A Complete Bibliography of Publications in the ICES Journal of Marine Science (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

09 April 2024 Version 1.22

Title word cross-reference

4 [SBD^{+}21]. 45 [Lov20]. ^{1}4 [LNW21, LNW21]. _{2} [BKR^{+}21, CCM^{+}21, CCJ^{+}21, EDC20, FDPD^{+}23, GS22, GSL22, Gal23, JWR^{+}21, LXBG21, Sta22]. F_{msy} [SBC^{+}21a, SBC^{+}21b].

-dimensional [SBD^{+}21]. -induced [CCM^{+}21]. -year [Lov20].

19 [BMM^{+}22, KGG^{+}23, PMTG^{+}23, SFW^{+}22]. 1940s [Hol21].

2020 [GGOP22]. 2030 [MKCA^{+}23].

3D [HK21].

62-year [LVT^{+}23].

abalone [AABM^{+}20, EVLK^{+}21, MDAMPE^{+}21]. abandon [MMN^{+}22]. able [SMK23]. above [PTD^{+}20]. abrupt [MAC21]. absolute [Mil24]. Abundance [KTK20, MDP^{+}21, BMB20, BAS^{+}21, DBT^{+}23, FFS^{+}22, FWC^{+}20, GKRB21, HJP22, HCB^{+}24, HCM24, KSHO23, KPS22, LVG^{+}20, LRGW23, LHL^{+}21, LMW^{+}22, PWCL21, PCCL21, SBW^{+}22, SWLW22, SACP^{+}21, SSF^{+}22, TFWE^{+}22, TCLB22, VSL22, VPB^{+}20, WBC^{+}21, WLR^{+}20, dlCRT^{+}22]. abundances [Bea22]. abundant [SAA^{+}22]. abused [PF21]. academic [BKF^{+}20, SCM^{+}23]. Acanthaster [DPD^{+}22a, DPD^{+}22b]. Acartia [VLUB^{+}22]. accelerates [KWS^{+}21]. accelerometers [PVGC^{+}22]. acceptance [TYF23]. access [vDHR^{+}22]. account [SBC^{+}21a, SBC^{+}21b, SSV^{+}22a]. Accounting [CBFS23, ZRW^{+}21, MACSR22, PDH^{+}21]. accounts [MTK^{+}21]. accuracy [ZYJZ23]. Accurate [MHKV20, NDE^{+}21]. achieve [BKF^{+}20, FJC^{+}20]. achieving [WZJ^{+}22]. acid [CLMB22]. acidification [AABM^{+}20, BKR^{+}21, DBGP22, DNW^{+}23, EVLK^{+}21, GPS^{+}21b, GPS^{+}21a, GGRB22, HMBDP21, LXBG21, MOC^{+}23, NKA^{+}21, PBC^{+}22, PBS^{+}23]. acids [MvdMS^{+}23]. acknowledgement [AGSS22]. Acoustic [BWJ^{+}20, HMWS22, PVGC^{+}22, AKL^{+}21, AKK^{+}21, BHQF20, BGH^{+}20, CSC^{+}23, Dom21, IOS21, Köp23, LDW23, LRHW21, MYW^{+}21, MTK^{+}21, POUH23, PMSK^{+}20, SOB23]. acoustically [CSC^{+}23]. across [AJC^{+}23, Ano23j, BPC^{+}20, BGR^{+}22b, CRLC23, CAE^{+}21, FND23, GHB^{+}23, HKZ^{+}23, HRB^{+}21, KEW^{+}21, MWS^{+}23, NAS^{+}20, OOB^{+}22, OHM^{+}23, OH23, SSS^{+}23, TBP20, VSS^{+}22, WRC^{+}21, YCF^{+}23]. action [DNW^{+}23, GT22, MKB^{+}23, MKCA^{+}23]. actions [GT22, WPP^{+}21]. active [BVS^{+}23, NBL^{+}21, SKA15, SKA22]. activities [BPC^{+}24, RHW20]. activity [Ano23i, BC23, CWC^{+}23, LQC^{+}22, MMMB^{+}23, PVGC^{+}22]. acutus [BBE^{+}20, GHB^{+}23]. adaptation [GBH^{+}24, HEHT^{+}22, RHO22, Woo22, WMB^{+}22, HBC^{+}22]. Adapting [SMSP^{+}23]. Adaption [NKA^{+}21]. adaptive [CdSSO^{+}21, MDAMPE^{+}21, NSJ^{+}22]. additive [NHL^{+}20]. address [BBB^{+}21]. addresses [OBK21]. Addressing [POUH23]. adoption [Jen23]. Adriatic [BFD^{+}23, CAE^{+}21]. Adult [GPS^{+}21b, GPS^{+}21a, AABM^{+}20, KDS^{+}22, SCP^{+}21]. advancement [Leg21]. advancements [GC21]. advances [AHP^{+}21, LWL^{+}22]. advantages [TR20]. Adventures [Day20]. advice [Ano23e, Hut22, JAB^{+}21, Kra23, TZE21, TNM^{+}23, WFG^{+}23]. Aegean [KLP^{+}24]. Aerial [CNWW22, BFFC20, GETvB21, MSWB22]. affect [AWN22, BHDD22, HRB^{+}21, HMWS22, LRHW21, MYMV20, Suu22]. affected [BVV^{+}20, Bra22, LLALF22]. affecting [KGV21, MPN20, MB21, THSZ21]. affective [JEPT23]. affects [CCM^{+}21]. affinities [HMBDP21]. Africa [BRG22, SFPP^{+}23, SHP^{+}21, dMBJ22]. African [BRG22, SHP^{+}21, SHP^{+}22]. after [BWBH21, BM20, CCJ^{+}21, HHL^{+}21, VO23, VNR^{+}23]. against [MOC^{+}23]. Age [BHWC^{+}20, GGA^{+}23, MDHB21, PHB^{+}20, WM20, DGM^{+}24, GSWBK20, HCH20, IEIT23, JBF^{+}20, NFN20, RGE^{+}23, TBH^{+}20, TMS23]. Age-related [GGA^{+}23]. Age-specific [WM20]. Age-structured [BHWC^{+}20, HCH20]. ageing [PHB^{+}20]. Aggregating [EHV^{+}21, MAI^{+}22, NSS^{+}23]. aggregation [EGMO^{+}20a, EGMO^{+}20b, LMvEM23, RDAOE20, SPBJ21]. aggregations [EDH^{+}22, GHÓ^{+}20, MBS^{+}23, SPJ23, SHW^{+}21]. agile [PJCS^{+}23a]. aging [JYK^{+}21]. agreement [CRV^{+}21]. Air [CSM^{+}23]. airgun [Ano23b, MSN^{+}23]. airguns [MMN^{+}22]. al [Bra22, BRG22, FDPD^{+}23, GS22, NBO^{+}20c, SFHP22, Sta22, SSV^{+}22a, JLA22]. al. [Ano23h, AAHT22, Gal23, HHN22, JG21, JHN^{+}21, SFHP22]. Alaska [Ano23j, CRLC23, WGCC^{+}22, ZB24]. albatross [KVS^{+}23]. albatrosses [BPC^{+}24, SWLW22]. albida [MGB^{+}23]. Alcyonacea [GICP^{+}22]. algae [BMN^{+}21, PBC^{+}22]. algorithm [AKL^{+}21, FDMK21b, FDMK21c]. Aligning [Woo22]. alleviate [CZ23]. allocation [AA21]. allometric [CZ23, UBD^{+}23]. allometrically [BCBJ^{+}20]. Allowable [PPMD22]. alone [MSW21]. along [BPC^{+}21, GSWM^{+}21, GSWM^{+}22, GPV^{+}22, HPS^{+}23, SSBV21, UVI^{+}21a, UVI^{+}21b]. altar [Abe20]. Altering [BPR^{+}23]. alternate [BMM^{+}21]. alternate-haul [BMM^{+}21]. Alternative [MGF^{+}21, BSH^{+}22, VCF^{+}20]. alternatives [MSWB22]. alters [CHZC22, LS20b]. ambition [Bir23]. Amblyraja [KKS^{+}20]. ameliorate [BBB^{+}22]. America [WBG^{+}22]. American [TSE^{+}21, Ano23i, BC23, GZSP22, HCCR20, MFP20, WRC^{+}21]. americanus [HCCR20, WRC^{+}21]. Ammodytes [BSW^{+}23, JLM^{+}23]. Among [dBABS^{+}22, JKF^{+}22, KBBGA22, MCH^{+}24, Pep23, PMM^{+}23, PTR22, PJP20, SL21, THSZ21, WPP^{+}21]. among-assessment [PJP20]. Among-individual [dBABS^{+}22]. amplicons [MKN^{+}21]. anadromous [KBBGA22, SBB^{+}22]. analogues [LTMG23]. analyses [MGO21, MVMdS21, ORGTO^{+}23, TDB^{+}21, TMG20]. analysis [BSBH22, BBP^{+}23, CASM23, CFB^{+}21, DSMM^{+}22, DRP^{+}23, EGMO^{+}20a, EGMO^{+}20b, FMF^{+}20, FSS^{+}23, HHNS21, KMK^{+}21b, KGG^{+}23, LMW^{+}22, MKN^{+}21, MMCS^{+}24, NFN20, NAB^{+}22, OdS21, PWCL21, PFP^{+}22, WPP^{+}21, WJB^{+}24, YY22, YMAH20]. analytical [KPS21, MCH^{+}24]. anchor [ON21]. anchored [MR20a]. anchovy [BPPKH20, SBPA^{+}21, SDGR^{+}20]. ancient [ADBC^{+}22]. angel [LPG^{+}20]. angler [MYAH21, ODHF^{+}23, SHG^{+}21, YMAH20]. anglers [SCCA20]. angling [SROR21]. Anguilla [Ano23i, BC23, DAB^{+}23, VBO^{+}21a, VBO^{+}21b]. anguillid [WJB^{+}24]. animal [WBWS21]. Annotating [BDM24]. annual [Ano22e, BSW22, DDL^{+}20, GGOP21, JRPHM^{+}22, KPM^{+}22, SH22, GGOP22]. Anoplopoma [Ano23j, CRLC23, ORGTO^{+}23]. antagonism [JAQ^{+}23]. Antarctic [WESWT21, BBE^{+}20, GT22, Man21, MTT^{+}23, SMP^{+}23, TFH^{+}21, TFWE^{+}22]. anthems [DeM22]. Anthropocene [PB21]. anthropogenic [BBB^{+}22, DFK^{+}22, PPV^{+}22, VKH^{+}20a, VKH^{+}20b]. anti [SWH^{+}23]. anti-trawl [SWH^{+}23]. apparent [ANB^{+}21, ANB^{+}22]. appealing [Leg21]. applicability [RUdPFRA^{+}21]. Application [AKL^{+}21, BGH^{+}20, FDMK21a, WBWS21, BHWC^{+}20, CWB^{+}23, Eay23, GPQ^{+}20, KSHO23, MVMdS21, PHB^{+}20, SdS20]. Applications [Ano20a, BXF23, MCH^{+}24, Ano20b, GHJ^{+}22, LSOD^{+}20, RBC^{+}23]. applied [BMR21, BP20, Gri20, PMO^{+}21a, PMO^{+}21b, QWT^{+}21, RDM^{+}21]. Applying [TB21]. approach [AGSS22, BGR^{+}22b, CCC^{+}21, CMG^{+}23, CFO^{+}22, FMBL23, GHÓ^{+}20, GGRB22, HPSF21, HLHC22, Kai20, KSHO23, KLF^{+}20a, KLF^{+}20b, KHHD20, LS20a, LvDM^{+}22, NBA^{+}23, PEWN20, PRD20, RDM^{+}21, dSRKC22, SRB^{+}23, SdS20, SOC^{+}20, SBA^{+}22, SBD^{+}21, Tho22, TFH^{+}21, WZJ^{+}22, ZMFS20, vDBF^{+}20]. approaches [BPKM21, BPKM22, FRSM22, GCG22, KOB^{+}22, MCH^{+}24, OTMN20, RCJ^{+}22, RKH^{+}21, Rin21, SBKO22, ZHL^{+}22]. apps [SHG^{+}21]. Aptenodytes [PRCM22]. Aquaculture [FNB^{+}20, AGSS22, CCC^{+}21, CFMB21, CFB^{+}21, FRSM22, FCF^{+}21, KF21, PB21, SDSMS21, SG21]. Aquaculture-driven [FNB^{+}20]. aquatic [BBL^{+}23, MCH^{+}24]. arabesque [YMH^{+}24]. Arabian [BHWC^{+}20]. Arabian/Persian [BHWC^{+}20]. archaeological [PAW22]. archipelago [HMOR22, MWJ^{+}24, YABM20, MAI^{+}22]. archival [KKS^{+}20]. Arctic [DSGG20, DHN^{+}21, DDL^{+}20, EVY^{+}21, HHA^{+}21, KTA^{+}23, LRHW21, MPH^{+}21, SBB^{+}22, VHT^{+}21]. Ardenna [LB23]. area [BFD^{+}23, CBB21, HKZ^{+}23, HAPW21, MPMGGAHR20, OHM^{+}23, VO23, ZMN^{+}21, dBABS^{+}22]. area-based [HAPW21]. Areas [BKF^{+}20, BWBH21, BAS^{+}21, BFJB21, CMG^{+}23, DA20, DPD^{+}22a, DPD^{+}22b, HMOR22, HHD^{+}23, JDJK22, SSBV21, SBW^{+}22, SPBJ21, SHK^{+}21, VPP^{+}23, WRC^{+}21, ZJH^{+}23, BMM^{+}22, RRR^{+}24, SCMT^{+}22]. argentea [PMSK^{+}20]. Argentina [TDC^{+}23]. argentinus [AWN22, CASM23]. arising [RHW20, TMS23]. armatus [MHJ^{+}20]. aromatic [LZZ^{+}21]. array [SSP^{+}23]. articles [Ano20a, Ano20c]. artificial [Ano20a, GPM21, KLF^{+}20a, KLF^{+}20b, SCW^{+}20, TMD^{+}20]. artisanal [AVC^{+}21, GBGA^{+}23, VPCM^{+}22]. asked [MBSR20]. aspera [WDPM23]. aspirations [ZMMF22]. assemblage [LTN21, SBW^{+}22]. assemblages [CRG^{+}23, DPD^{+}22a, DPD^{+}22b, HGS^{+}21, MHO^{+}20, NKA^{+}20, SCW^{+}20, WBB^{+}21]. assembled [EDC20]. assess [BGH^{+}20]. assessed [KLF^{+}20a, KLF^{+}20b]. Assessing [CBFS23, HvdMLS21, JDD^{+}21, RPB^{+}21, CMG^{+}23, CFO^{+}22, FMBL23, KNP21, NBA^{+}23, PCC23, SBM^{+}22]. assessment [AGSS22, Ano23e, AHP^{+}21, BDB^{+}21, BMR21, BNB21, CZS^{+}23, CMR^{+}20, CGH^{+}23, CMCF21, DA20, FSC^{+}23, GBGA^{+}23, JBJ^{+}22, KSK^{+}21, KBW22, LLLS20, LV23, MTCC21, MKS^{+}21, MCOÁ^{+}22, NHM^{+}21, PDL20, PGRP23, PCCL21, PWH23, PJP20, SBH^{+}21, SBKO22, SB21, SACP^{+}21, TRP^{+}21, TBM^{+}22, TB21, VBJ^{+}21, VNR^{+}23, WESWT21, WFG^{+}23, Woo22, ZJF^{+}21, dPMS^{+}22]. assessments [CFB^{+}23, CMS^{+}23, CFP20, FW22, FSS^{+}23, HLHC22, IEIT23, LVG^{+}20, MVMdS21, MKGC20, ÖHB^{+}23, PKT^{+}21, TPCM^{+}24]. assignment [BLM^{+}21]. assimilative [ANB^{+}21, ANB^{+}22]. associated [Ano24b, BSW^{+}23, KEW^{+}21, LH20, LAGC23, NSS^{+}23, RVR^{+}24, SCW^{+}20]. associations [SAM^{+}22, THSZ21, VRH^{+}23]. assumptions [DHB^{+}21, HMF^{+}21]. Asturias [RUdPFRA^{+}21]. Asymmetry [FRP^{+}23]. at-sea [SWLW22]. athwart [BMU^{+}23]. Atlantic [AMD^{+}23, BQBB^{+}21, DSF^{+}21, UPH^{+}21, dAdSM21, ADS^{+}23, ANA^{+}20, ADBC^{+}22, ANA^{+}23, Ano22d, Ano23b, Ano24c, AWN22, ASW^{+}23, BFFC20, BSBH22, BBP^{+}23, BJG^{+}21, BJPP^{+}20, BPR^{+}20, BPC^{+}20, BLM^{+}21, BSH^{+}22, BGR^{+}22b, CBT^{+}20, CRG^{+}23, CSH^{+}21, CRR^{+}22, CBB21, CGH^{+}23, DFK^{+}22, EMS^{+}21, FFH^{+}22, FNB^{+}20, GSWM^{+}21, GSWM^{+}22, GGLS22, GHÓ^{+}20, GPJ^{+}21, GHB^{+}23, GGW^{+}23, HMC^{+}20, HJP22, HMA^{+}23, HPS^{+}23, HBD^{+}20, HBD^{+}21, JG21, JSN^{+}22, JKF^{+}22, JHS^{+}21a, JHS^{+}21b, JLM^{+}23, KPM^{+}22, KEW^{+}21, KLSV21, LPG^{+}20, LAA^{+}21, LAGC23, LOÖK20, LVT^{+}23, LMW^{+}22, MTCC21, MRC^{+}21, MMH^{+}23, MMN^{+}22, MSN^{+}23, MYW^{+}21, MvdMS^{+}23, Moo23a, MKGC20, MPK^{+}22, NBO^{+}20a, NBO^{+}20b, NBL^{+}21, OOB^{+}22, OH23, OGMA20, PJCS^{+}23b, PBD^{+}21, PAW22, PE21, PFP^{+}22, PRRB^{+}20, QBF^{+}21, SLB^{+}20, SBM^{+}22, SdS20, SOC^{+}20, SG21, SCH^{+}21b, SBD^{+}21, SPBJ21, SPJ23, SBC^{+}21a]. Atlantic [SBC^{+}21b, SGM^{+}24, SHK^{+}21, TZE21, TSE^{+}21, VKH^{+}20a, VKH^{+}20b, VPP^{+}23, VBJ^{+}21, VET^{+}22, YCF^{+}23, ZJF^{+}21, ZJH^{+}23, Zho21]. atlanticus [DRP^{+}23]. Attracted [VCD^{+}20, CVJ^{+}20]. attraction [BPC^{+}24]. auction [ASV21]. audax [MGB^{+}23]. auratus [BBB^{+}23]. Australasian [SAM^{+}22]. Australia [WCCY20, BBB^{+}23, BMB20, CHZC22, PVHV20, PDH^{+}21]. Australian [Ano20b, BHdL^{+}23, FMPH24, LSOD^{+}20, TR20]. australis [TND^{+}20]. Austria [KTK20]. Automated [PMSK^{+}20, FMF^{+}20, QWT^{+}21, SMO22, SCN^{+}23]. Automatic [GPQ^{+}20, MBA^{+}20, SSS^{+}20, vEMvH^{+}21, YLC^{+}22]. autonomous [CSC^{+}23]. autumn [BØS^{+}21, HCCR20]. autumn-spawning [BØS^{+}21]. Auxiliary [MVMdS21]. Availability [SPBJ21, CVJ^{+}20, KJP^{+}22, PJCS^{+}23a, MTK^{+}21, ODT^{+}22, STS^{+}20]. avoidance [GLM^{+}23, Kra23, PM23]. avoiding [FCD^{+}23]. Azorean [FCD^{+}23]. Azores [PJCS^{+}23b].

BACI [VNR^{+}23, Met20]. backed [VCD^{+}20]. backscattering [KAK^{+}21]. bad [Kra23, MEOA23]. BAG [Met20]. bait [BR22]. Balanced [PBF^{+}21, BP20, SSF^{+}23]. Balancing [EGMO^{+}20a, EGMO^{+}20b, TMLH21]. Baltic [FKTM20, Ano22b, BW21, Bra20, Bra22, EDH^{+}21, GZW^{+}20, HWL^{+}22, HMOR22, JDJK22, MCD^{+}20, NBO^{+}20a, NBO^{+}20b, NBO^{+}20c, OOE^{+}22, RBF^{+}22, SPd^{+}22, SSV^{+}22a, SSV^{+}22b, VQN22]. ban [Bor21, CVJ^{+}20]. banana [PDH^{+}21]. bankruptcy [KMS20b]. Baranov [Sha21, Ken21]. barcode [BQBB^{+}21]. Barents [DBT^{+}23, SMLB^{+}22, YSL22, Ano23f, DSMM^{+}22, DDL^{+}20, DYS21, HvdMLS21, HHA^{+}21, HFT^{+}22, HHD^{+}23, KPS22, SMB^{+}21, SB21, ZMN^{+}21]. barotrauma [WPN^{+}21]. barriers [Suu22]. basal [JSR^{+}22]. based [ARHM21, ÁEPCL20, ASV21, BLT23, BCC^{+}24, Bir23, BFW^{+}21, BPKM21, BPKM22, BPPKH20, CZSR21, CZ23, CMR^{+}20, DGM^{+}21, DLL^{+}21, DHB^{+}21, FMBL23, GDI^{+}20, GT22, GJU^{+}23, GMT21, GFO^{+}20, HAPW21, HCH20, KMG22, KHHD20, LVG^{+}20, LZWH20, MD22, Mun20, MMMB^{+}23, NAB^{+}22, NSS^{+}23, OT22, ÖHB^{+}23, PLN20, PVHV20, PAW22, PMM^{+}23, RKH^{+}21, SSS^{+}20, TPCM^{+}24, TMD^{+}20, TLM^{+}20, TMG20, VDAT20, WOL^{+}21, Woo22, YARE23, YLC^{+}22, ZJF^{+}21, ZRW^{+}21, dPMS^{+}22, vDHR^{+}22]. baseline [SBB^{+}23]. baselines [BR21, DPH^{+}21, FSN^{+}22]. basin [HBMC20, JNT^{+}22, SBB^{+}23, Thi20]. basin-scale [Thi20]. Basque [IVU^{+}22]. bass [IOS21, MPN20, PRMS22, SCP^{+}21, TMLH21]. bassanus [dGW^{+}22]. bathypelagic [GSWM^{+}21, GSWM^{+}22]. Bay [JSR^{+}22, RRR^{+}24, RUdPFRA^{+}21, JBJ^{+}22, JJB^{+}22, BPPKH20, CCR^{+}24, IVU^{+}22, IOS21, Köp23, MKSM23, MMMB^{+}23, NAB^{+}22, PTR22, SBPA^{+}21, VLUB^{+}22]. Bayesian [KP24, PEWN20, PVHV20, dSRKC22, TB21, WPP^{+}21, ZMFS20, ZHL^{+}22]. be [ASV21, BR22, Bra20, GAP21, MR20a, NBO^{+}20c]. beach [MPQS20, PDP^{+}23]. beach-spawning [MPQS20]. beached [CNWW22]. beaches [MHO^{+}20, MGO^{+}23]. beaching [MAI^{+}22]. beam [BM20, BRvL^{+}23, BMU^{+}23, RDM^{+}21]. bearing [AKK^{+}21]. become [MR20b]. Becoming [Hol20, Sch21a]. Beebe [Dol20]. Before [VNR^{+}23, VO23]. Behavioral [SRM^{+}20]. Behaviour [PE21, Ano24c, BFFC20, BBE^{+}20, CKR^{+}21, FBF^{+}21, GGA^{+}23, KBBGA22, LLPdL22, LVL^{+}22, MYJ^{+}20, MYW^{+}21, NBL^{+}21, Pep23, SKH21, VRH^{+}23, WBWS21, dBABS^{+}22]. Behavioural [BFD^{+}23, MSN^{+}23, SBB^{+}22, MCH^{+}24, VPCM^{+}22, Ano23b]. behaviours [SRM^{+}20]. Being [SMK23, Leg21]. belief [PVHV20]. Belize [MKB^{+}23, SBB^{+}23]. below [PTD^{+}20]. beneficial [GPS^{+}21b, GPS^{+}21a]. Benefits [Cha21, Bac24, JJB^{+}22, MYAH21, dAPDDJ21]. Benguela [BCBJ^{+}20]. Benthic [CvdWC^{+}20, DBB^{+}20, BSW^{+}23, BM20, CRB21, GZW^{+}20, GGRB22, HCB^{+}24, JDD^{+}21, KLF^{+}20a, KLF^{+}20b, LBH^{+}21, PBC^{+}22, PASL21, SMO22, TLS21, vDBF^{+}20]. bentho [DMK^{+}21]. bentho-demersal [DMK^{+}21]. Bering [DFWR20, HYE^{+}22, IEIT23, ROH^{+}21, RBCD20, SCF^{+}21, VRH^{+}23, YFM^{+}20, YC20]. berrima [DGM^{+}24]. best [CFP20, FP20, JEPT23, PH21]. beta [dSRKC22]. better [HBFO20, Lov20, MKSM23, PRAR23, dlCRT^{+}22]. between [Ano23d, BCB^{+}22a, BØS^{+}21, BPP^{+}21, CRB21, CLMB22, FCF^{+}21, Gri20, HWS^{+}21, HAPW21, IVU^{+}22, KWEB21, LBW^{+}20, LOÖK20, MEOA23, MVNR^{+}21, MPMGGAHR20, NAS^{+}20, OTLH20, PBC^{+}22, PASL21, RTQW20, ZMMF22, lCRGCCA23]. Beverton [GC21, Ken21, Man21]. Beyond [SKH21, BFJB21, CMD^{+}23a, CRV^{+}21, HVDvD23]. bias [CFO^{+}22]. biases [LWH^{+}21]. Big [GCG22]. bigeye [LWZC23]. bigger [Bea22]. Bight [HMC^{+}20]. billfish [BHL^{+}23, LTK20]. bio [MYMV20, MMMB^{+}23]. bio-economic [MMMB^{+}23]. bio-irrigation [MYMV20]. bioacoustics [SOB23]. biochronology [VBO^{+}21b, VBO^{+}21a]. biodiversity [AJC^{+}23, BPKM21, BPKM22, CRV^{+}21, CvdWC^{+}20, CMD^{+}23b, EWD^{+}21, EWD^{+}22, HJS^{+}23, HAPW21, LS20b, VRM^{+}23, VPB^{+}20]. Bioeconomic [dAPDDJ21, BHWC^{+}20]. Bioenergetic [GGBC20, BBB^{+}21]. bioenergetics [BPPKH20]. biogenic [MSW21]. biogeochemical [GZW^{+}20]. biogeochemistry [MYMV20]. Biogeographic [BBB^{+}23, CWL^{+}20]. Biogeographical [GSWM^{+}21, GSWM^{+}22]. biogeography [BPKM21, BPKM22, KLSV21]. bioindicator [MOC^{+}23]. bioinspired [PMM^{+}23]. biological [AKK^{+}21, ASV21, CRV^{+}21, DFWR20, DMK^{+}21, GGRB22, Kai20, OTMN20, SSBV21, SBKO22, SGM22, SJX^{+}20, TMLH21]. biomass [CRB21, EWD^{+}21, EWD^{+}22, LPC^{+}23, MKS^{+}21, MKSM23, MBP^{+}20, NFN20, NDE^{+}21, PMO^{+}21a, PMO^{+}21b, STK^{+}20b, STK^{+}20a, SMJ^{+}21, STS^{+}20, SMSP^{+}23, SMP^{+}23, SBW^{+}22, UBD^{+}23, VPB^{+}20]. biophysical [BJPP^{+}20, CSM^{+}22]. Bird [TDC^{+}23]. Biscay [RUdPFRA^{+}21, BPPKH20, CCR^{+}24, IVU^{+}22, Köp23, MMMB^{+}23, SBPA^{+}21, VLUB^{+}22]. bite [RBH22]. black [SWLW22, SFPP^{+}23, VCD^{+}20, VSL22]. black-backed [VCD^{+}20]. black-browed [SWLW22]. blackfin [DRP^{+}23]. Bloch [TDB^{+}21]. blooms [FRP^{+}23, SJX^{+}20]. blue [AMD^{+}23, Bir23, DLLC^{+}22, FFT^{+}22, HMWS22, LVL^{+}22, MFB23, PMTG^{+}23, WSL^{+}23, MKB^{+}23]. bluefin [ADS^{+}23, ADBC^{+}22, ANA^{+}23, BFFC20, HBD^{+}20, HBD^{+}21, MKGC20, NFN20]. blueprint [PJCS^{+}23b]. blues [DeM22]. BMSY [SLRC21]. board [AMP^{+}23, CCR^{+}24]. body [BCB^{+}22a, DBT^{+}23, HRB^{+}21, KTY^{+}21, KDS^{+}22, LB23, MFV^{+}23, Pep23, RGE^{+}23]. bond [MFB23]. bones [ADBC^{+}22]. boosted [NHL^{+}20]. boreal [KMS^{+}20a]. borealis [HWS^{+}21]. Boreogadus [DDL^{+}20]. both [BHDD22]. Bottom [JBF^{+}20, BFJB21, BSH^{+}22, CCR^{+}24, DDP^{+}21, HAP^{+}21, KKS^{+}20, LTN21, LMH22, LAGC23, MTK^{+}21, MHMP^{+}23, PE21, RMF^{+}22, RHvD^{+}20, RDM^{+}21, SPJ23, SB21, TDC^{+}23, vDBF^{+}20]. bottom-associated [LAGC23]. bottom-set [BFJB21]. bottom-trawl [MHMP^{+}23]. bottom-up [LMH22]. bottoms [CRG^{+}23]. boundaries [BDB^{+}21, SCB^{+}23]. boundary [MMAE20, PASL21]. Bowdoin [NKA^{+}20]. box [KF21]. brachydactyla [RPVT^{+}21]. brachyuran [GPJ^{+}21]. Brander [SSV^{+}22a]. Brazil [LLALF22]. breaks [NAS^{+}20]. breeding [ADG^{+}20, GT22, KVS^{+}23, dGW^{+}22]. BRFSS [SRM^{+}20]. British [BPC^{+}21, GLM^{+}23]. broad [GHB^{+}23]. broad-scale [GHB^{+}23]. Broadband [YMH^{+}24, AKL^{+}21, KAK^{+}21]. broadcast [CCJ^{+}21]. browed [SWLW22]. brown [KBK^{+}22, KBBGA22, MDP^{+}21, SH22, SSH^{+}20]. BRUV [MSWB22]. bryozoan [MSW21]. bryozoan-dominated [MSW21]. Buccoo [PGRP23]. build [KOB^{+}22]. Building [PJCS^{+}23a, SCM^{+}23, MKCA^{+}23]. business [LKL^{+}21]. Butterworth [SHP^{+}22]. Bycatch [YSL22, AMC^{+}23, ASW^{+}23, Bar23, BJBB20, CCR^{+}24, CCH^{+}23, Eay23, FCD^{+}23, Jen23, JEPT23, PEWN20, PO24, SRBT23, Suu22, UBD^{+}23, WGL^{+}23, XCK^{+}23]. Bycatches [SB21].

C [LNW21]. C [?]inLarsson:2021:PPP. C-incubator [LNW21]. C.

[Ano22e, GGOP21, GGOP22]. Cadiz [dlCRT^{+}22]. cage [CFMB21]. Calanoides [BBE^{+}20]. Calanus [Ano22e, EDH^{+}22, GGOP21, GGOP22, JRDW22, JNT^{+}22, KPS22, SPBJ21, SPJ23, VLUB^{+}22]. calcareous [RPB^{+}21]. calcifying [PBC^{+}22]. calculate [OTMN20]. calculated [BCBJ^{+}20]. Caledonia [DPD^{+}22b, DPD^{+}22a]. Calibrating [BCC^{+}24]. California [CLS^{+}22, HKZ^{+}23, JRPHM^{+}22, KDS^{+}22, KJP^{+}22, MPQS20, PLN20, PCG^{+}22, SSF^{+}22, SSS^{+}23, SDGR^{+}20]. californianus [CLS^{+}22]. calls [LSR^{+}21]. Cambodia [SWH^{+}23]. camera [ARHM21, CNWW22, MHKV20, CRJ20]. campechanus [PHB^{+}20, PFP^{+}22]. can [ASV21, CFB^{+}21, MGF^{+}21, NDE^{+}21, PRAR23, Pro20]. Canada [HHN22, MAC21, PBD^{+}21, SFHP21, SFHP22]. cancellation [DLL^{+}21]. Cancer [GPJ^{+}21, MDP^{+}21]. Cantabrian [PPV^{+}22, VLUB^{+}22]. capacities [KOB^{+}22]. capacity [ANB^{+}21, ANB^{+}22, BVS^{+}23, JBJ^{+}22, KF21, NSJ^{+}22]. capelin [MAC21]. capital [CBFS23]. capture [BFD^{+}23, SHFK22, WJB^{+}21, WPN^{+}21]. carbo [KBJA20]. carbon [BBB^{+}21, FDPD^{+}23, Gal23, HMBDP21, JNT^{+}22, PBS^{+}23]. Carbonate [SSBV21, LBW^{+}20, TLS21]. Carcharhinus [FBF^{+}21, GSR^{+}20, WJB^{+}21]. care [Ano23h, AAHT22, JSLC21, JLA22]. career [AHC^{+}20, BB22, Fie21, JBDK23, Lov20, SFW^{+}22, SCM^{+}23, SCB^{+}23, SOO^{+}22, SOCN^{+}22]. Caribbean [PMTG^{+}23]. carneipes [LB23]. carnivorous [CRG^{+}23]. carrion [MGO^{+}23]. carry [EVLK^{+}21]. carry-over [EVLK^{+}21]. Carrying [JBJ^{+}22, CSC^{+}23, KF21]. cascades [Thi20]. case [BMM^{+}22, Bar23, FFR20a, FFR20b, FND23, GLM^{+}23, IEIT23, LLLS20, LvDM^{+}22, MKSM23, MFP20, MKGC20, OOE^{+}22, PRD20, SHFK22, SMSP^{+}23, SFPP^{+}23, VAS^{+}20, WOB^{+}23, dlCRT^{+}22]. Casting [Var21]. catadromy [DAB^{+}23]. catch [ARE^{+}22, ANA^{+}23, Ano23e, Ano23j, BR22, CMH^{+}23, CRLC23, CSH^{+}21, FDK20, FDMK21b, FDMK21c, GLM^{+}23, HCB^{+}24, Kra23, LOÖK20, MYAH21, NHL^{+}20, NSS^{+}23, NKS^{+}23, ODHF^{+}23, PCH^{+}24, PDL20, PM23, QWT^{+}21, RRJ22, dSRKC22, RHO22, TSD^{+}21, TZE21, Var21, WFG^{+}23, PPMD22, SFPP^{+}23]. catch-at-size [ANA^{+}23]. catch-rate [Ano23j, CRLC23]. Catch-regulated [PPMD22]. catchability [KP24, ZCX^{+}20]. catches [BPR^{+}23, FOC^{+}23, HHN22, PAW22, PE21, SFHP21, SFHP22, WBG^{+}22]. catchment [LVT^{+}23, SBB^{+}23]. categorizing [SFG22]. category [ASV21]. caught [BRvL^{+}23, TLM^{+}20]. causality [Gri20]. cause [BMB20, GAM^{+}22]. caused [NKA^{+}21]. causes [JMB^{+}22]. caution [HMF^{+}21]. Cayman [SHW^{+}21]. CCAMLR [GT22]. Celtic [DMK^{+}21, LAB^{+}22, SVML^{+}23, VLUB^{+}22]. central [AMC^{+}23, GHÓ^{+}20, MMAE20, WCL^{+}23, VQN22]. central/southern [MMAE20]. centred [PJCS^{+}23a]. centuries [HHN22, SFHP21, SFHP22]. century [BLM^{+}21, KTA^{+}23, MD21]. cephalopod [CASM23, MCC^{+}22]. cephalopods [AHP^{+}21, BBM^{+}22]. cetacean [BR21, CCR^{+}24, KSHO23]. cf [DPD^{+}22b]. cf. [DPD^{+}22a]. chain [LMH22]. chains [vdBDH21]. chalcogrammus [DLL^{+}21, IEIT23, LDW23]. challenge [TDB^{+}21, THSZ21]. Challenges [FCD^{+}23, HHO^{+}20, MFB23, AHP^{+}21, BD20, CFMB21, JMB^{+}22, KHSS21, Kos21, MMC^{+}23, MCX21, MPMGGAHR20, OMM^{+}23, PCC23, PM23, RCB^{+}21, SCB^{+}23]. chance [Fie21, PM20]. change [Abe20, ALLP^{+}22, BSBH22, BPV^{+}20, BMN^{+}21, BBL^{+}23, CHZC22, DA20, Eay23, FXG^{+}20, GKRB21, GZW^{+}20, HSS21a, HSS21b, HBC^{+}22, HEHT^{+}22, JEPT23, KWS^{+}21, MMC^{+}23, MKK^{+}24, NBO^{+}20a, NBO^{+}20b, NAB^{+}22, OOE^{+}22, PKAT20, ROH^{+}21, RSTL22, SH22, SWS^{+}21, SCF^{+}21, VKH^{+}20a, VKH^{+}20b, WMB^{+}22, YFM^{+}20]. changes [BGR^{+}22a, CT23, CdSSO^{+}21, LLLS20, LKL^{+}21, MAT^{+}23, MHJ^{+}20, NBA^{+}23, SBH^{+}21, STS^{+}20, SWLW22, TMG20, TLS21, WBWS21]. changing [DSR^{+}21, DFWR20, HMA^{+}23, KJP^{+}22, SVMBR22, TAPB22, TFH^{+}21, WBB^{+}23, WM20]. Channel [VKH^{+}20a, VKH^{+}20b, LV23, VLUB^{+}22, WOB^{+}23]. char [SBB^{+}22]. Characteristics [PAW22, RES^{+}23]. characterization [BØS^{+}21, LMvEM23]. characterize [BXF23]. Characterizing [MFP20, SSP^{+}23, TWC^{+}20, TSW^{+}21]. charter [LS20a]. Chasing [Moo23b]. chemical [DHN^{+}21, SASM22]. chemicals [SCG^{+}20]. chemistry [LBW^{+}20, SSBV21]. Chesapeake [IOS21, PTR22]. chick [PRCM22]. Chile [DASLRB22, MMAE20, Ano22c]. Chilean [SSBV21]. China [LZZ^{+}21, LQC^{+}22, STK^{+}20b, STK^{+}20a, yMlZlL^{+}21]. Chinese [XCK^{+}23]. chinned [BPC^{+}24]. Chinook [SAB^{+}20, SRBT23, YFM^{+}20]. Chionoecetes [Ano23f, DSMM^{+}22, HHA^{+}21, HFT^{+}22, HHD^{+}23, MB21, SMOD21]. choice [FSC^{+}23]. chondrichthyan [GRE^{+}20, SLB^{+}20]. chronologies [VDAT20]. Chrysophrys [BBB^{+}23]. chub [KTY^{+}21]. Chukchi [LRGW23]. ciguatera [AVC^{+}21]. Ciliate [SMS^{+}21]. circumpolar [MTT^{+}23]. citizen [MPQS20]. clam [Ano24b, dlTVLC^{+}21, RVR^{+}24]. class [POUH23, vEMvH^{+}21]. classic [WGI^{+}21]. classification [BWJ^{+}20, CKH^{+}21, LWL^{+}22, POUH23, PMSK^{+}20, ZYJZ23]. clausi [VLUB^{+}22]. clavata [MD22, SMSP^{+}23]. clearance [HMWS22]. Climate [BGR^{+}22a, BBL^{+}23, CHZC22, GBH^{+}24, HYE^{+}22, JMB^{+}22, KWS^{+}21, SCF^{+}21, Abe20, Ano23l, ALLP^{+}22, BJBB20, BB22, CBCM21, DA20, DSR^{+}21, EVY^{+}21, FMPH24, GA23, GKRB21, GZW^{+}20, HKZ^{+}23, HMA^{+}23, HBC^{+}22, HEHT^{+}22, LKL^{+}21, MAT^{+}23, MMC^{+}23, MWT^{+}21, MAE^{+}23, MKK^{+}24, MFP20, NAB^{+}22, OTMN20, OH23, ROH^{+}21, RSTL22, SH22, SMB^{+}21, SDSMS21, SG21, UVI^{+}21a, UVI^{+}21b, VLD^{+}21, WCLB23, WSL^{+}23, WM20, WGCC^{+}22, Zho21, vdWBL22, DHN^{+}21]. climate-driven [MAT^{+}23, MWT^{+}21]. climate-impact [BB22]. Climate-induced [BGR^{+}22a, Ano23l, LKL^{+}21, WCLB23]. climate-linked [OTMN20]. Climate-mediated [JMB^{+}22]. climate-ready [WGCC^{+}22]. climate-resilient [CBCM21]. Close [TCLB22]. Close-kin [TCLB22]. closed [HLHC22, ZJH^{+}23]. closed-loop [HLHC22]. Closing [Gri20]. closure [HMWS22]. closures [BRG22, CBB21, HMC^{+}20, SHP^{+}21]. Clupea [BBP^{+}23, KPM^{+}22, KÓP^{+}21, MDHB21, TNS^{+}21]. clustering [AKL^{+}21, OT22]. clusters [JLM^{+}23]. cluttered [vEMvH^{+}21]. CMSY [BMR21]. Cnidaria [GICP^{+}22]. CO [GS22, Gal23, Sta22, CMD^{+}23a, MAE^{+}23, PJCS^{+}23b, PAC^{+}23, RRJ22, DNW^{+}23, JJB^{+}22, JKF^{+}22, Köp23, PJCS^{+}23a, MEOA23, NKS^{+}23, BKR^{+}21, CCM^{+}21, CCJ^{+}21, FDPD^{+}23, GSL22, Ano23c]. co-cultivation [JJB^{+}22]. co-design [PJCS^{+}23a]. Co-designing [CMD^{+}23a, PJCS^{+}23b, PAC^{+}23, DNW^{+}23, MEOA23, Ano23c]. co-developed [NKS^{+}23]. co-developing [Köp23]. Co-existence [RRJ22]. co-occurring [JKF^{+}22]. Co-production [MAE^{+}23]. coal [Hol20, Hol21]. coast [CWB^{+}23, GPV^{+}22, HPS^{+}23, HHNS21, HBD^{+}20, IVU^{+}22, MD21, SSBV21, SBB^{+}23, VSS^{+}22, BPC^{+}21, BFW^{+}21, SAB^{+}20, STS^{+}20, WCCY20]. Coastal [LBW^{+}20, Ano22a, BCB^{+}22b, BBE^{+}20, CHZC22, DLA^{+}21, GGLS22, HMOR22, JBJ^{+}22, JJB^{+}22, LVG^{+}20, LMH22, LXBG21, LHN^{+}22, LBPC^{+}20, MWJ^{+}24, MSWB22, MYW^{+}21, NBL^{+}21, NBA^{+}23, ÖHB^{+}23, Pep23, PMM^{+}23, SSP^{+}23, SACP^{+}21, SHK^{+}21, VO23, VNBLZR^{+}23, WRC^{+}21, WBB^{+}23, WJB^{+}21, WBC^{+}22, dBABS^{+}22]. coastal-dependent [LVG^{+}20]. coastal-pelagic [CHZC22]. cod [Ano22b, Ano22d, Ano23b, BCB^{+}22b, BW21, Bra20, Bra22, CSH^{+}21, CRR^{+}22, DSGG20, DYS21, EVY^{+}21, FKTM20, GGLS22, HWL^{+}22, HPS^{+}23, HMNF22, HHN22, HHA^{+}21, Hut22, IHL21, JBF^{+}20, KTA^{+}23, KWEB21, LAGC23, LOÖK20, MCD^{+}20, MMN^{+}22, MSN^{+}23, MvdMS^{+}23, NBO^{+}20a, NBO^{+}20b, NBO^{+}20c, PAW22, PE21, RBF^{+}22, SPd^{+}22, SHFK22, SFHP21, SFHP22, SSP^{+}23, SOC^{+}20, SBW^{+}22, SBD^{+}21, SSV^{+}22a, SSV^{+}22b, SHK^{+}21, WGCC^{+}22, vLNC20]. coexisting [VNR^{+}23, dAPDDJ21]. Coherence [KBW22]. COI [BQBB^{+}21]. coincides [LHL^{+}21]. coioides [LHE20]. Cold [HSMK21]. Cold-water [HSMK21]. collaboration [MFA^{+}24, OdS21]. collaborative [ASW^{+}23, CCH^{+}23, DGM^{+}21, RRJ22, WBB^{+}23]. collapse [Cha21, MAC21]. Colleagues [Pop20]. collected [SK21]. collection [Ano23j, CRLC23]. collections [CRV^{+}21, GHB^{+}23]. colonization [GPJ^{+}21, TWC^{+}20, TSW^{+}21]. colonizing [TMD^{+}20]. Colour [BHQF20]. Columbia [GLM^{+}23]. column [MRC^{+}21, PASL21, RTQW20]. combination [BØS^{+}21, GZW^{+}20]. Combined [BSBH22, RFB^{+}22, ESK^{+}21, RMF^{+}22, SLRC21]. Combining [ARE^{+}22, MYW^{+}21, MSP^{+}20, NHL^{+}22, PIAF^{+}23, ODT^{+}22]. come [Mur21]. Comment [BRG22, HHN22, JG21, SFHP22, Sta22, Ano23h, AAHT22, Bra22, FDPD^{+}23, GS22, Gal23, Har21, JLA22, NBO^{+}20c, SSV^{+}22a]. Commercial [RHW20, ARE^{+}22, ÁEPCL20, Ano24b, BGR^{+}22a, DGM^{+}24, QMMC^{+}23, RVR^{+}24, SSH^{+}20, SWR^{+}22, TMLH21, WBC^{+}21, dCKS^{+}22]. commercially [GA23, RRJ22, TAPB22]. common [JAQ^{+}23, KMK^{+}21b, SMS^{+}22, SAM^{+}22, Ear21]. communication [Hut22, Moo23b, RDAOE20]. communities [BKR^{+}21, HHNS21, IVU^{+}22, ÖHB^{+}23, PKAT20, SDW^{+}21, SRM^{+}20, TMD^{+}20, ZMN^{+}21, vDBF^{+}20].

Community [MMAE20, BKF^{+}20, DMK^{+}21, EDC20, JNT^{+}22, KMK^{+}21b, KLF^{+}20a, KLF^{+}20b, OBJ^{+}22, VO23, VPB^{+}20, WBC^{+}22]. Comparative [BMR21, LL23, MKN^{+}21, SL21, BSBH22, JDD^{+}21, TRP^{+}21]. compared [HAC^{+}23, LNW21, VO23]. Comparing [CMD^{+}23b, LWH^{+}21, RCJ^{+}22, TBM^{+}22]. Comparison [OTMN20, Ano20b, BCBJ^{+}20, EMS^{+}21, HJP22, LSOD^{+}20, NHL^{+}20, PASL21, PVGC^{+}22, WRC^{+}21, ZHL^{+}22, vEMvH^{+}21]. Comparisons [HCH20]. compatibility [dBSQK21]. compensate [KVS^{+}23]. competition [KLP^{+}24, KWEB21]. compiled [VET^{+}22]. complement [MKN^{+}21]. complementary [FOC^{+}23]. complements [VRM^{+}23]. complex [BBP^{+}23, FOC^{+}23, MWJ^{+}24, SWS^{+}21, TFH^{+}21, VBO^{+}21a, VBO^{+}21b]. complexes [OT22]. complexities [OP22]. complexity [KF21]. compliance [BKF^{+}20, BMM^{+}22, MYJ^{+}20, SCH^{+}21b]. compliant [PKT^{+}21]. components [CZS^{+}23]. composition [AHK23, FFS^{+}22, GHÓ^{+}20, KLF^{+}20a, KLF^{+}20b, LTN21, LRHW21, LS20b, PASL21, TMS23]. compositions [MKN^{+}21]. comprehensive [Eay23]. compression [IOS21]. compulsory [BSH^{+}22]. conceal [RGE^{+}23]. conceived [GC21]. concepts [MMC^{+}23]. conceptual [DGM^{+}21, RKH^{+}21]. conceptualizations [CFB^{+}23]. conchilega [MYMV20]. Concurrent [TR20, ASV21]. condition [BCB^{+}22a, CWC^{+}23, DPH^{+}21, KTY^{+}21, LB23, LAGC23, LOÖK20, MFV^{+}23, UPH^{+}21]. conditions [Ano22b, BHL^{+}23, DSDA22, DLLC^{+}22, KVS^{+}23, LXBG21, MSVSA22, ODT^{+}22, PRCM22, SMO22, SRM^{+}20, SSV^{+}22b, vdWBL22]. conduct [WPM^{+}23]. conference [JBDK23]. confidence [NHM^{+}21]. confirms [BLM^{+}21]. conflict [Ano22a, DLA^{+}21]. Conflicting [Mun20, PWCL21, PCCL21]. Congruence [NAS^{+}20]. congruent [BBB^{+}23]. connecting [DHN^{+}21]. connection [MPMGGAHR20]. connectivity [BJPP^{+}20, BQBB^{+}21, CSM^{+}22, CASM23, CRD^{+}22, GHB^{+}23, HJ22, JAQ^{+}23, MMH^{+}23, MHO^{+}20, NSJ^{+}22, SGM22]. connectivity-driven [CRD^{+}22]. Consequences [HMF^{+}21, BDB^{+}21, CMCF21, FSC^{+}23, HS20, TZE21]. Conservation [BKF^{+}20, BMM^{+}22, RST^{+}21, AVC^{+}21, Ano23e, BBB^{+}22, BR21, BBL^{+}23, EDH^{+}21, EGMO^{+}20a, EGMO^{+}20b, GCG22, KMS^{+}20a, LPG^{+}20, Man21, MWS^{+}23, PJCS^{+}23b, SLB^{+}20, SWH^{+}23, WPN^{+}21, WFG^{+}23, WJB^{+}24]. consider [SCMT^{+}22]. considerations [Bac24]. considering [BD20, VKH^{+}20a, VKH^{+}20b]. Consistent [OHM^{+}23]. constraints [TNS^{+}21]. constructing [CFP20]. construction [WBG^{+}22]. consumer [CPD^{+}21]. consumption [Ano22a, Ano23f, DLA^{+}21, FCF^{+}21, HFT^{+}22, MGO^{+}23, SMLB^{+}22]. contamination [AMP^{+}23]. contemporary [KDS23, MCD^{+}20]. content [KPM^{+}22, Thi20]. contents [NDE^{+}21]. context [MAT^{+}23, OP22]. contexts [KOB^{+}22]. continental [Ano20b, LSOD^{+}20, MFP20]. contingent [VPP^{+}23]. continuous [OHM^{+}23, HJP22, OBJ^{+}22]. continuously [TDB^{+}21]. contractions [GA23]. Contrasting [LV23, NKA^{+}20, YC20, RBG^{+}20]. Contrastive [YLC^{+}22]. contrasts [BJG^{+}21]. contribute [BCP^{+}21, Pep23]. contributing [Leg21]. Contribution [PY22, BOLPHL21, SOO^{+}22]. contributions [Dol20, Ken21, LS20a, NFN20]. control [FW22, GDI^{+}20, GKRB21, MPN20, SLRC21, VNR^{+}23, ZB24]. controlling [SMB^{+}21, UBD^{+}23]. controls [VKH^{+}20a, VKH^{+}20b]. conventional [CMD^{+}23b]. convolutional [BWJ^{+}20, LTK20, TK20, ZYJZ23]. cool [KMS^{+}20a]. cooler [LHL^{+}21]. cooperation [GLM^{+}23]. Copepod [PTR22, HHNS21, IVU^{+}22, MPN20]. copepods [CP20, CLMB22, SPJ23, VLUB^{+}22]. Coping [FWA^{+}21]. coral [Abe20, DPD^{+}22a, DPD^{+}22b, LZZ^{+}21, MAT^{+}23, Rin21, WOL^{+}21, YABM20]. coralline [BMN^{+}21]. corallivorous [DPD^{+}22a, DPD^{+}22b]. Cormorant [HMOR22, KBJA20]. cormorants [Ano22a, DLA^{+}21]. Correction [Ano22a, Ano22b, Ano22c, Ano22d, Ano23b, Ano23d, Ano23c, Ano23l, Ano23a, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano24a, Ano24b, Ano24c, BMU^{+}23]. correlates [Ano23k, GBMM^{+}23]. Correlation [LOÖK20, lCRGCCA23, LIR^{+}21, Ano23d]. Corrigendum [ANB^{+}22, DPD^{+}22a, EWD^{+}22, KH21, PMO^{+}21a, SBC^{+}21a, UVI^{+}21a, VBO^{+}21a]. Corroborating [TSD^{+}21]. corrugata [MDAMPE^{+}21]. Corsican [GPV^{+}22]. Coryphaena [MCOÁ^{+}22]. costs [WPB20]. could [Moo23a]. Council [CFB^{+}23]. count [ARHM21]. Counterintuitive [NBL^{+}21]. Counting [SOO^{+}22, TK20]. Countries [RES^{+}23]. Coupled [STS^{+}20]. Coupling [VNBLZR^{+}23]. covariates [Mun20, TB21, dPMS^{+}22]. coverage [BAS^{+}21, VCF^{+}20, YARE23]. COVID [BMM^{+}22, KGG^{+}23, PMTG^{+}23, SFW^{+}22]. COVID-19 [BMM^{+}22, KGG^{+}23, SFW^{+}22]. crab [Ano23f, CZSR21, CBB21, DSMM^{+}22, DEL21, DSF^{+}21, GPJ^{+}21, HHA^{+}21, HFT^{+}22, HHD^{+}23, MDP^{+}21, MB21, NKS^{+}23, OTLH20, RPVT^{+}21, SMOD21, SB21, SCF^{+}21, WDPM23, ZMN^{+}21]. crabs [BPR^{+}23]. Cracking [PM23]. Crangon [SH22, SSH^{+}20]. Crassostrea [CCM^{+}21, DBGP22, JWR^{+}21, KWS^{+}21, NKA^{+}21]. credibility [dBSQK21]. criteria [MTCC21]. critical [KHSS21, KPS21]. critically [LPG^{+}20]. critique [JHN^{+}21]. croaker [LLLS20]. cross [AMP^{+}23, Ano20b, LSOD^{+}20]. cross-contamination [AMP^{+}23]. Crossing [SCB^{+}23]. crown [WDPM23]. crown-of-thorns [WDPM23]. Crustacea [KLSV21]. crustacean [DEL21, GZW^{+}20, MHJ^{+}20, VSS^{+}22, dLP21]. crustaceans [Ste21]. Crustose [BMN^{+}21]. cryptic [WDPM23]. ctenophore [VSL22]. cuckoo [MD22]. cues [JSN^{+}22]. cultivated [JJB^{+}22]. cultivation [ESK^{+}21, JBJ^{+}22, JJB^{+}22, TRP^{+}21]. cumulative [NBA^{+}23]. curiosity [MR20b]. Current [JRPHM^{+}22, KDS^{+}22, PCG^{+}22, SAA^{+}22, BJG^{+}21, BD20, GICP^{+}22, HMK^{+}20, PRCM22, BCBJ^{+}20, KJP^{+}22, SSF^{+}22, SSS^{+}23, SDGR^{+}20]. curves [Mil24]. Cuttlefish [BBB^{+}22, GPM21]. cycle [BGR^{+}22b, TRP^{+}21]. cygnus [LLPdL22]. cyprinid [PMSK^{+}20]. daily [Hol20, MYAH21, WGI^{+}21, Zho21]. Danish [ON21, SHFK22]. Dannevig [SRMS21]. Danube [KTK20]. Data [SCCA20, AKL^{+}21, ARE^{+}22, Ano23g, Ano23j, BSW22, BMR21, BWJ^{+}20, CRLC23, CKH^{+}21, CMR^{+}20, ECBJ22, FDK20, FDMK21a, FDMK21b, FDMK21c, FDMK22, FODD^{+}22, GCG22, GCT^{+}23, HK21, Har21, HJS^{+}23, HCB^{+}24, HCM24, KSHO23, KP24, PJCS^{+}23a, LWH^{+}21, LMW^{+}22, MGF^{+}21, MBA^{+}20, MHES20, MGB^{+}23, MCD^{+}20, MMCS^{+}24, NHL^{+}20, ODHF^{+}23, PIAF^{+}23, PEWN20, PCH^{+}24, PWH23, Pro20, PMSK^{+}20, QWT^{+}21, RCJ^{+}22, SRB^{+}23, SSH^{+}20, SBC^{+}21a, SBC^{+}21b, SHW^{+}21, TDB^{+}21, TBH^{+}20, TMS23, WOB^{+}23, YSL22, ZMFS20]. data-driven [MHES20]. data-limited [CMR^{+}20, FDK20, FDMK21a, FDMK21b, FDMK21c, FDMK22, Har21, KSHO23, RCJ^{+}22, WOB^{+}23]. data-moderate [BMR21]. data-poor [BMR21, ECBJ22, MGF^{+}21, PWH23]. data-rich [BMR21]. dataset [LSD^{+}20, VET^{+}22]. Davis [LBH^{+}21]. day [BMM^{+}21, RCB^{+}21]. dead [dCKS^{+}22]. debate [BP20]. debt [PRAR23]. Decadal [RES^{+}23, Ano22e, GGOP21, GGOP22, MTTG^{+}23, SWLW22]. decadal-scale [MTTG^{+}23]. Decade [PY22, VBO^{+}21a, VBO^{+}21b, MKCA^{+}23, SCM^{+}23]. Decades [MWS^{+}23, CdSSO^{+}21, PFH^{+}22, WRC^{+}21, dMBJ22]. Decapoda [GPJ^{+}21]. decapods [Bac24]. deciding [FSN^{+}22]. decision [WPP^{+}21]. decline [GGBC20, LB23, MD21, RBF^{+}22, SSV^{+}22a, WBC^{+}22]. declined [LPC^{+}23]. declines [GZSP22, LVT^{+}23, RHO22]. decommissioning [BD20, FP20, FJC^{+}20]. decorator [WDPM23]. Decrease [BMM^{+}22]. Deep [FMF^{+}20, LBH^{+}21, QWT^{+}21, ARHM21, ÁEPCL20, Ano23h, AAHT22, Bar22, BVS^{+}23, BDM24, BWJ^{+}20, CMD^{+}23b, Dol20, FCD^{+}23, GPQ^{+}20, GICP^{+}22, GHJ^{+}22, JSLC21, JLA22, PJCS^{+}23a, LCG^{+}22, LSD^{+}20, LTK20, MHA^{+}21, PJCS^{+}23b, POUH23, SSS^{+}20, SCN^{+}23, TK20, VKH^{+}20a, VKH^{+}20b, VNR^{+}23, WPN^{+}21, vEMvH^{+}21, vDHR^{+}22]. deep-dwelling [WPN^{+}21]. deep-ocean [LCG^{+}22, PJCS^{+}23b]. Deep-sea [LBH^{+}21, BVS^{+}23, Dol20, GICP^{+}22, VKH^{+}20a, VKH^{+}20b, vDHR^{+}22]. deep-water [FCD^{+}23, VNR^{+}23]. deepwater [GRE^{+}20, MSVSA22]. defence [MOC^{+}23]. defines [MRC^{+}21]. Defining [WOB^{+}23, MBS^{+}23]. delay [MAC21]. delimited [CMG^{+}23]. delineated [BBB^{+}23]. delphinid [KSHO23]. delta [BDL^{+}22]. Demersal [WBG^{+}22, ALLP^{+}22, BCB^{+}22a, Dom21, DMK^{+}21, GRE^{+}20, HMNF22, IHL21, KGV21, LTN21, MWT^{+}21, PPV^{+}22, PRRB^{+}20, VDAT20, VPB^{+}20]. demographic [Ano23k, GBMM^{+}23]. demography [GSR^{+}20, OH23, QCM^{+}20]. Density [ZEM21, BCB^{+}22a, KTY^{+}21, LLPdL22, MHJ^{+}20, Mun20, SBC^{+}21a, SBC^{+}21b, SHE^{+}22, Tho22, VDAT20, WESWT21, dLP21]. density-dependent [Mun20, SHE^{+}22, VDAT20, dLP21]. Density-independent [ZEM21]. Dentex [GPV^{+}22]. dependence [KTY^{+}21, MHJ^{+}20, SBC^{+}21a, SBC^{+}21b, Tho22]. dependent [Ano23j, BPR^{+}20, CRLC23, CRB21, CCJ^{+}21, ECBJ22, LVG^{+}20, Mun20, SHFK22, SBKO22, SHE^{+}22, VRH^{+}23, VDAT20, dLP21]. depends [GBH^{+}24]. depleted [SLRC21]. depletion [CZS^{+}23, MCOÁ^{+}22, RAH^{+}22, RUdPFRA^{+}21]. deposition [CCC^{+}21, GPM21]. depredation [BR22, EMS^{+}21]. depth [Ano23i, BC23, FKTM20, KPS21, KPS22]. derby [FFR20a, FFR20b]. derived [SBH^{+}21, SSH^{+}20]. desalination [WBB^{+}21]. describes [AJC^{+}23]. Descriptor [PKT^{+}21]. descriptors [Dom21]. design [BMM^{+}21, PJCS^{+}23a, OBK21, SMSP^{+}23, YARE23]. design-based [YARE23]. Designing [LCG^{+}22, CMD^{+}23a, DNW^{+}23, MEOA23, MFB23, MFA^{+}24, PJCS^{+}23b, PAC^{+}23, Ano23c]. designs [Met20, Mil24, PMM^{+}23, TMS23]. desirable [FSN^{+}22]. Destructive [DPD^{+}22b, DPD^{+}22a]. detect [MVMdS21]. detected [NFN20]. Detecting [PKAT20, TK20, SAA^{+}22]. Detection [CSC^{+}23, CZSR21, CNWW22, LSD^{+}20, MBA^{+}20, QWT^{+}21, SSS^{+}20, SMO22, SCN^{+}23]. determination [DLLC^{+}22, JBF^{+}20, vDHR^{+}22]. determine [TB21]. determined [PH21]. determines [ANB^{+}21, ANB^{+}22, FWA^{+}21]. Determining [JDJK22, NHL^{+}20, PVHV20]. Deterrents [SCH^{+}21b, RBH22]. detrimental [WBB^{+}21]. developed [NKS^{+}23, PCG^{+}22]. Developing [CSM^{+}23, LSD^{+}20, MFB23, Bir23, GFO^{+}20, Köp23, MKB^{+}23]. Development [CSH^{+}21, SCW^{+}20, Tho22, Ano24c, BVS^{+}23, CCM^{+}21, CCH^{+}23, Dom21, GGBC20, KHHD20, MD21, MR20a, MPN20, MPK^{+}22, RRJ22, SBH^{+}21, SBM^{+}22, ZRW^{+}21, PY22]. developments [MCX21]. Device [EHV^{+}21, MAI^{+}22, NSS^{+}23]. Device-associated [NSS^{+}23]. devices [BSH^{+}22, CSH^{+}21, GPM21, Jen23, TYF23]. Dexter [Gal23]. dFAD [MAI^{+}22]. diagnostic [MVMdS21]. diamond [Ano24a, SHK^{+}23]. Dicentrarchus [SCP^{+}21, dPHG^{+}23]. die-off [JSR^{+}22]. Diel [BSW^{+}23, AMD^{+}23]. Diet [ZMN^{+}21, BBM^{+}22, DBT^{+}23, KÓP^{+}21, LOÖK20, SSS^{+}23]. diets [BPR^{+}20, MPA^{+}23]. differ [PBC^{+}22, RPVT^{+}21, VSS^{+}22]. Differences [RPVT^{+}21, ALLP^{+}22, CAE^{+}21, GGA^{+}23, JKF^{+}22, MCD^{+}20, Pep23, SBD^{+}21, Thi20]. Different [OTLH20, RHvD^{+}20, AKL^{+}21, AMP^{+}23, HHO^{+}20, HS20, PBS^{+}23, PVGC^{+}22]. Differential [YFM^{+}20, RHvD^{+}20]. differentiation [HWS^{+}21, JAQ^{+}23]. differently [GVD^{+}23]. differing [KOB^{+}22, MHO^{+}20]. diffusion [Jen23]. digital [PMO^{+}21a, PMO^{+}21b]. digitata [KMS^{+}20a]. dilemmas [QMMC^{+}23]. dimensional [BGH^{+}20, SBD^{+}21]. dimensionality [BDB^{+}21]. dimensions [MTTG^{+}23, PGRP23, SRR21]. Diplodus [dBABS^{+}22]. directed [OGMA20]. Direction [SMOD21]. directional [NBL^{+}21]. directions [MCX21, VSS^{+}23, WJ22]. Directive [MTCC21, PKT^{+}21]. discard [BGH^{+}20, Bor21, CVJ^{+}20, KKS^{+}20, MHMP^{+}23, TZE21, vEMvH^{+}21]. discarded [FAV^{+}20, SCN^{+}23]. discarding [HMNF22, MV23]. Discards [BCB^{+}22b, FMF^{+}20, KVS^{+}23, SFPP^{+}23, WCCY20]. disciplinary [SCB^{+}23]. disclose [VDAT20]. discovered [GEBS^{+}23, LVT^{+}23]. Discoveries [Kos21, Mon22]. discrimination [CZ23]. disease [NKA^{+}21].

Disentangling [QCM^{+}20]. dispersal [HHD^{+}23, MWJ^{+}24, NAS^{+}20, PMS^{+}20, SBH^{+}21, SGM22]. display [BMN^{+}21]. dissolution [SASM22]. dissolved [PBS^{+}23, PTD^{+}20]. Dissostichus [EMS^{+}21]. distance [BPC^{+}21, BLM^{+}21, JYK^{+}21]. distances [RDAOE20]. distant [XCK^{+}23]. distinct [GVD^{+}23, JLM^{+}23, MHO^{+}20, SSS^{+}23]. distinguish [VKH^{+}20a, VKH^{+}20b]. distinguishing [MGB^{+}23]. distortion [BMU^{+}23]. distress [RBF^{+}22]. distributed [TDB^{+}21]. Distribution [SMP^{+}23, AMC^{+}23, ARE^{+}22, Ano22c, Ano23i, BLT23, BHL^{+}23, BC23, BHDD22, CCS^{+}24, CBB21, DASLRB22, DBT^{+}23, DYS21, EVY^{+}21, FNB^{+}20, GZW^{+}20, HBFO20, HCM24, IEIT23, KTK20, KPS22, LWZC23, MDP^{+}21, MTK^{+}21, MSP^{+}20, OT22, PIAF^{+}23, RMF^{+}22, RRJ22, ROH^{+}21, SSH^{+}20, STS^{+}20, SWLW22, SSF^{+}22, SCF^{+}21, Tho22, TFWE^{+}22, TRD^{+}22, WBA^{+}22, WESWT21, WM20, WDPM23, YC20, ZMN^{+}21, ZCX^{+}20, ZHL^{+}22, dlCRT^{+}22]. distributions [GZW^{+}20]. disturbance [HMWS22, SKA15, SKA22]. Diurnally [LXBG21]. divergence [KEW^{+}21]. diverse [LHN^{+}22, MFA^{+}24]. diversification [KLSV21, MPMGGAHR20]. diversities [MKN^{+}21]. diversity [AMC^{+}23, BQBB^{+}21, CPD^{+}21, CWL^{+}20, DPMZ21, MHA^{+}21, MGO^{+}23, SACP^{+}21, TRD^{+}22]. DNA [AJC^{+}23, AMC^{+}23, AMP^{+}23, HHNS21, MKN^{+}21, SAA^{+}22, VRM^{+}23]. DNA/RNA [HHNS21]. do [Ano23h, AAHT22, JLA22, MMN^{+}22, MSW21, TM20]. document [WBWS21]. documents [GMT21]. does [AWN22, DSDA22, HMWS22, KVS^{+}23, RST^{+}21]. doing [DGM^{+}21, Hol21]. dolphin [GHB^{+}23, SAM^{+}22]. dolphinfish [MCOÁ^{+}22]. dolphins [GAM^{+}22, Jen23, VPCM^{+}22]. dominant [KDS^{+}22]. dominated [MSW21]. domoic [CLMB22]. don’t [JSLC21, dCKS^{+}22]. dormancy [BBE^{+}20]. Doryteuthis [AWN22, SSF^{+}22]. down [KBJA20]. down-migrating [KBJA20]. downs [Bea22, Ste21]. draining [HFP^{+}23]. drastically [LPC^{+}23]. drawbacks [Bac24]. dredge [MDP^{+}21]. Dredging [MYMV20]. Dredging-induced [MYMV20]. drifting [EHV^{+}21, NSS^{+}23, WGKG21, MAI^{+}22]. drive [STS^{+}20, SSF^{+}22]. driven [BHL^{+}23, CRD^{+}22, FNB^{+}20, KJP^{+}22, LRGW23, LPC^{+}23, MHES20, MAT^{+}23, MWT^{+}21, PPV^{+}22]. driver [GZW^{+}20]. Drivers [HSS21a, JRDW22, RDAOE20, VSS^{+}22, BPPC^{+}20, CP20, DFK^{+}22, GPV^{+}22, GZSP22, HEHT^{+}22, Jen23, KBW22, LPA^{+}22, LAGC23, MAC21, PVHV20, RGE^{+}23, SJX^{+}20, SDR^{+}21, THSZ21, VKH^{+}20a, VKH^{+}20b, WBC^{+}22, ZJF^{+}21, HSS21b]. drives [LMH22]. drones [LMvEM23]. dubius [JLM^{+}23]. due [Bra20, NBO^{+}20c, WGKG21]. Dungeness [NKS^{+}23]. during [Ano24c, BMM^{+}22, BCB^{+}22a, ESK^{+}21, EDH^{+}22, GGLS22, GT22, JG21, JHS^{+}21a, JHS^{+}21b, JNT^{+}22, MPN20, NBL^{+}21, RAH^{+}22, SMP^{+}23, VSL22, WML^{+}21]. Dutch [JBJ^{+}22, JJB^{+}22, RDM^{+}21]. dwelling [WPN^{+}21]. Dynamic [PWCL21, BLT23, MBP^{+}20, MBSR20, OT22, SVMBR22, SVML^{+}23, WBA^{+}22]. Dynamics [HBMC20, BBP^{+}23, BPPKH20, GPJ^{+}21, HHL^{+}21, LL23, MKS^{+}21, MKSM23, MBSR20, MWS^{+}23, OTMN20, ON21, QCM^{+}20, SSBV21, SMJ^{+}21, SSF^{+}22, VO23].

Early [HCCR20, AHC^{+}20, BB22, JBDK23, KGG^{+}23, LHE20, PAW22, SFW^{+}22, SCM^{+}23, SCB^{+}23, SOO^{+}22, SOCN^{+}22, ZL22, ZEM21]. early-career [SFW^{+}22]. East [CGH^{+}23, FMPH24, WCCY20, STK^{+}20b, STK^{+}20a, HHA^{+}21]. Eastern [ADG^{+}20, ADBC^{+}22, Bra22, CHZC22, CCM^{+}21, GAM^{+}22, HHN22, MMAE20, MVMdS21, NBO^{+}20a, NBO^{+}20b, PBD^{+}21, SFHP22, SSV^{+}22a, YFM^{+}20, YC20, Ano22b, BHL^{+}23, LPG^{+}20, LV23, MWJ^{+}24, OHM^{+}23, ROH^{+}21, RBCD20, SFHP21, SSV^{+}22b, WML^{+}21]. easternmost [TRD^{+}22]. eat [TM20]. echograms [BHQF20]. echosounder [BWJ^{+}20, CKH^{+}21, HK21, WBWS21]. echosounders [LRGW23, WGKG21]. Ecoacoustic [MCX21]. Ecological [FP20, GBGA^{+}23, GRE^{+}20, Rin21, SAB^{+}20, Ano22b, BNV^{+}21, CP20, CT23, DPMZ21, FRSM22, FXG^{+}20, GCT^{+}21, KOB^{+}22, KDS23, MMC^{+}23, MTTG^{+}23, NAB^{+}22, ÖHB^{+}23, RBCD20, SROR21, SSV^{+}22b, TB21, VNBLZR^{+}23, VQN22, WMB^{+}22]. ecologically [BD20]. ecology [BPKM21, BPKM22, FJC^{+}20, GHJ^{+}22, GCG22, HHL^{+}21, NHL^{+}22, RBC^{+}23, TND^{+}20]. Economic [JYK^{+}21, MD21, LS20a, MMMB^{+}23, PCH^{+}24, SBM^{+}22, TMLH21, VQN22, WPB20]. economics [NHL^{+}22, TNM^{+}23]. economy [PMTG^{+}23]. Ecopath [RCJ^{+}22]. ecoregion [SVML^{+}23]. ecoregions [LAB^{+}22]. Ecosim [RCJ^{+}22]. Ecosystem [DPH^{+}21, WOL^{+}21, Ano23i, BLT23, BCC^{+}24, BPV^{+}20, BFW^{+}21, BC23, CFB^{+}23, DGM^{+}21, FMBL23, FSS^{+}23, GZW^{+}20, GMT21, GFO^{+}20, HVDvD23, HYE^{+}22, HSMK21, KPS20, LFH^{+}21b, LFH^{+}22, LL23, Lin21, MTT^{+}23, NAB^{+}22, NBA^{+}23, OOE^{+}22, OTL^{+}23, Pep23, PRD20, PPV^{+}22, RCJ^{+}22, RKH^{+}21, SRA^{+}21, SCMT^{+}22, SPP21, TPCM^{+}24, TFH^{+}21, WESWT21, Woo22, YFM^{+}20, FRP^{+}23, JRPHM^{+}22, KDS^{+}22, PCG^{+}22, LFH^{+}21a]. Ecosystem-based [WOL^{+}21, BLT23, BCC^{+}24, BFW^{+}21, DGM^{+}21, GMT21, GFO^{+}20, NAB^{+}22, RKH^{+}21, TPCM^{+}24, Woo22]. ecosystem-wide [NBA^{+}23]. ecosystems [BFJB21, GS22, GSL22, GJU^{+}23, LMH22, LZZ^{+}21, Lin21, LTMG23, MR20a, MCX21, PMM^{+}23, SVMBR22, Sta22, VNBLZR^{+}23, WBB^{+}23, WOL^{+}21, vDHR^{+}22]. edge [HCM24, OHM^{+}23]. eDNA [CMD^{+}23b, DPMZ21, HJS^{+}23, SACP^{+}21, UBD^{+}23]. education [KHHD20]. edulis [DLLC^{+}22]. eel [BDL^{+}22, DAB^{+}23, HFP^{+}23, LPA^{+}22, MD21, RST^{+}21, VBO^{+}21a, VBO^{+}21b, WJB^{+}24]. eelgrass [MWS^{+}23]. eels [Ano23i, BC23]. EFF [MPMGGAHR20]. Effect [SMS^{+}22, AHK23, CZ23, CdSSO^{+}21, JSR^{+}22, OHM^{+}23, PBS^{+}23, PPMD22, RAH^{+}22, SDSMS21, TBH^{+}20, Zho21]. effective [MOC^{+}23, RBH22, RDAOE20, TBH^{+}20]. Effects [CRG^{+}23, CBB21, EVY^{+}21, FFH^{+}22, GSWBK20, JWR^{+}21, KDS^{+}22, RBG^{+}20, SFPP^{+}23, ZJH^{+}23, Ano23l, AABM^{+}20, BSBH22, BKR^{+}21, BGH^{+}20, CCJ^{+}21, DBB^{+}20, DDL^{+}20, EDC20, EVLK^{+}21, HWL^{+}22, HUJ^{+}21, JYK^{+}21, KPS22, LTN21, MFV^{+}23, MHO^{+}20, MPA^{+}23, NAB^{+}22, NBA^{+}23, RRJ22, SH22, TSE^{+}21, VDAT20, VNR^{+}23, WBC^{+}21, WCLB23, WPN^{+}21, WBB^{+}21]. efficacy [RMF^{+}22]. efficiency [CWB^{+}23, CSH^{+}21, CSC^{+}23, DHB^{+}21, WGKG21, WCL^{+}23]. efficient [ASV21]. effort [AA21, CBB21, MPA^{+}23, PEWN20, PCH^{+}24, SRB^{+}23, TBH^{+}20, dAPDDJ21, SFPP^{+}23]. egg [GSWBK20, GPM21, HCCR20, WGI^{+}21]. Eight [CdSSO^{+}21]. elasmobranch [HCM24, MBS^{+}23, MSWB22, WJB^{+}21, WGL^{+}23]. elasmobranchs [Kai20]. elastic [KAK^{+}21]. electric [RBH22]. electrical [BRvL^{+}23, SMS^{+}22]. electron [BCBJ^{+}20]. electronic [ADS^{+}23, HMNF22, QWT^{+}21, TK20]. ELEFAN [ZHL^{+}22]. eleginoides [EMS^{+}21]. elevated [JWR^{+}21]. elongata [GICP^{+}22]. elongatum [BQBB^{+}21]. embayment [MHJ^{+}20]. embracing [KF21]. emergence [Gri20, MKK^{+}24]. emerging [MPMGGAHR20, Pro20]. Emiliania [LXBG21]. emission [HS20]. emissions [FDPD^{+}23, GS22, GSL22, Gal23, KGV21, Sta22]. emphasis [VLUB^{+}22]. empirical [Ano23a, FDK20, MBSR20, PEWN20, SVBM^{+}22]. empowerment [CMD^{+}23a]. enablers [SCB^{+}23]. enabling [Moo23a]. enclosed [PTD^{+}20]. encountered [IOS21]. end [SCN^{+}23]. end-to-end [SCN^{+}23]. endangered [BRG22, LPG^{+}20, MVNR^{+}21, SHP^{+}21]. endeavour [ZHL^{+}22]. endotherms [TM20]. energetics [JWR^{+}21]. energy [Ano24c, KDS23, MYW^{+}21, MPK^{+}22, NAB^{+}22, RBF^{+}22, SBM^{+}22]. engagement [SCCA20, TYF23]. engineering [Rin21]. England [GZSP22, GGW^{+}23]. English [LV23, VLUB^{+}22, WOB^{+}23]. enhance [SBW^{+}22]. enhanced [PGRP23]. enhancement [DEL21, HUJ^{+}21]. enhancer [CZSR21]. enhances [LXBG21]. Enhancing [ECBJ22, TFH^{+}21, WSL^{+}23]. enough [SBW^{+}22]. enriches [HJS^{+}23]. enrichment [ESK^{+}21, QSKL21]. ensemble [JAB^{+}21, LZWH20, SBC^{+}21a, SBC^{+}21b]. ensis [ZHL^{+}22]. entanglement [CBB21, Ste21]. enthusiasm [MR20b]. entrainment [Cha21]. entrance [BDL^{+}22, BPPC^{+}20, LPA^{+}22]. entry [MYW^{+}21]. envelope [Ano22b, SSV^{+}22b]. environment [CWL^{+}20, CdSSO^{+}21, HJS^{+}23, KDS23, MWJ^{+}24, MCH^{+}24, NFN20, OH23, PVGC^{+}22, QCM^{+}20, TFWE^{+}22]. Environmental [AJC^{+}23, CBFS23, CT23, KPS22, PPV^{+}22, SDR^{+}21, TNS^{+}21, VKH^{+}20a, VKH^{+}20b, VRM^{+}23, WBC^{+}22, AGSS22, Ano23l, Ano23k, BSBH22, BHL^{+}23, CP20, DLLC^{+}22, EDH^{+}21, FXG^{+}20, GETvB21, GPV^{+}22, GBMM^{+}23, GFO^{+}20, HAC^{+}23, JSN^{+}22, KEW^{+}21, LPA^{+}22, LWZC23, MSVSA22, MTS^{+}22, NBO^{+}20a, NBO^{+}20b, PRCM22, RGE^{+}23, Ste21, SAA^{+}22, TDB^{+}21, TRP^{+}21, WCLB23]. environments [CMD^{+}23b, vEMvH^{+}21]. epibenthic [SCW^{+}20]. Epifauna [LH20]. epifaunal [Ano23f, HFT^{+}22]. Epinephelus [BJPP^{+}20, LHE20, SHW^{+}21]. Epipelagic [MHA^{+}21, HHNS21]. epitaph [PF21]. equal [RST^{+}21]. equation [NSS^{+}23]. equations [Mun20]. equilibrium [Ano20c]. equipped [WGKG21]. Equity [PRAR23]. era [ANA^{+}23, BB20]. Erratum [ABB^{+}21a, BPKM22, GSWM^{+}22, GGOP22, HSS21b, LFH^{+}21a, SKA22]. Escape [MHMP^{+}23, SGM^{+}24]. escaped [DFK^{+}22]. escapement [HFP^{+}23]. especially [SAA^{+}22]. estimability [FSC^{+}23]. estimate [FOC^{+}23, KSHO23, PWH23, RCJ^{+}22, ZMFS20]. estimated [BSW22, LQC^{+}22, MKN^{+}21, PKAT20, PO24, TBM^{+}22]. estimates [AKK^{+}21, BCB^{+}22b, BHdL^{+}23, HJP22, Kai20, LDW23, LRHW21, PTR22, TSD^{+}21, TZE21]. Estimating [FWC^{+}20, HCB^{+}24, PMO^{+}21a, PMO^{+}21b, SBC^{+}21b, ZHL^{+}22, EWD^{+}21, EWD^{+}22, HPSF21, IEIT23, PEWN20, PDL20, PJP20, SPd^{+}22, SHFK22, SBKO22, SVMBR22, SBC^{+}21a]. Estimation [GGW^{+}23, RUdPFRA^{+}21, ÁEPCL20, CSM^{+}22, EMS^{+}21, KAK^{+}21, KP24, LZWH20, Mil24, MHKV20, PHB^{+}20, SMSP^{+}23, SMP^{+}23, SCN^{+}23, TCLB22, UBD^{+}23]. estimations [KF21]. estimators [HCH20]. estuaries [LL23, SWR^{+}22]. estuarine [BPR^{+}23, GZW^{+}20, HGS^{+}21, LS20b, NAS^{+}20]. estuary [PFT^{+}20, SOB23, MSVSA22]. ethological [SAM^{+}22]. ETS [BCBJ^{+}20]. EU [Ear21, Kra23, vDHR^{+}22]. Eubalaena [SPBJ21]. eukaryotic [KMK^{+}21b]. Euphausia [SMP^{+}23]. Euphausiacea [KLSV21]. euphausiid [Ano22c, BQBB^{+}21, DASLRB22]. euphausiids [PCG^{+}22]. Europe [BR21]. European [ANA^{+}20, BBM^{+}22, Bor21, BPPC^{+}20, CAE^{+}21, CGH^{+}23, DAB^{+}23, FAV^{+}20, GVD^{+}23, LvDM^{+}22, LVT^{+}23, Mac22, MD21, MCS^{+}20, OH23, RST^{+}21, SBPA^{+}21, SCP^{+}21, TSWM23, UVI^{+}21a, UVI^{+}21b, VBO^{+}21a, VBO^{+}21b, dPHG^{+}23, vdBDH21, MPMGGAHR20]. eutrophic [LMH22]. eutrophication [GZW^{+}20]. evaluate [BKF^{+}20, KKS^{+}20, TMG20]. Evaluating [AMP^{+}23, HAC^{+}23, LWZC23, LAGC23, RCJ^{+}22, SMO22, ZCX^{+}20, vDBF^{+}20, CMCF21, HLHC22, Kai20]. Evaluation [HUJ^{+}21, JBF^{+}20, KMG22, CMH^{+}23, CCR^{+}24, CMR^{+}20, HHK21, Kra23, MCC^{+}22, MD22, SPP21, ZB24, dMBJ22]. evaluations [WPM^{+}23]. event [QWT^{+}21]. events [DeM22, GLW21, MAI^{+}22, MAT^{+}23]. Evidence [ADS^{+}23, HBD^{+}21, Lin21, MOC^{+}23, BJPP^{+}20, KGV21, LLPdL22, MHJ^{+}20, TAPB22]. Evolution [Kra23, BPKM21, BPKM22, CRD^{+}22, FNB^{+}20, MACSR22]. evolving [DAB^{+}23]. Ex [SBPA^{+}21]. examination [YY22]. Examining [RBCD20, SRR21]. example [BBB^{+}21, HWL^{+}22, Kai20, SROR21]. excavations [PAW22]. exchange [JRDW22, RTQW20]. excluder [TYF23]. exerts [MHO^{+}20]. exhibit [LHE20]. existence [RRJ22]. existing [CFMB21]. exits [BW21]. expand [VLD^{+}21]. Expanding [TPCM^{+}24]. expansion [KWS^{+}21, STK^{+}20b, STK^{+}20a, SRR21]. experiences [MR20a]. experiment [MYJ^{+}20]. experimental [BMM^{+}21, BRG22, GAM^{+}22, Mil24, SHP^{+}21, TAPB22]. experimentally [SBB^{+}22]. Expert [SHG^{+}21, FSS^{+}23]. expertise [PCL^{+}21]. Explicit [MKS^{+}21, CZS^{+}23, FDMK21a, SCMT^{+}22]. Exploitation [ADBC^{+}22, AWN22, BSBH22, CdSSO^{+}21, GGW^{+}23, RAH^{+}22, WML^{+}21]. exploited [BJPP^{+}20, BSW22, DSMM^{+}22, ECR^{+}21, LTN21, RGE^{+}23, SSF^{+}23, TAPB22]. Exploration [CFB^{+}23]. Exploring [FDMK22, KMK21a, MAT^{+}23, SBD^{+}21, SSF^{+}23, YARE23, FXG^{+}20].

Exponential [WJ22]. exposed [MMN^{+}22, SG21]. Exposure [TAPB22, Ano23b, BPC^{+}24, CCJ^{+}21, GPS^{+}21b, GPS^{+}21a, HKZ^{+}23, MSN^{+}23, SASM22, vLNC20]. expression [MSVSA22]. extend [ADG^{+}20]. extending [SKM^{+}22]. external [JRDW22, LL23]. Extinction [LPG^{+}20]. extinctions [GA23]. extirpations [GA23]. extreme [MAT^{+}23]. eye [Ano24a, Lov20, SHK^{+}23].

face [HBC^{+}22, HEHT^{+}22, JAQ^{+}23]. facilitate [PJCS^{+}23a]. Factor [SRM^{+}20, PWCL21, UPH^{+}21]. Factors [KGV21, MB21, SMB^{+}21, SKH21, SMK23, XCK^{+}23, ZB21]. FADs [WGKG21]. fail [KH21]. failure [CvPTAJ22]. falciformis [FBF^{+}21, GSR^{+}20]. Falkland [ASW^{+}23]. fall [Mac22]. family [HHL^{+}21]. far [dCKS^{+}22]. farm [BHDD22, TSWM23, VCD^{+}20, WBG^{+}22]. farmed [DFK^{+}22, MvdMS^{+}23, SGM^{+}24]. farming [BBB^{+}21, SSBV21]. farms [BHDD22, CvdWC^{+}20, GKRB21, Met20, PRD20, SASM22, VLD^{+}21]. Faroe [OOB^{+}22, VKH^{+}20a, VKH^{+}20b]. fat [KPM^{+}22]. fate [vdWBL22]. fatty [MvdMS^{+}23]. fauna [SLB^{+}20]. Fear [JSLC21, KMS20b]. feasibility [PHB^{+}20]. features [MKSM23]. fecundity [CCJ^{+}21, GZSP22, RPVT^{+}21]. feed [MvdMS^{+}23]. Feeding [NBO^{+}20a, NBO^{+}20b, Pep23, ANA^{+}20, Ano22b, BPP^{+}21, GHÓ^{+}20, OOB^{+}22, OGMA20, SAM^{+}22, SSV^{+}22b, UPH^{+}21, dAdSM21]. Female [DSF^{+}21, CT23, RPVT^{+}21, WRC^{+}21]. fidelity [GGLS22, dPHG^{+}23]. fiducial [MHKV20]. Field [VCF^{+}20, PBC^{+}22, PVGC^{+}22]. fieldwork [CvPTAJ22]. Filbee [Gal23]. fill [MGF^{+}21]. fimbria [Ano23j, CRLC23, ORGTO^{+}23]. fin [SWR^{+}22]. financing [MFB23]. finding [SPd^{+}22]. fine [Ano23d, FPY^{+}22, NAS^{+}20, lCRGCCA23]. fine-scale [Ano23d, FPY^{+}22, NAS^{+}20, lCRGCCA23]. finfish [CFMB21, TLM^{+}20]. finmarchicus [Ano22e, EDH^{+}22, GGOP21, GGOP22, JRDW22, JNT^{+}22, KPS22]. Finnish [HHK21]. First [BM20, GGLS22, TWC^{+}20, WBG^{+}22, vdWBL22]. first-year [vdWBL22]. Fish [Ano22a, DLA^{+}21, EHV^{+}21, MAI^{+}22, PH21, VO23, AKK^{+}21, ARE^{+}22, ÁEPCL20, ANB^{+}21, ANB^{+}22, Bac24, BSW^{+}23, BBM^{+}22, BXF23, BCBJ^{+}20, BGR^{+}22a, BCB^{+}22a, BSW22, BGH^{+}20, BNB21, BPPC^{+}20, CRG^{+}23, Cha21, CSM^{+}22, CCC^{+}21, CSH^{+}21, CKR^{+}21, CATM22, CWC^{+}23, CCJ^{+}21, DA20, DSDA22, Dom21, DDL^{+}20, EGMO^{+}20a, EGMO^{+}20b, FMF^{+}20, GSWM^{+}21, GSWM^{+}22, GPQ^{+}20, HGS^{+}21, HAPW21, HPSF21, Hut22, JAQ^{+}23, KLP^{+}24, KH21, LTN21, LWL^{+}22, LPC^{+}23, LvDM^{+}22, LAGC23, MWT^{+}21, MFV^{+}23, MHKV20, MHO^{+}20, MPA^{+}23, MPH^{+}21, ÖHB^{+}23, OH23, PMO^{+}21a, PMO^{+}21b, Pep23, QCM^{+}20, RMF^{+}22, ROH^{+}21, RDAOE20, RGE^{+}23, SASM22, SSS^{+}20, SMJ^{+}21, SL21, SMO22, STS^{+}20, SdS20, SVML^{+}23, SBW^{+}22, SJS21, SCN^{+}23, SOB23, SBC^{+}21a, SBC^{+}21b, SB21, Ste21, SACP^{+}21, SAA^{+}22, SDR^{+}21, SWS^{+}21, TWC^{+}20, TSW^{+}21, TK20, VRM^{+}23, VDAT20, VLD^{+}21]. fish [WPN^{+}21, WBB^{+}21, WBG^{+}22, WLR^{+}20, dCKS^{+}22, dPMS^{+}22, NSS^{+}23]. fish-retention [CSH^{+}21]. fished [DSF^{+}21]. Fisher [Bjo21, BMM^{+}22, FKTM20, MAT^{+}23, RSTL22, SCCA20, SKH21, Suu22, TYF23]. Fisheries [Ear21, GVD^{+}23, HS20, MPMGGAHR20, NSJ^{+}22, SHP^{+}22, WOL^{+}21, ABB^{+}21a, ABB^{+}21b, AVC^{+}21, Ano20b, Ano23d, ALLP^{+}22, BLT23, BPC^{+}24, Bar23, BCB^{+}22b, BDB^{+}21, Bjo21, BHQF20, BW21, BTM^{+}21, BBL^{+}23, BRG22, Cha21, CMG^{+}23, CFO^{+}22, CBCM21, CMCF21, CCH^{+}23, ECR^{+}21, ECBJ22, EGMO^{+}20a, EGMO^{+}20b, FFR20a, FFR20b, FCD^{+}23, FSC^{+}23, FDMK21a, FDMK22, FAV^{+}20, FMF^{+}20, FMPH24, GBGA^{+}23, GDI^{+}20, GAP21, GBH^{+}24, GSR^{+}20, GGBC20, GFO^{+}20, HWL^{+}22, HHK21, Har20, Har21, HEHT^{+}22, HMNF22, Hol21, HHO^{+}20, Hut22, HMK^{+}20, IHL21, JMB^{+}22, JAB^{+}21, KLG^{+}23, KH22, KTA^{+}23, KOB^{+}22, KGG^{+}23, KGV21, KWEB21, KH21, LLLS20, LSR^{+}21, LWH^{+}21, LKL^{+}21, LBPC^{+}20, LSOD^{+}20, MGF^{+}21, MYJ^{+}20, Mac22, MMC^{+}23, MV23, MVNR^{+}21, MYAH21, Met20, MAE^{+}23, MSP^{+}20, MKGC20, MPH^{+}21, MMMB^{+}23, ODT^{+}22, OdS21, OP22, OMM^{+}23, PEWN20, PVHV20, PCH^{+}24]. fisheries [PBF^{+}21, PCC23, PRAR23, PWH23, PJP20, Pro20, RCB^{+}21, RHvD^{+}20, dSRKC22, ROH^{+}21, RHW20, SPd^{+}22, SMJ^{+}21, SMLB^{+}22, SHG^{+}21, SOO^{+}22, SWR^{+}22, SFG22, SCMT^{+}22, SLRC21, SSF^{+}23, SPP21, SRMS21, SHP^{+}21, TZE21, TBP20, TNM^{+}23, TMLH21, TFWE^{+}22, VCF^{+}20, WGKG21, WPB20, WBB^{+}23, WGCC^{+}22, WZJ^{+}22, WMB^{+}22, WCL^{+}23, YMAH20, ZL22, ZB21, dLP21, dCKS^{+}22, lCRGCCA23]. fisheries-independent [ODT^{+}22]. fishers [BKF^{+}20, Bar23, CMH^{+}23, Eay23, TLM^{+}20, TR20, VPCM^{+}22, dBSQK21]. fishery [Ano22d, Ano23j, AHP^{+}21, AWN22, ASW^{+}23, AA21, AHK23, BWBH21, BJBB20, BW21, BHdL^{+}23, BFW^{+}21, BSH^{+}22, CASM23, CRLC23, CRR^{+}22, DCL^{+}21, EMS^{+}21, ECR^{+}21, ECBJ22, EHV^{+}21, FFR20a, FFR20b, FOC^{+}23, FBF^{+}21, GAM^{+}22, GC21, Gra21, GLM^{+}23, HWL^{+}22, Har21, HSS21a, HS20, JYK^{+}21, KSHO23, KKS^{+}20, KVS^{+}23, LWZC23, LLALF22, LBH^{+}21, MDAMPE^{+}21, MDP^{+}21, MCOÁ^{+}22, MB21, MPK^{+}22, MHMP^{+}23, NKS^{+}23, OTLH20, PIAF^{+}23, PDH^{+}21, PPMD22, PO24, PRRB^{+}20, RCJ^{+}22, RDM^{+}21, RHO22, SHFK22, SBM^{+}22, SBKO22, SMS^{+}22, SROR21, SRBT23, SCH^{+}21b, SCP^{+}21, SB21, TSD^{+}21, TMG20, THSZ21, TFH^{+}21, VNR^{+}23, VQN22, WCCY20, WBC^{+}21, WPP^{+}21, XCK^{+}23, dAPDDJ21, HSS21b]. fishery-dependent [Ano23j, CRLC23, ECBJ22, SBKO22]. fishery-independent [MDP^{+}21]. fishery-targeted [KSHO23]. fishes [ARHM21, BOLPHL21, BRvL^{+}23, CLL21, CHZC22, KTK20, LRGW23, Sha21, ZYJZ23]. Fishing [RAH^{+}22, RGE^{+}23, AMP^{+}23, BBB^{+}22, BR22, BHL^{+}23, BFJB21, BP20, CRG^{+}23, CVJ^{+}20, CBB21, DPH^{+}21, FW22, FODD^{+}22, FXG^{+}20, GRE^{+}20, GGA^{+}23, GT22, JEPT23, KMK21a, KTK20, KMS20b, LTN21, LS20a, LQC^{+}22, LLPdL22, LHN^{+}22, MD21, MSW21, MPMGGAHR20, NHL^{+}22, NAB^{+}22, OTLH20, OH23, PPMD22, PFT^{+}20, QWT^{+}21, QCM^{+}20, RPVT^{+}21, RES^{+}23, SRB^{+}23, SRM^{+}20, Ste21, VO23, VSS^{+}23, WGKG21, WCL^{+}23, YABM20, YMAH20, ZJH^{+}23, dAPDDJ21]. Five [HHN22, SFHP21, SFHP22, MCOÁ^{+}22, SDGR^{+}20]. five-fleet [MCOÁ^{+}22]. Fjord [NKA^{+}20, KBK^{+}22, SJM^{+}20]. flat [DPD^{+}22a, DPD^{+}22b]. flatfish [BVV^{+}20, CWB^{+}23, RAH^{+}22]. flatfishes [YC20]. fledging [LB23]. fleet [HMNF22, LV23, LHN^{+}22, MCOÁ^{+}22, NHM^{+}21, PEWN20]. fleet- [PEWN20]. fleet-specific [NHM^{+}21]. fleets [CCR^{+}24, SRB^{+}23]. Flesh [LB23, KAK^{+}21]. Flesh-footed [LB23]. fleshy [HMBDP21]. Flexibility [BTM^{+}21, FWA^{+}21, GBH^{+}24]. Flexible [CKR^{+}21, KMK^{+}21b]. floor [RTQW20]. Florida [CSM^{+}22, JSR^{+}22]. flounder [YMH^{+}24]. flow [BØS^{+}21, BPR^{+}23, SGM22]. fluctuating [LXBG21, SMB^{+}21]. fluctuations [FFS^{+}22, LIR^{+}21, PTD^{+}20]. flumes [EDC20]. fluvial [Ano23i, BC23]. fluviatilis [HMOR22]. focal [CWC^{+}23]. focus [Ano23f, HFT^{+}22, Pec23]. focused [RCB^{+}21]. folk [DeM22]. followed [KBBGA22]. following [GPS^{+}21b, MAC21]. food [CVJ^{+}20, EDH^{+}21, HAC^{+}23, HHA^{+}21, JSR^{+}22, KMK^{+}21b, KJP^{+}22, LMH22, MTCC21, MVNR^{+}21, MPH^{+}21, PFH^{+}22, SCMT^{+}22, Thi20]. food-web [PFH^{+}22]. foods [WSL^{+}23]. footed [LB23]. footprint [FMPH24, ZL22]. forage [BSW^{+}23, SDR^{+}21, SWS^{+}21]. foraging [BFD^{+}23, BR22, CPD^{+}21, GGA^{+}23, HBD^{+}20, JDJK22, KVS^{+}23, SPBJ21, SPJ23]. foraminifera [TLS21]. forbesii [LAB^{+}22, SBA^{+}22]. forces [KMS20b]. forcing [GFO^{+}20]. fore [EDC20]. forecast [LTMG23, NHM^{+}21, SAB^{+}20]. Forecasting [FFR20a, FFR20b, RCH^{+}21, HMF^{+}21, WCL^{+}23]. forecasts [KNP21, NKS^{+}23]. Foreseeing [MACSR22]. forest [HK21]. forests [FDPD^{+}23, Gal23, HGS^{+}21, PMSK^{+}20]. fork [Mur21]. form [FSN^{+}22]. formal [Sha21]. formalizing [RKH^{+}21]. forthcoming [SWS^{+}21]. Forty [Sha21]. forward [DGM^{+}21, PE21]. fossil [JMR^{+}22]. four [SL21]. fragmentation [GA23]. frame [SPd^{+}22, VCF^{+}20]. Frameless [SPd^{+}22]. frames [VCF^{+}20]. Framework [MTCC21, PKT^{+}21, AVC^{+}21, BGR^{+}22b, GLW21, HLHC22, KNP21, KDS23, SLB^{+}20, SFG22, vDHR^{+}22]. frameworks [MFA^{+}24]. free [NFN20]. French [PRD20]. frequency [CKH^{+}21, LFH^{+}21a, LFH^{+}21b, LFH^{+}22, SHW^{+}21, ZB21, ZHL^{+}22]. Frequently [MBSR20]. friends [Pop20]. frontlines [Har20]. FT [PHB^{+}20]. FT-NIR [PHB^{+}20]. function [CZ23]. functional [ANB^{+}21, ANB^{+}22, MGO^{+}23, PMM^{+}23]. functioning [NAB^{+}22, NBA^{+}23]. Fund [MPMGGAHR20]. Fundy [MKSM23]. further [GMT21]. future [Abe20, BCB^{+}22b, BD20, BMN^{+}21, CFB^{+}21, DHN^{+}21, FCF^{+}21, GZW^{+}20, HMA^{+}23, HS20, HMK^{+}20, KHSS21, LSR^{+}21, LXBG21, MHES20, MACSR22, MCX21, MFA^{+}24, Mon22, MPH^{+}21, PB21, PDH^{+}21, PFT^{+}20, RCB^{+}21, Ste21, SCG^{+}20, SCF^{+}21, TNM^{+}23, VSS^{+}23, WBB^{+}23, WJ22, vdWBL22, MKB^{+}23]. future-projected [LXBG21].

gadids [Ano22a, DLA^{+}21]. Gadus [Ano22b, Ano22d, Ano23b, BCB^{+}22b, CSH^{+}21, CRR^{+}22, DLL^{+}21, DSGG20, DYS21, FKTM20, GGLS22, HPS^{+}23, HHA^{+}21, IEIT23, JBF^{+}20, KTA^{+}23, LDW23, LAGC23, LOÖK20, MMN^{+}22, MSN^{+}23, MvdMS^{+}23, NBO^{+}20a, NBO^{+}20b, PAW22, SHFK22, SBD^{+}21, SSV^{+}22b, vLNC20]. gag [GSWBK20]. gahi [AWN22]. Gallagher [Sta22, FDPD^{+}23]. game [KHHD20, dAPDDJ21]. game-based [KHHD20]. gammarus [CAE^{+}21, CGH^{+}23, RRJ22, TSWM23]. gannet [SAM^{+}22]. Gannets [CVJ^{+}20, dGW^{+}22]. gap [FCF^{+}21, Gri20, ZMMF22]. gaps [DBB^{+}20, MGF^{+}21]. gas [AKK^{+}21, HS20, KGV21, RHW20]. gas-bearing [AKK^{+}21]. gastropods [LBW^{+}20]. gateways [MPH^{+}21, DHN^{+}21]. Gauging [BPC^{+}24, PLN20]. gauntlet [VLD^{+}21]. Gaussian [BAS^{+}21, LHN^{+}22]. gear [BR22, BHdL^{+}23, BMM^{+}21, ECBJ22, HAP^{+}21, JEPT23, MTK^{+}21, RCJ^{+}22, SRB^{+}23, SMK23]. gear-in [ECBJ22]. gear-out [ECBJ22]. gears [Ano23j, CRLC23, GPM21, Mil24]. gelatinous [LHL^{+}21, TM20, YSL22]. Gender [JBDK23]. gene [BØS^{+}21, MSVSA22]. generalized [MCOÁ^{+}22, NHL^{+}20, RUdPFRA^{+}21]. generate [Thi20]. generating [LCG^{+}22, Mun20]. generation [DSR^{+}21]. generational [BJPP^{+}20]. generative [BDM24]. generosa [dlTVLC^{+}21]. Genetic [BPC^{+}21, BJG^{+}21, BJPP^{+}20, DSMM^{+}22, HWS^{+}21, JAQ^{+}23, MCS^{+}20, OOB^{+}22, ORGTO^{+}23, WSJ23, Ano22d, BØS^{+}21, BLM^{+}21, CRR^{+}22, FPY^{+}22, FDMK21b, FDMK21c, GHÓ^{+}20, HUJ^{+}21, KMS^{+}20a, KH21, NAS^{+}20, TRD^{+}22]. genetically [DeM22]. genetics [Gra21]. Genome [FPY^{+}22, DRP^{+}23, JLM^{+}23]. Genome-wide [FPY^{+}22, DRP^{+}23]. Genomic [MGB^{+}23, PFP^{+}22, CASM23, JAQ^{+}23, KEW^{+}21, MKN^{+}21]. genomically [BBB^{+}23]. genomics [WML^{+}21]. genotype [TDB^{+}21]. genus [MMH^{+}23, WPN^{+}21]. geoduck [dlTVLC^{+}21]. Geographic [DBT^{+}23, CAE^{+}21, CGH^{+}23]. geometry [ON21]. George [LLPdL22]. Georgia [EMS^{+}21, HHL^{+}21, TYF23, TFH^{+}21, TFWE^{+}22]. geospatial [SRB^{+}23]. German [Bar23]. get [DMK^{+}21]. giant [BPR^{+}23, RCB^{+}21]. gigas [DBGP22, JWR^{+}21, KWS^{+}21, NKA^{+}21]. Gillespie [SHP^{+}22]. Gillnet [FKTM20, Bar23, CCR^{+}24, GAM^{+}22, SOB23]. gillnets [SHFK22]. Gjerde [JHN^{+}21]. glacialis [SPBJ21]. glass [BDL^{+}22, LPA^{+}22]. glider [CSC^{+}23]. Global [GPV^{+}22, BBB^{+}22, BMN^{+}21, CRV^{+}21, CSM^{+}23, DNW^{+}23, MKCA^{+}23, NBA^{+}23, MEOA23]. go [CvPTAJ22]. goals [CFB^{+}21, TMLH21]. gonads [MvdMS^{+}23]. good [CMD^{+}23a, GDI^{+}20, HVDvD23, Kra23, MEOA23, MR20b, Pec23]. gorbuscha [DBT^{+}23]. gorgonian [SGM22]. governance [SMK23, SOCN^{+}22]. GPS [VCD^{+}20]. grabbing [QBF^{+}21]. gradient [UVI^{+}21a, UVI^{+}21b]. gradients [AJC^{+}23]. Grand [THSZ21]. Great [KBJA20, Ano22a, DLA^{+}21, Kos21]. greater [LHL^{+}21, OOB^{+}22, OMM^{+}23, LAB^{+}22]. green [dAdSM21]. greenhouse [HS20, KGV21]. Greenland [ANA^{+}20, GEBS^{+}23, LTN21, LBH^{+}21, NKA^{+}20, SCH^{+}21b, VHT^{+}21, VET^{+}22, WM20, YCF^{+}23]. greenling [YMH^{+}24]. grey [MTTG^{+}23, PMS^{+}20]. ground [AKK^{+}21, ADG^{+}20, OOB^{+}22, STK^{+}20b, STK^{+}20a]. ground-truthed [AKK^{+}21]. groundfish [GLM^{+}23, JKF^{+}22, KBW22, KKS^{+}20, ROH^{+}21, VRH^{+}23]. grounds [ANA^{+}20, HBD^{+}20, LAB^{+}22, OGMA20, VKH^{+}20a, VKH^{+}20b, dAdSM21]. group [SCCA20]. group-specific [SCCA20]. grouper [BJPP^{+}20, GSWBK20, LHE20, SHW^{+}21]. groups [FSS^{+}23, GVD^{+}23, OOB^{+}22]. Growth [CLL21, QSKL21, SSS^{+}23, VBO^{+}21b, ANA^{+}23, Bir23, Bra20, Bra22, BPP^{+}21, CMCF21, CMS^{+}23, DYS21, DGM^{+}24, FFH^{+}22, dlTVLC^{+}21, HPSF21, KTY^{+}21, LXBG21, LVT^{+}23, MHJ^{+}20, MSVSA22, MCD^{+}20, MSW21, NBO^{+}20a, NBO^{+}20b, NBO^{+}20c, PBS^{+}23, SBKO22, SdS20, SBD^{+}21, SHE^{+}22, TSE^{+}21, VDAT20, WJ22, YFM^{+}20, ZMFS20, Zho21, ZHL^{+}22, vdWBL22, VBO^{+}21a]. grunion [MPQS20]. Guadalupe [JRPHM^{+}22]. guide [CFB^{+}21, Eay23, SCM^{+}23]. Gulf [MSVSA22, SMS^{+}21, BHWC^{+}20, BCB^{+}22a, BQBB^{+}21, BPP^{+}21, FFR20a, FFR20b, JRDW22, KKS^{+}20, LL23, MHA^{+}21, MTS^{+}22, PFP^{+}22, SPJ23, WRC^{+}21, WGCC^{+}22, dlCRT^{+}22]. gull [GGA^{+}23]. gulls [VCD^{+}20].

Habitat [WDPM23, AMD^{+}23, BGR^{+}22a, CATM22, DPH^{+}21, DCL^{+}21, FKTM20, GPV^{+}22, GGLS22, GICP^{+}22, GVD^{+}23, HMA^{+}23, HAP^{+}21, HMF^{+}21, IOS21, JSN^{+}22, KTY^{+}21, KTK20, MRC^{+}21, MWT^{+}21, MKSM23, MSW21, MTT^{+}23, PTR22, PPMD22, SCW^{+}20, SWR^{+}22, SRA^{+}21, SWH^{+}23, TSWM23, THSZ21, VHT^{+}21, WCL^{+}23, vDBF^{+}20, CBFS23]. habitats [LBH^{+}21, MHO^{+}20, RFB^{+}22, RHvD^{+}20]. habits [MR20a]. hadal [WJ22]. haddock [BJG^{+}21, IHL21]. hake [SRBT23, SFPP^{+}23]. Half [KTA^{+}23, BLM^{+}21]. Halibut [KEW^{+}21, GEBS^{+}23, LBH^{+}21, LMW^{+}22, VHT^{+}21, VET^{+}22, WM20, YCF^{+}23]. Haliotis [AABM^{+}20, EVLK^{+}21, MDAMPE^{+}21]. hammerhead [LMvEM23]. handling [WJB^{+}21]. handling-related [WJB^{+}21]. Harald [SRMS21]. harbour [GETvB21]. harengus [BBP^{+}23, KPM^{+}22, KÓP^{+}21, TNS^{+}21]. Harnessing [CCH^{+}23]. Harvest [CBCM21, AHK23, DDP^{+}21, FDMK22, FODD^{+}22, GDI^{+}20, HAPW21, ODHF^{+}23, PBF^{+}21, RCH^{+}21, SPd^{+}22, SLRC21, SSF^{+}23, TLM^{+}20, WOB^{+}23, ZB24]. harvestable [HSMK21]. harvested [LTK20, TK20, VO23]. harvesting [BP20, SCCA20]. harveyi [NKA^{+}21]. hatchery [CCM^{+}21, TRP^{+}21]. Hatching [SH22, HCCR20]. haul [BMM^{+}21]. health [CFMB21, JDJK22, MAT^{+}23, Moo23a, SRM^{+}20]. hearing [PH21]. Heat [yMlZlL^{+}21, WGCC^{+}22]. heatwave [KDS^{+}22]. Heavy [AWN22, Ano24b, RVR^{+}24]. Helgoland [KMK^{+}21b]. helgolandicus [VLUB^{+}22]. helicina [MOC^{+}23]. help [FDPD^{+}23, Gal23, GT22, PRAR23]. helped [Hol20]. herbivory [ESK^{+}21]. here [Bea23]. Heritage [TSD^{+}21]. herring [Ano23k, BSBH22, BBP^{+}23, BØS^{+}21, CdSSO^{+}21, CSC^{+}23, GBMM^{+}23, HHK21, KPM^{+}22, KÓP^{+}21, MDHB21, PMO^{+}21a, PMO^{+}21b, SKA15, SKA22, SHE^{+}22, SPP21, TZE21, TNS^{+}21, ZB24]. Heterogeneity [BKR^{+}21, BCP^{+}21, HGS^{+}21, MTS^{+}22, PEWN20, PPMD22, SJX^{+}20]. heterotrophic [LFH^{+}21a, LFH^{+}21b, LFH^{+}22]. hidden [SDW^{+}21, TZR^{+}21, VAS^{+}20]. Hierarchical [KSHO23, KMS^{+}20a, CWB^{+}23, TMS23, ZMFS20]. High [CBT^{+}20, GGLS22, KBK^{+}22, LVL^{+}22, TRD^{+}22, BCP^{+}21, BFJB21, CASM23, CCJ^{+}21, EDC20, HK21, JAQ^{+}23, KTA^{+}23, MYW^{+}21, SK21, SCP^{+}21]. high-energy [MYW^{+}21]. high-latitude [KTA^{+}23]. High-resolution [LVL^{+}22, HK21]. highlight [MVMdS21, TR20]. Highlighting [SdS20, GT22]. highly [LHN^{+}22, LVL^{+}22, MMCS^{+}24, SSF^{+}23, Zho21]. hindcast [BPPKH20]. hippoglossoides [GEBS^{+}23, WM20]. hippurus [MCOÁ^{+}22]. historic [SWR^{+}22]. historical [GGBC20, VBJ^{+}21]. histories [CCS^{+}24]. History [PRRB^{+}20, ADBC^{+}22, BSBH22, BPR^{+}20, DBGP22, Dol20, Ear21, FDK20, JNT^{+}22, LLLS20, Mun20, OH23, PWH23, RCB^{+}21, SGM^{+}24, WBB^{+}23]. Hjort [KHSS21, SRMS21]. holistic [HHO^{+}20, MMMB^{+}23, SSV^{+}22a]. Holm [SFHP22]. Holt [ABB^{+}21a, ABB^{+}21b, GC21, Ken21, Man21, RCB^{+}21]. Homarus [CAE^{+}21, CGH^{+}23, HCCR20, RRJ22, TSWM23, WRC^{+}21]. hook [FCD^{+}23, FOC^{+}23]. hook-and-line [FCD^{+}23, FOC^{+}23]. hope [BB22]. horizontal [KTK20]. hotspot [GA23, MKK^{+}24]. hotspots [CRD^{+}22]. hubbsi [SdS20]. human [Ano22a, DLA^{+}21, PJCS^{+}23a, MTTG^{+}23, SRR21, vEMvH^{+}21]. human-centred [PJCS^{+}23a]. human-wildlife [Ano22a, DLA^{+}21]. humanities [TPCM^{+}24]. Humpback [ADG^{+}20]. Huron [HBMC20]. huxleyi [LXBG21]. Hybrid [Mil24, SSS^{+}20]. hydroacoustic [KTK20, RFB^{+}22, RMF^{+}22]. hydrocarbons [LZZ^{+}21]. hydrodynamic [MYW^{+}21]. hydrography [SWS^{+}21]. hydrophone [SSP^{+}23]. Hyne [PTD^{+}20]. hyperboreus [Ano22e, GGOP21, GGOP22]. hypercapnia [LHE20]. hyperspectral [CNWW22]. hypothesis [Bra22, FND23]. hypoxia [Bra20, Bra22, NBO^{+}20c, RBG^{+}20, vDBF^{+}20].

Iberian [dlCRT^{+}22, QCM^{+}20]. Ice [MTT^{+}23, DDL^{+}20, HYE^{+}22]. Iceland [Ano22e, CVJ^{+}20, GGOP21, GGOP22, KGV21, MWT^{+}21]. Icelandic [GPJ^{+}21, KÓP^{+}21, vLNC20]. ICES [Ano23a, FCF^{+}21, FSS^{+}23, SVBM^{+}22]. Ichthyophonus [Ano23k, GBMM^{+}23]. ichthyoplankton [YSL22]. ideal [DEL21]. Identification [TZR^{+}21, BPC^{+}21, DBB^{+}20, Dom21, OOB^{+}22, SDW^{+}21]. identifications [MKN^{+}21]. identified [WJ22]. identify [AKL^{+}21, ARHM21, BXF23, CMH^{+}23, KMG22, MSVSA22]. Identifying [Bar23, BNV^{+}21, GJU^{+}23, JBE^{+}21, KOB^{+}22, LTK20, MTCC21, MAC21, OT22, SCP^{+}21, BBP^{+}23, GGRB22, LvDM^{+}22, TK20, VSS^{+}23]. II [MPH^{+}21]. illegal [FW22]. Illex [AWN22, CASM23]. illuminate [KH21]. illumination [CZSR21]. illustrative [MKGC20]. Image [ÁEPCL20, CZSR21, GJU^{+}23, LSD^{+}20, TMD^{+}20, YLC^{+}22]. Image-based [ÁEPCL20, GJU^{+}23, TMD^{+}20]. images [ARHM21, YLC^{+}22]. imaging [GCT^{+}21, MCH^{+}24]. imbalance [POUH23, vEMvH^{+}21]. Impact [KLF^{+}20a, KLF^{+}20b, PO24, ZL22, Bor21, BFJB21, BB22, GPJ^{+}21, HMNF22, HKC^{+}23, HYE^{+}22, JDD^{+}21, LV23, MEOA23, MFP20, ODHF^{+}23, PRD20, RHvD^{+}20, RCH^{+}21, SASM22, VBJ^{+}21, VNR^{+}23]. Impacts [CFMB21, FND23, BPR^{+}23, GT22, HEHT^{+}22, HAC^{+}23, HLHC22, Kai20, KGG^{+}23, LV23, LWZC23, LLPdL22, LBH^{+}21, MTS^{+}22, RCJ^{+}22, RGE^{+}23, SFW^{+}22, SBM^{+}22, Ste21, SCMT^{+}22, TAPB22, VKH^{+}20a, VKH^{+}20b, VSS^{+}23, vDBF^{+}20]. imperilled [RST^{+}21]. Implementation [CFB^{+}23, BWBH21]. implications [Ano22d, Ano24c, BFFC20, BTM^{+}21, CASM23, CRR^{+}22, DEL21, GPM21, GZSP22, GSR^{+}20, GMT21, HKZ^{+}23, HMA^{+}23, HSS21a, HSS21b, KMS^{+}20a, MTCC21, MDAMPE^{+}21, PFT^{+}20, RUdPFRA^{+}21, ROH^{+}21, SSBV21, SWR^{+}22, SMK23, TFWE^{+}22, WPN^{+}21, ZJF^{+}21, ZB21]. importance [ALLP^{+}22, BBM^{+}22, BPC^{+}20, DGM^{+}24, OP22, VKH^{+}20a, VKH^{+}20b, VPP^{+}23, YMAH20]. important [HSMK21, RRJ22]. improve [LDW23, NKS^{+}23, SMSP^{+}23, SCH^{+}21b, dPMS^{+}22]. improved [LMW^{+}22, WCL^{+}23]. improves [Ano23g, GCT^{+}23, dlTVLC^{+}21, SDW^{+}21, UBD^{+}23]. Improving [BKF^{+}20, ZYJZ23, BCB^{+}22b, OMM^{+}23]. incentivizing [PM23]. incidence [CBT^{+}20]. incidental [dSRKC22]. including [VSS^{+}23]. inclusion [LVG^{+}20]. Incoherent [BDB^{+}21]. incorporate [LKL^{+}21, NHL^{+}22]. Incorporating [GFO^{+}20, IEIT23, MTS^{+}22, MTK^{+}21, OBK21, WBA^{+}22, WFG^{+}23, Ano23e]. incorporation [MKS^{+}21]. incorrect [FSC^{+}23]. increase [BWBH21, CSH^{+}21, Pro20, dSRKC22, WGKG21]. increased [HBD^{+}21, JMB^{+}22, TBH^{+}20]. Increases [OOE^{+}22, ANA^{+}23, ESK^{+}21, ZEM21]. Increasing [KLG^{+}23, WGI^{+}21, STK^{+}20b, STK^{+}20a, dBSQK21]. increment [CAE^{+}21, VDAT20]. increments [DGM^{+}24, Zho21]. incubator [LNW21]. independent [MDP^{+}21, ODT^{+}22, ZEM21]. index [GJU^{+}23, MAT^{+}23, MTK^{+}21, SFPP^{+}23, TAB^{+}20, WBC^{+}21, YARE23]. Indian [BPC^{+}24, FBF^{+}21, GRE^{+}20, SMJ^{+}21, SK21, YABM20]. indicate [DFWR20, GHB^{+}23, NSS^{+}23]. indicates [SWS^{+}21]. indications [BM20]. indicator [FMBL23, HLHC22, SBH^{+}21, SRA^{+}21, ZL22]. indicator-based [FMBL23]. indicators [BLT23, FXG^{+}20, HvdMLS21, KMG22, ÖHB^{+}23, SROR21]. indices [BAS^{+}21, LVG^{+}20, LMW^{+}22, MCX21, NDE^{+}21, PWCL21, PCCL21]. indicus [PDH^{+}21]. Individual [dlTVLC^{+}21, LRHW21, NSS^{+}23, AKK^{+}21, BPPKH20, HHK21, LAGC23, PMO^{+}21a, PMO^{+}21b, dBABS^{+}22]. Individual-based [NSS^{+}23, BPPKH20]. individuals [Pep23]. Indo [FPY^{+}22]. induced [Ano23l, BGR^{+}22a, BMU^{+}23, CCM^{+}21, HMWS22, JG21, JHS^{+}21a, JHS^{+}21b, LKL^{+}21, MYMV20, WCLB23]. induces [EVLK^{+}21]. industrial [ANA^{+}23, DCL^{+}21, GGBC20]. industrial-era [ANA^{+}23]. Industry [ASW^{+}23, SK21, GT22, KPM^{+}22, Mac22, SRM^{+}20, TYF23]. Industry-collaborative [ASW^{+}23]. Industry-collected [SK21]. Infaunal [HFT^{+}22, Ano23f]. infection [FFH^{+}22, JHN^{+}21, SDSMS21]. infections [Ano23k, GBMM^{+}23]. infestation [SJM^{+}20]. infested [SBB^{+}22]. Influence [BPR^{+}20, CWC^{+}23, RMF^{+}22, CT23, CVJ^{+}20, DLLC^{+}22, DHB^{+}21, GGBC20, MYJ^{+}20, QCM^{+}20, SKH21, ZCX^{+}20]. influences [CWL^{+}20, KMK21a, TNS^{+}21]. influencing [SMK23]. inform [AVC^{+}21, Bar23, BJBB20, CMD^{+}23b, HBC^{+}22, MSP^{+}20, RRR^{+}24, SROR21, SHG^{+}21, SCP^{+}21, TSD^{+}21, WJB^{+}24]. Information [BPV^{+}20, Ano23j, CMH^{+}23, CRLC23, ECBJ22, Mun20]. information-generating [Mun20]. Information-theoretic [BPV^{+}20]. infrared [CNWW22]. infrastructure [FJC^{+}20, RHW20, SCW^{+}20]. Ingredients [MR20b]. inhabited [MYMV20]. inherited [DeM22]. initial [BWBH21]. injuries [BRvL^{+}23, SMS^{+}22]. innate [KMS20b]. innovations [Jen23]. inorganic [HMBDP21, PBS^{+}23]. inputs [LRHW21]. inshore [HWS^{+}21, HSS21a, HSS21b]. insight [KÓP^{+}21]. Insights [WJB^{+}21, AHC^{+}20, ADBC^{+}22, BFD^{+}23, BW21, BPKM21, BPKM22, DMK^{+}21, HHO^{+}20, PCG^{+}22, SOC^{+}20, TND^{+}20, TCLB22]. inspire [Eay23]. institutions [SCM^{+}23]. insurance [MAT^{+}23, WSL^{+}23]. Integrated [CASM23, TDB^{+}21, AABM^{+}20, CFB^{+}23, CSC^{+}23, FSS^{+}23, PIAF^{+}23, SCN^{+}23, TPCM^{+}24, Woo22]. Integrating [FOC^{+}23, FODD^{+}22, GCT^{+}23, MKSM23, MTTG^{+}23, TNM^{+}23, CBFS23, ECBJ22, MGO21, SRR21, Ano23g]. Integration [BPPKH20]. integrity [LBW^{+}20]. intelligence [Ano20a, BB20, MHES20]. intelligent [LSD^{+}20]. intensity [RPVT^{+}21]. Intensive [PDP^{+}23]. intensively [BJPP^{+}20]. intentions [CMD^{+}23a]. Inter [Ano22e, GGOP21, AA21, KTY^{+}21, KPM^{+}22, Moo23b, SH22, SBD^{+}21, GGOP22]. Inter-annual [Ano22e, GGOP21, KPM^{+}22, SH22, GGOP22]. inter-population [SBD^{+}21]. inter-species [Moo23b]. inter-specific [KTY^{+}21]. inter-temporal [AA21]. interact [Thi20, VPCM^{+}22]. Interaction [ZB21, BHWC^{+}20, FBF^{+}21]. Interactions [BCB^{+}22a, CSM^{+}23, BW21, BHL^{+}23, ECR^{+}21, GGA^{+}23, GCT^{+}21, KMK^{+}21b, MCH^{+}24, PFH^{+}22, RCJ^{+}22, RHW20, SFG22]. Interannual [LHL^{+}21, PRMS22, Ano22c, DASLRB22, TAB^{+}20, VLUB^{+}22]. interdisciplinary [AHC^{+}20]. interference [RAH^{+}22]. intergenerational [CCM^{+}21]. Internal [BRvL^{+}23, JRDW22, SMS^{+}22]. International [JBF^{+}20, CASM23, JBDK23, LCG^{+}22, CFB^{+}23]. interpretation [PCL^{+}21]. interpreting [MGO21]. interspecific [KLP^{+}24]. Intra [AA21, KTY^{+}21, KPM^{+}22]. Intra- [AA21, KTY^{+}21, KPM^{+}22]. Intraspecific [GSR^{+}20, KLP^{+}24]. intrinsic [CP20]. introgression [DFK^{+}22]. invading [VSL22]. invasion [DA20]. invasions [DFWR20]. invasive [DBT^{+}23, KWS^{+}21, QMMC^{+}23, SKM^{+}22]. inversion [JAQ^{+}23, KAK^{+}21]. invertebrate [BM20, CRB21, TWC^{+}20, TSW^{+}21, WBG^{+}22]. invertebrates [GGRB22, HCB^{+}24, ZYJZ23]. investigate [LWH^{+}21]. investigation [AABM^{+}20]. Investigations [CLMB22, SRMS21]. investment [CdSSO^{+}21]. involving [IEIT23]. Iranian [ECR^{+}21]. Ireland [HBD^{+}20, HBD^{+}21]. iris [EVLK^{+}21]. Irish [ECBJ22]. irrigation [MYMV20]. irroratus [GPJ^{+}21].

Isidella [GICP^{+}22]. Island [MFB23, RES^{+}23, Bir23, PRCM22, VSS^{+}22]. Islands [SHW^{+}21, ASW^{+}23, MWJ^{+}24, OOB^{+}22, SMP^{+}23]. isotope [vLNC20]. isotopes [ANA^{+}20, Ano24a, CT23, JDJK22, LMH22, MCC^{+}22, SHK^{+}23]. Italian [CBFS23]. item [HHA^{+}21]. iterative [PJCS^{+}23a]. iteroparity [BPC^{+}20]. Izu [HHNS21].

Jamieson [Ano23h, AAHT22]. Jansen [JHN^{+}21]. Japan [HHNS21, KSHO23]. japonicus [KTY^{+}21]. jellyfish [CFMB21, RBCD20]. jeopardizing [dBSQK21]. Jersey [SACP^{+}21]. Johan [KHSS21, SRMS21]. johnii [TDB^{+}21]. Johnsen [JG21, JHN^{+}21]. joint [BTM^{+}21, CdSSO^{+}21, JBE^{+}21, Mil24, OT22]. jurisdiction [BFJB21, CRV^{+}21]. jurisdictions [HHO^{+}20]. just [Bea23]. juvenile [ANA^{+}23, DYS21, GGLS22, HPS^{+}23, KBJA20, MSVSA22, SCP^{+}21, YFM^{+}20, YC20, vdWBL22]. juveniles [GEBS^{+}23].

K. [MGB^{+}23]. Kajikia [MGB^{+}23]. Katsuwonus [WCL^{+}23]. Kattegat [SBW^{+}22]. Keeping [BB22]. kelp [KMS^{+}20a, TRP^{+}21]. Kep [SWH^{+}23]. Kerguelen [SCMT^{+}22]. key [DDL^{+}20, GZW^{+}20, GMT21, Jen23, MTT^{+}23, MOC^{+}23, ODHF^{+}23, PVHV20, PMM^{+}23, SAB^{+}20]. Keys [CSM^{+}22]. kin [TCLB22]. King [YABM20, DSF^{+}21, PRCM22, SB21]. Kingdom [HBD^{+}21, SCG^{+}20]. know [FJC^{+}20]. Knowledge [Bir23, AVC^{+}21, BD20, CFMB21, CCH^{+}23, DBB^{+}20, EDH^{+}21, FKTM20, Leg21, MAE^{+}23, MKCA^{+}23, SROR21, VNBLZR^{+}23]. Knowledge-based [Bir23]. krill [GT22, KDS^{+}22, LRHW21, MTT^{+}23, SMP^{+}23, TFH^{+}21, TFWE^{+}22].

L [VBO^{+}21a]. L. [Ano23b, BPC^{+}21, CGH^{+}23, MMN^{+}22, MSN^{+}23, NBO^{+}20a, NBO^{+}20b, PAW22, TNS^{+}21, VBO^{+}21b]. laboratory [FFH^{+}22, MYJ^{+}20, PBC^{+}22, SAA^{+}22]. Labrador [MB21]. labrax [SCP^{+}21, dPHG^{+}23]. Lagenorhynchus [GHB^{+}23]. lagoon [BBB^{+}21, CCC^{+}21, LPA^{+}22]. lake [HBMC20, PTD^{+}20, HBMC20, PMSK^{+}20]. Lambert [KPS21]. Laminaria [KMS^{+}20a]. lance [BSW^{+}23, JLM^{+}23, SWS^{+}21, YMH^{+}24]. land [GT22]. land-based [GT22]. landing [LV23, MHMP^{+}23]. landings [ÁEPCL20, ASV21, BPPC^{+}20, SSH^{+}20, SFPP^{+}23]. Lanice [MYMV20]. Large [SWR^{+}22, Ano23i, BC23, KTK20, Lin21, RHO22, SMOD21, dAPDDJ21]. Large-scale [SWR^{+}22, RHO22, dAPDDJ21]. largest [EHV^{+}21]. larvae [EVY^{+}21, GPS^{+}21b, GPS^{+}21a, HPSF21, HHD^{+}23, TNS^{+}21]. Larval [SGM22, BPP^{+}21, CSM^{+}22, CCM^{+}21, EVLK^{+}21, Pep23, SSS^{+}23, VRH^{+}23]. last [vLNC20]. latent [BAS^{+}21]. latissima [JBJ^{+}22, JJB^{+}22, TRP^{+}21]. latitude [KTA^{+}23]. latitudinal [UVI^{+}21a, UVI^{+}21b]. Lawrence [BPP^{+}21, MSVSA22, SPJ23]. layer [PASL21]. leaders [SCM^{+}23]. leads [GPS^{+}21a]. learned [TCLB22]. Learning [CZSR21, DGM^{+}21, dMBJ22, ARHM21, ÁEPCL20, Ano20a, BXF23, CZ23, GPQ^{+}20, GHJ^{+}22, LSD^{+}20, LZWH20, POUH23, QWT^{+}21, RBC^{+}23, SSS^{+}20, Sch21a, YLC^{+}22, vEMvH^{+}21]. Learning-based [CZSR21, ARHM21, YLC^{+}22]. learnt [MKB^{+}23]. leaves [RBG^{+}20]. leaving [OGMA20]. legacy [ABB^{+}21a, ABB^{+}21b, JSR^{+}22, KHSS21, RCB^{+}21]. legitimacy [dBSQK21]. Lenfest [HKC^{+}23]. length [CMR^{+}20, HCH20, KMG22, KDS^{+}22, LMH22, MD22, MHKV20, Pep23, SHFK22, SHW^{+}21, TBH^{+}20, TMS23, ZMFS20, ZHL^{+}22]. length-based [CMR^{+}20, HCH20, KMG22, MD22]. length-dependent [SHFK22]. length-frequency [SHW^{+}21]. lengths [DDL^{+}20]. lenses [Ano24a, SHK^{+}23]. Lessepsian [DA20, DSDA22]. lesser [OBJ^{+}22, VCD^{+}20]. lessoniana [CCS^{+}24]. Lessons [RRR^{+}24, BPR^{+}23, Har20, MKB^{+}23, QMMC^{+}23, SPd^{+}22, TCLB22]. Leucoraja [MD22]. Leuresthes [MPQS20]. level [KMK21a, PEWN20, SKA15, SKA22]. levels [FND23, TAPB22, TRD^{+}22]. Leveraging [PJP20]. lewini [LMvEM23]. lice [JG21, JHS^{+}21a, JHS^{+}21b, SJM^{+}20, SBH^{+}21, SDSMS21, SG21, SBB^{+}22]. lice-induced [JG21, JHS^{+}21a, JHS^{+}21b]. Life [DBGP22, LLLS20, MKCA^{+}23, OH23, Ano22c, BSBH22, BPR^{+}20, BGR^{+}22b, CPD^{+}21, CCJ^{+}21, DASLRB22, FDK20, Gra21, JNT^{+}22, LHE20, Mun20, PWH23, Sha21, TRP^{+}21, ZEM21]. Life-history [DBGP22, OH23, FDK20, Mun20]. life-history-dependent [BPR^{+}20]. life-stage [CPD^{+}21]. lifetime [CT23]. light [KMK21a, Kra23, SWLW22]. light-mantled [SWLW22]. lighting [MTTG^{+}23]. lights [LQC^{+}22]. like [GPM21]. likelihood [FSC^{+}23, TMS23]. likely [BGR^{+}22b, JAQ^{+}23]. Limacina [MOC^{+}23]. limbatus [WJB^{+}21]. limit [Ano23e, PE21, SRA^{+}21, TRD^{+}22, WFG^{+}23, dGW^{+}22]. limited [CMR^{+}20, FDK20, FDMK21a, FDMK21b, FDMK21c, FDMK22, GPS^{+}21a, Har21, KSHO23, PEWN20, RCJ^{+}22, WOB^{+}23, dPMS^{+}22]. limits [DFWR20, ODHF^{+}23, PJP20, YARE23]. line [BFJB21, DDP^{+}21, FCD^{+}23, FOC^{+}23]. linear [NSS^{+}23]. linearity [FXG^{+}20, SOC^{+}20]. lines [TDC^{+}23]. linkages [CCM^{+}21]. linked [HCCR20, JAQ^{+}23, OTMN20, STK^{+}20b, STK^{+}20a]. Linking [FDK20, JMR^{+}22, FODD^{+}22]. links [RTQW20]. lionfish [QMMC^{+}23]. lions [CLS^{+}22, BCB^{+}22a]. Lipid [CP20, LRHW21]. literature [CLL21]. litter [MKB^{+}23, SBB^{+}23]. live [CZSR21]. lived [AHP^{+}21, WOB^{+}23]. livelihoods [RES^{+}23]. liver [MvdMS^{+}23]. lives [ABB^{+}21a, ABB^{+}21b]. living [DSR^{+}21, EDH^{+}21, Man21]. loathing [JSLC21]. Lobster [WML^{+}21, CAE^{+}21, CGH^{+}23, FPY^{+}22, FWA^{+}21, GZSP22, HCCR20, LSR^{+}21, LLPdL22, MBA^{+}20, MTS^{+}22, RRJ22, TBP20, TSWM23, VNR^{+}23, WSJ23]. lobsters [SKM^{+}22, WRC^{+}21]. Local [SHK^{+}21, AVC^{+}21, DNW^{+}23, JAQ^{+}23, KH22, MKCA^{+}23, NBA^{+}23, RAH^{+}22, SROR21, SBW^{+}22, TFWE^{+}22, VNBLZR^{+}23, vDBF^{+}20]. localities [RPVT^{+}21]. Localized [SHP^{+}22]. location [ANA^{+}20]. loci [GHÓ^{+}20]. logbook [SSH^{+}20]. loggerhead [BFD^{+}23]. Loligo [LAB^{+}22, SBA^{+}22]. Long [BJBB20, DSGG20, LB23, RBH22, BPC^{+}21, BLM^{+}21, DPMZ21, EDC20, HHL^{+}21, MGB^{+}23, MBP^{+}20, WCCY20, WZJ^{+}22, YSL22]. long-distance [BPC^{+}21, BLM^{+}21]. Long-range [RBH22]. long-standing [MGB^{+}23]. Long-term [BJBB20, DSGG20, LB23, DPMZ21, MBP^{+}20, WCCY20, WZJ^{+}22, YSL22]. longer [MYAH21]. longline [BJBB20, BFJB21, LMW^{+}22, QWT^{+}21, dSRKC22, ZB21]. loop [HLHC22]. loss [BHdL^{+}23, CZ23, GICP^{+}22, HYE^{+}22, SWR^{+}22]. losses [GAM^{+}22, RHW20]. lost [VBJ^{+}21]. Lough [PTD^{+}20]. louse [FFH^{+}22, FNB^{+}20, GKRB21, HJ22, JHN^{+}21]. low [Ano24b, CZSR21, CCM^{+}21, GGLS22, KMK21a, RVR^{+}24]. low-illumination [CZSR21]. Lowestoft [Hol21]. lowland [KBJA20]. luck [MR20b]. luminous [KMK21a]. lumps [Hol20, Hol21]. Lutjanus [PHB^{+}20, PFP^{+}22, TDB^{+}21]. Lyme [RRR^{+}24].

Machine [MHES20, RBC^{+}23, Ano20a, BXF23, BB20, CZ23, LWL^{+}22]. mackerel [GHÓ^{+}20, KTY^{+}21, KÓP^{+}21, PMO^{+}21a, PMO^{+}21b, TZE21, VPP^{+}23]. Macquarie [PRCM22]. macro [JSN^{+}22, SBB^{+}23]. macro-habitat [JSN^{+}22]. macroalga [ESK^{+}21]. macroalgae [HMBDP21]. macroepibenthic [TMD^{+}20]. Macrophytes [CATM22]. macrozoobenthic [ZMN^{+}21]. made [BD20, WLR^{+}20]. magellanicus [ZJH^{+}23]. magnitude [WJB^{+}21]. main [HBMC20]. Maine [JRDW22, KKS^{+}20, MTS^{+}22, WRC^{+}21, WBB^{+}23]. Maja [RPVT^{+}21]. major [BHdL^{+}23, DHN^{+}21]. Makaira [AMD^{+}23]. make [Mon22]. makers [LCG^{+}22]. making [Var21]. mako [Kai20]. Maldives [YABM20]. male [MB21, RDAOE20]. male-only [MB21]. mammal [SMLB^{+}22]. mammals [Bar23, BR22, Moo23b, PO24]. man [BD20, WLR^{+}20]. man-made [BD20, WLR^{+}20]. manage [GAP21]. Management [SHP^{+}22, WOL^{+}21, ZB24, Ano20b, AHP^{+}21, BLT23, Bar23, BBP^{+}23, BCC^{+}24, BJG^{+}21, BDB^{+}21, BFW^{+}21, BTM^{+}21, BP20, CASM23, CRR^{+}22, CMG^{+}23, DGM^{+}21, DSR^{+}21, Eay23, EDH^{+}21, FFR20a, FFR20b, FDMK21a, FDMK22, GPM21, GDI^{+}20, GBH^{+}24, GSR^{+}20, Gri20, GMT21, GFO^{+}20, HWL^{+}22, Har20, HBC^{+}22, HAPW21, HFP^{+}23, JMB^{+}22, JAB^{+}21, KLG^{+}23, KMS^{+}20a, LSR^{+}21, LKL^{+}21, LSOD^{+}20, MDAMPE^{+}21, MCS^{+}20, MYAH21, MD22, MMMB^{+}23, NAB^{+}22, OTLH20, OTL^{+}23, OP22, OMM^{+}23, PHB^{+}20, PBF^{+}21, PGRP23, PFT^{+}20, PJP20, RRR^{+}24, RKH^{+}21, RPVT^{+}21, ROH^{+}21, SBH^{+}21, SHG^{+}21, SLRC21, SPP21, SDGR^{+}20, TDB^{+}21, TPCM^{+}24, TBP20, TFWE^{+}22, VQN22, WPM^{+}23, WPN^{+}21, WSJ23, WPP^{+}21, WJB^{+}24, Woo22, WMB^{+}22, WCL^{+}23, ZB21, dLP21, vDHR^{+}22, dMBJ22, Ano22d]. managers [NKS^{+}23]. Managing [BSH^{+}22, BGR^{+}22b, dAPDDJ21]. mangrove [CWL^{+}20, HGS^{+}21]. mantled [SWLW22]. map [ARE^{+}22, Eay23, SRB^{+}23]. Mapping [GICP^{+}22, HK21, AVC^{+}21, MMMB^{+}23, SWH^{+}23]. maps [BHQF20]. marbled [HHL^{+}21]. marginal [Ano24a, SHK^{+}23, SJX^{+}20]. mariculture [KGG^{+}23, yMlZlL^{+}21]. marina [QSKL21]. Marine [BR22, DPH^{+}21, FFT^{+}22, HMK^{+}20, MTCC21, MR20a, PGRP23, PKT^{+}21, QBF^{+}21, SMLB^{+}22, SWH^{+}23, AJC^{+}23, ANA^{+}20, ANB^{+}21, ANB^{+}22, AGSS22, AHC^{+}20, Ano20a, Ano23c, BWBH21, Bar23, BNV^{+}21, BCC^{+}24, BB20, BRvL^{+}23, BFJB21, BPKM21, BPKM22, BPPC^{+}20, CP20, CMD^{+}23a, CMG^{+}23, CFMB21, CRD^{+}22, CNWW22, CRV^{+}21, CvPTAJ22, DA20, DSR^{+}21, DHN^{+}21, ESK^{+}21, EGMO^{+}20a, EGMO^{+}20b, EWD^{+}21, EWD^{+}22, FOC^{+}23, FMBL23, FMPH24, GT22, GHJ^{+}22, GHB^{+}23, GGRB22, GCG22, Gri20, GJU^{+}23, HWL^{+}22, HKZ^{+}23, HJS^{+}23, HVDvD23, Hol20, JAQ^{+}23, JSN^{+}22, JBDK23, KDS^{+}22, KOB^{+}22, KPS20, KDS23, LFH^{+}21a, LFH^{+}21b, LFH^{+}22, LS20a, LvDM^{+}22, LAGC23, LLPdL22, Lin21, LTMG23, LLALF22, LVT^{+}23, Lov20, MHES20, MMC^{+}23, MHJ^{+}20, MVNR^{+}21, MFV^{+}23, MTTG^{+}23, MTT^{+}23, MCX21, MPMGGAHR20, MKB^{+}23, Moo23b, MSP^{+}20, MFP20, MPH^{+}21, MKCA^{+}23, NHL^{+}22, NAB^{+}22]. marine [ODHF^{+}23, OHM^{+}23, PBD^{+}21, PAC^{+}23, PVGC^{+}22, PB21, PY22, PTD^{+}20, PDP^{+}23, PO24, RCB^{+}21, RBC^{+}23, SMOD21, SMJ^{+}21, SFW^{+}22, SDW^{+}21, SCW^{+}20, SCM^{+}23, SCB^{+}23, SROR21, SG21, SHG^{+}21, SJS21, SOO^{+}22, SRA^{+}21, SACP^{+}21, TSD^{+}21, TM20, VO23, VRM^{+}23, VAS^{+}20, VCF^{+}20, WBB^{+}23, WESWT21, YLC^{+}22, YFM^{+}20, YMAH20, ZL22, ZMMF22, dPMS^{+}22, vDHR^{+}22, Ano20b, LSOD^{+}20, MKCA^{+}23, RRR^{+}24, SCMT^{+}22]. mark [BFD^{+}23, VET^{+}22]. marker [MHKV20]. market [JYK^{+}21, SSF^{+}22]. markets [QMMC^{+}23]. Markov [TZR^{+}21, VAS^{+}20]. marlin [AMD^{+}23, FFT^{+}22, LVL^{+}22, MGB^{+}23]. Mass [AKK^{+}21, GLW21]. Match [BMM^{+}21, FND23]. material [KMK21a, LRHW21, TLS21]. matrices [CFP20]. maturation [FFH^{+}22, MDHB21, YCF^{+}23]. mature [Ano23k, GBMM^{+}23]. maturity [AWN22, CGH^{+}23, MB21, WRC^{+}21, dLP21]. Maurolicus [CKR^{+}21]. maximize [MYAH21]. maximizing [AA21]. maxims [LAA^{+}21]. Maximum [Ear21, FSN^{+}22]. may [BR22, Bra20, FDPD^{+}23, GS22, GSL22, Gal23, LBPC^{+}20, MPN20, NBO^{+}20c, RGE^{+}23, Sta22]. Maybe [Pec23]. me [Bar22, KSK^{+}21, Wap20]. meadow [MWS^{+}23]. meadows [CBFS23, MACSR22]. mean [DDL^{+}20, HCH20]. measure [TBH^{+}20]. measurement [GPQ^{+}20]. measurements [AKK^{+}21, BCBJ^{+}20, KAK^{+}21, MDHB21, SBPA^{+}21]. measures [BPV^{+}20, GVD^{+}23, Kra23, NSJ^{+}22, PFT^{+}20, RPVT^{+}21, SLRC21]. Measuring [LS20a, SRM^{+}20, PEWN20]. mechanical [BRvL^{+}23]. mechanism [SMOD21]. mechanistic [THSZ21]. mediated [JMB^{+}22]. medieval [PAW22]. Mediterranean [FFS^{+}22, VPB^{+}20, ADBC^{+}22, BFD^{+}23, BFFC20, BDL^{+}22, CRG^{+}23, DA20, DPMZ21, GA23, GAM^{+}22, HEHT^{+}22, LPA^{+}22, LPG^{+}20, LBPC^{+}20, MACSR22, MCOÁ^{+}22, MHMP^{+}23, OHM^{+}23, PRRB^{+}20, SGM22, TRD^{+}22]. meet [CFB^{+}21]. meeting [WOL^{+}21]. meets [KPS21]. megafauna [BM20, GCG22, JDD^{+}21, PLN20, PDP^{+}23, VNR^{+}23]. meltwater [NKA^{+}20]. memes [DeM22]. mentella [MSVSA22]. Merluccius [SdS20, TND^{+}20]. mesh [ASW^{+}23]. meso [GSWM^{+}21, GSWM^{+}22, SBB^{+}23, VCD^{+}20]. meso- [GSWM^{+}21, GSWM^{+}22]. meso-scale [VCD^{+}20]. Mesopelagic [KAK^{+}21, AKL^{+}21, AKK^{+}21, ARHM21, BCBJ^{+}20, CLL21, CKR^{+}21, KLSV21]. mesozooplankton [BCP^{+}21, LHL^{+}21]. meta [YY22]. meta-analysis [YY22]. Metabarcoding [EWD^{+}21, KMK^{+}21b, SDW^{+}21, AJC^{+}23, AMC^{+}23, CMD^{+}23b, DPMZ21, HJS^{+}23, HHNS21, MHA^{+}21, MGO21, PCL^{+}21, PASL21, EWD^{+}22]. metabolism [MSVSA22]. Metapenaeus [ZHL^{+}22]. metapopulation [CRD^{+}22]. metatranscriptomics [MKN^{+}21]. Metazoan [DPMZ21]. method [ARHM21, CFP20, GGRB22, KTK20, Köp23, LOÖK20, SHFK22, SBA^{+}22, SVMBR22, WGI^{+}21]. methods [BBB^{+}22, BMR21, CMR^{+}20, EMS^{+}21, FOC^{+}23, KSK^{+}21, LNW21, MGO21, MPA^{+}23, PDL20, QWT^{+}21]. métiers [SKH21]. metrics [DHB^{+}21, JBE^{+}21, PTR22]. Mexico [FFR20a, FFR20b, LL23, MHA^{+}21, PFP^{+}22, QMMC^{+}23, SMS^{+}21]. Micro [JSN^{+}22, SBB^{+}23]. Micro- [JSN^{+}22]. microbial [BOLPHL21, Thi20]. microbiome [RPB^{+}21]. microplastics [MFV^{+}23]. microsatellite [GHÓ^{+}20]. microscopic [LSD^{+}20]. microzooplankton [SMS^{+}21]. mid [TDC^{+}23, HMC^{+}20, PJCS^{+}23b, ZJH^{+}23]. Mid-Atlantic [HMC^{+}20, PJCS^{+}23b, ZJH^{+}23]. mid-water [TDC^{+}23]. middle [NFN20]. middle-age [NFN20]. migrating [KBJA20]. Migration [VET^{+}22, Ano24a, BSW^{+}23, BPC^{+}21, BBP^{+}23, BPR^{+}20, BLM^{+}21, JG21, JHS^{+}21a, JHS^{+}21b, NBL^{+}21, SHK^{+}23, TND^{+}20, dPHG^{+}23]. migratory [Cha21, MYW^{+}21, Zho21]. mild [Bra20, Bra22, NBO^{+}20c]. Mind [FCF^{+}21]. Mining [LDW23, GMT21, SCCA20]. misaligned [BDB^{+}21]. mismatch [FND23]. misreported [PDL20]. misreporting [SFPP^{+}23]. Missing [RTQW20, VBJ^{+}21]. misspecification [CMCF21, MVMdS21]. mitigate [FDPD^{+}23, GS22, GSL22, Gal23, Sta22]. mitigation [CCH^{+}23, LV23, PEWN20, Ste21, WGL^{+}23]. Mixed [BBP^{+}23, HPS^{+}23, Ano22d, BTM^{+}21, CRR^{+}22, ECBJ22, GDI^{+}20, GAP21, MV23, MKGC20, NHL^{+}20, dSRKC22, WPP^{+}21, WZJ^{+}22]. mixed-fisheries [GDI^{+}20]. mixed-population [MKGC20]. Mixed-stock [BBP^{+}23, Ano22d, CRR^{+}22, WPP^{+}21]. mixing [Ano24a, SHK^{+}23]. mixture [LHN^{+}22]. mobile [CPD^{+}21, ECBJ22, KTK20]. mobility [PLN20]. mobilization [RDM^{+}21]. Mobulid [CCH^{+}23]. Model [NFN20, BHWC^{+}20, BBB^{+}21, BDM24, BAS^{+}21, BNB21, BPPKH20, CWB^{+}23, CRD^{+}22, Eay23, HMF^{+}21, JHN^{+}21, KSK^{+}21, Ken21, KAK^{+}21, KPS20, LRHW21, MR20a, MKS^{+}21, MVMdS21, MCOÁ^{+}22, NHM^{+}21, NSS^{+}23, PDL20, RDM^{+}21, SOC^{+}20, TZR^{+}21, TBM^{+}22, TB21, VHT^{+}21, ZJF^{+}21, dPMS^{+}22, MPK^{+}22]. model-based [ZJF^{+}21, dPMS^{+}22].

Model-free [NFN20]. Modelled [HHD^{+}23, HFP^{+}23]. Modelling [CMS^{+}23, FMPH24, JJB^{+}22, PM20, PPMD22, SASM22, BJPP^{+}20, CCC^{+}21, DGM^{+}21, DPH^{+}21, Har21, dlTVLC^{+}21, JBJ^{+}22, KSHO23, LL23, MYW^{+}21, MSP^{+}20, MBSR20, OGMA20, PCC23, SBH^{+}21, SMJ^{+}21, SVML^{+}23, SBD^{+}21, SCMT^{+}22, SPP21, TMLH21, WJ22, YCF^{+}23, ZCX^{+}20, vDBF^{+}20]. models [Ano23g, BLT23, BCC^{+}24, CSM^{+}22, CMCF21, DFWR20, FRSM22, FSC^{+}23, GC21, GZW^{+}20, GCT^{+}23, HBFO20, HCH20, JAB^{+}21, KLG^{+}23, LZWH20, LHN^{+}22, MKSM23, MBP^{+}20, NHL^{+}20, OTMN20, ODT^{+}22, OTL^{+}23, OT22, PIAF^{+}23, RKH^{+}21, RUdPFRA^{+}21, SRBT23, SFPP^{+}23, SRA^{+}21, SRR21, TAB^{+}20, Tho22, VAS^{+}20, VNBLZR^{+}23, WBA^{+}22, CBFS23]. moderate [BMR21]. moderately [FDMK22]. modern [PAW22]. modes [SHFK22]. modification [DCL^{+}21]. modified [SG21]. modifies [HGS^{+}21]. modify [MGO^{+}23]. molecular [BPKM21, BPKM22, KLF^{+}20a, KLF^{+}20b, MGO21]. molluscs [GA23]. mollusk [CWL^{+}20]. molt [MB21]. molt-type [MB21]. Molva [MMH^{+}23]. Monetary [PFT^{+}20]. monitor [PCH^{+}24, TMD^{+}20]. monitored [HFP^{+}23, MPQS20]. Monitoring [Met20, Ano22d, Ano24c, CCR^{+}24, CRR^{+}22, CMD^{+}23b, Dom21, FW22, FMF^{+}20, HJS^{+}23, HJP22, HMNF22, KPM^{+}22, LFH^{+}21a, LFH^{+}21b, LFH^{+}22, LBW^{+}20, MSWB22, QWT^{+}21, RRR^{+}24, SSH^{+}20, SOB23, TSD^{+}21, TK20, WBB^{+}23]. month [WCL^{+}23]. morhua [Ano22b, Ano22d, Ano23b, BCB^{+}22b, CSH^{+}21, CRR^{+}22, DSGG20, DYS21, FKTM20, GGLS22, HPS^{+}23, HHA^{+}21, JBF^{+}20, KTA^{+}23, LAGC23, LOÖK20, MMN^{+}22, MSN^{+}23, MvdMS^{+}23, NBO^{+}20a, NBO^{+}20b, PAW22, SHFK22, SBD^{+}21, SSV^{+}22b, vLNC20]. morio [BJPP^{+}20]. morphological [HHNS21, KLF^{+}20a, KLF^{+}20b, MKN^{+}21, MGO21]. morphology [DSGG20]. mortality [BBB^{+}22, BM20, BSW22, BGH^{+}20, FFH^{+}22, GGLS22, GLW21, HMOR22, HPSF21, HCH20, JG21, JHS^{+}21a, JHS^{+}21b, KKS^{+}20, LZWH20, PRMS22, TDC^{+}23, TZE21, TB21, ZEM21]. Morus [dGW^{+}22]. mosaic [TDB^{+}21, TBP20]. most [MSVSA22, Var21]. mostly [KH22]. motion [SSS^{+}20]. motivates [CMH^{+}23]. motivations [SCB^{+}23]. Moulding [DEL21]. moult [CAE^{+}21]. mounted [LRGW23]. mouth [Pep23]. Move [dCKS^{+}22]. Movement [FBF^{+}21, JBE^{+}21, WBWS21]. movements [FFT^{+}22, SCP^{+}21]. moving [PE21, SCB^{+}23]. MSP [PGRP23]. MSY [PF21, RUdPFRA^{+}21, TBM^{+}22, ZRW^{+}21]. MSY-based [ZRW^{+}21]. much [Pec23]. mud [BPR^{+}23]. muelleri [CKR^{+}21]. Multi [NHM^{+}21, SBA^{+}22, SWLW22, WESWT21, AHK23, CKH^{+}21, GDI^{+}20, HSS21a, HSS21b, OOB^{+}22, SOC^{+}20, TBH^{+}20, WCL^{+}23]. Multi-decadal [SWLW22]. Multi-fleet [NHM^{+}21]. multi-frequency [CKH^{+}21]. Multi-method [SBA^{+}22]. multi-model [SOC^{+}20]. multi-month [WCL^{+}23]. Multi-scale [WESWT21]. multi-sea [OOB^{+}22]. multi-species [AHK23, HSS21a, HSS21b]. multi-stage [TBH^{+}20]. multi-stock [GDI^{+}20]. multiannual [MCOÁ^{+}22]. multibeam [BMU^{+}23, HK21, WBWS21]. Multidecadal [LIR^{+}21, Zho21]. multidisciplinary [PJCS^{+}23b]. multifaceted [SRR21]. multifrequency [BWJ^{+}20]. multigear [SMJ^{+}21]. multiple [BGR^{+}22b, HKZ^{+}23, HEHT^{+}22, NAS^{+}20, OTMN20, ODT^{+}22, PFH^{+}22, RPVT^{+}21, WOL^{+}21]. Multispecies [MKK^{+}24, FSN^{+}22, KLG^{+}23, MHMP^{+}23, OBK21, SSF^{+}23, VQN22, WZJ^{+}22, ZCX^{+}20]. multivariate [TMS23]. multivariate-Tweedie [TMS23]. muscle [MvdMS^{+}23]. mussel [DLLC^{+}22]. mussels [HMWS22]. My [Gra21, KSK^{+}21]. myctophid [SSS^{+}23]. Mytilus [DLLC^{+}22].

naevus [MD22]. narrative [Abe20]. narratives [HKC^{+}23]. narrow [ZMFS20]. Narrowing [ZMMF22]. Nassau [SHW^{+}21]. natal [dAdSM21]. nation [FCF^{+}21]. National [FFT^{+}22, BFJB21, CRV^{+}21, MKB^{+}23]. Nations [PY22, PMTG^{+}23]. native [DSDA22, KWS^{+}21, SKM^{+}22]. Natural [BSW22, DFK^{+}22, LZWH20, CBFS23, CAE^{+}21, Dol20, KVS^{+}23, PM20, SSBV21, SCB^{+}23, TB21]. Nature [PMM^{+}23]. Nature-based [PMM^{+}23]. nauplii [MPN20]. Navigating [MMC^{+}23, Kos21]. Near [SPJ23, BMN^{+}21, RMF^{+}22]. Near-bottom [SPJ23]. nearshore [RFB^{+}22]. necessarily [RST^{+}21, SMK23]. necessity [PM20]. need [Bea22, BGR^{+}22b, FJC^{+}20, GT22, LTMG23, SSV^{+}22a]. needed [DBB^{+}20, HBC^{+}22]. needs [PF21, SHG^{+}21, TNM^{+}23]. negative [GT22]. negatively [CCM^{+}21, KMK21a]. neglected [Dol20]. nektonic [CASM23]. Nematoscelis [KLSV21]. Nephrops [FAV^{+}20, VNR^{+}23, WSJ23]. neritic [BFD^{+}23, IVU^{+}22]. net [AKK^{+}21, HJP22, Var21, ZL22]. netting [CMD^{+}23b, KMK21a]. Network [RKH^{+}21, CRV^{+}21, FRSM22, FSS^{+}23, HKZ^{+}23, HJ22, NAB^{+}22, PVHV20, SSS^{+}20, SCN^{+}23, SRR21, VNBLZR^{+}23, WJB^{+}24]. network-based [SSS^{+}20]. networks [BWJ^{+}20, FMF^{+}20, LTK20, OdS21, TK20, ZYJZ23]. neural [BWJ^{+}20, FMF^{+}20, LTK20, SSS^{+}20, SCN^{+}23, TK20, ZYJZ23]. Neutral [MDAMPE^{+}21]. Newfoundland [CWB^{+}23, HJP22, MB21, MAC21]. Next [DSR^{+}21]. Next-generation [DSR^{+}21]. niche [Day20, KÓP^{+}21]. night [EDH^{+}22, LQC^{+}22]. nigricans [AMD^{+}23]. Niño [FWA^{+}21]. NIR [PHB^{+}20]. No [KMS20b, MFV^{+}23, PBS^{+}23, WBB^{+}21, CCM^{+}21, CFP20, JSR^{+}22, LLPdL22, OHM^{+}23, PF21, SBW^{+}22, VO23, VNR^{+}23, dBABS^{+}22]. no-take [LLPdL22, OHM^{+}23, SBW^{+}22, VO23, VNR^{+}23, dBABS^{+}22]. noise [MMN^{+}22, VAS^{+}20]. noltei [Ano24b, RVR^{+}24]. Non [SOC^{+}20, TSE^{+}21, Ano20c, BM20, CWC^{+}23, DeM22, FXG^{+}20, KWS^{+}21, NSS^{+}23, ORGTO^{+}23, WBA^{+}22, ZRW^{+}21]. non-equilibrium [Ano20c]. non-focal [CWC^{+}23]. non-genetically [DeM22]. non-linear [NSS^{+}23]. Non-linearity [SOC^{+}20, FXG^{+}20]. non-native [KWS^{+}21]. non-panmictic [ORGTO^{+}23]. Non-stationary [TSE^{+}21, WBA^{+}22, ZRW^{+}21]. non-target [BM20]. Nondaily [DGM^{+}24]. nonlinear [Ano20c, MBSR20]. nonlocal [THSZ21]. Nordic [SMLB^{+}22]. Normalizing [CvPTAJ22]. North [BPC^{+}21, BJPP^{+}20, KMK^{+}21b, LAB^{+}22, PMS^{+}20, RDM^{+}21, Ano22e, GHÓ^{+}20, GGOP21, GGOP22, YFM^{+}20, AMD^{+}23, Ano24a, Ano24c, BVV^{+}20, CZS^{+}23, CBB21, CGH^{+}23, FND23, FFS^{+}22, FNB^{+}20, FJC^{+}20, HRB^{+}21, HFP^{+}23, HHA^{+}21, JKF^{+}22, JBF^{+}20, Kai20, KLF^{+}20a, KLF^{+}20b, LH20, LSR^{+}21, MEOA23, MDP^{+}21, Moo23a, MFP20, OTL^{+}23, RHvD^{+}20, SH22, SHK^{+}23, SDW^{+}21, SPBJ21, SPJ23, SCG^{+}20, TSE^{+}21, VET^{+}22, WBG^{+}22, WLR^{+}20, Zho21, vdWBL22]. North-east [HHA^{+}21]. northeast [MMC^{+}23, MWS^{+}23, SWS^{+}21, BSH^{+}22, DSGG20, EVY^{+}21, FRP^{+}23, KBW22, KTA^{+}23, LLALF22, MMH^{+}23, SBC^{+}21a, SBC^{+}21b, SDR^{+}21, SPP21, TZE21, UPH^{+}21]. Northeastern [GSWM^{+}21, GSWM^{+}22, SMS^{+}21]. Northern [dGW^{+}22, Ano22c, CZS^{+}23, DASLRB22, HWS^{+}21, HMOR22, HCM24, Hut22, JLM^{+}23, ORGTO^{+}23, PAW22, SDGR^{+}20, VPP^{+}23, CAE^{+}21, FAV^{+}20, WESWT21]. northwards [VLD^{+}21]. northwest [BBM^{+}22, BPC^{+}20, JLM^{+}23, SFPP^{+}23, ADS^{+}23, HJP22, KEW^{+}21, YCF^{+}23, ZJF^{+}21]. northwestern [BMB20, NKA^{+}20]. norvegicus [FAV^{+}20, VNR^{+}23, WSJ23]. Norway [JG21, CAE^{+}21, FOC^{+}23, JHS^{+}21a, JHS^{+}21b, LSR^{+}21, PAW22, VNR^{+}23, VCF^{+}20, WSJ23]. Norwegian [Ano22a, BCB^{+}22b, DLA^{+}21, DBT^{+}23, DFK^{+}22, IHL21, PFH^{+}22, SJM^{+}20, SDSMS21, SHE^{+}22, SGM^{+}24, TNS^{+}21, YSL22]. note [Ano20b, Ano22e]. Notothenia [HHL^{+}21]. Nototheniidae [HHL^{+}21]. Novel [SRMS21, CNWW22, CLS^{+}22, HPSF21, VCF^{+}20]. nudges [MYJ^{+}20, SCH^{+}21b]. numbers [PRCM22]. Numerical [Kai20]. numerically [KDS^{+}22]. nursery [LAB^{+}22]. nutrient [ESK^{+}21, KPS20, QSKL21, Thi20]. nutrient-phytoplankton [KPS20]. Nutritional [ANB^{+}21, SAM^{+}22, ANB^{+}22]. NW [BQBB^{+}21, TND^{+}20, VPB^{+}20].

OASIS [CSM^{+}23]. object [vEMvH^{+}21]. objectives [FODD^{+}22, WOL^{+}21]. obligation [LV23]. obscures [BKR^{+}21]. observation [PLN20]. observations [LCG^{+}22, dPMS^{+}22]. observatory [CRJ20, LBW^{+}20]. observed [LRGW23, SBD^{+}21]. observer [Ano23g, GCT^{+}23, PO24]. observing [PJCS^{+}23b, CSM^{+}23]. obtained [HK21]. occlusion [vEMvH^{+}21]. occupancy [SDR^{+}21]. occurrence [DFK^{+}22, HBD^{+}21, MPA^{+}23]. occurring [JKF^{+}22]. Ocean [DNW^{+}23, ESK^{+}21, EVLK^{+}21, HKC^{+}23, PY22, SCM^{+}23, dPMS^{+}22, Ano23c, AABM^{+}20, BJBB20, BKR^{+}21, BMN^{+}21, CMD^{+}23b, CSM^{+}23, DBGP22, DSR^{+}21, DFWR20, GPS^{+}21b, GPS^{+}21a, GZW^{+}20, GGRB22, HMBDP21, JSLC21, KHHD20, KJP^{+}22, PJCS^{+}23a, LCG^{+}22, LXBG21, MOC^{+}23, MKK^{+}24, MFA^{+}24, MHO^{+}20, MGO^{+}23, NKA^{+}21, NKS^{+}23, PJCS^{+}23b, PBC^{+}22, PBS^{+}23, PAC^{+}23, QBF^{+}21, Sch21a, SG21, SOCN^{+}22, WM20, AMD^{+}23, Ano23l, BPC^{+}24, BCBJ^{+}20, BHL^{+}23, BQBB^{+}21, CBT^{+}20, CRG^{+}23, DSF^{+}21, FPY^{+}22, FBF^{+}21, FFT^{+}22, GJU^{+}23, LWZC23, MVMdS21, MKCA^{+}23, SK21, SWLW22, UPH^{+}21, VBJ^{+}21, WCLB23, WCL^{+}23, YABM20, dAdSM21]. ocean-biogeochemical [GZW^{+}20]. oceanic [BLM^{+}21, Kos21, NKA^{+}20, OOB^{+}22, OGMA20]. oceanica [CBFS23, MACSR22, TRD^{+}22]. Oceanographic [BCP^{+}21, ODT^{+}22]. oceanography [DHN^{+}21, KTA^{+}23, dlCRT^{+}22]. oceans [HBC^{+}22, Moo23b, GRE^{+}20]. Octopus [DGM^{+}24, RUdPFRA^{+}21]. off [Ano22c, CWB^{+}23, CHZC22, DASLRB22, EGMO^{+}20a, EGMO^{+}20b, GSWBK20, HAPW21, HHNS21, HBD^{+}20, IVU^{+}22, JSR^{+}22, KSHO23, LWH^{+}21, MMAE20, MYAH21, ODHF^{+}23, SMP^{+}23, TND^{+}20, VSS^{+}22]. off-site [LWH^{+}21, ODHF^{+}23]. offs [OBK21, SPP21, WPP^{+}21]. Offshore [BHDD22, Ano20b, BVV^{+}20, BD20, DBB^{+}20, HWS^{+}21, LBPC^{+}20, LBH^{+}21, LSOD^{+}20, Met20, MPK^{+}22, PRD20, RRJ22, RHW20, SBM^{+}22, SFG22, SCG^{+}20, TSWM23, TWC^{+}20, TSW^{+}21, VCD^{+}20, WBG^{+}22, YABM20]. offspring [GPM21]. oil [MvdMS^{+}23, RHW20]. old [CvdWC^{+}20, MR20a]. Oman [ECR^{+}21]. omnidirectional [PMO^{+}21a, PMO^{+}21b]. omnivory [LMH22]. on-site [LWH^{+}21]. Oncorhynchus [DBT^{+}23, YFM^{+}20]. only [MB21, SRA^{+}21]. onset [dLP21]. Onshore [ASV21]. ontogeny [SMOD21]. opalescens [SSF^{+}22]. open [LQC^{+}22]. operating [SMOD21]. operation [WBG^{+}22]. operational [KNP21]. Operationalizing [DDP^{+}21, JAB^{+}21]. opercula [CT23]. opilio [Ano23f, DSMM^{+}22, HHA^{+}21, HFT^{+}22, HHD^{+}23, MB21, SMOD21]. opinion [SHG^{+}21]. opportunistic [GGA^{+}23, HCM24, OOE^{+}22]. Opportunities [OMM^{+}23, HMK^{+}20, MPMGGAHR20, PMTG^{+}23, TNM^{+}23, UPH^{+}21]. opportunity [CPD^{+}21, MR20b, YMAH20]. opposed [OTLH20]. Opposite [UVI^{+}21b, UVI^{+}21a]. Optimal [dSRKC22, AHK23, BP20]. Optimising [LFH^{+}21b, LFH^{+}22, LFH^{+}21a]. optimism [BMB20]. optimization [OBK21]. optimize [FDMK21b, FDMK21c, GETvB21, SAA^{+}22]. optimized [BMM^{+}21]. Optimizing [TMD^{+}20]. option [dLP21]. options [HWL^{+}22, HEHT^{+}22, NHM^{+}21, WMB^{+}22, vDHR^{+}22]. orange [LHE20, SK21]. orange-spotted [LHE20]. organic [PBS^{+}23]. organisms [MCH^{+}24]. origin [DSMM^{+}22, HPS^{+}23, WJB^{+}21]. originally [GC21]. origins [dAdSM21]. Orkney [SMP^{+}23]. ornate [FPY^{+}22]. ornatus [FPY^{+}22]. oscillations [BJBB20]. Otago [MSW21]. other [BBB^{+}22, DeM22, HAC^{+}23, LRHW21, RHW20, SRMS21]. Otolith [VDAT20, CZ23, DSGG20, VBO^{+}21a, VBO^{+}21b]. otoliths [BDM24, PAW22]. otter [WCCY20]. our [Mon22, Moo23b]. outbreaks [DPD^{+}22a, DPD^{+}22b]. outcomes [MB21, TMG20]. outdoor [EDC20]. outlook [GHJ^{+}22]. outperform [SLRC21]. output [DSF^{+}21]. ovarian [WRC^{+}21]. over-exploitation [WML^{+}21]. Overestimating [HFP^{+}23]. overexploitation [HHL^{+}21]. overexploited [EGMO^{+}20a, EGMO^{+}20b]. overfished [VNR^{+}23]. overfishing [Lin21, PJP20]. overlap [BMU^{+}23]. overlap-induced [BMU^{+}23]. overlapping [CCS^{+}24]. overlooking [ZEM21]. overrides [GZW^{+}20]. overview [CFMB21, GHJ^{+}22, PMM^{+}23, RBC^{+}23, VSS^{+}23, WGL^{+}23]. oxy [IOS21]. oxy-thermal [IOS21]. oxygen [MCC^{+}22, PLN20, PTR22, PTD^{+}20, vLNC20]. oyster [CCM^{+}21, DBGP22, GPS^{+}21b, GPS^{+}21a, JWR^{+}21, KWS^{+}21].

P [vdBDH21]. Pacific [BHL^{+}23, FPY^{+}22, GRE^{+}20, Ano23l, Ano23k, Ano24a, ADG^{+}20, BSW^{+}23, Bea22, CSC^{+}23, DBGP22, FFT^{+}22, GBMM^{+}23, JWR^{+}21, Kai20, KWS^{+}21, KJP^{+}22, LWZC23, MDHB21, MVMdS21, MWS^{+}23, NFN20, ORGTO^{+}23, RES^{+}23, SHK^{+}23, SRBT23, SWLW22, SRA^{+}21, SPP21, VSS^{+}22, WCLB23, WGCC^{+}22, WCL^{+}23, YMH^{+}24, ZB24]. pagurus [MDP^{+}21]. pair [CCR^{+}24, JBE^{+}21]. Paired [SOB23]. Palau [FFT^{+}22]. Pallas [MYMV20]. pallasii [MDHB21]. Pan [VHT^{+}21]. Pan-Arctic [VHT^{+}21]. Pandalus [HWS^{+}21]. pandemic [BMM^{+}22, SFW^{+}22]. panmictic [ORGTO^{+}23]. panmixia [WSJ23]. Panopea [dlTVLC^{+}21]. Panulirus [FPY^{+}22, LLPdL22]. paradigm [MBS^{+}23]. parallels [SRMS21]. parameters [KP24, LBW^{+}20, ZMFS20]. parasite [CRD^{+}22, GKRB21]. parasite-control [GKRB21]. Park [Ano20b, LSOD^{+}20, PGRP23]. partial [TND^{+}20]. partially [SBW^{+}22]. participants [SRM^{+}20]. participation [JBDK23, Mac22]. Participatory [MMMB^{+}23, CvPTAJ22]. Participatory-based [MMMB^{+}23]. particle [RTQW20]. particulates [CRJ20]. partitioning [KÓP^{+}21, MVNR^{+}21]. partners [JBE^{+}21]. partnership [BKF^{+}20]. passages [VAS^{+}20]. passive [SRB^{+}23, SOB23]. past [BCB^{+}22b, GICP^{+}22, KHSS21, MR20a, MACSR22, PDH^{+}21, SCG^{+}20]. Patagonia [TND^{+}20]. Patagonian [EMS^{+}21]. patagonicus [PRCM22]. patches [RAH^{+}22]. paternity [RPVT^{+}21]. path [DGM^{+}21]. pathogens [VLD^{+}21]. pathway [PMM^{+}23]. pathways [BDL^{+}22, CMD^{+}23a]. pattern [BHDD22, GCT^{+}21, VCD^{+}20, lCRGCCA23, Ano23d]. Patterns [BMB20, AMC^{+}23, Ano24a, Ano24c, BP20, CP20, CWL^{+}20, GPM21, GSWM^{+}21, GSWM^{+}22, GPV^{+}22, HLHC22, LOÖK20, NHL^{+}20, OHM^{+}23, OBJ^{+}22, PDP^{+}23, SHK^{+}23, SCCA20, SSH^{+}20, VET^{+}22, VPB^{+}20, ZJF^{+}21]. pCO [EDC20, JWR^{+}21, LXBG21]. peak [VSL22]. pelagic [ARHM21, BSW^{+}23, BJBB20, CHZC22, CWC^{+}23, CMD^{+}23b, GBGA^{+}23, GHB^{+}23, HYE^{+}22, JBE^{+}21, KWEB21, LBW^{+}20, LRGW23, MRC^{+}21, MTK^{+}21, MPA^{+}23, OH23, PFH^{+}22, RFB^{+}22]. pelamis [WCL^{+}23]. Penaeus [PDH^{+}21]. penguin [BRG22, PRCM22, SHP^{+}21]. Penguins [SHP^{+}22]. Peninsula [QCM^{+}20, dlCRT^{+}22, WESWT21]. People [AAHT22, Fie21, JSLC21, Ano23h, JLA22]. Per-recruit [Har21]. Perca [HMOR22]. perceived [JEPT23]. perceptions [MAT^{+}23, Mun20, OTLH20, PVHV20, PGRP23, RSTL22]. perch [HMOR22]. Performance [CMR^{+}20, MKGC20, BMR21, FDK20, LVG^{+}20, LLLS20, PEWN20, PVHV20, PCH^{+}24, PPMD22, SAB^{+}20, SL21]. period [VSL22]. periods [KHSS21, LVL^{+}22]. permanent [Ano22c, DASLRB22]. perseverance [MR20b]. Persian [BHWC^{+}20]. persistence [CHZC22]. Persistent [JKF^{+}22, MAC21]. persisting [JAQ^{+}23]. personal [RBH22]. personatus [BSW^{+}23]. perspective [FMPH24, SCG^{+}20, vdBDH21]. perspectives [LAA^{+}21, SFW^{+}22, SAM^{+}22, Suu22, WJ22]. perturbations [LL23]. Peruvian [KGG^{+}23]. Peterman [SVMBR22]. petrels [BPC^{+}24]. petroleum [TWC^{+}20, TSW^{+}21]. pH [CCM^{+}21, RPB^{+}21, TAPB22].

Phalacrocorax [KBJA20]. phase [Ano23i, BC23]. Phenological [CPD^{+}21, UVI^{+}21a, UVI^{+}21b]. Phenology [LPA^{+}22, FRP^{+}23, SH22, SJS21, TAB^{+}20]. phenotypes [KH21]. phenotypic [BØS^{+}21, DEL21, FWA^{+}21]. philippinarum [BBB^{+}21]. photic [Thi20]. photogrammetry [MHKV20]. photorespiration [RBG^{+}20]. photosynthesis [PBS^{+}23, RBG^{+}20]. physical [DHN^{+}21, SJX^{+}20, TFWE^{+}22]. Physiological [HMBDP21, CWC^{+}23, JWR^{+}21]. Phytoplankton [LNW21, HMWS22, KPS20, LSD^{+}20, SJX^{+}20, ZJF^{+}21]. picture [Bea22]. pinch [MR20b]. pink [DBT^{+}23, MDAMPE^{+}21]. pinnatifida [ESK^{+}21]. pinniped [RCH^{+}21]. pipelines [LH20, RHW20]. Placopecten [ZJH^{+}23]. plaice [BHDD22]. plan [MKB^{+}23]. Plankton [CRJ20, HJP22, OBJ^{+}22, KMK^{+}21b, SMB^{+}21, YLC^{+}22]. planning [AVC^{+}21, CFB^{+}21, NHL^{+}22, QBF^{+}21, SLB^{+}20, Woo22, ZMMF22]. Plastic [BHdL^{+}23, CNWW22]. Plasticity [BBE^{+}20, BFD^{+}23, DEL21]. plasticized [MFV^{+}23]. plastics [MFA^{+}24]. Plateau [SCMT^{+}22]. platessa [BHDD22]. platform [TWC^{+}20, TSW^{+}21]. platforms [CvdWC^{+}20]. players [HSMK21]. pleasure [Leg21]. Pleuronectes [BHDD22]. plume [NKA^{+}20]. plumes [MYMV20]. pointhead [YMH^{+}24]. points [Ano23a, HS20, MD22, OTMN20, SVBM^{+}22, SVMBR22, TBM^{+}22, ZRW^{+}21]. polar [EDH^{+}22]. polemics [Sha21]. Policy [Moo23a, GKRB21, HKZ^{+}23, KOB^{+}22, LCG^{+}22, RRR^{+}24, SMK23, vDHR^{+}22, Ear21]. policy-based [vDHR^{+}22]. political [Ear21]. pollock [DLL^{+}21, IEIT23, LDW23]. Pollution [LZZ^{+}21]. polycyclic [LZZ^{+}21]. polyvalent [SRB^{+}23]. pomfret [BHWC^{+}20]. Poor [UPH^{+}21, BMR21, ECBJ22, KVS^{+}23, MGF^{+}21, PWH23]. pop [KKS^{+}20]. pop-up [KKS^{+}20]. Population [BQBB^{+}21, GPJ^{+}21, GEBS^{+}23, KJP^{+}22, MPQS20, MMH^{+}23, WML^{+}21, BDB^{+}21, CCS^{+}24, DHB^{+}21, DRP^{+}23, GPM21, JYK^{+}21, JLM^{+}23, KBK^{+}22, LvDM^{+}22, LRHW21, MD21, MDAMPE^{+}21, MCS^{+}20, MKK^{+}24, MKGC20, OTMN20, PFP^{+}22, QCM^{+}20, RRJ22, RCH^{+}21, SKA15, SKA22, SBD^{+}21, SDGR^{+}20, TRD^{+}22, VBO^{+}21a, VBO^{+}21b, WSJ23, dPHG^{+}23]. population-dynamics [OTMN20]. population-level [SKA15, SKA22]. population-scale [MKK^{+}24]. populations [Ano23k, BØS^{+}21, BPC^{+}20, Cha21, DFK^{+}22, GRE^{+}20, GBMM^{+}23, HUJ^{+}21, HWS^{+}21, PBD^{+}21, ROH^{+}21, RGE^{+}23, SdS20, SOB23, SAA^{+}22, TDB^{+}21, VSL22, VNR^{+}23]. portfolio [AHK23]. ports [STS^{+}20]. Portugal [BPPC^{+}20]. Portunus [MHJ^{+}20]. pose [LBPC^{+}20]. Posidonia [CBFS23, MACSR22, TRD^{+}22]. position [BOLPHL21, LMH22]. positively [CCM^{+}21]. possibilities [DSMM^{+}22]. Possible [DHN^{+}21, MPH^{+}21, LVG^{+}20, MAC21]. Post [WPN^{+}21, JG21, JSN^{+}22, JHS^{+}21a, JHS^{+}21b, LVL^{+}22, UPH^{+}21]. Post-release [WPN^{+}21, LVL^{+}22]. post-smolt [JG21, JHS^{+}21a, JHS^{+}21b]. post-smolts [JSN^{+}22, UPH^{+}21]. pot [BHdL^{+}23]. Potential [HWL^{+}22, SKA15, BBB^{+}21, BFJB21, Cha21, CVJ^{+}20, DSF^{+}21, GT22, GHJ^{+}22, HHD^{+}23, KBW22, MVMdS21, NAS^{+}20, NKA^{+}21, ODHF^{+}23, PBF^{+}21, PMTG^{+}23, SSF^{+}23, TMG20, SKA22]. potentially [BVV^{+}20]. pots [CSH^{+}21]. practical [TCLB22]. practice [CFB^{+}23, FP20, RRJ22]. practices [Bar23, DCL^{+}21, JEPT23]. prawn [PDH^{+}21]. prawns [ZHL^{+}22]. Pre [TZE21, DPH^{+}21, LVG^{+}20, PMTG^{+}23]. Pre-catch [TZE21]. pre-COVID-19 [PMTG^{+}23]. pre-fishing [DPH^{+}21]. pre-recruit [LVG^{+}20]. precautionary [Ano23e, FDMK21a, WOB^{+}23, WFG^{+}23]. precision [UBD^{+}23, WGI^{+}21]. predation [HMOR22, KBJA20]. Predator [NDE^{+}21, GHB^{+}23, KMK^{+}21b, RCJ^{+}22, WDPM23]. predators [GT22, MRC^{+}21, RCH^{+}21]. predatory [ANB^{+}21, ANB^{+}22]. predict [DPH^{+}21]. Predicted [ROH^{+}21, JRPHM^{+}22, MDHB21, dlCRT^{+}22]. Predicting [BAS^{+}21, CLS^{+}22, SRBT23, SHE^{+}22]. Prediction [BNB21, SJM^{+}20, Gri20, MTTG^{+}23]. predictions [Ano23g, GCT^{+}23, NKS^{+}23, SGM22]. predictive [SRA^{+}21]. prefer [SKM^{+}22]. preference [HAP^{+}21]. preferred [RHO22]. preparing [WBB^{+}23]. Presence [SRA^{+}21, WPB20]. Presence-only [SRA^{+}21]. presences [GICP^{+}22]. present [KHSS21, RCB^{+}21, SCG^{+}20]. present-day [RCB^{+}21]. preservation [TRP^{+}21]. pressure [CRB21, FXG^{+}20, LLPdL22, SJM^{+}20, SDSMS21, SFW^{+}22]. pretty [GDI^{+}20]. prevalence [KBK^{+}22, RPVT^{+}21]. prevalent [BPC^{+}21]. previous [LDW23]. Prey [KVS^{+}23, Ano23f, CPD^{+}21, CLS^{+}22, DDL^{+}20, HFT^{+}22, KMK^{+}21b, KWEB21, NDE^{+}21, OBJ^{+}22, RCJ^{+}22, SKM^{+}22, SPBJ21, SWS^{+}21, TM20, dGW^{+}22, dlCRT^{+}22, MPH^{+}21]. Prey-switching [KVS^{+}23]. preying [BR22]. primary [LNW21, ZL22]. Prince [ZB24, CRJ20]. principal [GGW^{+}23]. principles [FDMK21a, GMT21, Man21]. prior [PCCL21]. priorities [LAA^{+}21]. prized [LVL^{+}22]. probability [CAE^{+}21, MHMP^{+}23, SASM22, ZEM21]. problem [KTA^{+}23]. problems [KH22]. procedures [AMP^{+}23]. procellariiform [WESWT21]. Process [FSC^{+}23, GCT^{+}21, TBH^{+}20]. processes [KH21, SdS20]. produce [PKT^{+}21]. produced [BXF23]. production [Ano23f, BNV^{+}21, BTM^{+}21, CFB^{+}21, FCF^{+}21, GSWBK20, HBMC20, HAC^{+}23, HFT^{+}22, JRDW22, JRPHM^{+}22, LNW21, LPC^{+}23, MMAE20, MAE^{+}23, PHB^{+}20, RDAOE20, SMB^{+}21, SFPP^{+}23, TLS21, VSS^{+}22, WGI^{+}21]. Productivity [SFW^{+}22, Ano22b, MKSM23, MBP^{+}20, SVMBR22, SVML^{+}23, SSV^{+}22b, SCF^{+}21, THSZ21, ZL22]. profiling [CRJ20, dlTVLC^{+}21]. profit [AA21]. profit-maximizing [AA21]. profitability [AHK23]. Program [HKC^{+}23]. programme [CCR^{+}24, CFO^{+}22, PJCS^{+}23b]. programs [Ano23j, CRLC23]. progress [CFB^{+}23, HFP^{+}23, TNM^{+}23]. Progressing [KLSV21]. projected [LXBG21]. Projecting [BWBH21, MWT^{+}21, HMF^{+}21]. projections [DSR^{+}21, KJP^{+}22, MFP20]. prolonged [WPN^{+}21]. promising [PMM^{+}23]. promote [MSW21]. properties [LRHW21, TLS21]. property [TSD^{+}21]. proportion [BVV^{+}20]. Proposed [LKL^{+}21, GLW21, TMG20]. prospective [PEWN20]. prospects [SHG^{+}21]. protect [GVD^{+}23, vDHR^{+}22].

Protected [RRR^{+}24, SCMT^{+}22, BWBH21, DA20, HKZ^{+}23, OHM^{+}23, SBW^{+}22, VO23]. protection [PMM^{+}23, VO23, dBABS^{+}22]. protists [LFH^{+}21a, LFH^{+}21b, LFH^{+}22]. protocol [PO24, TMD^{+}20]. protocols [Ano24c, CWB^{+}23, SAA^{+}22]. proven [JEPT23]. provide [NDE^{+}21]. provides [NHM^{+}21]. Province [SWH^{+}23]. provinces [BBB^{+}23]. proxies [MCC^{+}22, PVGC^{+}22]. pteropod [MOC^{+}23]. publications [OdS21]. published [Leg21]. Publisher [Ano22e, Ano20b]. Pulse [SHW^{+}21, BM20, Har21, HMWS22, SMS^{+}22]. pup [JRPHM^{+}22]. pups [PMS^{+}20]. purse [FBF^{+}21, WGKG21]. putative [KEW^{+}21, SHK^{+}21]. pyrethroid [FNB^{+}20].

qualitative [FRSM22, OTL^{+}23, SRR21]. Quantification [KLP^{+}24, PFH^{+}22]. quantify [SRB^{+}23]. Quantifying [EHV^{+}21, FW22, HAP^{+}21, MTT^{+}23, MPA^{+}23, ODHF^{+}23, WGKG21, WCCY20, XCK^{+}23, MCH^{+}24]. Quantitative [CSM^{+}22, EWD^{+}21, EWD^{+}22, OTL^{+}23, SROR21]. Queen [Gra21]. Queensland [WCCY20]. quest [LAA^{+}21]. question [GAP21]. Questionnaire [GAM^{+}22]. questions [MBSR20]. quinqueradiata [STK^{+}20b, STK^{+}20a]. quota [AHK23, HHK21, Hol20].

radiata [KKS^{+}20]. rainfall [Ano24b, RVR^{+}24]. Raja [MD22, SMSP^{+}23]. ranched [LVT^{+}23]. random [PMSK^{+}20]. Range [BLM^{+}21, GA23, CAE^{+}21, DRP^{+}23, GHB^{+}23, HCM24, KWS^{+}21, RBH22, SJS21, SKM^{+}22]. range-extending [SKM^{+}22]. Range-wide [BLM^{+}21, DRP^{+}23]. ranges [GDI^{+}20, SSBV21]. ranging [PDP^{+}23]. ranking [CMG^{+}23]. rapid [BGH^{+}20, STK^{+}20b, STK^{+}20a]. rapidly [SJS21]. rarely [KSHO23]. Rastrineobola [PMSK^{+}20]. rate [Ano23j, CRLC23, FDMK22, FRP^{+}23, GGW^{+}23, HMWS22]. rates [BCBJ^{+}20, FAV^{+}20, NKS^{+}23, PEWN20, PKAT20, PO24, RRJ22, TLM^{+}20, VBO^{+}21a, VBO^{+}21b, VSS^{+}22, VBJ^{+}21, WOB^{+}23]. ratios [PWH23, dAdSM21]. ray [MD22, GC21]. rays [AMC^{+}23]. Re [SSV^{+}22b, YY22, Ano22b]. Re-examination [YY22]. Re-thinking [SSV^{+}22b, Ano22b]. reached [SRA^{+}21]. readers [Leg21]. readiness [JEPT23]. ready [WGCC^{+}22]. Real [WCL^{+}23, BMR21]. Real-time [WCL^{+}23]. real-world [BMR21]. realist [HHK21]. realities [ZMMF22]. realized [Day20]. rebalancing [GBH^{+}24]. Rebuilding [SLRC21]. recapture [BFD^{+}23, MCD^{+}20, TCLB22, VET^{+}22]. receiver [CSC^{+}23]. recently [PMS^{+}20, PCG^{+}22]. Recognition [GMT21, VAS^{+}20, YLC^{+}22]. Recommendations [MGO21, CRV^{+}21, GMT21, MMC^{+}23]. recompression [BGH^{+}20]. reconcile [PWCL21]. Reconciling [PCCL21]. reconstructed [CT23]. Reconstructing [DPH^{+}21]. reconstruction [vLNC20]. record [JMR^{+}22]. Recorder [HJP22, OBJ^{+}22]. recover [VNR^{+}23]. recovered [PAW22, RCH^{+}21]. recovery [BRG22, CLS^{+}22, HHL^{+}21, LVL^{+}22, SCP^{+}21, SHW^{+}21, SHP^{+}21, WML^{+}21]. Recreational [LLPdL22, LBPC^{+}20, RSTL22, Ano20b, BKF^{+}20, BMM^{+}22, FFR20a, FFR20b, FOC^{+}23, FODD^{+}22, HWL^{+}22, HMK^{+}20, KMS20b, LS20a, LWH^{+}21, LVL^{+}22, LSOD^{+}20, MYJ^{+}20, MYAH21, MPMGGAHR20, NHL^{+}20, NHL^{+}22, ODHF^{+}23, PFT^{+}20, SPd^{+}22, SCCA20, SROR21, SWR^{+}22, TLM^{+}20, TR20, TSD^{+}21, TMG20, TMLH21, VCF^{+}20, YMAH20]. recruit [Har21, LVG^{+}20]. Recruitment [LPC^{+}23, AWN22, BVV^{+}20, BDL^{+}22, DSF^{+}21, FWA^{+}21, HRB^{+}21, JKF^{+}22, Kai20, KNP21, KTA^{+}23, Man21, MTS^{+}22, MPN20, Mun20, NFN20, PDH^{+}21, RUdPFRA^{+}21, STK^{+}20b, STK^{+}20a, SOC^{+}20, SVML^{+}23, SHW^{+}21, SHK^{+}21, TZR^{+}21, TNS^{+}21, TB21, YY22, ZRW^{+}21, ZEM21, vdWBL22]. Recruitment-driven [LPC^{+}23]. recurrent [LMvEM23]. Red [FFR20a, FFR20b, BJPP^{+}20, FWA^{+}21, PHB^{+}20, PFP^{+}22, SB21, WDPM23, ZHL^{+}22, Gra21]. redfish [BPP^{+}21, MSVSA22]. redistribution [JMB^{+}22]. redleg [PDH^{+}21]. reduce [ASW^{+}23, Eay23, GT22, dSRKC22, RCH^{+}21, TDC^{+}23]. Reduced [Bra20, NBO^{+}20c, BM20, SG21, UPH^{+}21, YARE23]. reduces [TLS21]. reducing [ODHF^{+}23, RBH22]. reduction [Jen23, JEPT23, Suu22, ZL22]. Reef [PGRP23, BKR^{+}21, BGH^{+}20, DPD^{+}22a, DPD^{+}22b, EDC20, LZZ^{+}21, MAT^{+}23, Rin21, SL21, WOL^{+}21, YABM20]. reefs [Abe20, CvdWC^{+}20]. reel [PRAR23]. reference [Ano23a, HS20, MD22, OTMN20, SVBM^{+}22, SVMBR22, TBM^{+}22, ZRW^{+}21]. refining [SPd^{+}22]. Reflecting [WBB^{+}23]. reflection [Köp23]. reflections [MEOA23, SOCN^{+}22]. regime [BCB^{+}22a, GMT21, PKAT20, TZR^{+}21, WSJ23]. regimes [BJG^{+}21, DBGP22, RBG^{+}20, SSS^{+}23, TBP20]. region [ADS^{+}23, MHA^{+}21, PJCS^{+}23b]. Regional [DFWR20, MCD^{+}20, OGMA20, SGM^{+}24, ALLP^{+}22, DSR^{+}21, KH22, TAPB22, vDBF^{+}20]. regionalization [SLB^{+}20]. regions [LPC^{+}23, VSS^{+}22]. registration [vEMvH^{+}21]. registry [TR20]. regression [NHL^{+}20, dSRKC22]. regulated [PPMD22]. regulates [HJ22]. regulation [SdS20, SCG^{+}20]. regulations [RCJ^{+}22, SB21, vDHR^{+}22]. regulatory [TMG20]. Reinhardtius [GEBS^{+}23, WM20]. related [GGA^{+}23, SMK23, WJB^{+}21]. relation [BNB21, BQBB^{+}21, GKRB21, JSN^{+}22, PRCM22, WLR^{+}20, YC20]. Relationship [LBW^{+}20, BPP^{+}21, CRB21, CLMB22, Man21, RUdPFRA^{+}21, VBO^{+}21a, VBO^{+}21b, ZEM21]. relationships [Kai20, KMK^{+}21b, MTS^{+}22, SOC^{+}20, UBD^{+}23, YY22, ZRW^{+}21]. relative [CWB^{+}23, CWL^{+}20, EWD^{+}21, EWD^{+}22, FDMK22, FWC^{+}20, Mil24, SACP^{+}21, WBG^{+}22]. release [LVL^{+}22, PBS^{+}23, WPN^{+}21, WGL^{+}23]. releases [SASM22]. relevance [BSH^{+}22, VBJ^{+}21]. remains [GPS^{+}21b]. remix [WGI^{+}21]. remote [LQC^{+}22]. removals [SMLB^{+}22]. renewable [NAB^{+}22, SCW^{+}20]. renewables [DBB^{+}20]. repeat [KBBGA22]. Reply [GS22, Gal23, JLA22, JHN^{+}21, NBO^{+}20c, SFHP22, SSV^{+}22a]. reported [JHN^{+}21]. reporting [ECBJ22, GLW21, HHO^{+}20, OOB^{+}22]. representation [SGM22, SCMT^{+}22]. reproducibility [SAA^{+}22]. reproduction [Ano22b, CLL21, CT23, MFV^{+}23, SSV^{+}22b]. reproductive [CdSSO^{+}21, CCM^{+}21, DSF^{+}21, FWA^{+}21, Moo23a, QSKL21]. requirements [HMA^{+}23]. resampling [LS20a]. research [AMP^{+}23, Ano20b, BBL^{+}23, DBB^{+}20, DPMZ21, DNW^{+}23, FP20, Hol21, HKC^{+}23, KTA^{+}23, LSOD^{+}20, OP22, PJCS^{+}23b, SCM^{+}23, SMSP^{+}23, SOCN^{+}22, VSS^{+}23, WBB^{+}23]. researchers [AHC^{+}20, SFW^{+}22, SCM^{+}23, SCB^{+}23, SOCN^{+}22]. reserve [VNR^{+}23]. reserves [LLPdL22, RBF^{+}22]. Residency [TSWM23]. resident [Ano23i, BC23]. residual [ZYJZ23]. resilience [BPV^{+}20, FMBL23, FWC^{+}20, KOB^{+}22, KPS20, MMC^{+}23, WSL^{+}23, WMB^{+}22, YABM20, ZL22]. resilient [CBCM21, MAE^{+}23]. resistance [CRD^{+}22, FNB^{+}20]. resolution [HK21, LVL^{+}22]. resolve [MGB^{+}23]. resolved [MMCS^{+}24, Zho21]. Resource [MVNR^{+}21, DSR^{+}21, JSR^{+}22, RAH^{+}22, WBB^{+}23]. resources [Ano23f, AHP^{+}21, EDH^{+}21, HFT^{+}22, HSMK21, LLALF22, Man21, Met20, MPH^{+}21, PCC23, VQN22]. respect [CBB21]. Respiration [BCBJ^{+}20, RBG^{+}20]. respiratory [BCBJ^{+}20]. Response [SHP^{+}22, ANB^{+}21, ANB^{+}22, Ano24b, DLL^{+}21, OOE^{+}22, RVR^{+}24, RCH^{+}21, SBH^{+}21, SG21, VCD^{+}20, WM20, YFM^{+}20]. Responses [DeM22, FXG^{+}20, PBC^{+}22, Ano23b, GPS^{+}21a, HMBDP21, MSN^{+}23, OH23, QSKL21, RPB^{+}21, RHO22, SJX^{+}20, SBB^{+}22, Suu22, THSZ21, TFWE^{+}22, UVI^{+}21a, UVI^{+}21b]. restocking [RST^{+}21]. restoration [LAA^{+}21, Rin21]. restored [PFT^{+}20]. restrictions [MSW21]. result [dGW^{+}22]. retained [MYAH21]. retention [CSH^{+}21, Mil24]. retrieval [YLC^{+}22]. retrospective [HLHC22]. returning [GGW^{+}23]. reveal [ANA^{+}23, Ano24a, BLT23, BP20, CCS^{+}24, FPY^{+}22, LMH22, LHN^{+}22, ORGTO^{+}23, SHK^{+}23, SOB23, TNS^{+}21, YSL22, Zho21]. revealed [DRP^{+}23, JLM^{+}23, KPM^{+}22, MCD^{+}20, OHM^{+}23, SHW^{+}21, VCD^{+}20, VET^{+}22, WPP^{+}21]. Revealing [BPP^{+}21]. reveals [AMC^{+}23, CASM23, CRD^{+}22, FKTM20, JAQ^{+}23, OOB^{+}22, SCCA20, SDW^{+}21, TDB^{+}21, VCF^{+}20]. review [Ano23a, Bac24, BBB^{+}22, FP20, MCX21, PMTG^{+}23, SVBM^{+}22, SDGR^{+}20, VPP^{+}23, WMB^{+}22, vEMvH^{+}21]. Revival [GC21]. rich [BMR21]. richness [MGO^{+}23, PASL21]. Ridge [HHNS21, PJCS^{+}23b]. right [Ano24c, CBB21, Moo23a, SPBJ21, SPJ23]. rights [PVHV20]. rights-based [PVHV20]. rigidity [GBH^{+}24]. rise [Mac22]. rising [HCCR20]. Risk [AGSS22, SRM^{+}20, AMP^{+}23, Ano23e, DA20, ESK^{+}21, FW22, GBGA^{+}23, Hut22, JMB^{+}22, LPG^{+}20, PLN20, QBF^{+}21, RBH22, WFG^{+}23, XCK^{+}23, ZB21]. Risks [HEHT^{+}22, GRE^{+}20]. river [Ano23e, BPR^{+}23, HFP^{+}23, KTK20, WFG^{+}23, SBB^{+}23]. rivers [DBT^{+}23, GGW^{+}23]. RNA [HHNS21]. road [Eay23, HHL^{+}21, Mur21]. Roads [KMK^{+}21b]. Robust [VQN22, DBGP22]. robustness [LHE20]. Rock [WML^{+}21, DeM22, GPJ^{+}21, LLPdL22, MBA^{+}20]. rockcod [HHL^{+}21]. Rockfish [BKF^{+}20, CT23, RFB^{+}22, BMM^{+}22]. rockfishes [WPN^{+}21]. rocky [CRG^{+}23, CvdWC^{+}20]. rod [GGW^{+}23]. role [Cha21, CZ23, DYS21, JEPT23, JNT^{+}22, PCL^{+}21, RBCD20, SBD^{+}21, TYF23, VPCM^{+}22]. roles [JRDW22, TPCM^{+}24]. rooted [MR20a]. ropes [ON21]. Ross [SHP^{+}22]. Ross-Gillespie [SHP^{+}22]. rossii [HHL^{+}21]. rostrata [Ano23i, BC23]. rotational [HMC^{+}20]. rotationally [ZJH^{+}23]. Rough [RRJ22]. roughy [SK21]. round [CRR^{+}22, Ano22d]. routine [KPM^{+}22]. Ruditapes [BBB^{+}21]. rule [FDK20, FDMK21b, FDMK21c, GDI^{+}20, SMOD21]. rules [LKL^{+}21, SLRC21, ZB24]. running [VLD^{+}21]. Russian [SB21].

sablefish [Ano23j, CRLC23, ORGTO^{+}23]. Saccharina [JBJ^{+}22, JJB^{+}22, TRP^{+}21]. sacrificing [Abe20]. Sagmariasus [WML^{+}21]. saida [DDL^{+}20]. sailfish [FFT^{+}22, LVL^{+}22]. salar [BGR^{+}22b, JSN^{+}22, LVT^{+}23, MYW^{+}21, SCH^{+}21b, TSE^{+}21]. salinity [Ano24b, RVR^{+}24]. Salmo [BPC^{+}21, BGR^{+}22b, JSN^{+}22, KBJA20, KBK^{+}22, KBBGA22, LVT^{+}23, MYW^{+}21, SCH^{+}21b, TSE^{+}21]. Salmon [JG21, JHS^{+}21a, JHS^{+}21b, ANA^{+}20, Ano23e, Bea22, BPR^{+}20, BPC^{+}20, BLM^{+}21, BGR^{+}22b, DBT^{+}23, DFK^{+}22, FFH^{+}22, FNB^{+}20, GKRB21, GGW^{+}23, HUJ^{+}21, HMA^{+}23, HJ22, JSN^{+}22, JHN^{+}21, JDJK22, LAA^{+}21, LVT^{+}23, MYW^{+}21, NBL^{+}21, OOB^{+}22, OGMA20, PBD^{+}21, SJM^{+}20, SBH^{+}21, SDSMS21, SAB^{+}20, SG21, SRBT23, SCH^{+}21b, SGM^{+}24, SBB^{+}22, TSE^{+}21, UPH^{+}21, WFG^{+}23, YFM^{+}20]. salmonids [MvdMS^{+}23, VLD^{+}21]. SAM [NHM^{+}21]. sample [TBH^{+}20]. sampled [JBF^{+}20]. samples [SBKO22]. sampling [AKK^{+}21, AMP^{+}23, ASV21, CFO^{+}22, FSC^{+}23, HJP22, JBF^{+}20, Köp23, LFH^{+}21a, LFH^{+}21b, LFH^{+}22, MPA^{+}23, SPd^{+}22, SBKO22, SMSP^{+}23, SOB23, TR20, TBH^{+}20, TMS23, VCF^{+}20]. Sanctuary [FFT^{+}22]. sand [BSW^{+}23, JLM^{+}23, SWS^{+}21, WLR^{+}20, YMH^{+}24]. sandeel [HRB^{+}21]. sandeels [OBJ^{+}22, SMS^{+}22]. sardine [KJP^{+}22]. Sargassum [HK21]. sargus [dBABS^{+}22]. satellite [KKS^{+}20]. save [Moo23a, SBW^{+}22]. saving [Day20]. scale [Ano20b, Ano23d, Ano23e, BHDD22, CCC^{+}21, FPY^{+}22, GHB^{+}23, JYK^{+}21, JNT^{+}22, LV23, LLALF22, LBPC^{+}20, LSOD^{+}20, MTTG^{+}23, MDHB21, MKK^{+}24, MWS^{+}23, MMMB^{+}23, NAS^{+}20, NSJ^{+}22, OP22, PBF^{+}21, PRAR23, RES^{+}23, RHO22, SWR^{+}22, TSD^{+}21, Thi20, TAPB22, VKH^{+}20a, VKH^{+}20b, VCD^{+}20, WPB20, WBC^{+}21, WESWT21, WFG^{+}23, dAPDDJ21, lCRGCCA23, vDBF^{+}20]. scale-up [TAPB22]. scales [BCP^{+}21, BGR^{+}22b, DNW^{+}23, JDJK22, MTCC21, MBP^{+}20]. Scaling [UBD^{+}23, Day20]. scallop [MKSM23, ZJH^{+}23]. scalloped [LMvEM23]. scallops [HMC^{+}20]. scamp [GSWBK20]. scaring [TDC^{+}23]. scats [CLS^{+}22]. scattering [LRHW21]. Scavenger [MGO^{+}23]. Scenario [CFB^{+}21, TMG20]. Scenario-based [TMG20]. Scenarios [CCC^{+}21, BMN^{+}21, BBL^{+}23, CMG^{+}23, DA20, DHN^{+}21, LV23, MPH^{+}21, OTL^{+}23]. schemes [CMH^{+}23]. Schijns [HHN22]. Schizophrys [WDPM23]. school [PMO^{+}21a, PMO^{+}21b]. schools [PMSK^{+}20]. Sciaena [GPV^{+}22]. Science [Ano20c, BD20, OdS21, PY22, SCM^{+}23, ABB^{+}21a, ABB^{+}21b, Ano20a, Ano23c, BB20, Bir23, CMD^{+}23a, CvPTAJ22, Har20, HBC^{+}22, Hut22, JBDK23, KH21, Lov20, Mac22, MEOA23, MHES20, PAC^{+}23, RCB^{+}21, SMK23, TNM^{+}23, THSZ21, WJ22]. science-related [SMK23]. sciences [SCB^{+}23, SOO^{+}22, TPCM^{+}24]. scientific [ARE^{+}22, FP20, JDD^{+}21, JAB^{+}21, Köp23, Kra23, PJP20, VRM^{+}23, VNBLZR^{+}23, dBSQK21]. scientist [Bjo21, BB22, Hol20, MR20b, Sch21a]. scientists [AHC^{+}20, MPQS20, RCB^{+}21, SOO^{+}22]. Scomber [GHÓ^{+}20, KTY^{+}21, KÓP^{+}21]. scombrus [GHÓ^{+}20, KÓP^{+}21]. scope [TPCM^{+}24]. Scotian [HJP22]. Scottish [CLMB22, HMNF22, LBW^{+}20, OGMA20]. Sea [CSM^{+}23, DDL^{+}20, GKRB21, Ano23h, AAHT22, BFD^{+}23, Bar22, BVS^{+}23, BPC^{+}21, CBT^{+}20, CLS^{+}22, Dol20, GGA^{+}23, GICP^{+}22, HMC^{+}20, HYE^{+}22, JSLC21, JLA22, JYK^{+}21, JDJK22, JBE^{+}21, JRPHM^{+}22, LBH^{+}21, LHL^{+}21, MKSM23, OOB^{+}22, PRAR23, RTQW20, SG21, SWLW22, SJX^{+}20, TMLH21, VKH^{+}20a, VKH^{+}20b, WLR^{+}20, ZJH^{+}23, vDHR^{+}22, ADS^{+}23, AMC^{+}23, Ano23f, BFD^{+}23, BVV^{+}20, BFFC20, BW21, Bra20, CZS^{+}23, CRG^{+}23, CFB^{+}23, DSMM^{+}22, DFWR20, DMK^{+}21, DDL^{+}20, DYS21, EDH^{+}21, ECR^{+}21, ECBJ22, FND23, FFS^{+}22, FJC^{+}20, FKTM20, GA23, GAM^{+}22, GZW^{+}20, HvdMLS21, HMOR22, HRB^{+}21, HFP^{+}23, HHA^{+}21, HFT^{+}22, HYE^{+}22, HHD^{+}23, IEIT23, JBF^{+}20, KMK^{+}21b, KLF^{+}20a, KLF^{+}20b, KLP^{+}24, KPS22, LH20, LIR^{+}21, LAB^{+}22, LPG^{+}20, LSR^{+}21, LRGW23, LZZ^{+}21, LQC^{+}22, MACSR22, MDP^{+}21, NBO^{+}20a, NBO^{+}20b, NBO^{+}20c, OOE^{+}22, OTL^{+}23, PMS^{+}20, PFH^{+}22, PPV^{+}22, RHvD^{+}20]. Sea [RDM^{+}21, ROH^{+}21, RBCD20, SH22, SMB^{+}21, STK^{+}20b, STK^{+}20a, SDW^{+}21, SB21, SCG^{+}20, SCF^{+}21, VLUB^{+}22, VRH^{+}23, VSL22, VPB^{+}20, WLR^{+}20, YFM^{+}20, YC20, ZMN^{+}21, vdWBL22]. sea-ice [HYE^{+}22]. Sea-louse [GKRB21]. seabass [GVD^{+}23, dPHG^{+}23]. seabed [GMT21, HVDvD23, ON21]. Seabird [TFWE^{+}22, dlCRT^{+}22, BJBB20, GBGA^{+}23, TDC^{+}23]. seabirds [Ano23d, Bar23, GLW21, NKA^{+}20, VSS^{+}23, WESWT21, ZB21, lCRGCCA23]. seabream [dBABS^{+}22]. seafloor [GJU^{+}23, LRGW23, RHvD^{+}20]. seafloor-mounted [LRGW23]. seafood [FCF^{+}21, Hut22, PRAR23]. seagrass [Ano24b, JSR^{+}22, QSKL21, RBG^{+}20, RVR^{+}24, TRD^{+}22]. Seal [BW21, JRPHM^{+}22, PMS^{+}20, TFWE^{+}22, WPB20]. seals [GETvB21]. searching [Day20]. seas [Ano24a, CBFS23, OH23, SHK^{+}23, SK21, DBT^{+}23, LAB^{+}22, PRRB^{+}20, SVML^{+}23, SMLB^{+}22, YSL22]. Seascape [HGS^{+}21, MHO^{+}20, MWS^{+}23, PLN20, SHK^{+}21]. seascape-scale [MWS^{+}23]. season [dSRKC22]. Seasonal [AMD^{+}23, BC23, HJ22, NKS^{+}23, TAB^{+}20, dPHG^{+}23, BCP^{+}21, FRP^{+}23, LRGW23, MBP^{+}20, SJX^{+}20, VSL22, Ano23i]. Seasonality [VLUB^{+}22, VPB^{+}20, DPMZ21, FKTM20, SACP^{+}21, ZJF^{+}21]. seasons [MYAH21]. seastar [WDPM23]. seaward [NBL^{+}21]. Seaweed [FDPD^{+}23, GS22, GSL22, Gal23, Sta22, NAS^{+}20, vdBDH21]. seaweeds [PBS^{+}23]. Sebastes [BPP^{+}21, MSVSA22, WPN^{+}21]. second [TSW^{+}21]. secondary [Ano23f, HFT^{+}22, SMSP^{+}23]. sector [LS20a, MPMGGAHR20]. sectors [OTLH20]. Sediment [RDM^{+}21, QSKL21]. sediments [MYMV20]. see [Bea22]. seen [KSHO23]. segmentation [GPQ^{+}20]. seine [FBF^{+}21, IHL21, ON21, WGKG21, NAB^{+}22]. seismic [Ano23b, MMN^{+}22, MSN^{+}23]. select [dCKS^{+}22]. selection [BSH^{+}22, CCC^{+}21, IHL21, JSN^{+}22, TB21]. selective [KWEB21, PRMS22, SMK23]. selectivity [BSH^{+}22]. self [CFO^{+}22, KMS20b, Köp23, TMS23]. self-reflection [Köp23]. self-sampling [CFO^{+}22]. self-subsidizing [KMS20b]. self-weighting [TMS23]. Semi [CKH^{+}21, Dom21, MTK^{+}21, PTD^{+}20, RFB^{+}22, SROR21]. semi-demersal [Dom21]. semi-enclosed [PTD^{+}20]. semi-pelagic [MTK^{+}21, RFB^{+}22]. semi-quantitative [SROR21]. Semi-supervised [CKH^{+}21]. sensible [SDGR^{+}20]. sensing [LQC^{+}22]. Sensitivity [SWS^{+}21, BMN^{+}21, DMK^{+}21, GGRB22, yMlZlL^{+}21, SGM22, SAA^{+}22]. sensors [ECBJ22]. separate [MGB^{+}23]. separation [KMK21a]. Sepioteuthis [CCS^{+}24]. sequencing [JLM^{+}23]. sequential [CZS^{+}23]. sequestration [BBB^{+}21]. Serendipity [Wap20]. serial [CCJ^{+}21]. series [JMR^{+}22, LVT^{+}23, NFN20, PCG^{+}22, PY22]. Seriola [STK^{+}20b, STK^{+}20a]. set [BFJB21, CCR^{+}24]. Setting [BB20, HVDvD23, dSRKC22]. settings [PBC^{+}22]. settlement [EVLK^{+}21, HHD^{+}23]. seven [MFP20, PMTG^{+}23, VBO^{+}21a, VBO^{+}21b]. seven-decade [VBO^{+}21a, VBO^{+}21b]. several [KBBGA22]. Sex [dAdSM21, DLLC^{+}22]. sexual [FFH^{+}22, JAQ^{+}23]. Seychelles [MAI^{+}22]. shallow [BKR^{+}21, CRG^{+}23]. shallowest [DPD^{+}22a, DPD^{+}22b]. shape [CZ23, CCM^{+}21, OH23]. shape-based [CZ23]. shaped [Fie21]. shaping [CWL^{+}20]. share [CMH^{+}23]. shared [MMC^{+}23]. sharing [CMH^{+}23, CRV^{+}21]. shark [CWC^{+}23, FCD^{+}23, GSR^{+}20, LMvEM23, RBH22, Zho21]. sharks [AMC^{+}23, BMB20, FBF^{+}21, LPG^{+}20, MVNR^{+}21]. sharp [Lov20]. shear [KAK^{+}21]. Shearwaters [LB23]. Shelf [FRP^{+}23, ZJF^{+}21, BBM^{+}22, JLM^{+}23, LHL^{+}21, MSW21, MFP20, PLN20, SDR^{+}21, SWS^{+}21, UVI^{+}21b, UVI^{+}21a]. shell [LBW^{+}20]. shellfish [BBB^{+}21, FRSM22, HSMK21, JJB^{+}22, SSBV21, TAPB22]. shells [TLS21]. shelves [HJP22]. Shetland [VKH^{+}20a, VKH^{+}20b]. shift [BCB^{+}22a, MB21]. Shifting [BR21, FSN^{+}22, VSL22, SJS21, SWS^{+}21]. Shifts [IVU^{+}22, IEIT23, MWT^{+}21, PKAT20, ROH^{+}21, TZR^{+}21, WM20]. Ship [BVS^{+}23, DLL^{+}21]. ship-based [DLL^{+}21]. Ship-to-shore [BVS^{+}23]. shocks [WSL^{+}23]. Shore [MWJ^{+}24, BVS^{+}23, TLM^{+}20]. shore-based [TLM^{+}20]. short [AHP^{+}21, CNWW22, WOB^{+}23]. short-lived [AHP^{+}21, WOB^{+}23]. short-wave [CNWW22]. shortages [dGW^{+}22]. shortfin [Kai20]. should [MR20a]. show [GAM^{+}22, SG21, VBO^{+}21a, VBO^{+}21b]. shows [BPC^{+}21, JJB^{+}22, SBA^{+}22]. shrimp [CZS^{+}23, HWS^{+}21, HHO^{+}20, LLALF22, SH22, SSH^{+}20]. shrimping [TYF23]. side [SRMS21]. sided [GHB^{+}23]. Sidney [ABB^{+}21a, ABB^{+}21b, GC21, Man21, RCB^{+}21]. SIDS [MKB^{+}23]. signals [MKK^{+}24]. significance [SPJ23]. significantly [CSH^{+}21]. silky [FBF^{+}21, GSR^{+}20]. silver [HFP^{+}23, PMSK^{+}20]. Simple [HBFO20]. simplex [SdS20]. simplify [PWH23]. Simulating [MWJ^{+}24]. Simulation [PDL20, TBH^{+}20, CFO^{+}22, HLHC22, SdS20, Tho22]. simulations [NSS^{+}23, SGM22]. Simultaneous [KP24]. since [BPPC^{+}20, FFS^{+}22, RBF^{+}22]. sinensis [KBJA20]. single [MHKV20, NHM^{+}21]. single-fleet [NHM^{+}21]. sinks [FDPD^{+}23, Gal23]. sit [Bea23]. site [CCC^{+}21, DPMZ21, GGLS22, LBW^{+}20, LWH^{+}21, MMN^{+}22, ODHF^{+}23, dPHG^{+}23]. site-fidelity [GGLS22]. sites [ASV21, PAW22, RDAOE20]. situ [AKL^{+}21, AKK^{+}21, CRJ20, KAK^{+}21, LNW21, PLN20, SBPA^{+}21, SHW^{+}21, SBB^{+}22, YLC^{+}22]. situations [Har21]. Six [PKT^{+}21, HCH20]. Sixty [SDGR^{+}20]. Sixty-five [SDGR^{+}20]. Size [KWEB21, MB21, ÖHB^{+}23, ÁEPCL20, ANA^{+}23, ASV21, BMB20, BSH^{+}22, CCM^{+}21, CGH^{+}23, CFP20, DBT^{+}23, FFS^{+}22, GSWBK20, GPQ^{+}20, HRB^{+}21, KP24, Pep23, PRMS22, RRJ22, RGE^{+}23, SMOD21, TBH^{+}20, WRC^{+}21, ZJH^{+}23, dLP21]. Size-at-maturity [MB21]. Size-based [ÖHB^{+}23]. Size-selective [KWEB21, PRMS22]. size-specific [KP24]. size-structured [CFP20]. size-transition [CFP20]. sizes [AWN22, CLS^{+}22]. skate [ASW^{+}23, KKS^{+}20]. skeleton [RPB^{+}21]. skill [KMG22, KNP21]. skipjack [WCL^{+}23]. Skipper [RHO22]. Slope [ADS^{+}23]. small [Bir23, BCP^{+}21, HMNF22, JYK^{+}21, LLLS20, LLPdL22, LLALF22, LBPC^{+}20, MPA^{+}23, MMMB^{+}23, NSJ^{+}22, OH23, PBF^{+}21, PRAR23, RES^{+}23, TSD^{+}21, WPB20, dAPDDJ21, MFB23]. small- [dAPDDJ21]. small-scale [JYK^{+}21, LLALF22, LBPC^{+}20, MMMB^{+}23, NSJ^{+}22, PBF^{+}21, PRAR23, RES^{+}23, TSD^{+}21, WPB20]. Smartphone [SHG^{+}21]. smolt [JG21, JHS^{+}21a, JHS^{+}21b]. smolts [JSN^{+}22, MYW^{+}21, OGMA20, UPH^{+}21]. Snapper [FFR20a, FFR20b, BBB^{+}23, PHB^{+}20, PFP^{+}22, TDB^{+}21]. Snow [HHA^{+}21, Ano23f, CBB21, DSMM^{+}22, HFT^{+}22, HHD^{+}23, MB21, ZMN^{+}21]. SNP [TDB^{+}21]. SNP-genotype [TDB^{+}21]. SNPs [DRP^{+}23, FPY^{+}22]. Social [FRSM22, WJB^{+}24, Bar23, CvPTAJ22, FSS^{+}23, KOB^{+}22, KF21, Köp23, MMC^{+}23, MTTG^{+}23, OdS21, SCCA20, SKH21, SCB^{+}23, SMK23, TPCM^{+}24]. Social-ecological [FRSM22, MMC^{+}23]. socio [WMB^{+}22]. socio-ecological [WMB^{+}22]. Socioeconomic [PGRP23, QMMC^{+}23]. solaris [DPD^{+}22a, DPD^{+}22b]. sole [SMS^{+}22, vdWBL22]. solute [RTQW20]. solution [KPS21]. Solutions [KH22, MVMdS21, PMM^{+}23]. somatic [CMCF21, SHE^{+}22]. some [SRMS21]. sometimes [HBFO20]. sonar [MCH^{+}24, SKA15, SKA22]. sonars [BMU^{+}23, PMO^{+}21a, PMO^{+}21b]. sooty [BPC^{+}24, SWLW22]. Sound [ZB24, BD20, HMWS22, RDAOE20, CRJ20]. sound-induced [HMWS22]. sounding [HK21]. sounds [BXF23]. source [SHK^{+}21]. sources [MGF^{+}21, SBC^{+}21a, SBC^{+}21b]. south [FMPH24, YFM^{+}20, ANA^{+}20, BRG22, EMS^{+}21, GRE^{+}20, HHL^{+}21, LZZ^{+}21, LQC^{+}22, MRC^{+}21, MEOA23, SMP^{+}23, SRA^{+}21, SHP^{+}21, TFH^{+}21, TFWE^{+}22, dMBJ22]. southeastern [BBB^{+}23, GSWBK20, HYE^{+}22, HCH20, IVU^{+}22, MSW21]. southern [DSF^{+}21, GZSP22, HHNS21, JRPHM^{+}22, KLF^{+}20a, KLF^{+}20b, MMAE20, SH22, SPJ23, SDGR^{+}20, dGW^{+}22, BPC^{+}24, BCBJ^{+}20, DSF^{+}21, GRE^{+}20, GJU^{+}23, PLN20, PMS^{+}20]. southwest [SWLW22, AWN22, ASW^{+}23, SLB^{+}20, dAdSM21]. southwestern [CBT^{+}20, VBJ^{+}21, SdS20]. sp [Ano23k, BPP^{+}21, GBMM^{+}23]. space [BNB21, CWL^{+}20, CMS^{+}23, NHM^{+}21, PDL20, SFPP^{+}23]. Spanish [CCR^{+}24, PRRB^{+}20]. spare [DPD^{+}22a, DPD^{+}22b]. Spatial [CMCF21, DASLRB22, LAB^{+}22, LMvEM23, SSH^{+}20, SJX^{+}20, BCP^{+}21, BHDD22, DYS21, EVY^{+}21, GC21, GCT^{+}21, GCG22, JJB^{+}22, LvDM^{+}22, LMW^{+}22, MTS^{+}22, MKS^{+}21, MDP^{+}21, NHL^{+}20, NHL^{+}22, OT22, PDP^{+}23, QBF^{+}21, SLB^{+}20, TBP20, VKH^{+}20a, VKH^{+}20b, VPB^{+}20, VCF^{+}20, WBA^{+}22, YC20, ZMMF22, dBABS^{+}22, Ano22c]. Spatialized [NAB^{+}22, NBA^{+}23]. Spatially [CZS^{+}23, SCMT^{+}22, ZJF^{+}21, CMG^{+}23, ZCX^{+}20]. Spatio [BLT23, BPC^{+}20, BDL^{+}22, MAI^{+}22, OBJ^{+}22, Ano23g, BAS^{+}21, GCT^{+}23, IEIT23, MKSM23, TAB^{+}20, TFWE^{+}22, WPP^{+}21]. Spatio-temporal [BLT23, BPC^{+}20, BDL^{+}22, MAI^{+}22, OBJ^{+}22, Ano23g, BAS^{+}21, GCT^{+}23, IEIT23, MKSM23, TAB^{+}20, TFWE^{+}22, WPP^{+}21]. Spatiotemporal [PCG^{+}22, YCF^{+}23, CCS^{+}24, LAGC23, MTK^{+}21, ODT^{+}22, OBK21, SRBT23, WBC^{+}21, YARE23]. spawner [NFN20]. Spawning [MMN^{+}22, SJS21, ADS^{+}23, Ano23b, BØS^{+}21, CCJ^{+}21, EVY^{+}21, EGMO^{+}20a, EGMO^{+}20b, GSWBK20, KBBGA22, LAB^{+}22, MPQS20, MSN^{+}23, MAC21, RUdPFRA^{+}21, RDAOE20, STK^{+}20b, STK^{+}20a, SHE^{+}22, SHW^{+}21, SHK^{+}21, TNS^{+}21, VPP^{+}23, VRH^{+}23, WML^{+}21, vdWBL22]. spearfishers [SCCA20]. special [VLUB^{+}22]. Species [MV23, Ano23d, Ano24b, AHK23, ASV21, BLT23, BPR^{+}23, BQBB^{+}21, CPD^{+}21, CZ23, CLS^{+}22, DGM^{+}24, GA23, GBGA^{+}23, GPJ^{+}21, GZW^{+}20, HBFO20, HSS21a, HSS21b, Hut22, JKF^{+}22, KSHO23, KMK21a, KDS^{+}22, KWS^{+}21, KLP^{+}24, LVG^{+}20, LvDM^{+}22, LBPC^{+}20, LTK20, LS20b, yMlZlL^{+}21, MGB^{+}23, MHJ^{+}20, MSWB22, MTT^{+}23, MTK^{+}21, Moo23b, Moo23a, MSP^{+}20, MFP20, ODHF^{+}23, OOE^{+}22, OBJ^{+}22, OT22, PIAF^{+}23, PLN20, Pep23, QMMC^{+}23, RVR^{+}24, SL21, SGM22, SDW^{+}21, SMSP^{+}23, SJS21, SCN^{+}23, SAA^{+}22, TDB^{+}21, Tho22, WBA^{+}22, lCRGCCA23]. specific [CPD^{+}21, KTY^{+}21, KMK^{+}21b, KP24, MCD^{+}20, NHM^{+}21, SCCA20, WM20]. specification [FSC^{+}23]. spectroscopy [PHB^{+}20]. spectrum [KP24]. speed [HMWS22]. Sphyrna [LMvEM23]. SPiCT [BMR21]. spider [RPVT^{+}21]. Spillover [HMC^{+}20]. spined [OOE^{+}22]. Spiny [SKM^{+}22, TBP20, FPY^{+}22, RPVT^{+}21]. sponge [RPB^{+}21, VKH^{+}20a, VKH^{+}20b]. sportfish [LVL^{+}22]. spotted [LHE20]. spp [SPBJ21, SPJ23]. sprat [LvDM^{+}22, MCS^{+}20, WOB^{+}23]. Sprattus [MCS^{+}20, WOB^{+}23]. spread [DSDA22, ZMFS20]. spring [BØS^{+}21, HCCR20, JNT^{+}22, MPN20, SHE^{+}22, TNS^{+}21]. spring- [BØS^{+}21]. spring-spawning [SHE^{+}22, TNS^{+}21]. squat [FWA^{+}21]. squid [Ano24a, AWN22, CASM23, GPM21, SHK^{+}23, SBA^{+}22, SSF^{+}22]. squids [Ano23l, LAB^{+}22, WCLB23]. St [BPP^{+}21, MSVSA22, SPJ23]. Stability [KPS20, DSGG20, LL23]. Stable [ANA^{+}20, LMH22, SHK^{+}23, vLNC20, CT23, JDJK22, Ano24a]. Stafford [GS22]. stage [BB20, CPD^{+}21, TBH^{+}20, THSZ21]. stage-structured [THSZ21]. stages [Ano22c, DASLRB22, LHE20, ZEM21]. staging [WRC^{+}21]. stakeholder [CMD^{+}23a, CCH^{+}23, PGRP23]. standard [KTK20]. standardization [MTK^{+}21, TAB^{+}20, YARE23]. standardized [CGH^{+}23]. Standardizing [Ano23j, CRLC23, TLM^{+}20, MMCS^{+}24]. standing [MGB^{+}23]. stanzas [YC20]. starfish [DPD^{+}22a, DPD^{+}22b]. stars [SOO^{+}22]. starvation [GPS^{+}21b]. state [BNB21, CMS^{+}23, HvdMLS21, HVDvD23, HMK^{+}20, NHM^{+}21, NKS^{+}23, PDL20, SFPP^{+}23]. state-space [BNB21, CMS^{+}23, NHM^{+}21, PDL20, SFPP^{+}23]. States [HMC^{+}20, SAB^{+}20, Bir23, GSWBK20, MMC^{+}23, MFB23]. static [ECBJ22]. static-gear [ECBJ22]. stationary [TSE^{+}21, WBA^{+}22, ZRW^{+}21]. statistical [LvDM^{+}22]. statolith [CASM23]. statoliths [MCC^{+}22]. Status [SMJ^{+}21, ANB^{+}21, ANB^{+}22, FWC^{+}20, JBDK23, KMG22, LZZ^{+}21, OTLH20, ÖHB^{+}23, RHvD^{+}20, SHG^{+}21, TBP20]. stay [ADG^{+}20]. STECF [Kra23]. steps [PKT^{+}21]. stick [Lov20]. stickleback [LIR^{+}21, OOE^{+}22]. stimulation [SMS^{+}22]. stimulations [BRvL^{+}23]. Stochastic [SVML^{+}23]. Stock [AHP^{+}21, Ano22d, Ano23e, BPC^{+}21, BBP^{+}23, BHWC^{+}20, BDB^{+}21, BD20, BMR21, BNB21, CZS^{+}23, CMR^{+}20, CRR^{+}22, CMCF21, CMS^{+}23, CFP20, DEL21, ECR^{+}21, EVY^{+}21, FSC^{+}23, FWC^{+}20, GDI^{+}20, GGW^{+}23, HUJ^{+}21, HLHC22, IEIT23, JMB^{+}22, Kai20, KSK^{+}21, KMG22, KBW22, Köp23, LVG^{+}20, Man21, MTS^{+}22, MKSM23, MCD^{+}20, MVMdS21, MCOÁ^{+}22, MKGC20, MAC21, ODT^{+}22, OOB^{+}22, PDL20, PCCL21, PKT^{+}21, RUdPFRA^{+}21, STK^{+}20b, STK^{+}20a, SMJ^{+}21, SOC^{+}20, SBA^{+}22, TZR^{+}21, TZE21, TB21, WFG^{+}23, WPP^{+}21, WML^{+}21, YY22, ZRW^{+}21, ZEM21, vLNC20]. stock-recruitment [MTS^{+}22, TZR^{+}21]. stock-specific [MCD^{+}20]. stocks [BGR^{+}22a, BBB^{+}23, BSW22, BMR21, BGR^{+}22b, DSF^{+}21, Dom21, HCH20, KBW22, QCM^{+}20, SMJ^{+}21, SAB^{+}20, STS^{+}20, SdS20, SBC^{+}21b, WOB^{+}23, SBC^{+}21a]. stomach [NDE^{+}21]. stomiiform [BOLPHL21]. storage [CP20]. stories [HKC^{+}23]. story [DAB^{+}23, Mac22]. Straddling [BFJB21]. strain [LXBG21]. Strait [LBH^{+}21]. Stranding [GHB^{+}23, CBT^{+}20, PDP^{+}23]. strategic [BHWC^{+}20]. strategies [BD20, BPR^{+}20, CBCM21, CCH^{+}23, DDP^{+}21, FODD^{+}22, GFO^{+}20, JBF^{+}20, LSR^{+}21, MAE^{+}23, PBS^{+}23, RCH^{+}21, SCP^{+}21]. Strategy [CSM^{+}23, MTCC21, PKT^{+}21, dMBJ22, AA21, AHK23, FDMK22, PJCS^{+}23a, MD22, PBF^{+}21, SSF^{+}23, SPP21, WPM^{+}23, WGL^{+}23, ZB24]. straying [KBK^{+}22, KBBGA22]. stream [KBJA20, BQBB^{+}21]. strength [LDW23, SK21, SBPA^{+}21, YMH^{+}24]. Strengthening [CRV^{+}21]. strengthens [NHM^{+}21]. stress [Ano24b, CT23, RVR^{+}24, TLS21, WJB^{+}21]. Stressed [PTD^{+}20]. stresses [RPB^{+}21]. stressors [HKZ^{+}23]. striatus [SHW^{+}21]. striking [VCF^{+}20]. striped [IOS21, MGB^{+}23, MPN20, PRMS22]. struck [VBJ^{+}21]. structural [KEW^{+}21]. Structure [WLR^{+}20, BCP^{+}21, CRG^{+}23, DRP^{+}23, EDC20, EVY^{+}21, FPY^{+}22, GEBS^{+}23, HK21, KLF^{+}20a, KLF^{+}20b, LTN21, MRC^{+}21, MDAMPE^{+}21, MMH^{+}23, MCS^{+}20, ORGTO^{+}23, PFP^{+}22, RGE^{+}23, SBA^{+}22, TRD^{+}22, WSJ23, WBWS21, ZJH^{+}23, dPHG^{+}23]. structured [BHWC^{+}20, CFP20, HCH20, MMAE20, THSZ21]. structures [BD20, SWH^{+}23, TMD^{+}20, WLR^{+}20]. structuring [BJG^{+}21, KMS^{+}20a, LvDM^{+}22]. struggle [dGW^{+}22]. studies [PCL^{+}21, SDGR^{+}20, TAB^{+}20, VAS^{+}20]. study [FFR20a, FFR20b, FND23, HCM24, IEIT23, JDD^{+}21, JAQ^{+}23, LLLS20, LvDM^{+}22, MKSM23, MYW^{+}21, MFP20, OGMA20, PRD20, RRJ22, RTQW20, SHFK22, SBB^{+}23, SMSP^{+}23, WOB^{+}23, dlCRT^{+}22]. studying [KDS23, PRD20]. Stylocheiron [BQBB^{+}21]. sub [CCC^{+}21, LRHW21, SCP^{+}21]. sub-adult [SCP^{+}21]. sub-Arctic [LRHW21]. sub-lagoon [CCC^{+}21]. Subarctic [Ano22e, DHN^{+}21, MPH^{+}21, GGLS22, GGOP21, GGOP22, VSS^{+}22]. subglacial [NKA^{+}20]. sublethal [WPN^{+}21]. subsea [LH20]. subsidizing [KMS20b]. substantial [GAM^{+}22]. substrates [BSW^{+}23]. subtropical [LL23]. success [FWA^{+}21, TNS^{+}21, VPCM^{+}22]. successful [LAA^{+}21]. suggest [ANA^{+}20]. suggestion [SDGR^{+}20]. suggests [BØS^{+}21, GZW^{+}20, KMK^{+}21b]. Suitability [CBFS23, DPH^{+}21, HMA^{+}23, HMF^{+}21, PTR22, SRA^{+}21, WBC^{+}21]. suitable [GPV^{+}22, MSVSA22, VHT^{+}21]. summer [MTT^{+}23, SPJ23]. superba [SMP^{+}23]. supervised [CKH^{+}21]. supply [SPBJ21]. Support [Bra22, Ano20c, BCC^{+}24, BD20, Bir23, CMD^{+}23a, Dom21, GDI^{+}20, Har20, LSD^{+}20, MAE^{+}23, MWS^{+}23, SWH^{+}23, TPCM^{+}24, Woo22, WMB^{+}22]. Supporting [AHC^{+}20, WOL^{+}21, SCM^{+}23]. supports [CPD^{+}21]. surf [MHO^{+}20, MGO^{+}23]. Surface [DCL^{+}21, EDH^{+}22, IHL21, DLL^{+}21, JYK^{+}21, JRPHM^{+}22, LHL^{+}21]. Surfacing [BFFC20]. surfclam [MPK^{+}22, SBM^{+}22]. surprise [Sch21a]. Surveillance [SRM^{+}20, FW22, FMF^{+}20, SROR21]. Survey [JBF^{+}20, SRM^{+}20, ARE^{+}22, Ano23g, Bac24, CWB^{+}23, DLL^{+}21, DHB^{+}21, FOC^{+}23, GCT^{+}23, HJS^{+}23, LVG^{+}20, LWH^{+}21, LMW^{+}22, MHA^{+}21, ODHF^{+}23, OBK21, PWCL21, PCCL21, PMSK^{+}20, RFB^{+}22, RMF^{+}22, YARE23]. survey-based [DHB^{+}21, LVG^{+}20]. surveying [SPd^{+}22, VCF^{+}20]. surveys [BFFC20, BSW22, DLL^{+}21, GETvB21, GAM^{+}22, JDD^{+}21, KTK20, LDW23, MSWB22, MDP^{+}21, MSP^{+}20, ODT^{+}22, OHM^{+}23, PDP^{+}23, SMSP^{+}23, TR20, VRM^{+}23, YSL22, dBSQK21]. Survival [FAV^{+}20, CCM^{+}21, DYS21, GPM21, HRB^{+}21, PBD^{+}21, SG21, TSE^{+}21, WPN^{+}21, vdWBL22]. suspects [BGR^{+}22b]. sustainability [BBB^{+}22, BPV^{+}20, DNW^{+}23, HAC^{+}23, HHO^{+}20, SBH^{+}21, WZJ^{+}22]. Sustainable [PY22, CFB^{+}21, Ear21, FSN^{+}22, HBC^{+}22, HAPW21, Hut22, KHHD20, PMM^{+}23, PRAR23, RCH^{+}21, TDB^{+}21, WOB^{+}23, WOL^{+}21, vdBDH21]. Svedäng [Bra22, SSV^{+}22a]. Sverdrup [KPS21]. SW [dlCRT^{+}22]. Swedish [BFW^{+}21, HPS^{+}23, MD21, SPd^{+}22, WPB20]. swimming [NBL^{+}21, OGMA20]. SWIR [CNWW22]. switching [KVS^{+}23]. Sydeman [BRG22]. syndromes [JDJK22]. synergy [PJCS^{+}23b]. synthesis [Bac24, CLL21, HMA^{+}23, SVML^{+}23]. synthetic [MBA^{+}20]. system [Ano22c, BCBJ^{+}20, DASLRB22, HHK21, KBK^{+}22, MMAE20, SSS^{+}20, TK20]. Systematic [MCC^{+}22, AVC^{+}21]. systems [BCC^{+}24, DHN^{+}21, Har21, HVDvD23, HAC^{+}23, MTTG^{+}23, MPH^{+}21, PM20, TMLH21].

TAC [WZJ^{+}22]. tactics [LHN^{+}22, RAH^{+}22]. tag [MCD^{+}20]. tag-recapture [MCD^{+}20]. tagged [CSC^{+}23]. tags [ADS^{+}23, KKS^{+}20]. take [LLPdL22, Mur21, OHM^{+}23, SBW^{+}22, SSV^{+}22a, VO23, VNR^{+}23, dBABS^{+}22]. taking [BD20]. talking [KSK^{+}21]. target [AKL^{+}21, BM20, CKH^{+}21, GAP21, LDW23, POUH23, dSRKC22, SK21, SBPA^{+}21, YMH^{+}24]. targeted [KSHO23]. targeting [LHN^{+}22, MV23]. taxa [SRA^{+}21]. taxonomic [FFS^{+}22, PCL^{+}21, PASL21]. taxonomy [Pep23]. technical [Kra23]. techniques [RBC^{+}23]. technologies [Pro20, Suu22]. technology [ECBJ22, LWL^{+}22]. TEDs [Jen23]. telemetry [BGH^{+}20, IOS21, PVGC^{+}22, VCD^{+}20]. telephone [TR20]. Telling [HKC^{+}23, Moo23b]. temperate [CRG^{+}23, LHL^{+}21, MHJ^{+}20, PBS^{+}23, SL21, VNBLZR^{+}23, WBB^{+}21]. Temperature [CCJ^{+}21, HRB^{+}21, SSF^{+}22, VRH^{+}23, CRB21, DSDA22, DDL^{+}20, HCCR20, HMBDP21, HJ22, JYK^{+}21, JRPHM^{+}22, KTY^{+}21, LIR^{+}21, LTN21, MPN20, RBG^{+}20, SH22, SMOD21, SBD^{+}21, VBO^{+}21a, VBO^{+}21b, VDAT20, vLNC20]. Temperature-dependent [CCJ^{+}21, VRH^{+}23, CRB21]. temperature-size [SMOD21]. temperatures [LHL^{+}21]. Temporal [FFS^{+}22, GBMM^{+}23, MHJ^{+}20, Ano23g, AA21, BLT23, BPC^{+}20, BDL^{+}22, BAS^{+}21, CHZC22, CMCF21, DSGG20, GCT^{+}23, IEIT23, LAB^{+}22, LMvEM23, MAI^{+}22, MKSM23, OBJ^{+}22, OT22, PDP^{+}23, SSH^{+}20, SGM^{+}24, TAB^{+}20, TFWE^{+}22, WPP^{+}21, Ano23k]. Ten [Har20]. Tensions [Hut22]. tenuis [MPQS20]. term [BJBB20, DSGG20, DPMZ21, LB23, MBP^{+}20, WCCY20, WZJ^{+}22, YSL22]. Terrestrial [MvdMS^{+}23]. territorial [HBD^{+}21]. Territories [RES^{+}23]. test [HBFO20]. Testing [MD22, OTL^{+}23, CSH^{+}21, PDL20, TBH^{+}20, Tho22]. their [ADG^{+}20, GGLS22, LIR^{+}21, LRHW21, MMN^{+}22, MCH^{+}24, OGMA20, RCH^{+}21, SCN^{+}23, SRA^{+}21, TLS21, dGW^{+}22]. them [LBH^{+}21]. theoretic [BPV^{+}20]. Theoretical [dLP21]. theory [Sha21]. There [CFP20]. Thermal [CCS^{+}24, TLS21, DBGP22, IOS21, MCC^{+}22, MWT^{+}21, RPB^{+}21, SL21, YC20]. thermohaline [VRH^{+}23]. things [CvPTAJ22]. think [Bea23]. Thinking [KF21, SSV^{+}22b, Ano22b]. thornback [MD22]. thorns [WDPM23]. thorny [KKS^{+}20]. thoughts [WBB^{+}23]. threat [BPC^{+}24, CBB21, LBPC^{+}20]. threats [BBB^{+}22]. Three [Hol21, Ano24b, BGH^{+}20, DSF^{+}21, FND23, GPJ^{+}21, Hol20, KDS^{+}22, OOE^{+}22, PBS^{+}23, RVR^{+}24, WRC^{+}21, dMBJ22]. three-dimensional [BGH^{+}20]. three-spined [OOE^{+}22]. threespine [LIR^{+}21]. thresholds [FXG^{+}20, HVDvD23, SAB^{+}20]. throughout [FPY^{+}22, FNB^{+}20, SJS21]. Thunnus [ADS^{+}23, BFFC20, DRP^{+}23]. thynnus [ADS^{+}23, BFFC20]. tidal [WBWS21]. Time [MBP^{+}20, CBB21, CMS^{+}23, JMR^{+}22, LVT^{+}23, MPN20, NFN20, OTMN20, PCG^{+}22, PY22, dSRKC22, SASM22, SSV^{+}22b, TBP20, WCL^{+}23, Ano22b]. time-area [CBB21]. Time-variant [MBP^{+}20]. time-varying [CMS^{+}23, OTMN20]. times [MD21]. timescale [BWBH21]. timing [MAC21, VSL22]. tissue [AMP^{+}23]. tolerance [PLN20]. too [Pec23]. tool [BBP^{+}23, EWD^{+}21, EWD^{+}22, JDJK22, PEWN20, WJB^{+}24]. tools [MCS^{+}20]. toothfish [EMS^{+}21]. topic [WJ22]. Total [PPMD22]. tourism [CWC^{+}23, MPMGGAHR20]. towed [SSP^{+}23]. tows [HJP22]. track [DSDA22]. Tracking [FSS^{+}23, HBD^{+}20, IOS21, MYW^{+}21, MMCS^{+}24, OMM^{+}23, vEMvH^{+}21]. Trade [SPP21, WPP^{+}21, EGMO^{+}20a, EGMO^{+}20b, HAPW21, OBK21, YABM20]. trade-off [EGMO^{+}20a, EGMO^{+}20b, HAPW21]. Trade-offs [SPP21, WPP^{+}21, OBK21]. Trading [MYAH21]. traditional [HJS^{+}23, MPMGGAHR20]. train [HMWS22]. training [BVS^{+}23]. trait [OBJ^{+}22]. traits [DeM22, DBGP22, DMK^{+}21, FDK20, FWA^{+}21, GGRB22, JNT^{+}22, OH23, SGM22]. trajectory [ECR^{+}21, PRCM22]. transboundary [ODT^{+}22]. Transdisciplinarity [SOCN^{+}22]. transect [GSWM^{+}21, GSWM^{+}22]. transfer [LZZ^{+}21]. transferability [HBFO20]. transferable [HHK21]. transferred [MvdMS^{+}23]. transformation [RHO22]. transformative [CMD^{+}23a, SOCN^{+}22]. transition [CFP20, NAS^{+}20, WMB^{+}22]. transitions [PPV^{+}22]. translating [LCG^{+}22]. transmitting [KKS^{+}20]. transparency [OMM^{+}23, Pro20]. transparent [CSH^{+}21]. Transport [LRGW23, BCBJ^{+}20]. Transport-driven [LRGW23]. trap [GPM21, HCB^{+}24, Ste21]. trap-catch [HCB^{+}24]. trap-like [GPM21]. traps [Bac24, Ste21]. Trawl [JBF^{+}20, SACP^{+}21, ARHM21, ASW^{+}23, AA21, BSW22, BSH^{+}22, CWB^{+}23, CCR^{+}24, DHB^{+}21, FAV^{+}20, GLM^{+}23, HJS^{+}23, KKS^{+}20, KP24, MDP^{+}21, Mil24, MTK^{+}21, MHMP^{+}23, PE21, RHvD^{+}20, RDM^{+}21, SMS^{+}22, SWH^{+}23, WCCY20, dCKS^{+}22]. trawlers [JBE^{+}21, TDC^{+}23]. trawling [AMC^{+}23, BM20, CRB21, DMK^{+}21, HAC^{+}23, HAP^{+}21, JDD^{+}21, LBH^{+}21, VRM^{+}23, vDBF^{+}20]. trawls [BRvL^{+}23, KMK21a, RDM^{+}21]. treatment [CRD^{+}22]. tree [LZWH20]. tree-based [LZWH20]. trees [NHL^{+}20]. Trend [LQC^{+}22]. Trends [BPPC^{+}20, PBD^{+}21, BPC^{+}20, CT23, KMG22, MPQS20, PCH^{+}24, STS^{+}20, SDR^{+}21, YSL22]. trials [ASW^{+}23, BMM^{+}21]. tribal [NKS^{+}23]. triggers [CSH^{+}21]. trip [ASV21]. trip-based [ASV21]. triple [DDP^{+}21, vdBDH21]. triplet [CZ23]. trophic [BOLPHL21, CRG^{+}23, FND23, KÓP^{+}21, LMH22, LZZ^{+}21, PFH^{+}22, Thi20, TND^{+}20]. Tropical [BHL^{+}23, EDC20, FBF^{+}21, HMBDP21, LLALF22, MVNR^{+}21, PBF^{+}21, QBF^{+}21, WGKG21, ADG^{+}20]. trout [BPC^{+}21, HBMC20, KBJA20, KBK^{+}22, KBBGA22]. trust [Pro20]. truthed [AKK^{+}21]. trutta [BPC^{+}21, KBJA20, KBK^{+}22, KBBGA22]. tshawytscha [YFM^{+}20]. tuberculata [AABM^{+}20]. tuna [ADS^{+}23, ADBC^{+}22, ANA^{+}23, BFFC20, CCH^{+}23, DRP^{+}23, DCL^{+}21, ECR^{+}21, EHV^{+}21, FBF^{+}21, FFT^{+}22, GBGA^{+}23, HBD^{+}20, HBD^{+}21, Jen23, LWZC23, LTK20, MVMdS21, MKGC20, NFN20, NSS^{+}23, RHO22, WGKG21, WCL^{+}23, YABM20]. turbid [MYMV20, SOB23]. Turbidity [LS20b]. turbine [BHDD22, WBWS21]. Turning [TBM^{+}22]. turnover [JNT^{+}22]. turtle [CBT^{+}20, TYF23]. Turtles [Jen23, BFD^{+}23, dAdSM21]. Tweedie [TMS23]. twelve [MRC^{+}21]. Two [JLM^{+}23, CWB^{+}23, DBGP22, DHN^{+}21, LPC^{+}23, LVL^{+}22, MTT^{+}23, RPVT^{+}21, VSL22, VAS^{+}20, YC20]. type [MB21]. types [AKL^{+}21, TK20]. Tyrrhenian [AMC^{+}23].

U.S. [Lin21, PFP^{+}22]. UAV [MSWB22]. ugly [Kra23, MEOA23]. UK [RRR^{+}24]. umbra [GPV^{+}22]. uncertain [LSR^{+}21]. uncertainties [HMF^{+}21, Kai20, SPP21]. uncertainty [CSM^{+}22, KBW22, MGB^{+}23, MFP20, OBK21, PJP20, SLRC21]. uncover [PDP^{+}23]. Uncovering [CWL^{+}20]. uncovers [CZS^{+}23]. Uncrewed [DLL^{+}21]. Undaria [ESK^{+}21]. under-coverage [VCF^{+}20]. underlying [KH21, ZEM21]. understand [Bea22, HKC^{+}23]. Understanding [ECR^{+}21, ODT^{+}22, SAM^{+}22, Suu22, AGSS22, JDJK22, MKSM23, PRAR23, PDH^{+}21]. Underwater [VAS^{+}20, CZSR21, CSC^{+}23, SSS^{+}20]. unexpected [DeM22]. unified [MBS^{+}23]. unintended [Bor21]. Union [BPPC^{+}20]. Unit [SFPP^{+}23]. United [HMC^{+}20, GSWBK20, HBD^{+}21, MMC^{+}23, PY22, SAB^{+}20, SCG^{+}20]. units [BBP^{+}23, BNV^{+}21]. unknown [Kos21, Sha21]. Unlocking [GHJ^{+}22]. unmanaged [BP20]. Unoccupied [MSWB22]. unregulated [BFJB21]. unsupervised [AKL^{+}21, ÁEPCL20]. untrawlable [RFB^{+}22]. unwanted [BBB^{+}22, GLM^{+}23, Kra23, PM23]. Upcalling [Ano24c]. upon [BR22]. ups [Bea22, Ste21]. Uptake [JEPT23, KLG^{+}23, PBS^{+}23, SMK23]. upwelling [Ano22c, DASLRB22, MMAE20, SSF^{+}22, SSS^{+}23]. urban [GGA^{+}23]. urchins [SKM^{+}22]. urgently [DBB^{+}20]. USA [JSR^{+}22, SACP^{+}21]. Use [EDH^{+}21, PCH^{+}24, AMD^{+}23, BFD^{+}23, DPH^{+}21, EHV^{+}21, FKTM20, GVD^{+}23, JSR^{+}22, KTK20, LVG^{+}20, OOB^{+}22, SWH^{+}23, TSWM23, WGKG21, WOL^{+}21]. used [HHK21, LQC^{+}22, MVMdS21, SMS^{+}22]. Using [CZ23, DMK^{+}21, FDMK21b, FDMK21c, GETvB21, HCM24, MSVSA22, SROR21, ÁEPCL20, AMP^{+}23, Bac24, BRvL^{+}23, BWJ^{+}20, CNWW22, FRSM22, GPQ^{+}20, GICP^{+}22, HvdMLS21, JDJK22, LS20a, LSD^{+}20, LZWH20, LVT^{+}23, LMvEM23, LTK20, MBA^{+}20, MCH^{+}24, ODHF^{+}23, OHM^{+}23, OTL^{+}23, PVHV20, PHB^{+}20, PKAT20, PMO^{+}21a, PMO^{+}21b, PMSK^{+}20, SRB^{+}23, SSP^{+}23, SRBT23, SHG^{+}21, SRM^{+}20, TK20, VAS^{+}20, VCF^{+}20, WRC^{+}21, WSL^{+}23, WJ22, ZYJZ23, vEMvH^{+}21, vDBF^{+}20]. USV [DLL^{+}21]. utility [SOB23, YMAH20]. Utilization [KKS^{+}20, EGMO^{+}20a, EGMO^{+}20b].

Validation [KSK^{+}21, DGM^{+}24]. valuable [GA23]. valuation [PFT^{+}20]. value [CBFS23, KNP21, SCP^{+}21, WBB^{+}23, vdBDH21]. values [Day20]. valve [HMWS22]. Vancouver [VSS^{+}22]. Variability [BHL^{+}23, KBBGA22, Ano22c, Ano22e, Ano23l, Ano23e, CMCF21, DASLRB22, GGOP21, GGOP22, HYE^{+}22, JRDW22, JJB^{+}22, JKF^{+}22, KPM^{+}22, LAB^{+}22, LWZC23, LAGC23, LHL^{+}21, LRHW21, MAI^{+}22, MKS^{+}21, OBK21, PRMS22, PCG^{+}22, PDH^{+}21, VLUB^{+}22, WCLB23, WBA^{+}22, WFG^{+}23, YC20, dLP21]. Variable [HKZ^{+}23, SMO22, TFH^{+}21]. variables [JBDK23, THSZ21]. variance [BNB21, FSC^{+}23]. variant [KEW^{+}21, MBP^{+}20]. variation [BSW22, BCP^{+}21, CGH^{+}23, GETvB21, IEIT23, KEW^{+}21, OOB^{+}22, OBJ^{+}22, OGMA20, PJP20, SGM^{+}24, TAB^{+}20, TFWE^{+}22, VPCM^{+}22, VSS^{+}22, dBABS^{+}22]. variations [BDL^{+}22, DDL^{+}20, JRPHM^{+}22, SH22]. varies [Ano23d, HMWS22, lCRGCCA23]. various [VPP^{+}23]. vary [PTR22]. varying [CMS^{+}23, HMBDP21, OTMN20, ZJF^{+}21, ZCX^{+}20]. vehicle [DLL^{+}21]. Venice [BBB^{+}21]. vents [BKR^{+}21]. verification [MPK^{+}22]. verreauxi [WML^{+}21]. versus [BRvL^{+}23, MPA^{+}23]. Vertebrae [ANA^{+}23]. vertical [BSW^{+}23, BFFC20, BHL^{+}23, HJP22, MRC^{+}21, MTK^{+}21]. vessel [MMCS^{+}24, OMM^{+}23, PEWN20, SSH^{+}20]. vessel-level [PEWN20]. vessels [AMP^{+}23, CVJ^{+}20, GGA^{+}23, LQC^{+}22, VAS^{+}20]. Vibrio [NKA^{+}21]. Victoria [PMSK^{+}20]. video [FMF^{+}20, JDD^{+}21, MSWB22, RFB^{+}22, RMF^{+}22]. videos [SSS^{+}20, TK20]. Vietnam [BNV^{+}21]. views [SSV^{+}22a]. virginica [CCM^{+}21]. viscosity [KAK^{+}21]. viscous [KAK^{+}21]. vision [LWL^{+}22]. vocalization [SSP^{+}23]. Voluntary [GT22, MSW21]. vs [HFP^{+}23, JDD^{+}21, KTK20, Met20]. vulgaris [LAB^{+}22, RUdPFRA^{+}21]. Vulnerability [LLALF22, ALLP^{+}22, GJU^{+}23]. Vulnerable [DPH^{+}21, BFJB21, GJU^{+}23, LBPC^{+}20, SRA^{+}21, vDHR^{+}22].

wait [Bea23]. Walbaum [GEBS^{+}23]. Wales [GGW^{+}23]. walleye [DLL^{+}21, LDW23]. want [Ano23c, PAC^{+}23]. warm [HCM24]. warmer [SDSMS21]. warming [ESK^{+}21, FRP^{+}23, GA23, GPS^{+}21b, GPS^{+}21a, HYE^{+}22, KJP^{+}22, MKK^{+}24, ZJH^{+}23, dGW^{+}22]. warning [ZL22]. Was [Fie21, PF21]. waste [CCC^{+}21, WBB^{+}21]. Water [MRC^{+}21, MPN20, FCD^{+}23, HCCR20, HCM24, HSMK21, MHA^{+}21, NKA^{+}20, PASL21, RTQW20, TDC^{+}23, VNR^{+}23, XCK^{+}23]. waters [Ano22a, Ano22e, BBE^{+}20, CLMB22, DLA^{+}21, FMPH24, GGOP21, GPJ^{+}21, GGOP22, HBD^{+}21, KÓP^{+}21, LBPC^{+}20, MWT^{+}21, MSWB22, MYW^{+}21, OGMA20, SB21, dGW^{+}22, QCM^{+}20]. wave [CNWW22, TWC^{+}20, TSW^{+}21]. Weak [DRP^{+}23]. weaned [PMS^{+}20]. weather [LTMG23]. web [EDH^{+}21, KMK^{+}21b, MTCC21, MVNR^{+}21, PFH^{+}22, SCMT^{+}22, Thi20]. webs [JSR^{+}22, MPH^{+}21]. weight [IEIT23, MPA^{+}23, SCN^{+}23]. weight-at-age [IEIT23]. weighting [TMS23]. West [BFW^{+}21, FPY^{+}22, SAB^{+}20, STS^{+}20, HPS^{+}23, HBD^{+}20, LBH^{+}21, MD21, VSS^{+}22, LTN21, VPP^{+}23]. West-Atlantic [VPP^{+}23]. Westermost [RRJ22]. Western [FFS^{+}22, MBA^{+}20, Ano24a, FFT^{+}22, HWL^{+}22, HHD^{+}23, LLPdL22, LBPC^{+}20, LVT^{+}23, SHK^{+}23, SPd^{+}22, WCL^{+}23, BJPP^{+}20, FBF^{+}21, FKTM20, RBF^{+}22, TR20, VLUB^{+}22]. Whale [EMS^{+}21, Moo23a, SPBJ21, SPJ23]. whales [Ano24c, ADG^{+}20, CBB21, VBJ^{+}21]. whaling [Man21, VBJ^{+}21]. where [HvdMLS21, LPC^{+}23]. white [BPC^{+}24, GHB^{+}23, MGB^{+}23, dBABS^{+}22, LIR^{+}21]. white-chinned [BPC^{+}24]. white-sided [GHB^{+}23]. who [RCB^{+}21, VPCM^{+}22]. whole [CCJ^{+}21, JLM^{+}23]. whole-life [CCJ^{+}21]. Wide [PMS^{+}20, BLM^{+}21, DRP^{+}23, FPY^{+}22, HMNF22, KH22, NBA^{+}23, PDP^{+}23, Var21]. wide-ranging [PDP^{+}23]. wideband [AKK^{+}21]. Widespread [JSR^{+}22]. width [VDAT20]. width-based [VDAT20]. Wild [SG21, VLD^{+}21, Ano23b, DFK^{+}22, HUJ^{+}21, KBK^{+}22, LVT^{+}23, Man21, MSN^{+}23, MvdMS^{+}23, SBB^{+}22]. wildlife [Ano22a, DLA^{+}21]. will [Bea23, Mon22]. William [CRJ20, ZB24, Dol20]. willing [SMK23]. wind [Ano24c, BHDD22, CvdWC^{+}20, Met20, MPK^{+}22, PRD20, RRJ22, SBM^{+}22, TSWM23, VCD^{+}20, WBG^{+}22]. windfarms [BVV^{+}20, SFG22]. winter [GGLS22, MPN20, OOB^{+}22, PRMS22]. within [ASV21, FFT^{+}22, FSC^{+}23, SCM^{+}23, SCMT^{+}22, TSWM23, TFH^{+}21, WESWT21]. without [BAS^{+}21, dBSQK21]. women [SCM^{+}23]. workflow [MMCS^{+}24]. World [TSD^{+}21, Ano20c, BMR21, Day20, EHV^{+}21, HBC^{+}22, KH22, SRMS21]. world-wide [KH22]. Wrasse [BFW^{+}21]. wrasses [Ano22a, DLA^{+}21]. wreck [GLW21]. writing [Leg21]. wrong [CvPTAJ22].

Year [CRR^{+}22, EDC20, LVT^{+}23, Lov20, PRMS22, SL21, vdWBL22, Ano22d]. Year-long [EDC20]. Year-round [CRR^{+}22, Ano22d]. years [BFD^{+}23, FNB^{+}20, HHL^{+}21, KBBGA22, Sha21, SDGR^{+}20, vLNC20]. yellow [Ano23i, BC23, LLLS20]. yellow-phase [Ano23i, BC23]. yelloweye [CT23]. yellowfin [ECR^{+}21, FFT^{+}22, MVMdS21]. yellowtail [STK^{+}20b, STK^{+}20a]. yield [BWBH21, Ear21, FSN^{+}22, GDI^{+}20]. young [CvdWC^{+}20, PRMS22, SL21]. young-of-the-year [PRMS22, SL21]. you’re [dCKS^{+}22]. YouTube [SCCA20].

Zalophus [CLS^{+}22]. Zealand [Ano23g, EVLK^{+}21, GCT^{+}23, MSW21]. zone [MTTG^{+}23, MGO^{+}23, NBL^{+}21, SBW^{+}22, Thi20]. zones [MHO^{+}20, NAS^{+}20, SASM22]. zoning [CCC^{+}21]. zooplankton [BPKM21, BPKM22, EWD^{+}21, EWD^{+}22, FFS^{+}22, GCT^{+}21, HJP22, JMR^{+}22, MKN^{+}21, MHA^{+}21, MGO21, MMAE20, OBJ^{+}22, PY22, SDW^{+}21, UVI^{+}21a, UVI^{+}21b, VSS^{+}22, WBC^{+}22, YSL22]. Zostera [Ano24b, QSKL21, RVR^{+}24].

References

Ashrafi:2021:IIT

[AA21] Tannaz Alizadeh Ashrafi and Keita Abe. Intra- and intertemporal effort allocation and profit-maximizing strategy of trawl fishery. ICES Journal of Marine Science, 78(8):2943– 2957, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/8/2943/6364350.

Avignon:2020:IIE

[AABM^{+}20] Solène Avignon, Stéphanie Auzoux-Bordenave, Sophie Martin, Philippe Dubois, Aïcha Badou, Manon Coheleach, Nicolas Richard, Sarah Di Giglio, Loïc Malet, Arianna Servili, Fanny Gaillard, Sylvain Huchette, and Sabine Roussel. An integrated investigation of the effects of ocean acidification on adult abalone (Haliotis tuberculata). ICES Journal of Marine Science, 77(2):757–772, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
757/5699268.

Armstrong:2022:PDC

[AAHT22] Claire W. Armstrong, Margrethe Aanesen, Stephen Hynes, and Rob Tinch. People do care about the deep sea. A comment on jamieson et al. (2020). ICES Journal of Marine Science, 79(8):2336–2339, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2336/6712687. See [JSLC21] and correction

[Ano23h].

Anderson:2021:ESH

[ABB^{+}21a] Emory D. Anderson, Valerio Bartolino, Silvana Birchenough, Howard I. Browman, Mark Gibbs, Manuel Hidalgo, Raúl Prellezo, and Katherine Yates. Erratum to: Sidney Holt’s legacy lives on in fisheries science. ICES Journal of Marine Science, 78(8):3016, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/

3016/6381243. See [ABB^{+}21b].

Anderson:2021:SHL

[ABB^{+}21b] Emory D. Anderson, Valerio Bartolino, Silvana Birchenough, Howard I. Browman, Mark Gibbs, Manuel Hidalgo, Raúl Prellezo, and Katherine Yates. Sidney Holt’s legacy lives on in fisheries science. ICES Journal of Marine Science, 78(6):2150–2154, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/

2150/6272137. See erratum [ABB^{+}21a].

Abelson:2020:WSF

[Abe20] Avigdor Abelson. Are we sacrificing the future of coral reefs on the altar of the “climate change” narrative? ICES Journal of Marine Science, 77(1):40–45, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/1/
40/5673597.

Andrews:2022:EHA

[ADBC^{+}22] Adam J. Andrews, Antonio Di Natale, Darío Bernal-Casasola, Veronica Aniceti, Vedat Onar, Tarek Oueslati, Tatiana Theodropoulou, Arturo Morales-Muñiz, Elisabetta Cilli, and Fausto Tinti. Exploitation history of Atlantic bluefin tuna in the eastern Atlantic and Mediterranean — insights from ancient bones. ICES Journal of Marine Science, 79(2):247–262, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/2/247/6511216.

Avila:2020:HWE

[ADG^{+}20] Isabel Cristina Avila, Carsten F. Dormann, Carolina García, Luis Fernando Payán, and María Ximena Zorrilla. Humpback whales extend their stay in a breeding ground in the Tropical Eastern Pacific. ICES Journal of Marine Science, 77(1):109–118, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/109/5688942.

Aalto:2023:EBT

[ADS^{+}23] Emilius A. Aalto, Simon Dedman, Michael J. W. Stokesbury, Robert J. Schallert, Michael Castleton, and Barbara A.

Block. Evidence of bluefin tuna (Thunnus thynnus) spawning in the Slope Sea region of the Northwest Atlantic from electronic tags. ICES Journal of Marine Science, 80(4):861– 877, May 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/4/861/7041968.

Alvarez-Ellacuria:2020:IBU

[ÁEPCL20] Amaya Álvarez-Ellacuría, Miquel Palmer, Ignacio A. Catalán, and Jose-Luis Lisani. Image-based, unsupervised estimation of fish size from commercial landings using deep learning. ICES Journal of Marine Science, 77(4):1330–1339, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/4/1330/5638881.

Andersen:2022:RUR

[AGSS22] Lasse Berg Andersen, Ellen Sofie Grefsrud, Terje Svåsand, and Nina Sandlund. Risk understanding and risk acknowledgement: a new approach to environmental risk assessment in marine aquaculture. ICES Journal of Marine Science, 79 (4):987–996, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/987/6547884.

Andrews:2020:SEC

[AHC^{+}20] Evan J. Andrews, Sarah Harper, Tim Cashion, Juliano Palacios-Abrantes, Jessica Blythe, Jack Daly, Sondra Eger, Carie Hoover, Nicolás Talloni-Alvarez, Louise Teh, Nathan Bennett, Graham Epstein, Christine Knott, Sarah L. Newell, and Charlotte K. Whitney. Supporting early career researchers: insights from interdisciplinary marine scientists. ICES Journal of Marine Science, 77(2):476–485, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/2/476/5704875.

Ashrafi:2023:EQP

[AHK23] Tannaz Alizadeh Ashrafi, Øystein Hermansen, and Sturla F. Kvamsdal. The effect of quota portfolio composition on optimal harvest strategy and profitability in a multi-species fishery. ICES Journal of Marine Science, 80(8):2099–2113, October 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/8/2099/7259324.

Arkhipkin:2021:SAM

[AHP^{+}21] Alexander I. Arkhipkin, Lisa C. Hendrickson, Ignacio Payá, Graham J. Pierce, Ruben H. Roa-Ureta, Jean-Paul Robin, and Andreas Winter. Stock assessment and management of cephalopods: advances and challenges for short-lived fishery resources. ICES Journal of Marine Science, 78(2):714–730, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/2/714/5828442.

Adams:2023:EDM

[AJC^{+}23] Clare I. M. Adams, Gert-Jan Jeunen, Hugh Cross, Helen R. Taylor, Antoine Bagnaro, Kim Currie, Chris Hepburn, Neil J. Gemmell, Lara Urban, Federico Baltar, Michael Stat, Michael Bunce, and Michael Knapp. Environmental DNA metabarcoding describes biodiversity across marine gradients. ICES Journal of Marine Science, 80(4):953–971, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/953/7058017.

Agersted:2021:MEI

[AKK^{+}21] Mette Dalgaard Agersted, Babak Khodabandeloo, Thor A. Klevjer, Eva García-Seoane, Espen Strand, Melanie J. Underwood, and Webjørn Melle. Mass estimates of individual gas-bearing mesopelagic fish from in situ wideband acoustic measurements ground-truthed by biological net sampling. ICES Journal of Marine Science, 78(10):3658–3673, December 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/10/3658/6425096.

Agersted:2021:AUC

[AKL^{+}21] Mette Dalgaard Agersted, Babak Khodabandeloo, Yi Liu, Webjørn Melle, and Thor A. Klevjer. Application of an unsupervised clustering algorithm on in situ broadband acoustic data to identify different mesopelagic target types. ICES Journal of Marine Science, 78(8):2907–2921, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2907/6360556.

Aragao:2022:IRD

[ALLP^{+}22] Guilherme Martins Aragão, Lucía López-López, Antonio Punzón, Elena Guijarro, Antonio Esteban, Encarnación García, José Manuel González-Irusta, Julia Polo, Miguel Vivas, and Manuel Hidalgo. The importance of regional differences in vulnerability to climate change for demersal fisheries. ICES Journal of Marine Science, 79(2):506–518, March 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/2/506/6324219.

Albonetti:2023:DMT

[AMC^{+}23] Linda Albonetti, Giulia Maiello, Alessia Cariani, Paolo Carpentieri, Alice Ferrari, Alice Sbrana, Peter Shum, Lorenzo Talarico, Tommaso Russo, and Stefano Mariani. DNA metabarcoding of trawling bycatch reveals diversity and distribution patterns of sharks and rays in the central Tyrrhenian Sea. ICES Journal of Marine Science, 80(4):664–674, May 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/4/664/7059508.

Andrzejaczek:2023:SDH

[AMD^{+}23] Samantha Andrzejaczek, Chloe S. Mikles, Jonathan J. Dale, Michael Castleton, and Barbara A. Block. Seasonal and diel habitat use of blue marlin Makaira nigricans in the North Atlantic Ocean. ICES Journal of Marine Science, 80(4): 1002–1015, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/4/1002/7067461.

Anderson:2023:EDC

[AMP^{+}23] Giulia Anderson, Jed I. Macdonald, Joanne Potts, Pierre Feutry, Peter M. Grewe, Marion Boutigny, Campbell R. Davies, Jeff A. Muir, Francois Roupsard, Caroline Sanchez, and Simon J. Nicol. Evaluating DNA cross-contamination risk using different tissue sampling procedures on board fishing and research vessels. ICES Journal of Marine Science, 80 (4):728–738, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/4/728/7025294.

Almodovar:2020:SIS

[ANA^{+}20] Ana Almodóvar, Graciela G. Nicola, Daniel Ayllón, Clive N. Trueman, Ian Davidson, Richard Kennedy, and Benigno Elvira. Stable isotopes suggest the location of marine feeding grounds of South European Atlantic salmon in Greenland. ICES Journal of Marine Science, 77(2):593–603, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/593/5716105.

Andrews:2023:VRI

[ANA^{+}23] Adam J. Andrews, Antonio Di Natale, Piero Addis, Federica Piattoni, Vedat Onar, Darío Bernal-Casasola, Veronica Aniceti, Gabriele Carenti, Verónica Gómez-Fernández, Fulvio Garibaldi, Arturo Morales-Muñiz, and Fausto Tinti. Vertebrae reveal industrial-era increases in Atlantic bluefin tuna catch-at-size and juvenile growth. ICES Journal of Marine Science, 80(4):836–847, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/4/836/7034402.

Andersen:2021:NSD

[ANB^{+}21] Niels Gerner Andersen, Stefan Neuenfeldt, Bjarte Bogstad, Ken Haste Andersen, and Jan Erik Beyer. Nutritional status determines apparent assimilative capacity and functional response of marine predatory fish. ICES Journal of Marine Science, 78(10):3615–3624, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/

10/3615/6386083. See corrigendum [ANB^{+}22].

Andersen:2022:CNS

[ANB^{+}22] Niels Gerner Andersen, Stefan Neuenfeldt, Bjarte Bogstad, Ken Haste Andersen, and Jan Erik Beyer. Corrigendum to: Nutritional status determines apparent assimilative capacity and functional response of marine predatory fish. ICES Journal of Marine Science, 79(1):243, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/243/6419776. See [ANB^{+}21].

Anonymous:2020:AML

[Ano20a] Anonymous. Applications of machine learning and artificial intelligence in marine science: all articles. ICES Journal of Marine Science, 77(4):1267–1455, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1267/5873749.

Anonymous:2020:PNC

[Ano20b] Anonymous. Publisher’s note: a cross continental scale comparison of Australian offshore recreational fisheries research and its applications to Marine Park and fisheries management. ICES Journal of Marine Science, 77(6):2206, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2206/5960369.

Anonymous:2020:SSN

[Ano20c] Anonymous. Science in support of a nonlinear nonequilibrium world: all articles. ICES Journal of Marine Science, 77(4):1456–1588, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1456/5873755.

Anonymous:2022:CFC

[Ano22a] Anonymous. Correction to: Fish consumption by great cormorants in Norwegian coastal waters — a human-wildlife conflict for wrasses, but not gadids. ICES Journal of Marine Science, 79(7):2162, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
7/2162/6651438.

Anonymous:2022:CRT

[Ano22b] Anonymous. Correction to: Re-thinking the “ecological envelope” of Eastern Baltic cod (Gadus morhua): conditions for productivity, reproduction, and feeding over time. ICES Journal of Marine Science, 79(10):2752, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/10/2752/6822702.

Anonymous:2022:CSI

[Ano22c] Anonymous. Correction to: Spatial and interannual variability in the distribution of euphausiid life stages in the permanent upwelling system off northern Chile. ICES Journal of Marine Science, 79(5):1709, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1709/6589470. See [DASLRB22].

Anonymous:2022:CYR

[Ano22d] Anonymous. Correction to: Year-round genetic monitoring of mixed-stock fishery of Atlantic cod (Gadus morhua); implications for management. ICES Journal of Marine Science, 79(6):1979, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
6/1979/6613267. See [CRR^{+}22].

Anonymous:2022:PNE

[Ano22e] Anonymous. Publisher’s note: Inter-annual and decadal variability of Calanus finmarchicus and C. hyperboreus in subarctic waters north of Iceland 1990–2020. ICES Journal of Marine Science, 79(3):829, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
3/829/6565774.

Anonymous:2023:CER

[Ano23a] Anonymous. Correction to: An empirical review of ICES reference points. ICES Journal of Marine Science, 80(4): 1156, May 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/4/1156/7033995. See [SVBM^{+}22].

Anonymous:2023:CBR

[Ano23b] Anonymous. Correction to: Behavioural responses of wild, spawning Atlantic cod (Gadus morhua L.) to seismic airgun exposure. ICES Journal of Marine Science, 80(6):1827–1828, August 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/6/1827/7173169. See [MSN^{+}23].

Anonymous:2023:CCD

[Ano23c] Anonymous. Correction to: Co-designing marine science for the ocean we want. ICES Journal of Marine Science, 80(2): 401, March 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/2/401/7071069. See [CMD^{+}23a,

PAC^{+}23].

Anonymous:2023:CCB

[Ano23d] Anonymous. Correction to: Correlation between seabirds and fisheries varies by species at fine-scale pattern. ICES Journal of Marine Science, 80(1):235, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/235/6969397.

Anonymous:2023:CIC

[Ano23e] Anonymous. Correction to: Incorporating conservation limit variability and stock risk assessment in precautionary salmon catch advice at the river scale. ICES Journal of Marine Science, 80(4):1159, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
4/1159/7078954. See [WFG^{+}23].

Anonymous:2023:CIE

[Ano23f] Anonymous. Correction to: Infaunal and epifaunal secondary production in the Barents Sea, with focus on snow crab (Chionoecetes opilio) prey resources and consumption. ICES Journal of Marine Science, 80(4):1157, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1157/7068453. See [HFT^{+}22].

Anonymous:2023:CIS

[Ano23g] Anonymous. Correction to: Integrating survey and observer data improves the predictions of New Zealand spatiotemporal models. ICES Journal of Marine Science, 80(9): 2456, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/9/2456/7283901. See [GCT^{+}23].

Anonymous:2023:CPD

[Ano23h] Anonymous. Correction to: People do care about the deep sea. A comment on Jamieson et al. (2020). ICES Journal of Marine Science, 80(4):1155, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
4/1155/6953698. See [AAHT22].

Anonymous:2023:CSA

[Ano23i] Anonymous. Correction to: Seasonal activity and depth distribution of resident yellow-phase American eels (Anguilla rostrata) in a large fluvial ecosystem. ICES Journal of Marine Science, 80(4):1160, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
4/1160/7085487. See [BC23].

Anonymous:2023:CSF

[Ano23j] Anonymous. Correction to: Standardizing fishery-dependent catch-rate information across gears and data collection programs for Alaska sablefish (Anoplopoma fimbria). ICES Journal of Marine Science, 80(4):1161, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
4/1161/7103486. See [CRLC23].

Anonymous:2023:CTE

[Ano23k] Anonymous. Correction to: Temporal, environmental, and demographic correlates of Ichthyophonus sp. infections in mature Pacific herring populations. ICES Journal of Marine Science, 80(9):2455, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
9/2455/7316694. See [GBMM^{+}23].

Anonymous:2023:CEC

[Ano23l] Anonymous. Correction to: The effects of climateinduced environmental variability on Pacific Ocean squids. ICES Journal of Marine Science, 80(4):1158, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1158/7077002. See [WCLB23].

Anonymous:2024:CSI

[Ano24a] Anonymous. Correction to: Stable isotopes in eye lenses reveal migration and mixing patterns of diamond squid in the western North Pacific and its marginal seas. ICES Journal of Marine Science, 81(1):204, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/81/
1/204/7460304. See [SHK^{+}23].

Anonymous:2024:CSR

[Ano24b] Anonymous. Correction to: The stress response of the seagrass Zostera noltei and three commercial clam species to low salinity associated with heavy rainfall. ICES Journal of Marine Science, 81(2):411, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/2/411/7585544. See [RVR^{+}24].

Anonymous:2024:CUB

[Ano24c] Anonymous. Correction to: Upcalling behaviour and patterns in North Atlantic right whales, implications for monitoring protocols during wind energy development. ICES Journal of Marine Science, 81(1):205, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/1/205/7491222.

Alglave:2022:CSS

[ARE^{+}22] Baptiste Alglave, Etienne Rivot, Marie-Pierre Etienne, Mathieu Woillez, James T. Thorson, and Youen Vermard. Combining scientific survey and commercial catch data to map fish distribution. ICES Journal of Marine Science, 79(4):1133–1149, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
4/1133/6543974.

Allken:2021:DLB

[ARHM21] Vaneeda Allken, Shale Rosen, Nils Olav Handegard, and Ketil Malde. A deep learning-based method to identify and count pelagic and mesopelagic fishes from trawl camera images. ICES Journal of Marine Science, 78(10):3780–3792, December 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/10/3780/6429121.

Azevedo:2021:OBS

[ASV21] Manuela Azevedo, Cristina Silva, and Jon Helge Vølstad. Onshore biological sampling of landings by species and size category within auction sites can be more efficient than trip-based concurrent sampling. ICES Journal of Marine Science, 78(8):2757–2773, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2757/6350842.

Arkhipkin:2023:ICM

[ASW^{+}23] Alexander Arkhipkin, Frane Skeljo, James Wallace, Cian Derbyshire, Ludovic Goyot, Toni Trevizan, and Andreas Winter. Industry-collaborative mesh trials to reduce bycatch in the Falkland Islands skate trawl fishery (Southwest Atlantic). ICES Journal of Marine Science, 80(3):578–590, April 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/3/578/6484966.

Andre:2021:FML

[AVC^{+}21] Laure Vaitiare André, Simon Van Wynsberge, Mireille Chinain, Clémence Mahana Iti Gatti, Alexandra Dempsey, and Serge Andréfouët. A framework for mapping local knowledge on ciguatera and artisanal fisheries to inform systematic conservation planning. ICES Journal of Marine Science, 78(4): 1357–1371, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1357/6168601.

Arkhipkin:2022:HFE

[AWN22] A. I. Arkhipkin, A. G. Winter, and C. M. Nigmatullin. Heavy fishery exploitation does not affect sizes at recruitment and maturity of squid Illex argentinus and Doryteuthis gahi, in the Southwest Atlantic. ICES Journal of Marine Science, 79(1):182–192, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/182/6482877.

Bacheler:2024:RSB

[Bac24] Nathan M. Bacheler. A review and synthesis of the benefits, drawbacks, and considerations of using traps to survey fish and decapods. ICES Journal of Marine Science, 81(1):1–21, January 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/1/1/7502967.

Barnhill:2022:DSM

[Bar22] Kelsey Archer Barnhill. The deep sea and me. ICES Journal of Marine Science, 79(7):1996–2002, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/1996/6674267.

Barz:2023:ISP

[Bar23] Fanny Barz. Identifying social practices to inform fisheries management — the case of bycatch practices of marine mammals and seabirds of German gillnet fishers. ICES Journal of Marine Science, 80(3):458–468, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/458/6874767.

Breivik:2021:PAI

[BAS^{+}21] Olav Nikolai Breivik, Fredrik Aanes, Guldborg Søvik, Asgeir Aglen, Sigbjørn Mehl, and Espen Johnsen. Predicting abundance indices in areas without coverage with a latent spatio-temporal Gaussian model. ICES Journal of Marine Science, 78(6):2031–2042, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2031/6298533.

Beyan:2020:SSM

[BB20] Cigdem Beyan and Howard I. Browman. Setting the stage for the machine intelligence era in marine science. ICES Journal of Marine Science, 77(4):1267–1273, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1267/5859738.

Bryndum-Buchholz:2022:KHE

[BB22] Andrea Bryndum-Buchholz. Keeping up hope as an early career climate-impact scientist. ICES Journal of Marine Science, 79(9):2345–2350, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
9/2345/6754317.

Bertolini:2021:BMA

[BBB^{+}21] C. Bertolini, I. Bernardini, D. Brigolin, V. Matozzo, M. Milan, and R. Pastres. A bioenergetic model to address carbon sequestration potential of shellfish farming: example from Ruditapes philippinarum in the Venice lagoon. ICES Journal of Marine Science, 78(6):2082–2091, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/2082/6294922.

Barrett:2022:CCG

[BBB^{+}22] C. J. Barrett, J. Bensbai, M. K. Broadhurst, P. Bustamante, R. Clark, G. M. Cooke, A. Di Cosmo, C. Drerup, O. Escolar, F. A. Fernández-Álvarez, K. Ganias, K. C. Hall, R. T. Hanlon, J. Hernández-Urcera, Q. Q. H. Hua, T. Lacoue-Labarthe, J. Lewis, F. Lishchenko, V. Maselli, H. Moustahfid, R. Nakajima, C. E. O’Brien, L. Parkhouse, S. Pengelly, G. J. Pierce, J. G. Ramírez, J-P Robin, K. K. Sajikumar, G. Sasikumar, C. L. Smith, R. Villanueva, and D. T. H. Y´ ên. Cuttlefish conservation: a global review of methods to ameliorate unwanted fishing mortality and other anthropogenic threats to sustainability. ICES Journal of Marine Science, 79(10):2579–2596, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/10/2579/6844536.

Bertram:2023:BPG

[BBB^{+}23] A. Bertram, J. Bell, C. J. Brauer, A. Fowler, P. Hamer, J. Sandoval-Castillo, J. Stewart, M. Wellenreuther, and L. B. Beheregaray. Biogeographic provinces and genomically delineated stocks are congruent in snapper (Chrysophrys auratus) from southeastern Australia. ICES Journal of Marine Science, 80(5):1422–1430, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1422/7147045.

Biggs:2020:PDB

[BBE^{+}20] Tristan E. G. Biggs, Corina P. D. Brussaard, Claire Evans, Hugh J. Venables, and David W. Pond. Plasticity in dormancy behaviour of Calanoides acutus in Antarctic coastal waters. ICES Journal of Marine Science, 77(5):1738–1751, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1738/5815632.

Burgess:2023:CCS

[BBL^{+}23] M. G. Burgess, S. L. Becker, R. E. Langendorf, A. Fredston, and C. M. Brooks. Climate change scenarios in fisheries and aquatic conservation research. ICES Journal of Marine Science, 80(5):1163–1178, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1163/7103492.

Barrett:2022:ICD

[BBM^{+}22] C. J. Barrett, P. Barry, E. MacLeod, S. Stott, R. Vieira, and V. Laptikhovsky. The importance of cephalopods in the diet of fish on the northwest European shelf. ICES Journal of Marine Science, 79(5):1675–1686, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1675/6589473.

Bekkevold:2023:MSA

[BBP^{+}23] Dorte Bekkevold, Florian Berg, Patrick Polte, Valerio Bartolino, Henn Ojaveer, Henrik Mosegaard, Edward D. Farrell, Jelena Fedotova, Jakob Hemmer-Hansen, Bastian Huwer, Vanessa Trijoulet, Christoffer Moesgaard Albertsen, Angela P. Fuentes-Pardo, Tomas Gröhsler, Mats Pettersson, Teunis Jansen, Arild Folkvord, and Leif Andersson. Mixedstock analysis of Atlantic herring (Clupea harengus): a tool for identifying management units and complex migration dynamics. ICES Journal of Marine Science, 80(1):173–184, January 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/1/173/6969398.

Burliuk:2023:SAD

[BC23] Colleen M. M. Burliuk and John M. Casselman. Seasonal activity and depth distribution of resident yellow-phase American eels (Anguilla rostrata) in a large fluvial ecosystem. ICES Journal of Marine Science, 80(4):923–940, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/923/7049338. See correction [Ano23i].

Bensebaini:2022:IBD

[BCB^{+}22a] C. M. Bensebaini, G. Certain, N. Billet, A. Jadaud, S. Gourguet, T. Hattab, and J. M. Fromentin. Interactions between demersal fish body condition and density during the regime shift of the Gulf of Lions. ICES Journal of Marine Science, 79(6):1765–1776, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1765/6610013.

Berg:2022:DCE

[BCB^{+}22b] Hilde Sofie Fantoft Berg, Thomas L. Clegg, Geir Blom, Jeppe Kolding, Kotaro Ono, and Kjell Nedreaas. Discards of cod (Gadus morhua) in the Norwegian coastal fisheries: improving past and future estimates. ICES Journal of Marine Science, 79(5):1548–1560, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1548/6585625.

Belcher:2020:RMF

[BCBJ^{+}20] Anna Belcher, Kathryn Cook, Daniel Bondyale-Juez, Gabriele Stowasser, Sophie Fielding, Ryan A. Saunders, Daniel J. Mayor, and Geraint A. Tarling. Respiration of mesopelagic fish: a comparison of respiratory electron transport system (ETS) measurements and allometrically calculated rates in the Southern Ocean and Benguela Current. ICES Journal of Marine Science, 77(5):1672–1684, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1672/5805416.

Bentley:2024:CEM

[BCC^{+}24] Jacob W Bentley, David Chagaris, Marta Coll, Johanna J Heymans, Natalia Serpetti, Carl J Walters, and Villy Christensen. Calibrating ecosystem models to support ecosystem-based management of marine systems. ICES Journal of Marine Science, 81(2):260–275, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/2/260/7587678.

Brandao:2021:OSS

[BCP^{+}21] Manoela C. Brandão, Thierry Comtet, Patrick Pouline, Caroline Cailliau, Aline Blanchet-Aurigny, Marc Sourisseau, Raffaele Siano, Laurent Memery, Frédérique Viard, and Flávia Nunes. Oceanographic structure and seasonal variation contribute to high heterogeneity in mesozooplankton over small spatial scales. ICES Journal of Marine Science, 78(9):3288–3302, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3288/6318567.

Birchenough:2020:SSE

[BD20] Silvana N. R. Birchenough and Steven Degraer. Science in support of ecologically sound decommissioning strategies for offshore man-made structures: taking stock of current knowledge and considering future challenges. ICES Journal of Marine Science, 77(3):1075–1078, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1075/5820932.

Berger:2021:IDF

[BDB^{+}21] Aaron M. Berger, Jonathan J. Deroba, Katelyn M. Bosley, Daniel R. Goethel, Brian J. Langseth, Amy M. Schueller, and Dana H. Hanselman. Incoherent dimensionality in fisheries management: consequences of misaligned stock assessment and population boundaries. ICES Journal of Marine Science, 78(1):155–171, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
155/6043739.

Bouchard:2022:STV

[BDL^{+}22] Colin Bouchard, Hilaire Drouineau, Patrick Lambert, Olivier Boutron, and Delphine Nicolas. Spatio-temporal variations in glass eel recruitment at the entrance pathways of a Mediterranean delta. ICES Journal of Marine Science, 79(6):1874–1887, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
6/1874/6633760.

Bojesen:2024:AOD

[BDM24] Troels Arnfred Bojesen, Côme Denechaud, and Ketil Malde. Annotating otoliths with a deep generative model. ICES Journal of Marine Science, 81(1):55–65, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/1/55/7341839.

Beamish:2022:NSB

[Bea22] Richard Beamish. The need to see a bigger picture to understand the ups and downs of Pacific salmon abundances. ICES Journal of Marine Science, 79(4):1005–1014, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1005/6550345.

Beamish:2023:TWJ

[Bea23] Richard Beamish. I think that I will just sit here and wait. ICES Journal of Marine Science, 80(4):710–718, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/710/7085489.

Baldi:2023:BPU

[BFD^{+}23] Giulia Baldi, Giovanni Furii, Massimo Del Vecchio, Pasquale Salvemini, Carola Vallini, Valeria Angelini, Sauro Pari, Kamyla Lombardi Moraes, Chiara Profico, Vincenzo Olivieri, Dimitris Margaritoulis, Alan F. Rees, Anxhela Çurri, Sandra Hochscheid, Daniela Freggi, Bojan Lazar, Paolo Luschi, and Paolo Casale. Behavioural plasticity in the use of a neritic foraging area by loggerhead sea turtles: insights from 37 years of capture–mark–recapture in the Adriatic Sea (Mediterranean Sea). ICES Journal of Marine Science, 80(1):210–217, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/1/210/6956801.

Bauer:2020:SVB

[BFFC20] Robert Klaus Bauer, Fabien Forget, Jean-Marc Fromentin, and Manuela Capello. Surfacing and vertical behaviour of Atlantic bluefin tuna (Thunnus thynnus) in the Mediterranean Sea: implications for aerial surveys. ICES Journal of Marine Science, 77(5):1979–1991, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1979/5862935.

Brewin:2021:SLH

[BFJB21] Paul E. Brewin, Thomas J. Farrugia, Chris Jenkins, and Paul Brickle. Straddling the line: high potential impact on vulnerable marine ecosystems by bottom-set longline fishing in unregulated areas beyond national jurisdiction. ICES Journal of Marine Science, 78(6):2132–2145, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/2132/5875731.

Bourlat:2021:WFS

[BFW^{+}21] Sarah J. Bourlat, Ellika Faust, Håkan Wennhage, Andreas Wikström, Kristie Rigby, Maria Vigo, Paul Kraly, Erik Selander, and Carl André. Wrasse fishery on the Swedish West Coast: towards ecosystem-based management. ICES Journal of Marine Science, 78(4):1386–1397, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1386/6168020.

Bohaboy:2020:ATD

[BGH^{+}20] Erin Collings Bohaboy, Tristan L. Guttridge, Neil Hammerschlag, Maurits P. M. Van Zinnicq Bergmann, and William F. Patterson III. Application of three-dimensional acoustic telemetry to assess the effects of rapid recompression on reef fish discard mortality. ICES Journal of Marine Science, 77(1):83–96, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/77/1/83/5614227.

Bell:2022:CIH

[BGR^{+}22a] Richard J. Bell, Brian Grieve, Marta Ribera, John Manderson, and Dave Richardson. Climate-induced habitat changes in commercial fish stocks. ICES Journal of Marine Science, 79(8):2247–2264, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
8/2247/6710433.

Bull:2022:LSF

[BGR^{+}22b] C. D. Bull, S. D. Gregory, E. Rivot, T. F. Sheehan, D. Ensing, G. Woodward, and W. Crozier. The likely suspects framework: the need for a life cycle approach for managing Atlantic salmon (Salmo salar) stocks across multiple scales. ICES Journal of Marine Science, 79(5):1445– 1456, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1445/6604492.

Buyse:2022:OWF

[BHDD22] Jolien Buyse, Kris Hostens, Steven Degraer, and Annelies De Backer. Offshore wind farms affect the spatial distribution pattern of plaice Pleuronectes platessa at both the turbine and wind farm scale. ICES Journal of Marine Science, 79(6): 1777–1786, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/6/1777/6610008.

Bornt:2023:PGL

[BHdL^{+}23] Katrina Bornt, Jason How, Simon de Lestang, Kathryn Linge, Renae Hovey, and Tim Langlois. Plastic gear loss estimates from a major Australian pot fishery. ICES Journal of Marine Science, 80(1):158–172, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/158/6966540.

Blondin:2023:VBV

[BHL^{+}23] H. E. Blondin, D. E. Haulsee, R. Logan, M. Shivji, E. R. Hoffmayer, J. H. Walker, J. M. Dean, E. L. Hazen, and L. B. Crowder. Variability in billfish vertical distribution and fishing interactions driven by environmental conditions in the Eastern Tropical Pacific Ocean. ICES Journal of Marine Science, 80(6):1629–1642, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1629/7198278.

Blackwell:2020:CMF

[BHQF20] Robert E. Blackwell, Richard Harvey, Bastien Y. Queste, and Sophie Fielding. Colour maps for fisheries acoustic echograms. ICES Journal of Marine Science, 77(2):826–834, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/826/5684813.

Ben-Hasan:2020:ASB

[BHWC^{+}20] A. Ben-Hasan, C. Walters, V. Christensen, G. Munro, U. R. Sumaila, and A. Al-Baz. Age-structured bioeconomic model for strategic interaction: an application to pomfret stock in the Arabian/Persian Gulf. ICES Journal of Marine Science, 77(5):1787–1795, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1787/5825339.

Birchenough:2023:KBS

[Bir23] S. N. R. Birchenough. Knowledge-based science in support of the blue growth ambition for small island developing states. ICES Journal of Marine Science, 80(8):2166–2170, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/8/2166/7281163.

Bi:2020:LTC

[BJBB20] Rujia Bi, Yan Jiao, Haakon Bakka, and Joan A. Browder. Long-term climate ocean oscillations inform seabird bycatch from pelagic longline fishery. ICES Journal of Marine Science, 77(2):668–679, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/2/668/5704436.

Berg:2021:GSA

[BJG^{+}21] Paul R. Berg, Per E. Jorde, Kevin A. Glover, Geir Dahle, John B. Taggart, Knut Korsbrekke, Gjert E. Dingsør, Jon E. Skjæraasen, Peter J. Wright, Steven X. Cadrin, Halvor Knutsen, and Jon-Ivar Westgaard. Genetic structuring in Atlantic haddock contrasts with current management regimes. ICES Journal of Marine Science, 78(1):1–13, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/1/1/6032773.

Bjordal:2021:FFS

[Bjo21] Åsmund Bjordal. Fisher or fisheries scientist? ICES Journal of Marine Science, 78(3):848–854, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
848/6145327.

Bernard:2020:GBM

[BJPP^{+}20] Andrea M. Bernard, Matthew W. Johnston, Rocío Pérez-Portela, Marjorie F. Oleksiak, Felicia C. Coleman, and Mahmood S. Shivji. Genetic and biophysical modelling evidence of generational connectivity in the intensively exploited, Western North Atlantic red grouper (Epinephelus morio). ICES Journal of Marine Science, 77(1):359–370, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/1/359/5614358.

Ban:2020:ICR

[BKF^{+}20] Natalie C. Ban, Katrina Kushneryk, Jenna Falk, Alanna Vachon, and Laurel Sleigh. Improving compliance of recreational fishers with rockfish conservation areas: community– academic partnership to achieve and evaluate conservation. ICES Journal of Marine Science, 77(6):2308–2318, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2308/5543080.

Blain:2021:HAC

[BKR^{+}21] Caitlin O. Blain, Sara Kulins, Craig A. Radford, Mary A. Sewell, and Nick T. Shears. Heterogeneity around CO_{2} vents obscures the effects of ocean acidification on shallow reef communities. ICES Journal of Marine Science, 78(9):3162– 3175, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/78/9/3162/6374671.

Bradbury:2021:RWG

[BLM^{+}21] I. R. Bradbury, S. J. Lehnert, A. Messmer, S. J. Duffy, E. Verspoor, T. Kess, J. Gilbey, V. Wennevik, M. Robertson, G. Chaput, T. Sheehan, P. Bentzen, J. B. Dempson, and D. Reddin. Range-wide genetic assignment confirms longdistance oceanic migration in Atlantic salmon over half a century. ICES Journal of Marine Science, 78(4):1434–1443, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1434/6169450.

Badger:2023:STS

[BLT23] J. J. Badger, S. I. Large, and J. T. Thorson. Spatiotemporal species distribution models reveal dynamic indicators for ecosystem-based fisheries management. ICES Journal of Marine Science, 80(7):1949–1962, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1949/7241088.

Bergman:2020:FIR

[BM20] Magda J. N. Bergman and Erik H. Meesters. First indications for reduced mortality of non-target invertebrate benthic megafauna after pulse beam trawling. ICES Journal of Marine Science, 77(2):846–857, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
846/5709124.

Braccini:2020:PAS

[BMB20] Matias Braccini, Brett Molony, and Nick Blay. Patterns in abundance and size of sharks in northwestern Australia: cause for optimism. ICES Journal of Marine Science, 77(1): 72–82, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/1/72/5584949.

Browne:2021:MDO

[BMM^{+}21] D. Browne, C. Minto, M. McHugh, S. Murphy, M. Oliver, and R. Cosgrove. Match of the day: optimized experimental design in alternate-haul gear trials. ICES Journal of Marine Science, 78(6):1988–1998, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
1988/6295279.

Ban:2022:DRF

[BMM^{+}22] Natalie C. Ban, Chessi Miltner, Charlotte Matthews, Madeleine Ankenman, Sarah Stelte, Dana Haggarty, Hailey L. Davies, Paul A. Venturelli, and Francis Juanes. Decrease in recreational fisher compliance during the COVID-19 pandemic: the case of Rockfish Conservation Areas. ICES Journal of Marine Science, 79(8):2277–2285, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2277/6711582.

Britton:2021:CCA

[BMN^{+}21] Damon Britton, Craig N. Mundy, Fanny Noisette, Christina M. McGraw, and Catriona L. Hurd. Crustose coralline algae display sensitivity to near future global ocean change scenarios. ICES Journal of Marine Science, 78(10):3748–3756, December 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/10/3748/6425786.

Bouch:2021:CPD

[BMR21] Paul Bouch, Cóilín Minto, and Dave G. Reid. Comparative performance of data-poor CMSY and data-moderate SPiCT stock assessment methods when applied to data-rich, real-world stocks. ICES Journal of Marine Science, 78(1): 264–276, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/1/264/6032767.

Boyra:2023:CBO

[BMU^{+}23] Guillermo Boyra, Udane Martínez, Jon Uranga, Gala Moreno, and Héctor Peña. Correction of beam overlapinduced athwart distortion in multibeam sonars. ICES Journal of Marine Science, 80(1):197–209, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/197/6964912.

Breivik:2021:PVR

[BNB21] Olav Nikolai Breivik, Anders Nielsen, and Casper W. Berg. Prediction–variance relation in a state-space fish stock assessment model. ICES Journal of Marine Science, 78(10):3650–3657, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
10/3650/6412908.

Bell:2021:IME

[BNV^{+}21] James B. Bell, Nguy˜ ên Văn Nguyên, Hà Vũ Vi{ê \atop .}t, Minh Hoàng Nguy˜ ên, Hùng Thanh Bùi, Tuân Văn Trâng, Paul McIlwaine, Andrew Kenny, and Bát Kh´ ăc Nguy˜ ên. Identifying marine ecological production units in Vietnam. ICES Journal of Marine Science, 78(4):1241–1252, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/4/1241/6169453.

Bode:2021:MCT

[BOLPHL21] Antonio Bode, M. Pilar Olivar, Cristina López-Pérez, and Santiago Hernández-León. The microbial contribution to the trophic position of stomiiform fishes. ICES Journal of Marine Science, 78(9):3245–3253, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3245/6377542.

Borges:2021:UIE

[Bor21] Lisa Borges. The unintended impact of the European discard ban. ICES Journal of Marine Science, 78(1):134–141, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/1/134/6026103.

Berg:2021:CGP

[BØS^{+}21] Florian Berg, Hedda D. Østgaard, Aril Slotte, Leif Andersson, and Arild Folkvord. A combination of genetic and phenotypic characterization of spring- and autumn-spawning herring suggests gene flow between populations. ICES Journal of Marine Science, 78(2):694–703, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/2/
694/5817457.

Burgess:2020:WUF

[BP20] Matthew G. Burgess and Michael J. Plank. What unmanaged fishing patterns reveal about optimal management: applied to the balanced harvesting debate. ICES Journal of Marine Science, 77(3):901–910, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
901/5740017.

Bordeleau:2020:STT

[BPC^{+}20] X. Bordeleau, S. A. Pardo, G. Chaput, J. April, B. Dempson, M. Robertson, A. Levy, R. Jones, J. A. Hutchings, F. G. Whoriskey, and G. T. Crossin. Spatio-temporal trends in the importance of iteroparity across Atlantic salmon populations of the northwest Atlantic. ICES Journal of Marine Science, 77(1):326–344, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/326/5588516.

Bekkevold:2021:GSI

[BPC^{+}21] Dorte Bekkevold, Adam Piper, Ronald Campbell, Philip Rippon, Ros M. Wright, Charles Crundwell, Klaus Wysujack, Jamie R. Stevens, R. Andrew King, Kim Aarestrup, and Alistair Maltby. Genetic stock identification of sea trout (Salmo trutta L.) along the British North Sea Coast shows prevalent long-distance migration. ICES Journal of Marine Science, 78(3):952–966, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/3/952/6104103.

Banda:2024:GTE

[BPC^{+}24] Shamiso Banda, Pierre Pistorius, Julien Collet, Alexandre Corbeau, Henri Weimerskirch, Adrien Pajot, Danielle Z Keys, and Florian Orgeret. Gauging the threat: exposure and attraction of sooty albatrosses and white-chinned petrels to fisheries activities in the Southern Indian Ocean. ICES Journal of Marine Science, 81(1):75–85, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/1/75/7379035.

Bucklin:2021:NIB

[BPKM21] Ann Bucklin, Katja T. C. A. Peijnenburg, Ksenia Kosobokova, and Ryuji J. Machida. New insights into biodiversity, biogeography, ecology, and evolution of marine zooplankton based on molecular approaches. ICES Journal of Marine Science, 78(9):3281–3287, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/

9/3281/6403476. See erratum [BPKM22].

Bucklin:2022:ENI

[BPKM22] Ann Bucklin, Katja T. C. A. Peijnenburg, Ksenia Kosobokova, and Ryuji J. Machida. Erratum to: New insights into biodiversity, biogeography, ecology, and evolution of marine zooplankton based on molecular approaches. ICES Journal of Marine Science, 79(1):245, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
1/245/6453067. See [BPKM21].

Burns:2021:RRB

[BPP^{+}21] Corinne M. Burns, Pierre Pepin, Stéphane Plourde, Guillaume Veillet, Pascal Sirois, and Dominique Robert. Revealing the relationship between feeding and growth of larval redfish (Sebastes sp.) in the Gulf of St. Lawrence. ICES Journal of Marine Science, 78(10):3757–3766, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3757/6423499.

Bueno-Pardo:2020:TDM

[BPPC^{+}20] Juan Bueno-Pardo, Graham J. Pierce, Edna Cabecinha, Catarina Grilo, Jorge Assis, Vasilis Valavanis, Cristina Pita, Jesús Dubert, Francisco Leitão, and Henrique Queiroga. Trends and drivers of marine fish landings in Portugal since its entrance in the European Union. ICES Journal of Marine Science, 77(3):988–1001, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/3/988/5733766.

Bueno-Pardo:2020:IBI

[BPPKH20] Juan Bueno-Pardo, Pierre Petitgas, Susan Kay, and Martin Huret. Integration of bioenergetics in an individualbased model to hindcast anchovy dynamics in the Bay of Biscay. ICES Journal of Marine Science, 77(2):655–667, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/655/5671744.

Boe:2020:ILH

[BPR^{+}20] Kristin Bøe, Michael Power, Martha J. Robertson, Corey J. Morris, J. Brian Dempson, Christopher C. Parrish, and Ian A. Fleming. Influence of life-history-dependent migration strategies on Atlantic salmon diets. ICES Journal of Marine Science, 77(1):345–358, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/1/
345/5589084.

Blamey:2023:ARF

[BPR^{+}23] Laura K. Blamey, Éva E. Plagányi, Julie Robins, Rob Kenyon, Roy A. Deng, Justin Hughes, and Shaun Kim. Altering river flow impacts estuarine species and catches: lessons from giant mud crabs. ICES Journal of Marine Science, 80(9):2295–2312, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
9/2295/7321881.

Boschetti:2020:ITM

[BPV^{+}20] Fabio Boschetti, Karine Prunera, Mathew A. Vanderklift, Damian P. Thomson, Russell C. Babcock, Christopher Doropoulos, Anna Cresswell, and Hector Lozano-Montes. Information-theoretic measures of ecosystem change, sustainability, and resilience. ICES Journal of Marine Science, 77(4):1532–1544, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1532/5520595.

Bucklin:2021:PCE

[BQBB^{+}21] Ann Bucklin, Jennifer M. Questel, Leocadio Blanco-Bercial, Alexandra Frenzel, Sara B. Smolenack, and Peter H. Wiebe. Population connectivity of the euphausiid, Stylocheiron elongatum, in the Gulf Stream (NW Atlantic Ocean) in relation to COI barcode diversity of Stylocheiron species.

ICES Journal of Marine Science, 78(9):3464–3476, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3464/6358061.

Bearzi:2021:SBC

[BR21] Giovanni Bearzi and Randall R. Reeves. Shifting baselines of cetacean conservation in Europe. ICES Journal of Marine Science, 78(7):2337–2341, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2337/6324790.

Bearzi:2022:MMF

[BR22] Giovanni Bearzi and Randall R. Reeves. Marine mammals foraging around fishing gear or preying upon fishing catch and bait: it may not be “depredation”. ICES Journal of Marine Science, 79(8):2178–2183, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2178/6724454.

Brander:2020:RGB

[Bra20] Keith Brander. Reduced growth in Baltic Sea cod may be due to mild hypoxia. ICES Journal of Marine Science, 77(5):2003–2005, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
2003/5818946.

Brander:2022:SHG

[Bra22] Keith Brander. Support for the hypothesis that growth of eastern Baltic cod is affected by mild hypoxia. a comment on Svedäng et al. (2022). ICES Journal of Marine Science, 79(7):2155–2156, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/

7/2155/6571568. See [SSV^{+}22b] and reply [SSV^{+}22a].

Butterworth:2022:CSA

[BRG22] D. S. Butterworth and A. Ross-Gillespie. Comment on “South Africa’s experimental fisheries closures and recovery of the endangered African penguin” by Sydeman et al.

(2021). ICES Journal of Marine Science, 79(6):1965–1971, August 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/6/1965/6640476. See [SHP^{+}21] and

response [SHP^{+}22].

Boute:2023:IIM

[BRvL^{+}23] P. G. Boute, A. D. Rijnsdorp, J. L. van Leeuwen, R. P. M. Pieters, and M. J. Lankheet. Internal injuries in marine fishes caught in beam trawls using electrical versus mechanical stimulations. ICES Journal of Marine Science, 80(5): 1367–1381, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/5/1367/7186190.

Beaudry-Sylvestre:2022:CEE

[BSBH22] Manuelle Beaudry-Sylvestre, Hugues P. Benoît, and Jeffrey A. Hutchings. Combined effects of exploitation and environmental change on life history: a comparative analysis on Atlantic herring. ICES Journal of Marine Science, 79(10):2728–2740, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
10/2728/6852668.

Brinkhof:2022:MSS

[BSH^{+}22] Jesse Brinkhof, Manu Sistiaga, Bent Herrmann, Eduardo Grimaldo, and Roger B. Larsen. Managing size selectivity: the relevance of compulsory and alternative selection devices in the Northeast Atlantic bottom trawl fishery. ICES Journal of Marine Science, 79(9):2399–2412, November 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/9/2399/6761001.

Bjornsson:2022:NME

[BSW22] Björn Björnsson, Jón Sólmundsson, and Pamela J. Woods. Natural mortality in exploited fish stocks: annual variation estimated with data from trawl surveys. ICES Journal of Marine Science, 79(5):1569–1582, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1569/6590896.

Baker:2023:DVM

[BSW^{+}23] Matthew R. Baker, T. S. Smeltz, Kresimir Williams, Casey Greufe, Megan Ewing, Jonathan Chapman, Julia Glassy, Eva Hasegawa, Kathleen P. Cieri, Sofia Matson, and Rick Towler. Diel vertical migration in Pacific sand lance (Ammodytes personatus) — a pelagic forage fish associated with benthic substrates. ICES Journal of Marine Science, 80(6): 1758–1772, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/6/1758/7217009.

Briton:2021:FJP

[BTM^{+}21] Florence Briton, Olivier Thébaud, Claire Macher, Caleb Gardner, and Lorne Richard Little. Flexibility of joint production in mixed fisheries and implications for management. ICES Journal of Marine Science, 78(5):1599–1613, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/5/1599/6225992.

Barnhill:2023:SST

[BVS^{+}23] Kelsey Archer Barnhill, Beatriz Vinha, Alycia J. Smith, Daniëlle S. W. de Jonge, Daniela Y. Gaurisas, Roger Mocholí Segura, Pedro Madureira, Mónica Albuquerque, Veerle A. I. Huvenne, Covadonga Orejas, and Vikki Gunn. Shipto-shore training for active deep-sea capacity development. ICES Journal of Marine Science, 80(6):1619–1628, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1619/7187734.

Barbut:2020:PFR

[BVV^{+}20] Léo Barbut, Berthe Vastenhoud, Laurence Vigin, Steven Degraer, Filip A. M. Volckaert, and Geneviève Lacroix. The proportion of flatfish recruitment in the North Sea potentially affected by offshore windfarms. ICES Journal of Marine Science, 77(3):1227–1237, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1227/5445401.

Blomquist:2021:SIE

[BW21] Johan Blomquist and Staffan Waldo. Seal interactions and exits from fisheries: insights from the Baltic Sea cod fishery. ICES Journal of Marine Science, 78(8):2958–2966, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2958/6363516.

Barcelo:2021:PTI

[BWBH21] Caren Barceló, J. Wilson White, Louis W. Botsford, and Alan Hastings. Projecting the timescale of initial increase in fishery yield after implementation of marine protected areas. ICES Journal of Marine Science, 78(5):1860–1871, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/5/1860/6282521.

Brautaset:2020:ACM

[BWJ^{+}20] Olav Brautaset, Anders Ueland Waldeland, Espen Johnsen, Ketil Malde, Line Eikvil, Arnt-Børre Salberg, and Nils Olav Handegard. Acoustic classification in multifrequency echosounder data using deep convolutional neural networks. ICES Journal of Marine Science, 77(4):1391–1400, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/4/1391/5712978.

Barroso:2023:AML

[BXF23] V. R. Barroso, F. C. Xavier, and C. E. L. Ferreira. Applications of machine learning to identify and characterize the sounds produced by fish. ICES Journal of Marine Science, 80(7):1854–1867, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
7/1854/7246570.

Coleman:2021:ANN

[CAE^{+}21] Matthew T. Coleman, Anne-Lisbeth Agnalt, Jack Emmerson, Martial Laurens, Joanne S. Porter, and Michael C. Bell. From the Adriatic to Northern Norway — geographic differences in moult increment and moult probability of the European lobster (Homarus gammarus), across the natural range. ICES Journal of Marine Science, 78(2):611–620, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/2/611/5957471.

Chemshirova:2023:ISG

[CASM23] I. Chemshirova, A. Arkhipkin, P. W. Shaw, and N. J. McKeown. Integrated statolith and genomic analysis reveals high connectivity in the nektonic squid Illex argentinus: implications for management of an international cephalopod fishery. ICES Journal of Marine Science, 80(7):1976–1990, September 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/7/1976/7241617.

Christie:2022:MHF

[CATM22] Hartvig Christie, Guri Sogn Andersen, Lise Ann Tveiten, and Frithjof Emil Moy. Macrophytes as habitat for fish. ICES Journal of Marine Science, 79(2):435–444, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/435/6521671.

Cole:2021:ETA

[CBB21] Alexandra K. Cole, Sean W. Brillant, and Stephanie A. Boudreau. Effects of time-area closures on the distribution of snow crab fishing effort with respect to entanglement threat to North Atlantic right whales. ICES Journal of Marine Science, 78(6):2109–2119, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2109/6294923.

Collie:2021:HSC

[CBCM21] Jeremy S. Collie, Richard J. Bell, Samuel B. Collie, and Cóilín Minto. Harvest strategies for climate-resilient fisheries. ICES Journal of Marine Science, 78(8):2774–2783, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2774/6355111.

Catucci:2023:ANC

[CBFS23] Elena Catucci, Elvira Buonocore, Pier Paolo Franzese, and Michele Scardi. Assessing the natural capital value of Posidonia oceanica meadows in the Italian seas by integrating Habitat Suitability and Environmental Accounting Models. ICES Journal of Marine Science, 80(4):739–750, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/739/6549753.

Cantor:2020:HIS

[CBT^{+}20] Mauricio Cantor, André Silva Barreto, Renata M. Taufer, Bruno Giffoni, Pedro V. Castilho, Andrea Maranho, Carla Beatriz, Christiane Kolesnikovas, Daniela Godoy, Daniel W. Rogério, Jeferson L. Dick, Karina R. Groch, Liana Rosa, Marta J. Cremer, Pâmela E. Cattani, Rodrigo R. Valle, and Camila Domit. High incidence of sea turtle stranding in the southwestern Atlantic Ocean. ICES Journal of Marine Science, 77(5):1864–1878, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1864/5836006.

Chary:2021:SFW

[CCC^{+}21] Killian Chary, Myriam D. Callier, Denis Covès, Joël Aubin, Julien Simon, and Annie Fiandrino. Scenarios of fish waste deposition at the sub-lagoon scale: a modelling approach for aquaculture zoning and site selection. ICES Journal of Marine Science, 78(3):922–939, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
922/6066594.

Cronin:2023:HSK

[CCH^{+}23] Melissa R. Cronin, Donald A. Croll, Martin A. Hall, Nerea Lezama-Ochoa, Jon Lopez, Hilario Murua, Jefferson Murua, Victor Restrepo, Stefany Rojas-Perea, Joshua D. Stewart, Jennifer L. Waldo, and Gala Moreno. Harnessing stakeholder knowledge for the collaborative development of Mobulid bycatch mitigation strategies in tuna fisheries. ICES Journal of Marine Science, 80(3):620–634, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/620/6604949.

Concannon:2021:TDE

[CCJ^{+}21] Callie A. Concannon, Emma L. Cross, Lucas F. Jones, Christopher S. Murray, Catherine M. Matassa, Richard S.

McBride, and Hannes Baumann. Temperature-dependent effects on fecundity in a serial broadcast spawning fish after whole-life high CO_{2} exposure. ICES Journal of Marine Science, 78(10):3724–3734, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
10/3724/6419777.

Clements:2021:CIL

[CCM^{+}21] Jeff C. Clements, Claire E. Carver, Martin A. Mallet, Luc A. Comeau, and André L. Mallet. CO_{2}-induced low pH in an eastern oyster (Crassostrea virginica) hatchery positively affects reproductive development and larval survival but negatively affects larval shape and size, with no intergenerational linkages. ICES Journal of Marine Science, 78(1):349–359, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/1/349/5862572.

Castro:2024:ECB

[CCR^{+}24] Jose Castro, Lucia Cañás, Jose Rodríguez, Graham John Pierce, and Camilo Saavedra. An evaluation of the cetacean bycatch monitoring programme on board the Spanish set gillnet and bottom pair trawl fleets in the Bay of Biscay. ICES Journal of Marine Science, 81(2):307–316, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/2/307/7471698.

Chiang:2024:THR

[CCS^{+}24] Chun-I Chiang, Ming-Tsung Chung, Jen-Chieh Shiao, Pei-Ling Wang, and Chia-Hui Wang. Thermal histories reveal spatiotemporal distribution and population overlapping of Sepioteuthis lessoniana. ICES Journal of Marine Science, 81(1):162–177, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/81/1/162/7460293.

Claireaux:2021:EDA

[CdSSO^{+}21] Marion Claireaux, Thassya C. dos Santos Schmidt, Esben Moland Olsen, Aril Slotte, Øystein Varpe, Mikko Heino, and Katja Enberg. Eight decades of adaptive changes in herring reproductive investment: the joint effect of environment and exploitation. ICES Journal of Marine Science, 78(2): 631–639, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/2/631/5901668.

Couture:2021:SAC

[CFB^{+}21] Jessica L. Couture, Halley E. Froehlich, Bela H. Buck, Keith R. Jeffery, Gesche Krause, James A. Morris, Jr., Montse Pérez, Grant D. Stentiford, Harri Vehviläinen, and Benjamin S. Halpern. Scenario analysis can guide aquaculture planning to meet sustainable future production goals. ICES Journal of Marine Science, 78(3):821–831, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/821/6133890.

Clay:2023:IIE

[CFB^{+}23] Patricia M. Clay, Johanna Ferretti, Jennifer L. Bailey, Leyre Goti, Dorothy J. Dankel, Marina Santurtun, Jessica Fuller, Sebastian Linke, Jörn Schmidt, Kåre Nolde Nielsen, David Goldsborough, Rolf Groeneveld, Ana Rita Fraga, Isa Elegbede, and Christine Röckmann. Implementation of integrated ecosystem assessments in the International Council for the Exploration of the Sea — conceptualizations, practice, and progress. ICES Journal of Marine Science, 80(5): 1516–1528, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/5/1516/7176243.

Clinton:2021:IJM

[CFMB21] Morag Clinton, David E. K. Ferrier, Samuel A. M. Martin, and Andrew S. Brierley. Impacts of jellyfish on marine cage aquaculture: an overview of existing knowledge and the challenges to finfish health. ICES Journal of Marine Science, 78(5):1557–1573, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1557/6209450.

Clegg:2022:SAA

[CFO^{+}22] Thomas L. Clegg, Edvin Fuglebakk, Kotaro Ono, Jon Helge Vølstad, and Kjell Nedreaas. A simulation approach to assessing bias in a fisheries self-sampling programme. ICES Journal of Marine Science, 79(1):76–87, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/76/6453069.

Cronin-Fine:2020:TNB

[CFP20] Lee Cronin-Fine and André E. Punt. There is no best method for constructing size-transition matrices for sizestructured stock assessments. ICES Journal of Marine Science, 77(1):136–147, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/136/5637824.

Coleman:2023:SAG

[CGH^{+}23] Matthew T. Coleman, Matthew Garratt, Natalie Hold, Isobel S. M. Bloor, Stuart R. Jenkins, Joanne S. Porter, Oliver Tully, and Michael C. Bell. A standardized assessment of geographic variation in size at maturity of European lobster (Homarus gammarus L.) in the North East Atlantic. ICES Journal of Marine Science, 80(4):911–922, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/911/7049336.

Chambers:2021:BMF

[Cha21] Mark S. Chambers. Benefits to migratory fish populations of entrainment and its potential role in fisheries collapse. ICES Journal of Marine Science, 78(1):36–44, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/1/36/6026104.

Champion:2022:CCA

[CHZC22] Curtis Champion, Alistair J. Hobday, Xuebin Zhang, and Melinda A. Coleman. Climate change alters the temporal persistence of coastal-pelagic fishes off eastern Australia. ICES Journal of Marine Science, 79(4):1083–1097, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1083/6532052.

Choi:2021:SST

[CKH^{+}21] Changkyu Choi, Michael Kampffmeyer, Nils Olav Handegard, Arnt-Børre Salberg, Olav Brautaset, Line Eikvil, and Robert Jenssen. Semi-supervised target classification in multi-frequency echosounder data. ICES Journal of Marine Science, 78(7):2615–2627, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2615/6348794.

Christiansen:2021:FBM

[CKR^{+}21] Svenja Christiansen, Thor A. Klevjer, Anders Røstad, Dag L. Aksnes, and Stein Kaartvedt. Flexible behaviour in a mesopelagic fish (Maurolicus muelleri). ICES Journal of Marine Science, 78(5):1623–1635, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1623/6225993.

Caiger:2021:GRM

[CLL21] Paul E. Caiger, Lyndsey S. Lefebve, and Joel K. Llopiz. Growth and reproduction in mesopelagic fishes: a literature synthesis. ICES Journal of Marine Science, 78(3):765– 781, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/765/6103223.

Cook:2022:IRB

[CLMB22] Kathryn B. Cook, Jean-Pierre Lacaze, Margarita Machairopoulou, and Eileen Bresnan. Investigations into the relationship between domoic acid and copepods in Scottish waters. ICES Journal of Marine Science, 79(3):963–973, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/963/6512392.

Curtis:2022:PPR

[CLS^{+}22] K. Alexandra Curtis, Mark S. Lowry, Joelle M. Sweeney, Anthony J. Orr, and James T. Harvey. Predicting prey recovery from scats of California sea lions (Zalophus californianus) for novel prey species and sizes. ICES Journal of Marine Science, 79(1):193–203, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/193/6484961.

Correa:2021:STV

[CMCF21] Giancarlo M. Correa, Carey R. McGilliard, Lorenzo Ciannelli, and Claudio Fuentes. Spatial and temporal variability in somatic growth in fisheries stock assessment models: evaluating the consequences of misspecification. ICES Journal of Marine Science, 78(5):1900–1908, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1900/6287349.

Caze:2023:CDM

[CMD^{+}23a] Cosma Cazé, Camille Mazé, Anatole Danto, Hanieh Saeedi, Dan Lear, Saara Suominen, Joape Ginigini, Gilianne Brodie, Isoa Korovulavula, and Isabel Sousa Pinto. Co-designing marine science beyond good intentions: support stakeholders’ empowerment in transformative pathways. ICES Journal of Marine Science, 80(2):374–377, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/2/374/6698551. See correction [Ano23c].

Cote:2023:CEM

[CMD^{+}23b] D. Cote, B. McClenaghan, J. Desforges, N. A. Fahner, M. Hajibabaei, J. Chawarski, S. Roul, G. Singer, C. Aubry, and M. Geoffroy. Comparing eDNA metabarcoding and conventional pelagic netting to inform biodiversity monitoring in deep ocean environments. ICES Journal of Marine Science, 80(10):2545–2562, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/10/2545/7394702.

Clare:2023:AAR

[CMG^{+}23] D. S. Clare, R. Martinez, F. Goodsir, M. A. Spence, E. Tyllianakis, S. Ware, and M. Schratzberger. An approach for assessing and ranking fisheries management scenarios in spatially delimited marine areas. ICES Journal of Marine Science, 80(5):1291–1302, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1291/7107387.

Calderwood:2023:EIS

[CMH^{+}23] Julia Calderwood, C. Tara Marshall, Karl Haflinger, Joanna Alfaro-Shigueto, Jeffrey C. Mangel, and David G. Reid. An evaluation of information sharing schemes to identify what motivates fishers to share catch information. ICES Journal of Marine Science, 80(3):556–577, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/556/6482872.

Chong:2020:PED

[CMR^{+}20] Lisa Chong, Tobias K. Mildenberger, Merrill B. Rudd, Marc H. Taylor, Jason M. Cope, Trevor A. Branch, Matthias Wolff, and Moritz Stäbler. Performance evaluation of datalimited, length-based stock assessment methods. ICES Journal of Marine Science, 77(1):97–108, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/1/
97/5621359.

Correa:2023:MTV

[CMS^{+}23] Giancarlo M. Correa, Cole C. Monnahan, Jane Y. Sullivan, James T. Thorson, and André E. Punt. Modelling time-varying growth in state-space stock assessments. ICES Journal of Marine Science, 80(7):2036–2049, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/2036/7249995.

Cocking:2022:ADB

[CNWW22] Jennifer Cocking, Bhavani E. Narayanaswamy, Claire M. Waluda, and Benjamin J. Williamson. Aerial detection of beached marine plastic using a novel, hyperspectral shortwave infrared (SWIR) camera. ICES Journal of Marine Science, 79(3):648–660, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/3/648/6524754.

Cavallo:2020:LSP

[CP20] Alessandro Cavallo and Lloyd S. Peck. Lipid storage patterns in marine copepods: environmental, ecological, and intrinsic drivers. ICES Journal of Marine Science, 77(5):1589–1601, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1589/5827681.

Chamberlin:2021:PDP

[CPD^{+}21] Joshua Chamberlin, Eleni Petrou, Will Duguid, Russel Barsh, Francis Juanes, Jessica Qualley, and Lorenz Hauser. Phenological diversity of a prey species supports life-stage specific foraging opportunity for a mobile consumer. ICES Journal of Marine Science, 78(9):3089–3100, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3089/6374190.

Clare:2021:TDR

[CRB21] David S. Clare, Leonie A. Robinson, and Silvana N. R. Birchenough. A temperature-dependent relationship between benthic invertebrate biomass and trawling pressure. ICES Journal of Marine Science, 78(1):82–88, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/1/82/6032766.

Coates:2022:MMR

[CRD^{+}22] Andrew Coates, Nick Robinson, Tim Dempster, Francisca Samsing, Ingrid Johnsen, and Ben L. Phillips. A metapopulation model reveals connectivity-driven hotspots in treatment resistance evolution in a marine parasite. ICES Journal of Marine Science, 79(10):2682–2696, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/10/2682/6854446.

Cardona:2023:EFT

[CRG^{+}23] Luis Cardona, Olga Reñones, Adam Gouraguine, Fabiana Saporiti, Asunción Borrell, Alex Aguilar, and Joan Moranta. Effects of fishing on the trophic structure of carnivorous fish assemblages from shallow rocky bottoms of the Mediterranean Sea and the temperate Atlantic Ocean. ICES Journal of Marine Science, 80(4):751–765, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/751/6962286.

Campbell:2020:PWS

[CRJ20] R. W. Campbell, P. L. Roberts, and J. Jaffe. The Prince William Sound Plankton Camera: a profiling in situ observatory of plankton and particulates. ICES Journal of Marine Science, 77(4):1440–1455, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1440/5811106.

Cheng:2023:SFD

[CRLC23] M. L. H. Cheng, C. J. Rodgveller, J. A. Langan, and C. J. Cunningham. Standardizing fishery-dependent catchrate information across gears and data collection programs for Alaska sablefish (Anoplopoma fimbria). ICES Journal of Marine Science, 80(4):1028–1042, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1028/7074460. See correction [Ano23j].

Christensen:2022:YRG

[CRR^{+}22] Helle Torp Christensen, Frank Rigét, Anja Retzel, Erik Haar Nielsen, Einar Eg Nielsen, and Rasmus Berg Hedeholm. Year-round genetic monitoring of mixed-stock fishery of Atlantic cod (Gadus morhua); implications for management. ICES Journal of Marine Science, 79(5):1515–1529, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1515/6581599. See correction [Ano22d].

Collins:2021:SGN

[CRV^{+}21] Jane Eva Collins, Muriel Rabone, Thomas Vanagt, Diva J. Amon, Judith Gobin, and Isabelle Huys. Strengthening the global network for sharing of marine biological collections: recommendations for a new agreement for biodiversity beyond national jurisdiction. ICES Journal of Marine Science, 78(1):305–314, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/305/6032765.

Cypher:2023:DEA

[CSC^{+}23] Alysha D. Cypher, Hank Statscewich, Robert Campbell, Seth L. Danielson, John Eiler, and Mary Anne Bishop. Detection efficiency of an autonomous underwater glider carrying an integrated acoustic receiver for acoustically tagged Pacific herring. ICES Journal of Marine Science, 80(2):329– 341, March 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/2/329/7003480.

Chladek:2021:DTF

[CSH^{+}21] Jérôme Chladek, Daniel Stepputtis, Andreas Hermann, Peter Ljungberg, Paco Rodriguez-Tress, Juan Santos, and Jon Christian Svendsen. Development and testing of fishretention devices for pots: transparent triggers significantly increase catch efficiency for Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 78(1):199–219, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/1/199/6032366.

Chaput:2022:QUE

[CSM^{+}22] Romain Chaput, Pierre Sochala, Philippe Miron, Vassiliki H. Kourafalou, and Mohamed Iskandarani. Quantitative uncertainty estimation in biophysical models of fish larval connectivity in the Florida Keys. ICES Journal of Marine Science, 79(3):609–632, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/3/609/6541347.

Cronin:2023:DOA

[CSM^{+}23] M. F. Cronin, S. Swart, C. A. Marandino, C. Anderson, P. Browne, S. Chen, W. R. Joubert, U. Schuster, R. Venkatesan, C. I. Addey, O. Alves, F. Ardhuin, S. Battle, M. A. Bourassa, Z. Chen, M. Chory, C. Clayson, R. B. de Souza, M. du Plessis, M. Edmondson, J. B. Edson, S. T. Gille, J. Hermes, V. Hormann, S. A. Josey, M. Kurz, T. Lee, F. Maicu, E. H. Moustahfid, S-A Nicholson, E. S. Nyadjro, J. Palter, R. G. Patterson, S. G. Penny, L. P. Pezzi, N. Pinardi, J. E. J. Reeves Eyre, N. Rome, A. C. Subramanian, C. Stienbarger, T. Steinhoff, A. J. Sutton, H. Tomita, S. M. Wills, C. Wilson, and L. Yu. Developing an Observing Air–Sea Interactions Strategy (OASIS) for the global ocean. ICES Journal of Marine Science, 80(2):367–373, March 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/2/367/6724458.

Charapata:2023:EEC

[CT23] Patrick Charapata and Stephen Trumble. Environmental and ecological changes influence lifetime trends of reproduction, stress, and stable isotopes reconstructed from female yelloweye rockfish opercula. ICES Journal of Marine Science, 80(5):1500–1515, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1500/7181088.

Coolen:2020:BBO

[CvdWC^{+}20] Joop W. P. Coolen, Babeth van der Weide, Joël Cuperus, Maxime Blomberg, Godfried W. N. M. Van Moorsel, Marco A. Faasse, Oscar G. Bos, Steven Degraer, and Han J. Lindeboom. Benthic biodiversity on old platforms, young wind farms, and rocky reefs. ICES Journal of Marine Science, 77(3):1250–1265, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1250/5057660.

Clark:2020:GAF

[CVJ^{+}20] Bethany L. Clark, Freydís Vigfúsdóttir, Mark J. Jessopp, Julian M. Burgos, Thomas W. Bodey, and Stephen C. Votier. Gannets are not attracted to fishing vessels in Iceland — potential influence of a discard ban and food availability. ICES Journal of Marine Science, 77(2):692–700, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/692/5652242.

Cvitanovic:2022:NFW

[CvPTAJ22] C. Cvitanovic, E. I. van Putten, P. Tuohy, and R. Annand-Jones. Normalizing failure: when things go wrong in participatory marine social science fieldwork. ICES Journal of Marine Science, 79(8):2184–2195, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2184/6705562.

Cadigan:2023:HMR

[CWB^{+}23] N. G. Cadigan, S. J. Walsh, H. P. Benoît, P. M. Regular, and L. J. Wheeland. A hierarchical model of the relative efficiency of two trawl survey protocols, with application to flatfish off the coast of Newfoundland. ICES Journal of Marine Science, 80(4):1087–1102, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1087/7080897.

Clarke:2023:IST

[CWC^{+}23] Thomas M. Clarke, Sasha K. Whitmarsh, Curtis Champion, Hugh Pederson, Lauren Meyer, Joshua D. Dennis, Ross G. Dwyer, and Charlie Huveneers. Influence of shark tourism on the activity and physiological condition of a non-focal pelagic fish. ICES Journal of Marine Science, 80(6):1670– 1682, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/6/1670/7208601.

Chen:2020:URI

[CWL^{+}20] Guogui Chen, Wenqing Wang, Yi Liu, Yamian Zhang, Wei Ma, Kun Xin, and Mao Wang. Uncovering the relative influences of space and environment in shaping the biogeographic patterns of mangrove mollusk diversity. ICES Journal of Marine Science, 77(1):30–39, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/77/1/30/5614895.

Chen:2023:UML

[CZ23] Yuwen Chen and Guoping Zhu. Using machine learning to alleviate the allometric effect in otolith shape-based species discrimination: the role of a triplet loss function. ICES Journal of Marine Science, 80(5):1277–1290, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1277/7100608.

Cardinale:2023:SES

[CZS^{+}23] M. Cardinale, F. Zimmermann, G. Søvik, C. A. Griffiths, M. Bergenius Nord, and H. Winker. Spatially explicit stock assessment uncovers sequential depletion of northern shrimp stock components in the North Sea. ICES Journal of Marine Science, 80(7):1868–1880, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1868/7231251.

Cao:2021:LBL

[CZSR21] Shuo Cao, Dean Zhao, Yueping Sun, and Chengzhi Ruan. Learning-based low-illumination image enhancer for underwater live crab detection. ICES Journal of Marine Science, 78(3):979–993, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/3/979/6104240.

DAmen:2020:LFI

[DA20] M. D’Amen and E. Azzurro. Lessepsian fish invasion in Mediterranean marine protected areas: a risk assessment under climate change scenarios. ICES Journal of Marine Science, 77(1):388–397, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/388/5634224.

Durif:2023:ESC

[DAB^{+}23] C. M. F. Durif, M. Arts, F. Bertolini, A. Cresci, F. Daverat, E. Karlsbakk, J. Koprivnikar, E. Moland, E. M. Olsen, C. Parzanini, M. Power, M. Rohtla, A. B. Skiftesvik, E. Thorstad, L. A. Vøllestad, and H. I. Browman. The evolving story of catadromy in the European eel (Anguilla anguilla). ICES Journal of Marine Science, 80(9):2253–2265, November 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/9/2253/7283903.

deAlmeida:2021:SRN

[dAdSM21] João Paulo Felix Augusto de Almeida, Robson Guimarães dos Santos, and Tamí Mott. Sex ratios and natal origins of green turtles from feeding grounds in the Southwest Atlantic Ocean. ICES Journal of Marine Science, 78(5):1840–1848, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/5/1840/6278037.

deAzevedo:2021:BBM

[dAPDDJ21] Eric Zettermann Dias de Azevedo, Pedro Pintassilgo, David Valença Dantas, and Fábio Gonçalves Daura-Jorge. Bioeconomic benefits of managing fishing effort in a coexisting small- and large-scale fishery game. ICES Journal of Marine Science, 78(7):2486–2495, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2486/6308528.

Diaz-Astudillo:2022:SIV

[DASLRB22] Macarena Díaz-Astudillo, Gonzalo S. Saldías, Jaime Letelier, and Ramiro Riquelme-Bugueño. Spatial and interannual variability in the distribution of euphausiid life stages in the permanent upwelling system off northern Chile. ICES Journal of Marine Science, 79(1):61–75, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/61/6453072. See correction [CRR^{+}22].

Dayton:2020:ASR

[Day20] Paul K. Dayton. Adventures scaling the realized niche, saving the world, and searching for values. ICES Journal of Marine Science, 77(5):1648–1657, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1648/5837188.

deBenito-Abello:2022:AIV

[dBABS^{+}22] Carmela de Benito-Abelló, Luís Bentes, Inês Sousa, Marie Pedaccini, David Villegas-Ríos, Esben Moland Olsen, Jorge M. S. Gonçalves, and Bárbara Horta e Costa. Amongindividual variation in white seabream (Diplodus sargus) spatial behaviour and protection in a coastal no-take area. ICES Journal of Marine Science, 79(8):2265–2276, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2265/6708393.

Dannheim:2020:BEO

[DBB^{+}20] Jennifer Dannheim, Lena Bergström, Silvana N. R. Birchenough, Radosław Brzana, Arjen R. Boon, Joop W. P.

Coolen, Jean-Claude Dauvin, Ilse De Mesel, Jozefien Derweduwen, Andrew B. Gill, Zoë L. Hutchison, Angus C. Jackson, Urszula Janas, Georg Martin, Aurore Raoux, Jan Reubens, Liis Rostin, Jan Vanaverbeke, Thomas A. Wilding, Dan Wilhelmsson, and Steven Degraer. Benthic effects of offshore renewables: identification of knowledge gaps and urgently needed research. ICES Journal of Marine Science, 77(3): 1092–1108, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/3/1092/5368123.

DiPoi:2022:LHT

[DBGP22] Carole Di Poi, Nicolas Brodu, Frédéric Gazeau, and Fabrice Pernet. Life-history traits in the Pacific oyster Crassostrea gigas are robust to ocean acidification under two thermal regimes. ICES Journal of Marine Science, 79(10):2614–2629, December 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/10/2614/6808347.

deBoois:2021:CFS

[dBSQK21] Ingeborg J. de Boois, Nathalie A. Steins, Floor J. Quirijns, and Marloes Kraan. The compatibility of fishers and scientific surveys: increasing legitimacy without jeopardizing credibility. ICES Journal of Marine Science, 78(5):1769– 1780, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/5/1769/6276505.

DiazPauli:2023:GDA

[DBT^{+}23] Beatriz Diaz Pauli, Henrik H. Berntsen, Eva B. Thorstad, Eydna ì Homrum, Susan Mærsk Lusseau, Vidar Wennevik, and Kjell Rong Utne. Geographic distribution, abundance, diet, and body size of invasive pink salmon (Oncorhynchus gorbuscha) in the Norwegian and Barents Seas, and in Norwegian rivers. ICES Journal of Marine Science, 80(1):76–90, January 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/1/76/6916929.

deCastro:2022:MYD

[dCKS^{+}22] Francisco de Castro, Sarah B. M. Kraak, Samuel Shephard, Alexander Sadykov, David G. Reid, and Keith D.

Farnsworth. Move and you’re dead: commercial trawl fisheries select for fish that don’t move far. ICES Journal of Marine Science, 79(6):1742–1752, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1742/6610010.

Dupaix:2021:SHM

[DCL^{+}21] Amaël Dupaix, Manuela Capello, Christophe Lett, Marco Andrello, Nicolas Barrier, Gaëlle Viennois, and Laurent Dagorn. Surface habitat modification through industrial tuna fishery practices. ICES Journal of Marine Science, 78(9):3075–3088, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3075/6368416.

Dupont:2020:SIT

[DDL^{+}20] Nicolas Dupont, Joël M. Durant, Øystein Langangen, Harald Gjøsæter, and Leif Christian Stige. Sea ice, temperature, and prey effects on annual variations in mean lengths of a key Arctic fish, Boreogadus saida, in the Barents Sea. ICES Journal of Marine Science, 77(5):1796–1805, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1796/5828349.

Dichmont:2021:OTB

[DDP^{+}21] Catherine M. Dichmont, Natalie A. Dowling, Sean Pascoe, Toni Cannard, Rachel J. Pears, Sian Breen, Tom Roberts, George M. Leigh, and Marc Mangel. Operationalizing triple bottom line harvest strategies. ICES Journal of Marine Science, 78(2):731–742, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/2/731/5812750.

Daly:2021:MIC

[DEL21] Benjamin J. Daly, Ginny L. Eckert, and W. Christopher Long. Moulding the ideal crab: implications of phenotypic plasticity for crustacean stock enhancement. ICES Journal of Marine Science, 78(1):421–434, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
421/5818947.

DeMartini:2022:RUE

[DeM22] Edward E. DeMartini. Responses to unexpected events; folk–rock–blues anthems; memes and other non-genetically inherited traits. ICES Journal of Marine Science, 79(3):593– 599, April 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/3/593/6535188.

Diserud:2022:NAD

[DFK^{+}22] O. H. Diserud, P. Fiske, S. Karlsson, K. A. Glover, T. Næsje, T. Aronsen, G. Bakke, B. T. Barlaup, J. Erkinaro, B. Florø-Larsen, A. Foldvik, M. Heino, Ø Kanstad-Hanssen, H. Lo, R. A. Lund, R. Muladal, E. Niemelä, F. Økland, G. M. Østborg, H. Otterå, Ø Skaala, H. Skoglund, I. Solberg, M. F. Solberg, V. P. Sollien, H. Sægrov, K. Urdal, V. Wennevik, and K. Hindar. Natural and anthropogenic drivers of escaped farmed salmon occurrence and introgression into wild Norwegian Atlantic salmon populations. ICES Journal of Marine Science, 79(4):1363–1379, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1363/6571569.

Droghini:2020:ROM

[DFWR20] A. Droghini, A. S. Fischbach, J. T. Watson, and J. P. Reimer. Regional ocean models indicate changing limits to biological invasions in the Bering Sea. ICES Journal of Marine Science, 77(3):964–974, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/3/964/5753900.

DePiper:2021:LDC

[DGM^{+}21] Geret DePiper, Sarah Gaichas, Brandon Muffley, Greg Ardini, Jeffrey Brust, Jessica Coakley, Kiley Dancy, G. Warren Elliott, Dustin C. Leaning, Douglas Lipton, Jason McNamee, Charles Perretti, Kirby Rootes-Murdy, and Michael J. Wilberg. Learning by doing: collaborative conceptual modelling as a path forward in ecosystem-based management. ICES Journal of Marine Science, 78(4):1217–1228, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/4/1217/6207633.

Durante:2024:NGI

[DGM^{+}24] Erica D. Durante, Gretchen L. Grammer, Jasmin C. Martino, Justin L. Payne, and Zoë A. Doubleday. Nondaily growth increments in the commercial species, Octopus berrima, and the importance of age validation. ICES Journal of Marine Science, 81(2):293–306, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/2/293/7457023.

dEntremont:2022:NGM

[dGW^{+}22] Kyle J. N. d’Entremont, Leanne M. Guzzwell, Sabina I. Wilhelm, Vicki L. Friesen, Gail K. Davoren, Carolyn J. Walsh, and William A. Montevecchi. Northern gannets (Morus bassanus) breeding at their southern limit struggle with prey shortages as a result of warming waters. ICES Journal of Marine Science, 79(1):50–60, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/50/6453071.

Dean:2021:ITE

[DHB^{+}21] Micah J. Dean, William S. Hoffman, Nicholas C. Buchan, Steven X. Cadrin, and Jonathan H. Grabowski. The influence of trawl efficiency assumptions on survey-based population metrics. ICES Journal of Marine Science, 78(8):2858– 2874, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/8/2858/6363577.

Drinkwater:2021:PFS

[DHN^{+}21] Kenneth F. Drinkwater, Naomi Harada, Shigeto Nishino, Melissa Chierici, Seth L. Danielson, Randi B. Ingvaldsen, Trond Kristiansen, George L. Hunt, Franz Mueter, and Jan Erik Stiansen. Possible future scenarios for two major Arctic Gateways connecting subarctic and Arctic marine systems: I. Climate and physical–chemical oceanography. ICES Journal of Marine Science, 78(9):3046–3065, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3046/6382408.

Dehnhard:2021:FCG

[DLA^{+}21] Nina Dehnhard, Magdalene Langset, Asgeir Aglen, Svein-Håkon Lorentsen, and Tycho Anker-Nilssen. Fish consumption by great cormorants in Norwegian coastal waters — a human-wildlife conflict for wrasses, but not gadids. ICES Journal of Marine Science, 78(3):1074–1089, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/1074/6132742.

delaCruz:2022:SDB

[dlCRT^{+}22] Andrés de la Cruz, Fernando Ramos, Jorge Tornero, Margarita María Rincón, Ma. Paz Jiménez, and Gonzalo Muñoz Arroyo. Seabird distribution is better predicted by abundance of prey than oceanography. A case study in the Gulf of Cadiz (SW, Iberian Peninsula). ICES Journal of Marine Science, 79(1):204–217, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/204/6479775.

DeRobertis:2021:USV

[DLL^{+}21] Alex De Robertis, Mike Levine, Nathan Lauffenburger, Taina Honkalehto, James Ianelli, Cole C. Monnahan, Rick Towler, Darin Jones, Sarah Stienessen, and Denise McKelvey. Uncrewed surface vehicle (USV) survey of walleye pollock, Gadus chalcogrammus, in response to the cancellation of ship-based surveys. ICES Journal of Marine Science, 78(8):2797–2808, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2797/6356484.

Dalpe:2022:IEC

[DLLC^{+}22] Andréanne Dalpé, Ariane Lalonde-Larue, Thierry Choquette, Jade Cimmino, Émélie Leroux, Christelle Leung, Donald T. Stewart, Bernard Angers, and Sophie Breton. The influence of environmental conditions on sex determination in the blue mussel Mytilus edulis. ICES Journal of Marine Science, 79(2):394–402, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/394/6503930.

daSilvaSantana:2021:TSO

[dLP21] Cesar Augusto da Silva Santana, Colm Lordan, and Anne Marie Power. Theoretical size at the onset of maturity and its density-dependent variability as an option in crustacean fisheries management. ICES Journal of Marine Science, 78(4):1421–1433, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1421/6168569.

Hidalgo-de-la-Toba:2021:IGP

[dlTVLC^{+}21] José Angel Hidalgo de-la Toba, Brent Vadopalas, Daniel Bernardo Lluch-Cota, Enrique Morales-Bojórquez, J. Jesús Bautista-Romero, and Sergio Scarry González-Peláez. Individual growth profiling improves growth modelling in the geoduck clam Panopea generosa. ICES Journal of Marine Science, 78(1):112–124, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/112/6026108.

deMoor:2022:LTD

[dMBJ22] Carryn L. de Moor, Douglas S. Butterworth, and Susan Johnston. Learning from three decades of Management Strategy Evaluation in South Africa. ICES Journal of Marine Science, 79(6):1843–1852, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1843/6645268.

Dupaix:2021:UBT

[DMK^{+}21] Amaël Dupaix, Laurène Mérillet, Dorothée Kopp, Maud Mouchet, and Marianne Robert. Using biological traits to get insights into the bentho-demersal community sensitivity to trawling in the Celtic Sea. ICES Journal of Marine Science, 78(3):1063–1073, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
1063/6130110.

Dobson:2023:OAR

[DNW^{+}23] K. L. Dobson, J. A. Newton, S. Widdicombe, K. L. Schoo, M. P. Acquafredda, G. Kitch, A. Bantelman, K. Lowder, A. Valauri-Orton, K. Soapi, K. Azetsu-Scott, and K. Isensee. Ocean acidification research for sustainability: co-designing global action on local scales. ICES Journal of Marine Science, 80(2):362–366, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/2/362/6698552.

Dolan:2020:NCW

[Dol20] John R. Dolan. The neglected contributions of William Beebe to the natural history of the deep-sea. ICES Journal of Marine Science, 77(5):1617–1628, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1617/5851565.

Domokos:2021:DAD

[Dom21] Réka Domokos. On the development of acoustic descriptors for semi-demersal fish identification to support monitoring stocks. ICES Journal of Marine Science, 78(3):1117– 1130, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/1117/6134884.

Dumas:2022:CDO

[DPD^{+}22a] Pascal Dumas, Christophe Peignon, Mahé Dumas, Bertrand Bourgeois, Hugues Gossuin, and Sylvie Fiat. Corrigendum to: Destructive outbreaks of the corallivorous starfish Acanthaster cf. solaris spare coral assemblages in the shallowest reef flat areas in New Caledonia. ICES Journal of Marine Science, 79(7):2161, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2161/6542747. See [DPD^{+}22b].

Dumas:2022:DOC

[DPD^{+}22b] Pascal Dumas, Christophe Peignon, Mahé Dumas, Bertrand Bourgeois, Hugues Gossuin, and Sylvie Fiat. Destructive outbreaks of the corallivorous starfish Acanthaster cf. solaris spare coral assemblages in the shallowest reef flat areas in New Caledonia. ICES Journal of Marine Science, 79(2): 350–361, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/2/350/6512396. See corri-

gendum [DPD^{+}22a].

Downie:2021:RBU

[DPH^{+}21] Anna-Leena Downie, Nils Piechaud, Kerry Howell, Christopher Barrio Froján, Mar Sacau, and Andrew Kenny. Reconstructing baselines: use of habitat suitability modelling to predict pre-fishing condition of a vulnerable marine ecosystem. ICES Journal of Marine Science, 78(8):2784–2796, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2784/6352558.

dePontual:2023:SMS

[dPHG^{+}23] Helene de Pontual, Karine Heerah, Jolien Goossens, François Garren, Stephane Martin, Loic Le Ru, Didier Le Roy, and Mathieu Woillez. Seasonal migration, site fidelity, and population structure of European seabass (Dicentrarchus labrax). ICES Journal of Marine Science, 80(6):1606–1618, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1606/7190301.

duPontavice:2022:OMB

[dPMS^{+}22] Hubert du Pontavice, Timothy J. Miller, Brian C. Stock, Zhuomin Chen, and Vincent S. Saba. Ocean model-based covariates improve a marine fish stock assessment when observations are limited. ICES Journal of Marine Science, 79(4): 1259–1273, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/1259/6567589.

DiCapua:2021:MDS

[DPMZ21] Iole Di Capua, Roberta Piredda, Maria Grazia Mazzocchi, and Adriana Zingone. Metazoan diversity and seasonality through eDNA metabarcoding at a Mediterranean long-term ecological research site. ICES Journal of Marine Science, 78(9):3303–3316, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3303/6226984.

Dimens:2023:WRW

[DRP^{+}23] Pavel Dimens, Lionel Reynal, Cedric Pau, Freddy Arocha, Fabio Hazin, Pollyana Roque, Nancie J. Cummings, James S. Franks, Kenneth L. Jones, and Eric A. Saillant. Weak range-wide population structure in the blackfin tuna (Thunnus atlanticus) revealed by analysis of genome-wide SNPs. ICES Journal of Marine Science, 80(4):941–952, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/941/7058011.

DAmen:2022:SLF

[DSDA22] Manuela D’Amen, Sonia Smeraldo, Antonio Di Franco, and Ernesto Azzurro. The spread of Lessepsian fish does not track native temperature conditions. ICES Journal of Marine Science, 79(6):1864–1873, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1864/6633758.

DiSalvatore:2021:FRO

[DSF^{+}21] Pablo Di Salvatore, Hernán J. Sacristán, Olga Florentín, Martín Varisco, and Gustavo A. Lovrich. Female reproductive output and potential recruitment of three fished southern king crab stocks from the Southern Atlantic Ocean. ICES Journal of Marine Science, 78(7):2628–2642, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/7/2628/6350833.

Denechaud:2020:LTT

[DSGG20] Côme Denechaud, Szymon Smoliński, Audrey J. Geffen, and Jane A. Godiksen. Long-term temporal stability of Northeast Arctic cod (Gadus morhua) otolith morphology. ICES Journal of Marine Science, 77(3):1043–1054, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/3/1043/5770870.

Dahle:2022:GAE

[DSMM^{+}22] Geir Dahle, Bernard Sainte-Marie, Sarah L. Mincks, Eva Farestveit, Knut E. Jørstad, Ann Merete Hjelset, and Ann-Lisbeth Agnalt. Genetic analysis of the exploited snow crab (Chionoecetes opilio) in the Barents Sea — possibilities of origin. ICES Journal of Marine Science, 79(9):2389–2398, November 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/9/2389/6753243.

Drenkard:2021:NGR

[DSR^{+}21] Elizabeth J. Drenkard, Charles Stock, Andrew C. Ross, Keith W. Dixon, Alistair Adcroft, Michael Alexander, Venkatramani Balaji, Steven J. Bograd, Momme Butenschön, Wei Cheng, Enrique Curchitser, Emanuele Di Lorenzo, Raphael Dussin, Alan C. Haynie, Matthew Harrison, Albert Hermann, Anne Hollowed, Kirstin Holsman, Jason Holt, Michael G. Jacox, Chan Joo Jang, Kelly A. Kearney, Barbara A. Muhling, Mercedes Pozo Buil, Vincent Saba, Anne Britt Sandø, Désirée Tommasi, and Muyin Wang. Next-generation regional ocean projections for living marine resource management in a changing climate. ICES Journal of Marine Science, 78(6):1969–1987, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
1969/6307379.

Rodrigues:2022:OST

[dSRKC22] Lucas dos Santos Rodrigues, Paul Gerhard Kinas, and Luis Gustavo Cardoso. Optimal setting time and season increase the target and reduce the incidental catch in longline fisheries: a Bayesian beta mixed regression approach. ICES Journal of Marine Science, 79(4):1245–1258, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1245/6563423.

Durant:2021:RSD

[DYS21] Joël M. Durant, Natalia Yaragina, and Leif Christian Stige. The role of spatial distribution for growth and survival of juvenile cod Gadus morhua in the Barents Sea. ICES Journal of Marine Science, 78(8):2700–2708, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/8/2700/6352557.

Earle:2021:MSY

[Ear21] Michael Earle. Maximum sustainable yield in the EU’s Common Fisheries Policy — a political history. ICES Journal of Marine Science, 78(6):2173–2181, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2173/6168021.

Eayrs:2023:RMC

[Eay23] Stephen Eayrs. A road map to change: application of a comprehensive change management model to guide and inspire fishers to reduce bycatch. ICES Journal of Marine Science, 80(3):446–457, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/3/446/6612121.

Emmerson:2022:EFD

[ECBJ22] J. A. Emmerson, M. T. Coleman, I. S. M. Bloor, and S. R. Jenkins. Enhancing fishery-dependent information in datapoor fisheries; integrating gear-in–gear-out sensors and mobile reporting technology in a mixed Irish Sea static-gear fishery. ICES Journal of Marine Science, 79(7):2126–2137, September 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/7/2126/6674270.

Eighani:2021:UFI

[ECR^{+}21] Morteza Eighani, Jason M. Cope, Paria Raoufi, Reza Abbaspour Naderi, and Pascal Bach. Understanding fishery interactions and stock trajectory of yellowfin tuna exploited by Iranian fisheries in the Sea of Oman. ICES Journal of Marine Science, 78(7):2420–2431, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2420/6299664.

Edmunds:2020:YLE

[EDC20] Peter J. Edmunds, Steve S. Doo, and Robert C. Carpenter. Year-long effects of high pCO_{2} on the community structure of a tropical fore reef assembled in outdoor flumes. ICES Journal of Marine Science, 77(3):1055–1065, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/3/1055/5762457.

Eero:2021:UFW

[EDH^{+}21] Margit Eero, Jan Dierking, Christoph Humborg, Emma Undeman, Brian R. MacKenzie, Henn Ojaveer, Tiina Salo, and Friedrich Wilhelm Köster. Use of food web knowledge in environmental conservation and management of living resources in the Baltic Sea. ICES Journal of Marine Science, 78(8):2645–2663, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2645/6355112.

Espinasse:2022:SAE

[EDH^{+}22] Boris Espinasse, Malin Daase, Elisabeth Halvorsen, Marit Reigstad, Jørgen Berge, and Sünnje L. Basedow. Surface aggregations of Calanus finmarchicus during the polar night. ICES Journal of Marine Science, 79(3):803–814, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/803/6549746.

Erisman:2020:BCUa

[EGMO^{+}20a] Brad E. Erisman, Arnaud Grüss, Ismael Mascareñas-Osorio, Hector Lícon-González, Andrew F. Johnson, and Catalina López-Sagástegui. Balancing conservation and utilization in spawning aggregation fisheries: a trade-off analysis of an overexploited marine fish. ICES Journal of Marine Science, 77(1):148–161, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/148/5601456.

Erisman:2020:BCUb

[EGMO^{+}20b] Brad E. Erisman, Arnaud Grüss, Ismael Mascareñas-Osorio, Hector Lícon-González, Andrew F. Johnson, and Catalina López-Sagástegui. Balancing conservation and utilization in spawning aggregation fisheries: a trade-off analysis of an overexploited marine fish. ICES Journal of Marine Science, 77(5):2010–2011, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
2010/5815627.

Escalle:2021:QDF

[EHV^{+}21] Lauriane Escalle, Steven R. Hare, Tiffany Vidal, Maurice Brownjohn, Paul Hamer, and Graham Pilling. Quantifying drifting fish aggregating device use by the world’s largest tuna fishery. ICES Journal of Marine Science, 78(7):2432– 2447, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/7/2432/6307380.

Earl:2021:WDS

[EMS^{+}21] Timothy Earl, Eleanor MacLeod, Marta Söffker, Nico Gasco, Félix Massiot-Granier, Paul Tixier, and Christopher Darby. Whale depredation in the South Georgia Patagonian toothfish (Dissostichus eleginoides) fishery in the South Atlantic: a comparison of estimation methods. ICES Journal of Marine Science, 78(10):3817–3833, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3817/6444887.

Endo:2021:OWC

[ESK^{+}21] Hikaru Endo, Yoichi Sato, Kenji Kaneko, Daisuke Takahashi, Kazue Nagasawa, Yutaka Okumura, and Yukio Agatsuma. Ocean warming combined with nutrient enrichment increases the risk of herbivory during cultivation of the marine macroalga Undaria pinnatifida. ICES Journal of Marine Science, 78(1):402–409, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
402/5851564.

Espinel-Velasco:2021:OAI

[EVLK^{+}21] Nadjejda Espinel-Velasco, Miles Lamare, Anna Kluibenschedl, Graeme Moss, and Vonda Cummings. Ocean acidification induces carry-over effects on the larval settlement of the New Zealand abalone, Haliotis iris. ICES Journal of Marine Science, 78(1):340–348, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
340/5859739.

Endo:2021:ECS

[EVY^{+}21] Clarissa Akemi Kajiya Endo, Frode B. Vikebø, Natalia A. Yaragina, Solfrid Sætre Hjøllo, and Leif Christian Stige. Effects of climate and spawning stock structure on the spatial distribution of Northeast Arctic cod larvae. ICES Journal of Marine Science, 78(2):666–679, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/2/
666/5835270.

Ershova:2021:MQT

[EWD^{+}21] E. A. Ershova, O. S. Wangensteen, R. Descoteaux, C. Barth-Jensen, and K. Præbel. Metabarcoding as a quantitative tool for estimating biodiversity and relative biomass of marine zooplankton. ICES Journal of Marine Science, 78(9):3342– 3355, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/78/9/3342/6360557. See cor-

rigendum [EWD^{+}22].

Ershova:2022:CMQ

[EWD^{+}22] E. A. Ershova, O. S. Wangensteen, R. Descoteaux, C. Barth-Jensen, and K. Præbel. Corrigendum to: Metabarcoding as a quantitative tool for estimating biodiversity and relative biomass of marine zooplankton. ICES Journal of Marine Science, 79(1):242, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/242/6425782.

See [EWD^{+}21].

Fox:2020:SRN

[FAV^{+}20] Clive J. Fox, Amaya Albalat, Daniel Valentinsson, Hans C. Nilsson, Frank Armstrong, Peter Randall, and Thomas Catchpole. Survival rates for Nephrops norvegicus discarded from Northern European trawl fisheries. ICES Journal of Marine Science, 77(5):1698–1710, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1698/5807909.

Filmalter:2021:MBF

[FBF^{+}21] John D. Filmalter, Robert K. Bauer, Fabien Forget, Paul D. Cowley, and Laurent Dagorn. Movement behaviour and fishery interaction of silky sharks (Carcharhinus falciformis) in the tropical tuna purse seine fishery in the Western Indian Ocean. ICES Journal of Marine Science, 78(7):2474–2485, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/7/2474/6307376.

Fauconnet:2023:CAD

[FCD^{+}23] Laurence Fauconnet, Diana Catarino, Diya Das, Eva Giacomello, José Manuel Gonzalez-Irusta, Pedro Afonso, and Telmo Morato. Challenges in avoiding deep-water shark bycatch in Azorean hook-and-line fisheries. ICES Journal of Marine Science, 80(3):605–619, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/605/6775934.

Froehlich:2021:MGB

[FCF^{+}21] Halley E. Froehlich, Jessica Couture, Lynne Falconer, Gesche Krause, James A. Morris, Montse Perez, Grant D. Stentiford, Harri Vehviläinen, and Benjamin S. Halpern. Mind the gap between ICES nations’ future seafood consumption and aquaculture production. ICES Journal of Marine Science, 78(1):468–477, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
468/5828449.

Fischer:2020:LPD

[FDK20] Simon H. Fischer, José A. A. De Oliveira, and Laurence T. Kell. Linking the performance of a data-limited empirical catch rule to life-history traits. ICES Journal of Marine Science, 77(5):1914–1926, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1914/5856265.

Fischer:2021:AEP

[FDMK21a] Simon H. Fischer, José A. A. De Oliveira, John D. Mumford, and Laurence T. Kell. Application of explicit precautionary principles in data-limited fisheries management. ICES Journal of Marine Science, 78(8):2931–2942, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2931/6364352.

Fischer:2021:UGAa

[FDMK21b] Simon H. Fischer, José A. A. De Oliveira, John D. Mumford, and Laurence T. Kell. Using a genetic algorithm to optimize a data-limited catch rule. ICES Journal of Marine Science, 78(4):1311–1323, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1311/6161236.

Fischer:2021:UGAb

[FDMK21c] Simon H. Fischer, José A. A. De Oliveira, John D. Mumford, and Laurence T. Kell. Using a genetic algorithm to optimize a data-limited catch rule. ICES Journal of Marine Science, 78(8):3013–3014, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
3013/6382646.

Fischer:2022:ERH

[FDMK22] Simon H. Fischer, José A. A. De Oliveira, John D. Mumford, and Laurence T. Kell. Exploring a relative harvest rate strategy for moderately data-limited fisheries management. ICES Journal of Marine Science, 79(6):1730–1741, August 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/6/1730/6611681.

Filbee-Dexter:2023:SFC

[FDPD^{+}23] Karen Filbee-Dexter, Albert Pessarrodona, Carlos M. Duarte, Dorte Krause-Jensen, Kasper Hancke, Daniel Smale, and Thomas Wernberg. Seaweed forests are carbon sinks that may help mitigate CO_{2} emissions: a comment on Gallagher et al. (2022). ICES Journal of Marine Science, 80(6):1814–1819, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/

6/1814/7220133. See [GSL22] and reply [FDPD^{+}23].

Fjelldal:2022:ELS

[FFH^{+}22] Per Gunnar Fjelldal, Thomas W. K. Fraser, Tom J. Hansen, Ørjan Karlsen, and Samantha Bui. Effects of laboratory salmon louse infection on mortality, growth, and sexual maturation in Atlantic salmon. ICES Journal of Marine Science, 79(5):1530–1538, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
5/1530/6581593.

Farmer:2020:FRFa

[FFR20a] Nicholas A. Farmer, John T. Froeschke, and David L. Records. Forecasting for recreational fisheries management: a derby fishery case study with Gulf of Mexico red snapper. ICES Journal of Marine Science, 77(3):1248, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/3/1248/5780425.

Farmer:2020:FRFb

[FFR20b] Nicholas A. Farmer, John T. Froeschke, and David L. Records. Forecasting for recreational fisheries management: a derby fishery case study with Gulf of Mexico red snapper. ICES Journal of Marine Science, 77(6):2265–2284, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2265/5687231.

Feuilloley:2022:TFZ

[FFS^{+}22] Guillaume Feuilloley, Jean-Marc Fromentin, Claire Saraux, Jean-Olivier Irisson, Laetitia Jalabert, and Lars Stemmann. Temporal fluctuations in zooplankton size, abundance, and taxonomic composition since 1995 in the North Western Mediterranean Sea. ICES Journal of Marine Science, 79 (3):882–900, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/3/882/6383441.

Filous:2022:MYT

[FFT^{+}22] Alexander Filous, Alan M. Friedlander, Melvin Toribiong, Robert J. Lennox, Geory Mereb, and Yimnang Golbuu. The movements of yellowfin tuna, blue marlin, and sailfish within the Palau National Marine Sanctuary and the western Pacific Ocean. ICES Journal of Marine Science, 79(2):445–456, March 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/2/445/6519834.

Field:2021:WCS

[Fie21] John G. Field. Was this career shaped by people or by chance? ICES Journal of Marine Science, 78(10):3552–3561, December 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/10/3552/6430632.

Fowler:2020:EID

[FJC^{+}20] A. M. Fowler, A.-M. Jørgensen, J. W. P. Coolen, D. O. B. Jones, J. C. Svendsen, R. Brabant, B. Rumes, and S. Degraer. The ecology of infrastructure decommissioning in the North Sea: what we need to know and how to achieve it. ICES Journal of Marine Science, 77(3):1109–1126, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/3/1109/5543459.

Funk:2020:GFK

[FKTM20] Steffen Funk, Uwe Krumme, Axel Temming, and Christian Möllmann. Gillnet fishers’ knowledge reveals seasonality in depth and habitat use of cod (Gadus morhua) in the Western Baltic Sea. ICES Journal of Marine Science, 77(5):1816– 1829, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/5/1816/5831971.

Flensborg:2023:IBA

[FMBL23] L. C. Flensborg, A. A. Maureaud, D. N. Bravo, and M. Lindegren. An indicator-based approach for assessing marine ecosystem resilience. ICES Journal of Marine Science, 80(5): 1487–1499, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/5/1487/7175313.

French:2020:DNN

[FMF^{+}20] Geoff French, Michal Mackiewicz, Mark Fisher, Helen Holah, Rachel Kilburn, Neil Campbell, and Coby Needle. Deep neural networks for analysis of fisheries surveillance video and automated monitoring of fish discards. ICES Journal of Marine Science, 77(4):1340–1353, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1340/5542623.

Fulton:2024:MPC

[FMPH24] Elizabeth A Fulton, Nastaran Mazloumi, Aaron Puckeridge, and Roshan Hanamseth. Modelling perspective on the climate footprint in south east Australian marine waters and its fisheries. ICES Journal of Marine Science, 81(1):130–144, January 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/1/130/7458235.

Fjortoft:2020:ADE

[FNB^{+}20] Helene Børretzen Fjørtoft, Frank Nilsen, Francois Besnier, Per Gunnar Espedal, Anne Stene, Ann-Kristin Tveten, Pål Arne Bjørn, Vidar Teis Aspehaug, and Kevin Alan Glover. Aquaculture-driven evolution: distribution of pyrethroid resistance in the salmon louse throughout the North Atlantic in the years 2000–2017. ICES Journal of Marine Science, 77(5):1806–1815, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1806/5831179.

Ferreira:2023:IMM

[FND23] A. Sofia A. Ferreira, Anna B. Neuheimer, and Joël M. Durant. Impacts of the match–mismatch hypothesis across three trophic levels — a case study in the North Sea. ICES Journal of Marine Science, 80(2):308–316, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/2/308/6988656.

Ferter:2023:ICS

[FOC^{+}23] Keno Ferter, Håkon Otterå, Mary Christman, Alf Ring Kleiven, Marc Simon Weltersbach, Sofie Gundersen, Christine Djønne, Otte Bjelland, Bruce Hartill, Jeremy Lyle, Kieran Hyder, Trude Borch, and Jon Helge Vølstad. Integrating complementary survey methods to estimate catches in Norway’s complex marine recreational hook-and-line fishery. ICES Journal of Marine Science, 80(1):107–121, January 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/1/107/6958703.

Fowler:2022:IRF

[FODD^{+}22] Ashley M. Fowler, Faith A. Ochwada-Doyle, Natalie A. Dowling, Heath Folpp, Julian M. Hughes, Michael B. Lowry, Jeremy M. Lyle, Tim P. Lynch, Nathan G. Miles, and Rowan C. Chick. Integrating recreational fishing into harvest strategies: linking data with objectives. ICES Journal of Marine Science, 79(2):285–307, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/285/6507800.

Fortune:2020:EBP

[FP20] I. S. Fortune and D. M. Paterson. Ecological best practice in decommissioning: a review of scientific research. ICES Journal of Marine Science, 77(3):1079–1091, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1079/5107800.

Farhadi:2022:GWS

[FPY^{+}22] Ahmad Farhadi, Florian Pichlmueller, Bhargavi Yellapu, Shane Lavery, and Andrew Jeffs. Genome-wide SNPs reveal fine-scale genetic structure in ornate spiny lobster Panulirus ornatus throughout Indo–West Pacific Ocean. ICES Journal of Marine Science, 79(6):1931–1941, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1931/6646037.

Friedland:2023:ARW

[FRP^{+}23] Kevin D. Friedland, Nicholas R. Record, Daniel E. Pendleton, William M. Balch, Karen Stamieszkin, John R. Moisan, and Damian C. Brady. Asymmetry in the rate of warming and the phenology of seasonal blooms in the Northeast US Shelf Ecosystem. ICES Journal of Marine Science, 80 (4):775–786, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/4/775/7028404.

Ferriss:2022:SEA

[FRSM22] Bridget E. Ferriss, Jonathan C. P. Reum, Beth L. Sanderson, and P. Sean McDonald. Social-ecological approaches to shellfish aquaculture using qualitative network models. ICES Journal of Marine Science, 79(4):1289–1301, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1289/6565269.

Fisch:2023:PSV

[FSC^{+}23] N. Fisch, K. Shertzer, E. Camp, M. Maunder, and R. Ahrens. Process and sampling variance within fisheries stock assessment models: estimability, likelihood choice, and the consequences of incorrect specification. ICES Journal of Marine Science, 80(8):2125–2149, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/8/2125/7260034.

Fulton:2022:SBD

[FSN^{+}22] E. A. Fulton, K. Sainsbury, P. Noranarttragoon, D. Leadbitter, D. J. Staples, J. Porobic, Y. Ye, R. Phoonsawat, and N. Kulanujaree. Shifting baselines and deciding on the desirable form of multispecies maximum sustainable yield. ICES Journal of Marine Science, 79(7):2138–2154, September 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/7/2138/6698554.

Fuller:2023:TIE

[FSS^{+}23] Jessica L. Fuller, Harry V. Strehlow, Jörn O. Schmidt, Örjan Bodin, and Dorothy J. Dankel. Tracking integrated ecosystem assessments in the ICES network: a social network analysis of the ICES expert groups. ICES Journal of Marine Science, 80(2):282–294, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/2/282/6987542.

Ford:2022:QRA

[FW22] Jessica H. Ford and Chris Wilcox. Quantifying risk assessments for monitoring control and surveillance of illegal fishing. ICES Journal of Marine Science, 79(4):1113–1119, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1113/6542748.

Flores:2021:CNP

[FWA^{+}21] Andrés Flores, Rodrigo Wiff, Mauricio Ahumada, Dante Queirolo, and Pedro Apablaza. Coping with El Niño: phenotypic flexibility of reproductive traits in red squat lobster determines recruitment success. ICES Journal of Marine Science, 78(10):3709–3723, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
10/3709/6425789.

Froese:2020:ESS

[FWC^{+}20] Rainer Froese, Henning Winker, Gianpaolo Coro, Nazli Demirel, Athanassios C. Tsikliras, Donna Dimarchopoulou, Giuseppe Scarcella, Maria Lourdes Deng Palomares, Manuel Dureuil, and Daniel Pauly. Estimating stock status from relative abundance and resilience. ICES Journal of Marine Science, 77(2):527–538, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/2/527/5682447.

Fu:2020:REI

[FXG^{+}20] Caihong Fu, Yi Xu, Arnaud Grüss, Alida Bundy, Lynne Shannon, Johanna J. Heymans, Ghassen Halouani, Ekin Akoglu, Christopher P. Lynam, Marta Coll, Elizabeth A. Fulton, Laure Velez, and Yunne-Jai Shin. Responses of ecological indicators to fishing pressure under environmental change: exploring non-linearity and thresholds. ICES Journal of Marine Science, 77(4):1516–1531, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1516/5575246.

Gallagher:2023:RCFa

[GA23] K. M. Gallagher and P. G. Albano. Range contractions, fragmentation, species extirpations, and extinctions of commercially valuable molluscs in the Mediterranean Sea — a climate warming hotspot. ICES Journal of Marine Science, 80(5):1382–1398, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1382/7136035.

Gallagher:2023:RCFb

[Gal23] John B. Gallagher. Reply to the comment by Filbee– Dexter et al. (2023) “Seaweed forests are carbon sinks that may help mitigate CO_{2} emissions”. ICES Journal of Marine Science, 80(6):1820–1826, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1820/7230593. See [GSL22, FDPD^{+}23].

Garagouni:2022:QES

[GAM^{+}22] Maria Garagouni, Georgia Avgerinou, Foivos-Alexandros Mouchlianitis, George Minos, and Konstantinos Ganias. Questionnaire and experimental surveys show that dolphins cause substantial losses to a gillnet fishery in the eastern Mediterranean Sea. ICES Journal of Marine Science, 79(9):2552–2561, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
9/2552/6795159.

Garcia:2021:TQM

[GAP21] Dorleta Garcia, Inmaculada Arostegui, and Raúl Prellezo. To be or not to be a target: that is the question to manage mixed fisheries. ICES Journal of Marine Science, 78(7):2562–2578, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2562/6327202.

Garcia-Baron:2023:ERA

[GBGA^{+}23] Isabel García-Barón, Igor Granado, Amaia Astarloa, Guillermo Boyra, Anna Rubio, José A. Fernandes-Salvador, Lucía Zarauz, Iñigo Onandia, Estanis Mugerza, and Maite Louzao. Ecological risk assessment of a pelagic seabird species in artisanal tuna fisheries. ICES Journal of Marine Science, 80(9):2441–2454, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
9/2441/6671890.

Golden:2024:CAD

[GBH^{+}24] Abigail S. Golden, Marissa L. Baskett, Dan Holland, Arielle Levine, Kathy Mills, and Timothy Essington. Climate adaptation depends on rebalancing flexibility and rigidity in US fisheries management. ICES Journal of Marine Science, 81 (2):252–259, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/81/2/252/7459506.

Groner:2023:TED

[GBMM^{+}23] M. L. Groner, E. D. Bravo-Mendosa, A. H. MacKenzie, J. L. Gregg, C. M. Conway, J. T. Trochta, and P. K. Hershberger. Temporal, environmental, and demographic correlates of Ichthyophonus sp. infections in mature Pacific herring populations. ICES Journal of Marine Science, 80(9):2342– 2355, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/9/2342/7285197. See cor-

rection [Ano23k].

Goethel:2021:RRA

[GC21] Daniel R. Goethel and Steven X. Cadrin. Revival and recent advancements in the spatial fishery models originally conceived by Sidney Holt and Ray Beverton. ICES Journal of Marine Science, 78(6):2298–2315, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2298/6288531.

Gremillet:2022:BDA

[GCG22] David Grémillet, Damien Chevallier, and Christophe Guinet. Big data approaches to the spatial ecology and conservation of marine megafauna. ICES Journal of Marine Science, 79 (4):975–986, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/975/6564869.

Greer:2021:SPE

[GCT^{+}21] Adam T. Greer, Luciano M. Chiaverano, Laura M. Treible, Christian Briseño-Avena, and Frank J. Hernandez. From spatial pattern to ecological process through imaging zooplankton interactions. ICES Journal of Marine Science, 78(8):2664–2674, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2664/6350848.

Gruss:2023:ISO

[GCT^{+}23] A. Grüss, A. R. Charsley, J. T. Thorson, O. F. Anderson, R. L. O'Driscoll, B. Wood, O. N. Breivik, and C. A. O’Leary. Integrating survey and observer data improves the predictions of New Zealand spatio-temporal models. ICES Journal of Marine Science, 80(7):1991–2007, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1991/7246967. See correction [Ano23g].

Garcia:2020:MSH

[GDI^{+}20] Dorleta Garcia, Paul J. Dolder, Ane Iriondo, Claire Moore, Raúl Prellezo, and Agurtzane Urtizberea. A multi-stock harvest control rule based on “pretty good yield” ranges to support mixed-fisheries management. ICES Journal of Marine Science, 77(1):119–135, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/1/
119/5585654.

Gislason:2023:PSD

[GEBS^{+}23] Daví Gíslason, Daniel Estévez-Barcia, Sæmundur Sveinsson, Agneta Hansen, Denis Roy, Margaret Treble, Jesper Boje, Mikko Vihtakari, Bjarki ór Elvarsson, Kevin Hedges, Elvar H. Hallfredsson, and Torild Johansen. Population structure discovered in juveniles of Greenland halibut (Reinhardtius hippoglossoides Walbaum, 1792). ICES Journal of Marine Science, 80(4):889–896, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/889/7048517.

Galatius:2021:UEV

[GETvB21] Anders Galatius, Søs Gerster Engbo, Jonas Teilmann, and Floris M. van Beest. Using environmental variation to optimize aerial surveys of harbour seals. ICES Journal of Marine Science, 78(4):1500–1507, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1500/6186734.

Guo:2020:IEF

[GFO^{+}20] Chuanbo Guo, Caihong Fu, Norm Olsen, Yi Xu, Arnaud Grüss, Huizhu Liu, Philippe Verley, and Yunne-Jai Shin. Incorporating environmental forcing in developing ecosystembased fisheries management strategies. ICES Journal of Marine Science, 77(2):500–514, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
500/5682448.

Gimeno:2023:ARD

[GGA^{+}23] Míriam Gimeno, José Antonio García, Isabel Afán, Raül Aymí, Tomás Montalvo, and Joan Navarro. Age-related differences in foraging behaviour at sea and interactions with fishing vessels in an opportunistic urban gull. ICES Journal of Marine Science, 80(9):2405–2413, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/9/2405/6623969.

Guiet:2020:BIH

[GGBC20] J. Guiet, E. D. Galbraith, D. Bianchi, and W. W. L. Cheung. Bioenergetic influence on the historical development and decline of industrial fisheries. ICES Journal of Marine Science, 77(5):1854–1863, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1854/5835400.

Geissinger:2022:HSF

[GGLS22] Emilie A. Geissinger, Robert S. Gregory, Benjamin J. Laurel, and Paul V. R. Snelgrove. High site-fidelity and low mortality of juvenile Atlantic cod (Gadus morhua) in subarctic coastal habitat during their first winter. ICES Journal of Marine Science, 79(4):1408–1418, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1408/6568583.

Gislason:2021:IAD

[GGOP21] Astthor Gislason, Kristinn Gudmundsson, Solveig R. Olafsdottir, and Hildur Petursdottir. Inter-annual and decadal variability of Calanus finmarchicus and C. hyperboreus in subarctic waters north of Iceland 1990–2020. ICES Journal of Marine Science, 78(10):3735–3747, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3735/6429122. See erratum [GGOP22].

Gislason:2022:EIA

[GGOP22] Astthor Gislason, Kristinn Gudmundsson, Solveig R. Olafsdottir, and Hildur Petursdottir. Erratum to: Inter-annual and decadal variability of Calanus finmarchicus and C. hyperboreus in subarctic waters north of Iceland 1990– 2020. ICES Journal of Marine Science, 79(1):246, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/246/6470667. See [GGOP21].

Gray:2022:MIS

[GGRB22] P. Gray, C. Garcia, C. Robinson, and J. Bremner. A method for identifying sensitivity of marine benthic invertebrates to ocean acidification through a biological traits approach. ICES Journal of Marine Science, 79(7):2117–2125, September 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/7/2117/6674736.

Gregory:2023:ERA

[GGW^{+}23] Stephen D. Gregory, Jonathan P. Gillson, Katie Whitlock, Jon Barry, Peter Gough, Robert J. Hillman, David Mee, Graeme Peirson, Brian A. Shields, Lawrence Talks, Simon Toms, Alan M. Walker, Ben Wilson, and Ian C. Davidson. Estimation of returning Atlantic salmon stock from rod exploitation rate for principal salmon rivers in England & Wales. ICES Journal of Marine Science, 80(10):2504–2519, December 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/10/2504/7335584.

Gose:2023:SCI

[GHB^{+}23] M-A Gose, E. Humble, A. Brownlow, B. Mikkelsen, C. Loftus, D. Wall, E. Rogan, M. ten Doeschate, N. Davison, and R. Ogden. Stranding collections indicate broad-scale connectivity across the range of a pelagic marine predator, the Atlantic white-sided dolphin (Lagenorhynchus acutus). ICES Journal of Marine Science, 80(4):1120–1128, May 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/4/1120/7091392.

Goodwin:2022:UPD

[GHJ^{+}22] Morten Goodwin, Kim Tallaksen Halvorsen, Lei Jiao, Kristian Muri Knausgård, Angela Helen Martin, Marta Moyano, Rebekah A. Oomen, Jeppe Have Rasmussen, Tonje Knutsen Sørdalen, and Susanna Huneide Thorbjørnsen. Unlocking the potential of deep learning for marine ecology: overview, applications, and outlook. ICES Journal of Marine Science, 79(2):319–336, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/2/319/6507793.

Gislason:2020:GCF

[GHÓ^{+}20] David Gíslason, Sarah J. Helyar, Gudmundur J. Óskarsson, Gudbjörg Ólafsdóttir, Aril Slotte, Teunis Jansen, Jan Arge Jacobsen, Kristinn Ólafsson, Sigurlaug Skirnisdottir, Geir Dahle, Helle Siegstad, Hóraldur Joensen, Kiersten L. Curti, François Grégoire, Jacques Masse, Sæmundur Sveinsson, Anna Kristín Daníelsdóttir, and Christophe Pampoulie. The genetic composition of feeding aggregations of the Atlantic mackerel (Scomber scombrus) in the central north Atlantic: a microsatellite loci approach. ICES Journal of Marine Science, 77(2):604–612, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/2/604/5721484.

Gonzalez-Irusta:2022:MHL

[GICP^{+}22] José Manuel González-Irusta, Joan E. Cartes, Antonio Punzón, David Díaz, Luis Gil de Sola, and Alberto Serrano. Mapping habitat loss in the deep-sea using current and past presences of Isidella elongata (Cnidaria: Alcyonacea). ICES Journal of Marine Science, 79(6):1888–1901, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1888/6640479.

Gros:2023:IVM

[GJU^{+}23] Charley Gros, Jan Jansen, Candice Untiedt, Tabitha R. R. Pearman, Rachel Downey, David K. A. Barnes, David A.

Bowden, Dirk C. Welsford, and Nicole A. Hill. Identifying vulnerable marine ecosystems: an image-based vulnerability index for the Southern Ocean seafloor. ICES Journal of Marine Science, 80(4):972–986, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/972/7058020.

Godwin:2021:SLA

[GKRB21] Sean C. Godwin, Martin Krkosek, John D. Reynolds, and Andrew W. Bateman. Sea-louse abundance on salmon farms in relation to parasite-control policy and climate change. ICES Journal of Marine Science, 78(1):377–387, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/1/377/5983852.

Gronbaek:2023:AUC

[GLM^{+}23] L. Grønbæk, M. Lindroos, G. Munro, P. Pintassilgo, and B. Turris. The avoidance of unwanted catch and cooperation: the case of the British Columbia groundfish trawl fishery. ICES Journal of Marine Science, 80(3):483–491, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/483/6524923.

Glencross:2021:PFR

[GLW21] Jacqueline S. Glencross, Jennifer L. Lavers, and Eric J. Woehler. A proposed framework for reporting mass mortality (wreck) events of seabirds. ICES Journal of Marine Science, 78(6):1935–1942, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
1935/6304978.

Guilhon:2021:REB

[GMT21] Maila Guilhon, Francesc Montserrat, and Alexander Turra. Recognition of ecosystem-based management principles in key documents of the seabed mining regime: implications and further recommendations. ICES Journal of Marine Science, 78(3):884–899, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/3/884/6050568.

Gislason:2021:PDT

[GPJ^{+}21] Sindri Gíslason, Snæbjörn Pálsson, Jónas P. Jónasson, Hermann Dreki Guls, Jörundur Svavarsson, and Halldór P. Halldórsson. Population dynamics of three brachyuran crab species (Decapoda) in Icelandic waters: impact of recent colonization of the Atlantic rock crab (Cancer irroratus). ICES Journal of Marine Science, 78(2):534–544, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/2/534/5827682.

Ganias:2021:CSE

[GPM21] Konstantinos Ganias, Rozalia Perri, and Anastasios Moditsis. Cuttlefish and squid egg deposition patterns on artificial devices and trap-like gears: implications for offspring survival and population management. ICES Journal of Marine Science, 78(5):1614–1622, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1614/6225976.

Garcia:2020:ASF

[GPQ^{+}20] Rafael Garcia, Ricard Prados, Josep Quintana, Alexander Tempelaar, Nuno Gracias, Shale Rosen, Håvard Vågstøl, and Kristoffer Løvall. Automatic segmentation of fish using deep learning with application to fish size measurement. ICES Journal of Marine Science, 77(4):1354–1366, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/4/1354/5602457.

Gibbs:2021:AEOb

[GPS^{+}21a] Mitchell C. Gibbs, Laura M. Parker, Elliot Scanes, Maria Byrne, Wayne A. O’Connor, and Pauline M. Ross. Adult exposure to ocean acidification and warming leads to limited beneficial responses for oyster larvae. ICES Journal of Marine Science, 78(6):2017–2030, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2017/6306814.

Gibbs:2021:AEOa

[GPS^{+}21b] Mitchell C. Gibbs, Laura M. Parker, Elliot Scanes, Maria Byrne, Wayne A. O’Connor, and Pauline M. Ross. Adult exposure to ocean acidification and warming remains beneficial for oyster larvae following starvation. ICES Journal of Marine Science, 78(5):1587–1598, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1587/6224218.

Garcia:2022:GPE

[GPV^{+}22] J. Garcia, V. Pasqualini, L. Vanalderweireldt, P. A. Bisgambiglia, M. Marengo, P. Lejeune, A. Aiello, and E. D. H. Durieux. Global patterns and environmental drivers of suitable habitat for Dentex dentex and Sciaena umbra along the Corsican coast. ICES Journal of Marine Science, 79(9):2461–2472, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
9/2461/6762983.

Grant:2021:MLR

[Gra21] W. Stewart Grant. My life with the Red Queen in fishery genetics. ICES Journal of Marine Science, 78(7):2351–2358, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/7/2351/6297550.

Georgeson:2020:ERD

[GRE^{+}20] L. Georgeson, C. L. Rigby, T. J. Emery, M. Fuller, J. Hartog, A. J. Williams, A. J. Hobday, C. A. J. Duffy, C. A. Simpfendorfer, T. Okuda, I. C. Stobutzki, and S. J. Nicol. Ecological risks of demersal fishing on deepwater chondrichthyan populations in the Southern Indian and South Pacific Oceans. ICES Journal of Marine Science, 77(5):1711–1727, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1711/5811377.

Griffith:2020:CGB

[Gri20] Gary P. Griffith. Closing the gap between causality, prediction, emergence, and applied marine management. ICES Journal of Marine Science, 77(4):1456–1462, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/4/1456/5857784.

Gallagher:2022:RSC

[GS22] John Barry Gallagher and Victor Shelamoff. Reply to Stafford’s (2022) comment on “Seaweed ecosystems may not mitigate CO_{2} emissions” by Gallagher et al. (2022). ICES Journal of Marine Science, 79(5):1703–1704, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1703/6589471. See [GSL22, Sta22].

Gallagher:2022:SEM

[GSL22] John Barry Gallagher, Victor Shelamoff, and Cayne Layton. Seaweed ecosystems may not mitigate CO_{2} emissions. ICES Journal of Marine Science, 79(3):585–592, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.oup.

com/icesjms/article/79/3/585/6525671. See comments

[Sta22, FDPD^{+}23] and replies [GS22, Gal23].

Grant:2020:IDS

[GSR^{+}20] Michael I. Grant, Jonathan J. Smart, Cassandra L. Rigby, William T. White, Andrew Chin, Leontine Baje, and Colin A. Simpfendorfer. Intraspecific demography of the silky shark (Carcharhinus falciformis): implications for fisheries management. ICES Journal of Marine Science, 77(1): 241–255, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/1/241/5614226.

Gamboa-Salazar:2020:EAS

[GSWBK20] Keilin R. Gamboa-Salazar, David M. Wyanski, Walter J. Bubley, and Nikolai Klibansky. Effects of age and size on spawning and egg production in gag and scamp grouper off the southeastern United States. ICES Journal of Marine Science, 77(1):290–299, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/290/5581803.

Garcia-Seoane:2021:BPM

[GSWM^{+}21] Eva García-Seoane, Rupert Wienerroither, Kjell Arne Mork, Melanie J. Underwood, and Webjørn Melle. Biogeographical patterns of meso- and bathypelagic fish along a Northeastern Atlantic transect. ICES Journal of Marine Science, 78(4): 1444–1457, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1444/6169451. See erra-

tum [GSWM^{+}22].

Garcia-Seoane:2022:EBP

[GSWM^{+}22] Eva García-Seoane, Rupert Wienerroither, Kjell Arne Mork, Melanie J. Underwood, and Webjørn Melle. Erratum to: Biogeographical patterns of meso- and bathypelagic fish along a Northeastern Atlantic transect. ICES Journal of Marine Science, 79(7):2159–2160, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2159/6649360. See [GSWM^{+}21].

Godo:2022:VAA

[GT22] Olav Rune Godø and Philip Trathan. Voluntary actions by the Antarctic krill fishing industry help reduce potential negative impacts on land-based marine predators during breeding, highlighting the need for CCAMLR action. ICES Journal of Marine Science, 79(5):1457–1466, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1457/6596342.

Goossens:2023:FMP

[GVD^{+}23] Jolien Goossens, Damian Villagra, Georges De Putter, Pieterjan Verhelst, Els Torreele, Tom Moens, and Jan Reubens. Fisheries measures protect European seabass groups with distinct habitat use differently. ICES Journal of Marine Science, 80(7):1899–1910, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1899/7241087.

Goldstein:2022:RDA

[GZSP22] Jason S. Goldstein, Katrina A. Zarrella-Smith, and Tracy L. Pugh. Recent declines in American lobster fecundity in southern New England: drivers and implications. ICES Journal of Marine Science, 79(5):1662–1674, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1662/6594655.

Gogina:2020:CSD

[GZW^{+}20] Mayya Gogina, Michael L. Zettler, Irene Wåhlström, Helén Andersson, Hagen Radtke, Ivan Kuznetsov, and Brian R. MacKenzie. A combination of species distribution and oceanbiogeochemical models suggests that climate change overrides eutrophication as the driver of future distributions of a key benthic crustacean in the estuarine ecosystem of the Baltic Sea. ICES Journal of Marine Science, 77(6):2089– 2105, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/6/2089/5880824.

Hilborn:2023:ESE

[HAC^{+}23] R. Hilborn, R. Amoroso, J. Collie, J. G. Hiddink, M. J. Kaiser, T. Mazor, R. A. McConnaughey, A. M. Parma, C. R. Pitcher, M. Sciberras, and P. Suuronen. Evaluating the sustainability and environmental impacts of trawling compared to other food production systems. ICES Journal of Marine Science, 80(6):1567–1579, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1567/7226311.

Hintzen:2021:QHP

[HAP^{+}21] N. T. Hintzen, G. Aarts, J. J. Poos, K. J. Van der Reijden, and A. D. Rijnsdorp. Quantifying habitat preference of bottom trawling gear. ICES Journal of Marine Science, 78(1): 172–184, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/1/172/6026099.

Hilborn:2021:TBB

[HAPW21] Ray Hilborn, Caitlin Allen Akselrud, Henry Peterson, and George A. Whitehouse. The trade-off between biodiversity and sustainable fish harvest with area-based management. ICES Journal of Marine Science, 78(6):2271–2279, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/2271/5891348.

Hare:2020:TLF

[Har20] Jonathan A. Hare. Ten lessons from the frontlines of science in support of fisheries management. ICES Journal of Marine Science, 77(3):870–877, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/3/870/5803784.

Hart:2021:PRM

[Har21] Deborah R. Hart. Per-recruit modelling of pulse fisheries: comment on “Modelling pulse fishery systems in data-limited situations”. ICES Journal of Marine Science, 78(6):2146– 2149, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/6/2146/6306810.

Hidalgo:2022:LAS

[HBC^{+}22] Manuel Hidalgo, Valerio Bartolino, Marta Coll, Mary E. Hunsicker, Morgane Travers-Trolet, and Howard I. Browman. ‘Adaptation science’ is needed to inform the sustainable management of the world’s oceans in the face of climate change. ICES Journal of Marine Science, 79(2):457–462, March 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/2/457/6529728.

Horton:2020:TAB

[HBD^{+}20] Thomas W. Horton, Barbara A. Block, Alan Drumm, Lucy A. Hawkes, Macdara O’Cuaig, Niall Ó Maoiléidigh, Ross O’Neill, Robert J. Schallert, Michael J. W. Stokesbury, and Matthew J. Witt. Tracking Atlantic bluefin tuna from foraging grounds off the west coast of Ireland. ICES Journal of Marine Science, 77(6):2066–2077, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2066/5878157.

Horton:2021:EIO

[HBD^{+}21] Thomas W. Horton, Barbara A. Block, Rachel Davies, Lucy A. Hawkes, Duncan Jones, Hannah Jones, Keith Leeves, Niall Ó Maoiléidigh, David Righton, Jeroen van der Kooij, Dave Wall, and Matthew J. Witt. Evidence of increased occurrence of Atlantic bluefin tuna in territorial waters of the United Kingdom and Ireland. ICES Journal of Marine Science, 78(5):1672–1683, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1672/6231587.

Haulsee:2020:SSB

[HBFO20] Danielle E. Haulsee, Matthew W. Breece, Dewayne A. Fox, and Matthew J. Oliver. Simple is sometimes better: a test of the transferability of species distribution models. ICES Journal of Marine Science, 77(5):1752–1761, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1752/5817084.

He:2020:DLT

[HBMC20] Ji X. He, James R. Bence, Charles P. Madenjian, and Randall M. Claramunt. Dynamics of lake trout production in the main basin of Lake Huron. ICES Journal of Marine Science, 77(3):975–987, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/3/975/5802541.

Hiddink:2024:EAB

[HCB^{+}24] Jan Geert Hiddink, Matthew T Coleman, Stephen Brouwer, Isobel S M Bloor, and Stuart R Jenkins. Estimating the abundance of benthic invertebrates from trap-catch data. ICES Journal of Marine Science, 81(1):86–96, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/1/86/7379034.

Haarr:2020:ESE

[HCCR20] Marthe Larsen Haarr, Michel Comeau, Jöel Chassé, and Rémy Rochette. Early spring egg hatching by the American lobster (Homarus americanus) linked to rising water temperature in autumn. ICES Journal of Marine Science, 77(5):1685–1697, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1685/5807920.

Huynh:2020:CML

[HCH20] Quang C. Huynh, Nancie J. Cummings, and John M. Hoenig. Comparisons of mean length-based mortality estimators and age-structured models for six southeastern US stocks. ICES Journal of Marine Science, 77(1):162–173, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/1/162/5607277.

Hiddink:2024:UOD

[HCM24] Jan Geert Hiddink, Ryan Charles, and Alec B M Moore. Using opportunistic data to study the distribution and abundance of a warm water elasmobranch at the northern edge of its range. ICES Journal of Marine Science, 81(1):108–118, January 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/1/108/7458872.

Hidalgo:2022:RAO

[HEHT^{+}22] M. Hidalgo, A. E. El-Haweet, A. C. Tsikliras, E. M. Tirasin, T. Fortibuoni, F. Ronchi, V. Lauria, O. Ben Abdallah, E. Arneri, L. Ceriola, N. Milone, S. Lelli, P. Hernández, M. Bernal, and M. Vasconcellos. Risks and adaptation options for the Mediterranean fisheries in the face of multiple climate change drivers and impacts. ICES Journal of Marine Science, 79(9):2473–2488, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/9/2473/6775938.

Hohne:2023:OMP

[HFP^{+}23] Leander Höhne, Marko Freese, Jan-Dag Pohlmann, Markus Diekmann, Erik Fladung, Jeroen B. J. Huisman, Reinhold Hanel, and Lasse Marohn. Overestimating management progress — modelled vs. monitored silver eel escapement in a North Sea draining river. ICES Journal of Marine Science, 80(7):1936–1948, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1936/7239532.

Holte:2022:IES

[HFT^{+}22] Børge Holte, Mona M. Fuhrmann, Anne Helene S. Tandberg, Carsten Hvingel, and Ann Merete Hjelset. Infaunal and epifaunal secondary production in the Barents Sea, with focus on snow crab (Chionoecetes opilio) prey resources and consumption. ICES Journal of Marine Science, 79(9):2524– 2539, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/9/2524/6795160. See cor-

rection [Ano23f].

Henderson:2021:SHM

[HGS^{+}21] Christopher J. Henderson, Ben L. Gilby, Edward Stone, Hayden P. Borland, and Andrew D. Olds. Seascape heterogeneity modifies estuarine fish assemblages in mangrove forests. ICES Journal of Marine Science, 78(3):1108–1116, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/1108/6133740.

Holt:2021:SCC

[HHA^{+}21] Rebecca E. Holt, Carsten Hvingel, Ann-Lisbeth Agnalt, Andrey V. Dolgov, Ann Merete Hjelset, and Bjarte Bogstad. Snow crab (Chionoecetes opilio), a new food item for North-east Arctic cod (Gadus morhua) in the Barents Sea. ICES Journal of Marine Science, 78(2):491–501, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/2/491/6255454.

Huserbraaten:2023:MDS

[HHD^{+}23] M. Huserbråten, A. M. Hjelset, H. E. H. Danielsen, C. Hvingel, and A-L. Agnalt. Modelled dispersal of snow crab (Chionoecetes opilio) larvae and potential settlement areas in the western Barents Sea. ICES Journal of Marine Science, 80(5): 1342–1350, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/5/1342/7127569.

Hansten:2021:REI

[HHK21] Magnus Hanstén, Päivi Haapasaari, and Sakari Kuikka. A realist evaluation of the individual transferable quota system used in Finnish herring fisheries. ICES Journal of Marine Science, 78(10):3603–3614, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3603/6383435.

Hollyman:2021:LRR

[HHL^{+}21] Philip R. Hollyman, Simeon L. Hill, Vladimir V. Laptikhovsky, Mark Belchier, Susan Gregory, Alice Clement, and Martin A. Collins. A long road to recovery: dynamics and ecology of the marbled rockcod (Notothenia rossii, family: Nototheniidae) at South Georgia, 50 years after overexploitation. ICES Journal of Marine Science, 78(8):2745– 2756, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/8/2745/6354501.

Holm:2022:CFC

[HHN22] Poul Holm, Patrick W. Hayes, and John Nicholls. Comment on “Five centuries of cod catches in eastern Canada,” by Schijns et al. ICES Journal of Marine Science, 79(5):1705– 1707, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1705/6593175. See [SFHP21] and

reply [SFHP22].

Hirai:2021:DRM

[HHNS21] Junya Hirai, Kiyotaka Hidaka, Satoshi Nagai, and Yugo Shimizu. DNA/RNA metabarcoding and morphological analysis of epipelagic copepod communities in the Izu Ridge off the southern coast of Japan. ICES Journal of Marine Science, 78(9):3444–3456, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3444/6282522.

Hornborg:2020:CIH

[HHO^{+}20] Sara Hornborg, Alistair J. Hobday, Emily Ogier, Aysha Fleming, Linda Thomas, and Jason R. Hartog. Challenges and insights from holistic sustainability reporting for shrimp fisheries in different jurisdictions. ICES Journal of Marine Science, 77(6):2022–2032, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2022/5870404.

Huserbraaten:2022:STR

[HJ22] M. B. O. Huserbråten and I. A. Johnsen. Seasonal temperature regulates network connectivity of salmon louse. ICES Journal of Marine Science, 79(4):1075–1082, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1075/6532050.

Head:2022:PMN

[HJP22] Erica J. H. Head, Catherine L. Johnson, and Pierre Pepin. Plankton monitoring in the Northwest Atlantic: a comparison of zooplankton abundance estimates from vertical net tows and Continuous Plankton Recorder sampling on the Scotian and Newfoundland shelves, 1999– 2015. ICES Journal of Marine Science, 79(3):901–916, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/901/6414827.

He:2023:EME

[HJS^{+}23] X. He, N. W. Jeffery, R. R. E. Stanley, L. C. Hamilton, E. M. Rubidge, and C. L. Abbott. eDNA metabarcoding enriches traditional trawl survey data for monitoring biodiversity in the marine environment. ICES Journal of Marine Science, 80(5):1529–1538, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1529/7181086.

Hamana:2021:MSS

[HK21] Masahiro Hamana and Teruhisa Komatsu. Mapping 3D structure of a Sargassum forest with high-resolution sounding data obtained by multibeam echosounder. ICES Journal of Marine Science, 78(4):1458–1469, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1458/6169513.

Hudson:2023:TSU

[HKC^{+}23] Charlotte G. Hudson, Emily Knight, Sarah L. Close, Jason P. Landrum, Angela Bednarek, and Ben Shouse. Telling stories to understand research impact: narratives from the Lenfest Ocean Program. ICES Journal of Marine Science, 80 (2):394–400, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/2/394/6747651.

Hamilton:2023:VEM

[HKZ^{+}23] S. L. Hamilton, E. G. Kennedy, M. Zulian, T. M. Hill, B. Gaylord, E. Sanford, A. M. Ricart, M. Ward, A. K. Spalding, and K. Kroeker. Variable exposure to multiple climate stressors across the California marine protected area network and policy implications. ICES Journal of Marine Science, 80(7):1923–1935, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1923/7231250.

Huynh:2022:CLS

[HLHC22] Quang C. Huynh, Christopher M. Legault, Adrian R. Hordyk, and Tom R. Carruthers. A closed-loop simulation framework and indicator approach for evaluating impacts of retrospective patterns in stock assessments. ICES Journal of Marine Science, 79(7):2003–2016, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2003/6574661.

Henderson:2023:SUA

[HMA^{+}23] M. E. Henderson, K. E. Mills, M. A. Alexander, M. Barajas, M. J. Collins, M. Dzaugis, D. Kircheis, and T. F. Sheehan. A synthesis of US Atlantic salmon habitat requirements and implications for future suitability under a changing climate. ICES Journal of Marine Science, 80(8):2051–2073, October 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/8/2051/7272634.

Ho:2021:PRT

[HMBDP21] Maureen Ho, James McBroom, Ellie Bergstrom, and Guillermo Diaz-Pulido. Physiological responses to temperature and ocean acidification in tropical fleshy macroalgae with varying affinities for inorganic carbon. ICES Journal of Marine Science, 78(1):89–100, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
89/6032771.

Hart:2020:SSS

[HMC^{+}20] Deborah R. Hart, Daphne M. Munroe, Joseph C. Caracappa, Dale Haidvogel, Burton V. Shank, David B. Rudders, John M. Klinck, Eileen E. Hofmann, and Eric N. Powell. Spillover of sea scallops from rotational closures in the Mid-Atlantic Bight (United States). ICES Journal of Marine Science, 77(5):1992–2002, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1992/5864487.

Hodgdon:2021:CMA

[HMF^{+}21] Cameron T. Hodgdon, Mackenzie D. Mazur, Kevin D. Friedland, Nathan Willse, and Yong Chen. Consequences of model assumptions when projecting habitat suitability: a caution of forecasting under uncertainties. ICES Journal of Marine Science, 78(6):2092–2108, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2092/6294921.

Hyder:2020:MRF

[HMK^{+}20] Kieran Hyder, Christos D. Maravelias, Marloes Kraan, Zachary Radford, and Raul Prellezo. Marine recreational fisheries — current state and future opportunities. ICES Journal of Marine Science, 77(6):2171–2180, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/6/2171/5924560.

Holah:2022:IEM

[HMNF22] Helen Holah, C. Tara Marshall, Coby Needle, and Rob Fryer. The impact of electronic monitoring on fleet wide discarding of small cod in Scottish demersal fisheries. ICES Journal of Marine Science, 79(6):1753–1764, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1753/6619463.

Heikinheimo:2022:CPM

[HMOR22] Outi Heikinheimo, Timo J. Marjomäki, Mikko Olin, and Pekka Rusanen. Cormorant predation mortality of perch (Perca fluviatilis) in coastal and archipelago areas, northern Baltic Sea. ICES Journal of Marine Science, 79(2):337–349, March 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/2/337/6491039.

Hubert:2022:ADB

[HMWS22] Jeroen Hubert, Rosalie Moens, Rob Witbaard, and Hans Slabbekoorn. Acoustic disturbance in blue mussels: soundinduced valve closure varies with pulse train speed but does not affect phytoplankton clearance rate. ICES Journal of Marine Science, 79(9):2540–2551, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/9/2540/6779724.

Holt:2020:BMS

[Hol20] Sidney J. Holt. Becoming a marine scientist: helped by a daily quota of three lumps of coal. ICES Journal of Marine Science, 77(2):463–468, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
463/5722013.

Holt:2021:TLC

[Hol21] Sidney J. Holt. Three lumps of coal: doing fisheries research in Lowestoft in the 1940s. ICES Journal of Marine Science, 78(6):2155–2165, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2155/6032774.

Henriksson:2023:MOJ

[HPS^{+}23] Simon Henriksson, Ricardo T. Pereyra, Marte Sodeland, Olga Ortega-Martinez, Halvor Knutsen, Håkan Wennhage, and Carl André. Mixed origin of juvenile Atlantic cod (Gadus morhua) along the Swedish west coast. ICES Journal of Marine Science, 80(1):145–157, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/145/6908767.

Hinchliffe:2021:NAE

[HPSF21] Charles Hinchliffe, Pierre Pepin, Iain M. Suthers, and Daniel S. Falster. A novel approach for estimating growth and mortality of fish larvae. ICES Journal of Marine Science, 78(8):2684–2699, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2684/6364353.

Henriksen:2021:TBS

[HRB^{+}21] Ole Henriksen, Anna Rindorf, Mollie E. Brooks, Martin Lindegren, and Mikael van Deurs. Temperature and body size affect recruitment and survival of sandeel across the North Sea. ICES Journal of Marine Science, 78(4):1409–1420, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1409/6168659.

Hornborg:2020:FFG

[HS20] Sara Hornborg and Anthony D. M. Smith. Fisheries for the future: greenhouse gas emission consequences of different fishery reference points. ICES Journal of Marine Science, 77(5):1666–1671, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1666/5827678.

Hvingel:2021:CWS

[HSMK21] Carsten Hvingel, Bernard Sainte-Marie, and Gordon H. Kruse. Cold-water shellfish as harvestable resources and important ecosystem players. ICES Journal of Marine Science, 78(2):479–490, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/2/479/6130080.

Henly:2021:DIC

[HSS21a] Lauren Henly, James E. Stewart, and Stephen D. Simpson. Drivers and implications of change in an inshore multispecies fishery. ICES Journal of Marine Science, 78(5):1815– 1825, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/5/1815/6278034. See erra-

tum [HSS21b].

Henly:2021:EDI

[HSS21b] Lauren Henly, James E. Stewart, and Stephen D. Simpson. Erratum to: Drivers and implications of change in an inshore multi-species fishery. ICES Journal of Marine Science, 78(8):3015, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/

3015/6372893. See [HSS21a].

Hagen:2021:EGE

[HUJ^{+}21] Ingerid J. Hagen, Ola Ugedal, Arne J. Jensen, Håvard Lo, Espen Holthe, Bjørn Bjøru, Bjørn Florø-Larsen, Harald Sægrov, Helge Skoglund, and Sten Karlsson. Evaluation of genetic effects on wild salmon populations from stock enhancement. ICES Journal of Marine Science, 78(3):900– 909, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/900/6050573.

Hutchings:2022:TCS

[Hut22] Jeffrey A. Hutchings. Tensions in the communication of science advice on fish and fisheries: northern cod, species at risk, sustainable seafood. ICES Journal of Marine Science, 79(2):308–318, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/2/308/6502540.

Hansen:2021:ASB

[HvdMLS21] Cecilie Hansen, Gro I. van der Meeren, Harald Loeng, and Morten D. Skogen. Assessing the state of the Barents Sea using indicators: how, when, and where? ICES Journal of Marine Science, 78(8):2983–2998, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2983/6290288.

Hiddink:2023:STG

[HVDvD23] J. G. Hiddink, S. Valanko, A. J. Delargy, and P. D. van Denderen. Setting thresholds for good ecosystem state in marine seabed systems and beyond. ICES Journal of Marine Science, 80(4):698–709, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/4/698/7070114.

Haase:2022:PEM

[HWL^{+}22] Kevin Haase, Marc Simon Weltersbach, Wolf-Christian Lewin, Christopher Zimmermann, and Harry Vincent Strehlow. Potential effects of management options on marine recreational fisheries — the example of the western Baltic cod fishery. ICES Journal of Marine Science, 79(3):661– 676, April 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/3/661/6529070.

Hansen:2021:GDB

[HWS^{+}21] Agneta Hansen, Jon-Ivar Westgaard, Guldborg Søvik, Tanja Hanebrekke, Einar Magnus Nilssen, Per Erik Jorde, Jon Albretsen, and Torild Johansen. Genetic differentiation between inshore and offshore populations of northern shrimp (Pandalus borealis). ICES Journal of Marine Science, 78(9):3135–3146, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3135/6372894.

Hunt:2022:CWL

[HYE^{+}22] George L. Hunt, Jr., Ellen M. Yasumiishi, Lisa B. Eisner, Phyllis J. Stabeno, and Mary Beth Decker. Climate warming and the loss of sea ice: the impact of sea-ice variability on the southeastern Bering Sea pelagic ecosystem. ICES Journal of Marine Science, 79(3):937–953, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/937/6046591.

Indivero:2023:IDS

[IEIT23] Julia Indivero, Timothy E. Essington, James N. Ianelli, and James T. Thorson. Incorporating distribution shifts and spatio-temporal variation when estimating weight-at-age for stock assessments: a case study involving the Bering Sea pollock (Gadus chalcogrammus). ICES Journal of Marine Science, 80(2):258–271, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/2/258/6971590.

Ingolfsson:2021:SSH

[IHL21] Ólafur Arnar Ingólfsson, Odd-Børre Humborstad, and Svein Løkkeborg. Surface selection of haddock and cod in the Norwegian demersal seine fisheries. ICES Journal of Marine Science, 78(4):1508–1518, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1508/6199614.

Itakura:2021:TOT

[IOS21] Hikaru Itakura, Michael H. P. O’Brien, and David Secor. Tracking oxy-thermal habitat compression encountered by Chesapeake Bay striped bass through acoustic telemetry. ICES Journal of Marine Science, 78(3):1049–1062, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/3/1049/6129638.

Iriarte:2022:SNC

[IVU^{+}22] Arantza Iriarte, Fernando Villate, Ibon Uriarte, Gorka Bidegain, and Ziortza Barroeta. Shifts in neritic copepod communities off the Basque coast (southeastern Bay of Biscay) between 1998 and 2015. ICES Journal of Marine Science, 79 (3):830–843, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/3/830/6514217.

Jardim:2021:OEM

[JAB^{+}21] Ernesto Jardim, Manuela Azevedo, Jon Brodziak, Elizabeth N. Brooks, Kelli F. Johnson, Nikolai Klibansky, Colin P. Millar, Cóilín Minto, Iago Mosqueira, Richard D. M. Nash, Paraskevas Vasilakopoulos, and Brian K. Wells. Operationalizing ensemble models for scientific advice to fisheries management. ICES Journal of Marine Science, 78(4):1209–1216, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1209/6154492.

Jansson:2023:GSR

[JAQ^{+}23] Eeva Jansson, Carl André, María Quintela, Kim T. Halvorsen, François Besnier, Fernando Ayllon, Ellika Faust, Halvor Knutsen, Åsa Strand, and Kevin A. Glover. Genetic study reveals local differentiation persisting in the face of high connectivity and a genomic inversion likely linked with sexual antagonism in a common marine fish. ICES Journal of Marine Science, 80(4):1103–1112, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
4/1103/7083380.

Johannesen:2023:GEC

[JBDK23] E. Johannesen, F. Barz, D. J. Dankel, and S. B. M. Kraak. Gender and early career status: variables of participation at an international marine science conference. ICES Journal of Marine Science, 80(4):1016–1027, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1016/7071071.

Joo:2021:IPS

[JBE^{+}21] Rocío Joo, Nicolas Bez, Marie-Pierre Etienne, Pablo Marin, Nicolas Goascoz, Jérôme Roux, and Stéphanie Mahévas. Identifying partners at sea from joint movement metrics of pelagic pair trawlers. ICES Journal of Marine Science, 78(5):1758–1768, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1758/6276504.

Jourdain:2020:ESS

[JBF^{+}20] N. O. A. S. Jourdain, O. Breivik, E. Fuglebakk, S. Aanes, and J. H. Vølstad. Evaluation of sampling strategies for age determination of cod (Gadus morhua) sampled at the North Sea international bottom trawl survey. ICES Journal of Marine Science, 77(3):859–869, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
859/5770878.

Jiang:2022:CCE

[JBJ^{+}22] Long Jiang, Lander Blommaert, Henrice M. Jansen, Ole Jacob Broch, Klaas R. Timmermans, and Karline Soetaert. Carrying capacity of Saccharina latissima cultivation in a Dutch coastal bay: a modelling assessment. ICES Journal of Marine Science, 79(3):709–721, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/709/6535875.

Jac:2021:AIT

[JDD^{+}21] Cyrielle Jac, Nicolas Desroy, Jean-Claude Duchêne, Aurélie Foveau, Céline Labrune, Lyvia Lescure, and Sandrine Vaz. Assessing the impact of trawling on benthic megafauna: comparative study of video surveys vs. scientific trawling. ICES Journal of Marine Science, 78(5):1636–1649, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/5/1636/6231553.

Jones:2022:DBS

[JDJK22] Douglas Jones, Elin Dahlgren, Philip Jacobson, and Agnes M. L. Karlson. Determining Baltic salmon foraging areas at sea using stable isotopes in scales — a tool for understanding health syndromes. ICES Journal of Marine Science, 79(1): 158–168, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/1/158/6461168.

Jenkins:2023:TTT

[Jen23] Lekelia D. Jenkins. Turtles, TEDs, tuna, dolphins, and diffusion of innovations: key drivers of adoption of bycatch reduction devices. ICES Journal of Marine Science, 80(3): 417–436, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/3/417/6964916.

Jenkins:2023:UPB

[JEPT23] Lekelia D. Jenkins, Stephen Eayrs, Michael V. Pol, and Katie R. Thompson. Uptake of proven bycatch reduction fishing gear: perceived best practices and the role of affective change readiness. ICES Journal of Marine Science, 80 (3):437–445, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/3/437/6643483.

Jansen:2021:CSL

[JG21] Peder A. Jansen and Bjarne Gjerde. Comment on “Salmon lice-induced mortality of Atlantic salmon post-smolt during migration in Norway” by Johnsen et al. (2021). ICES Journal of Marine Science, 78(10):3847–3851, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3847/6440303. See [JHS^{+}21b, JHN^{+}21].

Johnsen:2021:RJG

[JHN^{+}21] Ingrid A. Johnsen, Alison Harvey, Pål Næverlid Sævik, Anne D. Sandvik, Ola Ugedal, Bjørn Ådlandsvik, Vidar Wennevik, Kevin A. Glover, and Ørjan Karlsen. Reply to Jansen and Gjerde’s (2021) critique of the salmon louse infection model reported in Johnsen et al. (2021). ICES Journal of Marine Science, 78(10):3852–3857, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3852/6440307. See [JHS^{+}21b, JG21].

Johnsen:2021:SLIa

[JHS^{+}21a] Ingrid A. Johnsen, Alison Harvey, Pål Næverlid Sævik, Anne D. Sandvik, Ola Ugedal, Bjørn Ådlandsvik, Vidar Wennevik, Kevin A. Glover, and Ørjan Karlsen. Salmon liceinduced mortality of Atlantic salmon during post-smolt migration in Norway. ICES Journal of Marine Science, 78(1): 142–154, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/1/142/6026111.

Johnsen:2021:SLIb

[JHS^{+}21b] Ingrid A. Johnsen, Alison Harvey, Pål Næverlid Sævik, Anne D. Sandvik, Ola Ugedal, Bjørn Ådlandsvik, Vidar Wennevik, Kevin A. Glover, and Ørjan Karlsen. Salmon lice-induced mortality of Atlantic salmon during post-smolt migration in Norway. ICES Journal of Marine Science, 78 (4):1555, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1555/6304977. See com-

ment [JG21] and reply [JHN^{+}21].

Jiang:2022:MSV

[JJB^{+}22] Long Jiang, Henrice M. Jansen, Ole Jacob Broch, Klaas R. Timmermans, and Karline Soetaert. Modelling spatial variability of cultivated Saccharina latissima in a Dutch coastal bay shows benefits of co-cultivation with shellfish. ICES Journal of Marine Science, 79(8):2324–2335, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2324/6726653.

Johannesen:2022:PDR

[JKF^{+}22] Edda Johannesen, Dave Keith, Michael J. Fogarty, Nancy Shackell, and Kenneth T. Frank. Persistent differences in recruitment variability among co-occurring North Atlantic groundfish species. ICES Journal of Marine Science, 79(9):2430–2441, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
9/2430/6758131.

Jamieson:2022:RPD

[JLA22] Alan J. Jamieson, Thomas D. Linley, and Prema Arasu. Reply to: People do care about the deep sea. a comment on jamieson et al. (2020). ICES Journal of Marine Science, 79(8):2340–2343, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
8/2340/6712685.

Jones:2023:TDP

[JLM^{+}23] Lucas F. Jones, R. Nicolas Lou, Christopher S. Murray, Dominique Robert, Christina M. Bourne, Caroline Bouchard, Marek Kučka, Y. Frank Chan, David B. Carlon, David N. Wiley, Nina O. Therkildsen, and Hannes Baumann. Two distinct population clusters of northern sand lance (Ammodytes dubius) on the northwest Atlantic shelf revealed by whole genome sequencing. ICES Journal of Marine Science, 80(1): 122–132, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/1/122/6872753.

Jacobsen:2022:CMS

[JMB^{+}22] Nis S. Jacobsen, Kristin N. Marshall, Aaron M. Berger, Chris Grandin, and Ian G. Taylor. Climate-mediated stock redistribution causes increased risk and challenges for fisheries management. ICES Journal of Marine Science, 79(4):1120– 1132, May 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/4/1120/6537151.

Jonkers:2022:LZT

[JMR^{+}22] Lukas Jonkers, Julie Meilland, Marina C. Rillo, Thibault de Garidel Thoron, John A. Kitchener, and Michal Kucera. Linking zooplankton time series to the fossil record. ICES Journal of Marine Science, 79(3):917–924, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/917/6317417.

Jonasdottir:2022:ECF

[JNT^{+}22] Sigrún H. Jónasdóttir, Lars Naustvoll, Frederik Wolff Teglhus, Mette Dalgaard Agersted, Julie C. Grenwald, Webjørn Melle, and Torkel Gissel Nielsen. Calanus finmarchicus basin scale life history traits and role in community carbon turnover during spring. ICES Journal of Marine Science, 79 (3):785–802, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/3/785/6532051.

Ji:2022:DVE

[JRDW22] Rubao Ji, Jeffrey A. Runge, Cabell S. Davis, and Peter H. Wiebe. Drivers of variability of Calanus finmarchicus in the Gulf of Maine: roles of internal production and external exchange. ICES Journal of Marine Science, 79(3):775– 784, April 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/3/775/6350837.

Juarez-Ruiz:2022:GSP

[JRPHM^{+}22] Ariadna Juárez-Ruiz, Mario A. Pardo, Julio C. Hernández-Montoya, Fernando R. Elorriaga-Verplancken, María de los Ángeles Milanés-Salinas, Tenaya Norris, Emilio Beier, and Gisela Heckel. Guadalupe fur seal pup production predicted from annual variations of sea surface temperature in the southern California Current Ecosystem. ICES Journal of Marine Science, 79(5):1637–1648, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1637/6604493.

Jamieson:2021:FLD

[JSLC21] Alan J. Jamieson, Glenn Singleman, Thomas D. Linley, and Susan Casey. Fear and loathing of the deep ocean: why don’t people care about the deep sea? ICES Journal of Marine Science, 78(3):797–809, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/

797/6042988. See reply [AAHT22].

Jensen:2022:MMH

[JSN^{+}22] J. L. A. Jensen, J. F. Strøm, A. Nikolopoulos, R. Primicerio, J. Skardhamar, B. J. Atencio, J. E. T. Strand, P. A. Bjørn, and T. Bøhn. Micro- and macro-habitat selection of Atlantic salmon, Salmo salar, post-smolts in relation to marine environmental cues. ICES Journal of Marine Science, 79(4): 1394–1407, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/1394/6571570.

James:2022:WSN

[JSR^{+}22] W. Ryan James, Rolando O. Santos, Jonathan R. Rodemann, Ryan J. Rezek, Zachary W. Fratto, Bradley T. Furman, Margaret O. Hall, Christopher R. Kelble, Jennifer S. Rehage, and James A. Nelson. Widespread seagrass die-off has no legacy effect on basal resource use of seagrass food webs in Florida Bay, USA. ICES Journal of Marine Science, 79(6):1831–1842, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1831/6611676.

Jiang:2021:EEP

[JWR^{+}21] Weiwei Jiang, Xiaoqin Wang, Samuel P. S. Rastrick, Junwei Wang, Yitao Zhang, Øivind Strand, Jianguang Fang, and Zengjie Jiang. Effects of elevated pCO_{2} on the physiological energetics of Pacific oyster, Crassostrea gigas. ICES Journal of Marine Science, 78(7):2579–2590, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2579/6332555.

Jang:2021:EES

[JYK^{+}21] Ho Geun Jang, Satoshi Yamazaki, Shoichi Kiyama, Keisaku Higashida, and Dugald Tinch. Economic effects of sea surface temperature, aging population, and market distance on a small-scale fishery. ICES Journal of Marine Science, 78(3): 1038–1048, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/3/1038/6120382.

Kai:2020:NAE

[Kai20] Mikihiko Kai. Numerical approach for evaluating impacts of biological uncertainties on estimates of stock–recruitment relationships in elasmobranchs: example of the North Pacific shortfin mako. ICES Journal of Marine Science, 77(1): 200–215, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/1/200/5614896.

Khodabandeloo:2021:MFS

[KAK^{+}21] Babak Khodabandeloo, Mette Dalgaard Agersted, Thor A. Klevjer, Geir Pedersen, and Webjørn Melle. Mesopelagic flesh shear viscosity estimation from in situ broadband backscattering measurements by a viscous–elastic model inversion. ICES Journal of Marine Science, 78(9):3147–3161, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3147/6378058.

Kallo:2022:VSB

[KBBGA22] Kristi Källo, Henrik Baktoft, Kim Birnie-Gauvin, and Kim Aarestrup. Variability in straying behaviour among repeat spawning anadromous brown trout (Salmo trutta) followed over several years. ICES Journal of Marine Science, 79(9):2453–2460, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
9/2453/6764497.

Kallo:2020:GCP

[KBJA20] Kristi Källo, Henrik Baktoft, Niels Jepsen, and Kim Aarestrup. Great cormorant (Phalacrocorax carbo sinensis) predation on juvenile down-migrating trout (Salmo trutta) in a lowland stream. ICES Journal of Marine Science, 77(2):721– 729, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/721/5652233.

Kallo:2022:HPS

[KBK^{+}22] Kristi Källo, Henrik Baktoft, Martin Lykke Kristensen, Kim Birnie-Gauvin, and Kim Aarestrup. High prevalence of straying in a wild brown trout (Salmo trutta) population in a fjord system. ICES Journal of Marine Science, 79(5): 1539–1547, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/5/1539/6582865.

Kerr:2022:CPD

[KBW22] L. Kerr, M. Barajas, and J. Wiedenmann. Coherence and potential drivers of stock assessment uncertainty in Northeast US groundfish stocks. ICES Journal of Marine Science, 79(8):2217–2230, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2217/6713983.

Killeen:2022:EMH

[KDS^{+}22] Helen Killeen, Jeffrey Dorman, William Sydeman, Connor Dibble, and Steven Morgan. Effects of a marine heatwave on adult body length of three numerically dominant krill species in the California Current Ecosystem. ICES Journal of Marine Science, 79(3):761–774, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/761/6425095.

Kressler:2023:FSE

[KDS23] Molly M. Kressler, Sasha R. X. Dall, and Richard B. Sherley. A framework for studying ecological energy in the contemporary marine environment. ICES Journal of Marine Science, 80(6):1580–1593, August 2023.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1580/7190725.

Kenchington:2021:BCB

[Ken21]  Trevor J. Kenchington.  Baranov’s contributions to

the Beverton–Holt model. ICES Journal of Marine Science, 78(6):2166–2172, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2166/6169452.

Kess:2021:PSV

[KEW^{+}21] Tony Kess, Anthony L. Einfeldt, Brendan Wringe, Sarah J. Lehnert, Kara K. S. Layton, Meghan C. McBride, Dominique Robert, Jonathan Fisher, Arnault Le Bris, Cornelia den Heyer, Nancy Shackell, Daniel E. Ruzzante, Paul Bentzen, and Ian R. Bradbury. A putative structural variant and environmental variation associated with genomic divergence across the Northwest Atlantic in Atlantic halibut. ICES Journal of Marine Science, 78(7):2371–2384, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/7/2371/6333511.

Kluger:2021:TOB

[KF21] Lotta Clara Kluger and Ramón Filgueira. Thinking outside the box: embracing social complexity in aquaculture carrying capacity estimations. ICES Journal of Marine Science, 78(1):435–442, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/435/5854844.

Kluger:2023:AEI

[KGG^{+}23] L. C. Kluger, M. Garteizgogeascoa, I. E. Gonzales, L. A. Odar, M. Flitner, and G. Damonte. An analysis of the early impacts of COVID-19 on Peruvian fisheries and mariculture. ICES Journal of Marine Science, 80(9):2280–2294, November 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/9/2280/7287043.

Kristofersson:2021:FAG

[KGV21] Dadi Kristofersson, Stefan Gunnlaugsson, and Hreidar Valtysson. Factors affecting greenhouse gas emissions in fisheries: evidence from Iceland’s demersal fisheries. ICES Journal of Marine Science, 78(7):2385–2394, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/7/2385/6294918.

Kuparinen:2021:CWP

[KH21] Anna Kuparinen and Jeffrey A. Hutchings. Corrigendum to: When phenotypes fail to illuminate underlying genetic processes in fish and fisheries science. ICES Journal of Marine Science, 78(4):1554, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1554/6298534.

Kearney:2022:SWW

[KH22] Bob Kearney and Ray Hilborn. Solutions to world-wide fisheries problems are mostly local or regional. ICES Journal of Marine Science, 79(4):997–1004, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/997/6545039.

Koenigstein:2020:GBE

[KHHD20] Stefan Koenigstein, Lisa-Henrike Hentschel, Lena Christin Heel, and Catherine Drinkorn. A game-based education approach for sustainable ocean development. ICES Journal of Marine Science, 77(5):1629–1638, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1629/5815628.

Kjesbu:2021:LJH

[KHSS21] Olav Sigurd Kjesbu, Jennifer Hubbard, Iain Suthers, and Vera Schwach. The legacy of Johan Hjort: challenges and critical periods — past, present, and future. ICES Journal of Marine Science, 78(2):621–630, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/2/
621/6047267.

Koenigstein:2022:PPP

[KJP^{+}22] Stefan Koenigstein, Michael G. Jacox, Mercedes Pozo Buil, Jerome Fiechter, Barbara A. Muhling, Stephanie Brodie, Peter T. Kuriyama, Toby D. Auth, Elliott L. Hazen, Steven J. Bograd, and Desiree Tommasi. Population projections of Pacific sardine driven by ocean warming and changing food availability in the California Current. ICES Journal of Marine Science, 79(9):2510–2523, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/9/2510/6779722.

Knotek:2020:UPS

[KKS^{+}20] Ryan Knotek, Jeff Kneebone, James Sulikowski, Tobey Curtis, Joseph Jurek, and John Mandelman. Utilization of popup satellite archival transmitting tags to evaluate thorny skate (Amblyraja radiata) discard mortality in the Gulf of Maine groundfish bottom trawl fishery. ICES Journal of Marine Science, 77(1):256–266, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/1/
256/5599855.

Klunder:2020:IASa

[KLF^{+}20a] Lise Klunder, Marc S. S. Lavaleye, Amalia Filippidi, Judith D. L. van Bleijswijk, Gert-Jan Reichart, Henk W. van der Veer, Gerard C. A. Duineveld, and Furu Mienis. Impact of an artificial structure on the benthic community composition in the southern North Sea: assessed by a morphological and molecular approach. ICES Journal of Marine Science, 77(3): 1167–1177, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/3/1167/5096671.

Klunder:2020:IASb

[KLF^{+}20b] Lise Klunder, Marc S. S. Lavaleye, Amalia Filippidi, Judith D. L. van Bleijswijk, Gert-Jan Reichart, Henk W. van der Veer, Gerard C. A. Duineveld, and Furu Mienis. Impact of an artificial structure on the benthic community composition in the southern North Sea: assessed by a morphological and molecular approach. ICES Journal of Marine Science, 77(3): 1247, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/3/1247/5702176.

Karp:2023:IUM

[KLG^{+}23] Melissa A. Karp, Jason S. Link, Max Grezlik, Steve Cadrin, Gavin Fay, Patrick Lynch, Howard Townsend, Richard D. Methot, Grant D. Adams, Kristan Blackhart, Caren Barceló, Andre Buchheister, Matthew Cieri, David Chagaris, Villy Christensen, J. Kevin Craig, Jonathan Cummings, Matthew D. Damiano, Mark Dickey-Collas, Bjarki ór Elvarsson, Sarah Gaichas, Melissa A. Haltuch, Janne B. Haugen, Daniel Howell, Isaac C. Kaplan, Willem Klajbor, Scott I. Large, Michelle Masi, Jason McNamee, Brandon Muffley, Sarah Murray, Éva Plagányi, David Reid, Anna Rindorf, Skyler R. Sagarese, Amy M. Schueller, Robert Thorpe, James T. Thorson, Maciej T. Tomczak, Vanessa Trijoulet, and Rudi Voss. Increasing the uptake of multispecies models in fisheries management. ICES Journal of Marine Science, 80(2):243–257, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/2/243/6997897.

Koutsidi:2024:QII

[KLP^{+}24] Martha Koutsidi, Alexis Lazaris, Panagiota Peristeraki, George Tserpes, and Evangelos Tzanatos. Quantification of intraspecific and interspecific competition in fish species of the Aegean Sea. ICES Journal of Marine Science, 81(2): 334–347, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/81/2/334/7491223.

Kulagin:2021:PDB

[KLSV21] D. N. Kulagin, A. A. Lunina, U. V. Simakova, and A. L. Vereshchaka. Progressing diversification and biogeography of the mesopelagic Nematoscelis (Crustacea: Euphausiacea) in the Atlantic. ICES Journal of Marine Science, 78(9):3457– 3463, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/78/9/3457/6145328.

Kell:2022:ESL

[KMG22] Laurence T. Kell, Cóilín Minto, and Hans D. Gerritsen. Evaluation of the skill of length-based indicators to identify stock status and trends. ICES Journal of Marine Science, 79(4): 1202–1216, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/1202/6554688.

Karlsen:2021:ENN

[KMK21a] Junita Diana Karlsen, Valentina Melli, and Ludvig Ahm Krag. Exploring new netting material for fishing: the low light level of a luminous netting negatively influences species separation in trawls. ICES Journal of Marine Science, 78(8):2818–2829, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2818/6359259.

Kase:2021:MAS

[KMK^{+}21b] Laura Käse, Katja Metfies, Alexandra C. Kraberg, Stefan Neuhaus, Cédric L. Meunier, Karen H. Wiltshire, and Maarten Boersma. Metabarcoding analysis suggests that flexible food web interactions in the eukaryotic plankton community are more common than specific predator–prey relationships at Helgoland Roads, North Sea. ICES Journal of Marine Science, 78(9):3372–3386, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3372/6219380.

King:2020:HGS

[KMS^{+}20a] Nathan G. King, Niall J. McKeown, Dan A. Smale, Sunny Bradbury, Thomas Stamp, Alexander Jüterbock, Hrönn Egilsdóttir, Emily A. Groves, and Pippa J. Moore. Hierarchical genetic structuring in the cool boreal kelp, Laminaria digitata: implications for conservation and management. ICES Journal of Marine Science, 77(5):1906–1913, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1906/5847977.

Kleiven:2020:NFB

[KMS20b] Alf Ring Kleiven, Even Moland, and U. Rashid Sumaila. No fear of bankruptcy: the innate self-subsidizing forces in recreational fishing. ICES Journal of Marine Science, 77(6):2304– 2307, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/6/2304/5530702.

Kiaer:2021:FAS

[KNP21] Christian Kiaer, Stefan Neuenfeldt, and Mark R. Payne. A framework for assessing the skill and value of operational recruitment forecasts. ICES Journal of Marine Science, 78(10):3581–3591, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
10/3581/6409471.

Kleisner:2022:IPA

[KOB^{+}22] Kristin M. Kleisner, Elena Ojea, Willow Battista, Merrick Burden, Erica Cunningham, Rod Fujita, Kendra Karr, Samuel Amorós, Julia Mason, Doug Rader, Nicolas Rovegno, and Alice Thomas-Smyth. Identifying policy approaches to build social–ecological resilience in marine fisheries with differing capacities and contexts. ICES Journal of Marine Science, 79(2):552–572, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/2/552/6303758.

Kvaavik:2021:NIT

[KÓP^{+}21] Cecilia Kvaavik, Gudmundur J. Óskarsson, Hildur Pétursdóttir, Gudrun Marteinsdóttir, and Anna Kristín Daníelsdóttir. New insight into trophic niche partitioning and diet of mackerel (Scomber scombrus) and herring (Clupea harengus) in Icelandic waters. ICES Journal of Marine Science, 78(4): 1485–1499, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1485/6184589.

Kopsel:2023:SSS

[Köp23] Vera Köpsel. A social scientific self-reflection of codeveloping an acoustic stock sampling method in the Bay of Biscay. ICES Journal of Marine Science, 80(2):378–389, March 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/2/378/6653506.

Koslow:2021:DCN

[Kos21] Julian Anthony Koslow. Discoveries and challenges: navigating the oceanic great unknown. ICES Journal of Marine Science, 78(7):2342–2350, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2342/6295518.

Krumsick:2024:SBE

[KP24] Kyle J. Krumsick and Eric J. Pedersen. Simultaneous Bayesian estimation of size-specific catchability and size spectrum parameters from trawl data. ICES Journal of Marine Science, 81(1):145–161, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/81/
1/145/7459507.

Kenyon:2022:IIA

[KPM^{+}22] Susan Kenyon, Martin Pastoors, Steven Mackinson, Thomas Cornulier, and C. Tara Marshall. Intra- and inter-annual variability in the fat content of Atlantic herring (Clupea harengus) as revealed by routine industry monitoring. ICES Journal of Marine Science, 79(1):88–99, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/88/6459961.

Kovac:2020:SRN

[KPS20] Žarko Kovač, Trevor Platt, and Shubha Sathyendranath. Stability and resilience in a nutrient-phytoplankton marine ecosystem model. ICES Journal of Marine Science, 77(4): 1556–1572, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/4/1556/5864492.

Kovac:2021:SML

[KPS21] Žarko Kovač, Trevor Platt, and Shubha Sathyendranath. Sverdrup meets Lambert: analytical solution for Sverdrup’s critical depth. ICES Journal of Marine Science, 78(4):1398– 1408, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1398/6166287.

Kvile:2022:EEE

[KPS22] Kristina Ø Kvile, Irina P. Prokopchuk, and Leif C. Stige. Environmental effects on Calanus finmarchicus abundance and depth distribution in the Barents Sea. ICES Journal of Marine Science, 79(3):815–828, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/815/6354504.

Kraak:2023:EET

[Kra23] Sarah B. M. Kraak. Evolution of EU technical measures for the avoidance of unwanted catch in the light of scientific evaluation and advice from the STECF; the good, the bad, and the ugly. ICES Journal of Marine Science, 80(3):635– 646, April 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/3/635/6550344.

Kanaji:2023:HMA

[KSHO23] Y. Kanaji, H. Sasaki, T. Hakamada, and H. Okamura. Hierarchical modelling approach to estimate the abundance of data-limited cetacean species and its application to fisherytargeted and rarely seen delphinid species off Japan. ICES Journal of Marine Science, 80(6):1643–1657, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1643/7192207.

Kell:2021:VSA

[KSK^{+}21] Laurence T. Kell, Rishi Sharma, Toshihide Kitakado, Henning Winker, Iago Mosqueira, Massimiliano Cardinale, and Dan Fu. Validation of stock assessment methods: is it me or my model talking? ICES Journal of Marine Science, 78(6):2244–2255, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2244/6296435.

Kjesbu:2023:HCH

[KTA^{+}23] O. S. Kjesbu, I. Tiedemann, M. Alix, A. Thorsen, and S. Sundby. Half a century of high-latitude fisheries oceanography research on the “recruitment problem” in Northeast Arctic cod (Gadus morhua). ICES Journal of Marine Science, 80(5):1179–1201, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1179/7176247.

Kerschbaumer:2020:ADH

[KTK20] Philip Kerschbaumer, Michael Tritthart, and Hubert Keckeis. Abundance, distribution, and habitat use of fishes in a large river (Danube, Austria): mobile, horizontal hydroacoustic surveys vs. a standard fishing method. ICES Journal of Marine Science, 77(5):1966–1978, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/5/1966/5862934.

Kamimura:2021:IIS

[KTY^{+}21] Yasuhiro Kamimura, Makoto Taga, Ryuji Yukami, Chikako Watanabe, and Sho Furuichi. Intra- and inter-specific density dependence of body condition, growth, and habitat temperature in chub mackerel (Scomber japonicus). ICES Journal of Marine Science, 78(9):3254–3264, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3254/6380055.

Kuepfer:2023:PSF

[KVS^{+}23] Amanda Kuepfer, Stephen C. Votier, Richard B. Sherley, Francesco Ventura, Rafael Matias, Orea Anderson, Paul Brickle, Alexander Arkhipkin, and Paulo Catry. Preyswitching to fishery discards does not compensate for poor natural foraging conditions in breeding albatross. ICES Journal of Marine Science, 80(9):2414–2426, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/9/2414/6582866.

Kulatska:2021:SSC

[KWEB21] Nataliia Kulatska, Pamela J. Woods, Bjarki Tór Elvarsson, and Valerio Bartolino. Size-selective competition between cod and pelagic fisheries for prey. ICES Journal of Marine Science, 78(5):1872–1886, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1872/6290973.

King:2021:CCA

[KWS^{+}21] Nathan G. King, Sophie B. Wilmes, David Smyth, Jonathan Tinker, Peter E. Robins, Jamie Thorpe, Laurence Jones, and Shelagh K. Malham. Climate change accelerates range expansion of the invasive non-native species, the Pacific oyster, Crassostrea gigas. ICES Journal of Marine Science, 78(1): 70–81, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/1/70/6032772.

Lennox:2021:QSA

[LAA^{+}21] Robert J. Lennox, Carlos M. Alexandre, Pedro R. Almeida, Kevin M. Bailey, Bjørn T. Barlaup, Kristin Bøe, André Breukelaar, Jaakko Erkinaro, Torbjørn Forseth, Sven-Erik Gabrielsen, Edmund Halfyard, Erlend M. Hanssen, Sten Karlsson, Stephanie Koch, Anders Koed, Roy M. Langåker, Håvard Lo, Martyn C. Lucas, Shad Mahlum, Charles Perrier, Ulrich Pulg, Timothy Sheehan, Helge Skoglund, Martin Svenning, Eva B. Thorstad, Gaute Velle, Frederick G. Whoriskey, and Knut Wiik Vollset. The quest for successful Atlantic salmon restoration: perspectives, priorities, and maxims. ICES Journal of Marine Science, 78(10):3479–3497, December 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/10/3479/6425092.

Laptikhovsky:2022:STV

[LAB^{+}22] Vladimir Laptikhovsky, A. Louise Allcock, Leigh Barnwall, Christopher Barrett, Gavan Cooke, Christian Drerup, Christopher Firmin, Sophie Lozach, Eleanor MacLeod, Daniel Oesterwind, Michael Petroni, Jean-Paul Robin, Edel Sheerin, Anne-Marie Power, and Graham J. Pierce. Spatial and temporal variability of spawning and nursery grounds of Loligo forbesii and Loligo vulgaris squids in ecoregions of Celtic Seas and Greater North Sea. ICES Journal of Marine Science, 79(6):1918–1930, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1918/6633838.

Lindmark:2023:EDS

[LAGC23] M. Lindmark, S. C. Anderson, M. Gogina, and M. Casini. Evaluating drivers of spatiotemporal variability in individual condition of a bottom-associated marine fish, Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 80(5): 1539–1550, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/5/1539/7186976.

Lavers:2023:LTD

[LB23] J. L. Lavers and A. L. Bond. Long-term decline in fledging body condition of flesh-footed shearwaters (Ardenna carneipes). ICES Journal of Marine Science, 80(4):1113– 1119, May 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/4/1113/7083385.

Long:2021:DSB

[LBH^{+}21] Stephen Long, Martin E. Blicher, Nanette Hammeken Arboe, Mona Fuhrmann, Michael Darling, Kirsty M. Kemp, Rasmus Nygaard, Karl Zinglersen, and Chris Yesson. Deepsea benthic habitats and the impacts of trawling on them in the offshore Greenland halibut fishery, Davis Strait, west Greenland. ICES Journal of Marine Science, 78(8):2724– 2744, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/8/2724/6357213.

Lloret:2020:RSS

[LBPC^{+}20] J. Lloret, S. Biton-Porsmoguer, A. Carreño, A. Di Franco, R. Sahyoun, P. Melià, J. Claudet, C. Sève, A. Ligas, M. Belharet, A. Calò, P. Carbonara, M. Coll, X. Corrales, G. Lembo, P. Sartor, I. Bitetto, D. Vilas, C. Piroddi, G. Prato, E. Charbonnel, O. Bretton, V. Hartmann, L. Prats, and T. Font. Recreational and small-scale fisheries may pose a threat to vulnerable species in coastal and offshore waters of the western Mediterranean. ICES Journal of Marine Science, 77(6):2255–2264, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2255/5486184.

Leon:2020:RBS

[LBW^{+}20] Pablo León, Nina Bednaršek, Pam Walsham, Kathryn Cook, Susan E. Hartman, Deborah Wall-Palmer, Jennifer Hindson, Kevin Mackenzie, Lynda Webster, and Eileen Bresnan. Relationship between shell integrity of pelagic gastropods and carbonate chemistry parameters at a Scottish Coastal Observatory monitoring site. ICES Journal of Marine Science, 77(1):436–450, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/436/5584407.

Levin:2022:DGT

[LCG^{+}22] Lisa A. Levin, Laura Cimoli, Kristina Gjerde, Harriet Harden-Davies, Patrick Heimbach, Diana LaScala-Gruenewald, Maria Pachiadaki, Helen R. Pillar, Leslie M. Smith, Karen Stocks, Justin E. Stopa, and Dawn J. Wright. Designing, generating, and translating deep-ocean observations for and with international policy makers. ICES Journal of Marine Science, 79(7):1992–1995, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/1992/6661567.

laCruz:2023:CBS

[lCRGCCA23] Andrés De la Cruz, Carlos Rodríguez-García, Remedios Cabrera-Castro, and Gonzalo M. Arroyo. Correlation between seabirds and fisheries varies by species at fine-scale pattern. ICES Journal of Marine Science, 80(9):2427–2440, November 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/9/2427/6758115.

Lauffenburger:2023:MPA

[LDW23] N. Lauffenburger, A. De Robertis, and K. Williams. Mining previous acoustic surveys to improve walleye pollock (Gadus chalcogrammus) target strength estimates. ICES Journal of Marine Science, 80(6):1683–1696, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1683/7201931.

Legendre:2021:PWB

[Leg21] Louis Legendre. The pleasure of writing, being published, appealing to readers, and contributing to the advancement of knowledge. ICES Journal of Marine Science, 78(6):1943– 1955, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/6/1943/6294919.

Lehtiniemi:2021:EOS

[LFH^{+}21a] Maiju Lehtiniemi, Elaine Fileman, Heidi Hällfors, Harri Kuosa, Sirpa Lehtinen, Inga Lips, Outi Setälä, Sanna Suikkanen, Jarno Tuimala, and Claire Widdicombe. Erratum to: Optimising sampling frequency for monitoring heterotrophic protists in a marine ecosystem. ICES Journal of Marine Science, 78(9):3477, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3477/6357218. See [LFH^{+}21b].

Lehtiniemi:2021:OSF

[LFH^{+}21b] Maiju Lehtiniemi, Elaine Fileman, Heidi Hällfors, Harri Kuosa, Sirpa Lehtinen, Inga Lips, Outi Setälä, Sanna Suikkanen, Jarno Tuimala, and Claire Widdicombe. Optimising sampling frequency for monitoring heterotrophic protists in a marine ecosystem. ICES Journal of Marine Science, 78(3): 1–12, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). See erratum [LFH^{+}21a].

Lehtiniemi:2022:OSF

[LFH^{+}22] Maiju Lehtiniemi, Elaine Fileman, Heidi Hällfors, Harri Kuosa, Sirpa Lehtinen, Inga Lips, Outi Setälä, Sanna Suikkanen, Jarno Tuimala, and Claire Widdicombe. Optimising sampling frequency for monitoring heterotrophic protists in a marine ecosystem. ICES Journal of Marine Science, 79 (3):925–936, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/3/925/6324792.

Lacey:2020:EAS

[LH20] Nichola C. Lacey and Peter Hayes. Epifauna associated with subsea pipelines in the North Sea. ICES Journal of Marine Science, 77(3):1137–1147, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1137/5315644.

Lonthair:2020:ELS

[LHE20] Joshua Lonthair, Pung-Pung Hwang, and Andrew J. Esbaugh. The early life stages of the orange-spotted grouper, Epinephelus coioides, exhibit robustness to hypercapnia. ICES Journal of Marine Science, 77(3):1066–1074, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/3/1066/5762494.

Long:2021:IVG

[LHL^{+}21] A. P. Long, D. Haberlin, O. Lyashevska, D. Brophy, Brendan O’ Hea, C. O’Donnell, R. G. Scarrott, C. Lawton, and T. K. Doyle. Interannual variability of gelatinous mesozooplankton in a temperate shelf sea: greater abundance coincides with cooler sea surface temperatures. ICES Journal of Marine Science, 78(4):1372–1385, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1372/6168604.

Liu:2022:GMM

[LHN^{+}22] Xiaozi Liu, Cecilie Hansen, Kjell Nedreaas, Hans Hagen Stockhausen, and Mikko Heino. Gaussian mixture models reveal highly diverse targeting tactics in a coastal fishing fleet. ICES Journal of Marine Science, 79(10):2709–2727, December 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/10/2709/6839999.

Link:2021:EEO

[Lin21] Jason S. Link. Evidence of ecosystem overfishing in U.S. large marine ecosystems. ICES Journal of Marine Science, 78(9):3176–3201, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3176/6381244.

Lajus:2021:MFT

[LIR^{+}21] Dmitry Lajus, Tatiana Ivanova, Elena Rybkina, Julia Lajus, and Mikhail Ivanov. Multidecadal fluctuations of threespine stickleback in the White Sea and their correlation with temperature. ICES Journal of Marine Science, 78(2):653–665, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/2/653/6000680.

Link:2021:PBR

[LKL^{+}21] J. S. Link, M. A. Karp, P. Lynch, W. E. Morrison, and J. Peterson. Proposed business rules to incorporate climate-induced changes in fisheries management. ICES Journal of Marine Science, 78(10):3562–3580, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3562/6425783.

Li:2023:CEM

[LL23] C. Li and H. Liu. Comparative ecosystem modelling of dynamics and stability of subtropical estuaries under external perturbations in the Gulf of Mexico. ICES Journal of Marine Science, 80(5):1303–1318, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1303/7109675.

Lira:2022:VMR

[LLALF22] Alex Souza Lira, François Le Loc’h, Humber Agrelli Andrade, and Flávia Lucena-Frédou. Vulnerability of marine resources affected by a small-scale tropical shrimp fishery in Northeast Brazil. ICES Journal of Marine Science, 79(3): 633–647, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/3/633/6526682.

Lee:2020:LHC

[LLLS20] Qi Lee, Alice Lee, Zunlei Liu, and Cody S. Szuwalski. Life history changes and fisheries assessment performance: a case study for small yellow croaker. ICES Journal of Marine Science, 77(2):645–654, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/2/645/5680031.

Lindstedt:2022:RFP

[LLPdL22] Daniel G. Lindstedt, Tim Langlois, Jane Prince, and Simon de Lestang. Recreational fishing pressure impacts the density and behaviour of the western rock lobster (Panulirus cygnus, George): evidence from small, no-take marine reserves. ICES Journal of Marine Science, 79(9):2413–2421, November 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/9/2413/6754312.

Lerner:2022:SIR

[LMH22] Jacob E. Lerner, Christian Marchese, and Brian P. V. Hunt. Stable isotopes reveal that bottom-up omnivory drives food chain length and trophic position in eutrophic coastal ecosystems. ICES Journal of Marine Science, 79(8):2311–2323, October 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/8/2311/6748271.

Lopez:2023:STC

[LMvEM23] N. A. López, R. B. McAuley, S. van Elden, and J. J. Meeuwig. Spatial and temporal characterization of a recurrent scalloped hammerhead shark Sphyrna lewini aggregation using drones. ICES Journal of Marine Science, 80(9):2356–2367, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
9/2356/7301220.

Luo:2022:SAL

[LMW^{+}22] Jiaxin Luo, Raphaël R. McDonald, Brendan F. Wringe, Cornelia E. den Heyer, Bruce Smith, Yuan Yan, and Joanna Mills Flemming. A spatial analysis of longline survey data for improved indices of Atlantic halibut abundance. ICES Journal of Marine Science, 79(6):1954–1964, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1954/6646719.

Larsson:2021:PPP

[LNW21] Ulf Larsson, Svante Nyberg, and Jakob Walve. Phytoplankton primary production: ^{14}C-in situ and ^{1}4Cincubator methods compared. ICES Journal of Marine Science, 78(10):3592–3602, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
10/3592/6407845.

Ljungberg:2020:CBC

[LOÖK20] Peter Ljungberg, Maria Ovegård, Kristin Öhman, and Sara Königson. Correlation between catch method, condition, and diet patterns in Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 77(1):267–277, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/1/
267/5565270.

Love:2020:YCM

[Lov20] Milton S. Love. A 45-year career in marine science — better than a sharp stick in the eye. ICES Journal of Marine Science, 77(2):469–475, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/2/469/5728770.

Lagarde:2022:PED

[LPA^{+}22] Raphaël Lagarde, Jason Peyre, Elsa Amilhat, François Bourrin, François Prellwitz, Alexandre Perrault, Gaël Simon, and Elisabeth Faliex. Phenology and environmental drivers of glass eel entrance in a Mediterranean lagoon. ICES Journal of Marine Science, 79(7):2107–2116, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2107/6682847.

Liang:2023:RDF

[LPC^{+}23] Cui Liang, Daniel Pauly, Villy Christensen, Weiwei Xian, and Carl Walters. Recruitment-driven fish production in two regions where fish biomass has drastically declined. ICES Journal of Marine Science, 80(4):1043–1051, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1043/7075373.

Lawson:2020:ERC

[LPG^{+}20] Julia M. Lawson, Riley A. Pollom, Cat A. Gordon, Joanna Barker, Eva K. M. Meyers, Heike Zidowitz, Jim R. Ellis, Álex Bartolí, Gabriel Morey, Sarah L. Fowler, David Jiménez Alvarado, Sonja V. Fordham, Rowland Sharp, Ali R. Hood, and Nicholas K. Dulvy. Extinction risk and conservation of critically endangered angel sharks in the Eastern Atlantic and Mediterranean Sea. ICES Journal of Marine Science, 77 (1):12–29, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/1/12/5675589.

Li:2022:TFA

[LQC^{+}22] Jiajun Li, Yongsong Qiu, Yancong Cai, Kui Zhang, Peng Zhang, Zhiyou Jing, Qiaer Wu, Shengwei Ma, Huaxue Liu, and Zuozhi Chen. Trend in fishing activity in the open South China Sea estimated from remote sensing of the lights used at night by fishing vessels. ICES Journal of Marine Science, 79(1):230–241, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/230/6485077.

Levine:2023:TDS

[LRGW23] R. M. Levine, A. De Robertis, D. Grünbaum, and C. D. Wilson. Transport-driven seasonal abundance of pelagic fishes in the Chukchi Sea observed with seafloor-mounted echosounders. ICES Journal of Marine Science, 80(4):987– 1001, May 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/4/987/7058021.

Lucca:2021:IVS

[LRHW21] Brandyn M. Lucca, Patrick H. Ressler, H. Rodger Harvey, and Joseph D. Warren. Individual variability in sub-Arctic krill material properties, lipid composition, and other scattering model inputs affect acoustic estimates of their population. ICES Journal of Marine Science, 78(4):1470–1484, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1470/6179526.

Lew:2020:MEC

[LS20a] Daniel K. Lew and Chang K. Seung. Measuring economic contributions of the marine recreational charter fishing sector using a resampling approach. ICES Journal of Marine Science, 77(6):2285–2294, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2285/5372749.

Lunt:2020:TAE

[LS20b] Jessica Lunt and Delbert L. Smee. Turbidity alters estuarine biodiversity and species composition. ICES Journal of Marine Science, 77(1):379–387, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/1/
379/5620392.

Li:2020:DMI

[LSD^{+}20] Qiong Li, Xin Sun, Junyu Dong, Shuqun Song, Tongtong Zhang, Dan Liu, Han Zhang, and Shuai Han. Developing a microscopic image dataset in support of intelligent phytoplankton detection using deep learning. ICES Journal of Marine Science, 77(4):1427–1439, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1427/5572160.

Lynch:2020:CCS

[LSOD^{+}20] T. P. Lynch, C. B. Smallwood, F. A. Ochwada-Doyle, J. Lyle, J. Williams, K. L. Ryan, C. Devine, B. Gibson, and A. Jordan. A cross continental scale comparison of Australian offshore recreational fisheries research and its applications to Marine Park and fisheries management. ICES Journal of Marine Science, 77(3):1190–1205, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1190/5513323.

Letschert:2021:UFN

[LSR^{+}21] Jonas Letschert, Nicole Stollberg, Henrike Rambo, Alexander Kempf, Jörg Berkenhagen, and Vanessa Stelzenmüller. The uncertain future of the Norway lobster fisheries in the North Sea calls for new management strategies. ICES Journal of Marine Science, 78(10):3639–3649, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3639/6414428.

Lu:2020:ISH

[LTK20] Yi-Chin Lu, Chen Tung, and Yan-Fu Kuo. Identifying the species of harvested tuna and billfish using deep convolutional neural networks. ICES Journal of Marine Science, 77(4):1318–1329, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1318/5509966.

Link:2023:WWN

[LTMG23] J. S. Link, S. Thur, G. Matlock, and M. Grasso. Why we need weather forecast analogues for marine ecosystems. ICES Journal of Marine Science, 80(8):2087–2098, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/8/2087/7280043.

Lekanda:2021:EBT

[LTN21] Aitor Lekanda, Nick Tolimieri, and Adriana Nogueira. The effects of bottom temperature and fishing on the structure and composition of an exploited demersal fish assemblage in West Greenland. ICES Journal of Marine Science, 78(8):2895–2906, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2895/6359258.

Lehuta:2023:CIL

[LV23] S. Lehuta and Y. Vermard. Contrasting impacts of the landing obligation at fleet scale: impact assessment of mitigation scenarios in the Eastern English Channel. ICES Journal of Marine Science, 80(3):518–531, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/518/6724459.

Lindegren:2022:SSA

[LvDM^{+}22] Martin Lindegren, Mikael van Deurs, Aurore Maureaud, James T. Thorson, and Dorte Bekkevold. A spatial statistical approach for identifying population structuring of marine fish species: European sprat as a case study. ICES Journal of Marine Science, 79(2):423–434, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/423/6520723.

LePape:2020:UPS

[LVG^{+}20] Olivier Le Pape, Youen Vermard, Jérome Guitton, Elliot J. Brown, Karen E. van de Wolfshaar, Romuald N. Lipcius, Josianne G. Støttrup, and Kenneth A. Rose. The use and performance of survey-based pre-recruit abundance indices for possible inclusion in stock assessments of coastal-dependent species. ICES Journal of Marine Science, 77(5):1953–1965, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1953/5861726.

Logan:2022:HRP

[LVL^{+}22] Ryan K. Logan, Jeremy J. Vaudo, Christopher G. Lowe, Bradley M. Wetherbee, and Mahmood S. Shivji. Highresolution post-release behaviour and recovery periods of two highly prized recreational sportfish: the blue marlin and sailfish. ICES Journal of Marine Science, 79(7):2055–2068, September 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/7/2055/6660731.

Long:2023:RMG

[LVT^{+}23] A. P. Long, L. Vaughan, E. Tray, K. Thomas, N. Ó Maoiléidigh, R. Poole, D. Cotter, A. Doogan, and D. Brophy. Recent marine growth declines in wild and ranched Atlantic salmon Salmo salar from a western European catchment discovered using a 62-year time series. ICES Journal of Marine Science, 80(6):1697–1709, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1697/7205010.

Lewin:2021:CSS

[LWH^{+}21] Wolf-Christian Lewin, Marc Simon Weltersbach, Kevin Haase, Carsten Riepe, Christian Skov, Casper Gundelund, and Harry V. Strehlow. Comparing on-site and off-site survey data to investigate survey biases in recreational fisheries data. ICES Journal of Marine Science, 78(7):2528–2546, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/7/2528/6321555.

Li:2022:RAM

[LWL^{+}22] Daoliang Li, Qi Wang, Xin Li, Meilin Niu, He Wang, and Chunhong Liu. Recent advances of machine vision technology in fish classification. ICES Journal of Marine Science, 79(2):263–284, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/2/263/6511219.

Lin:2023:EIE

[LWZC23] H. Lin, J. Wang, J. Zhu, and X. Chen. Evaluating the impacts of environmental and fishery variability on the distribution of bigeye tuna in the Pacific Ocean. ICES Journal of Marine Science, 80(10):2642–2656, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/10/2642/7335585.

Li:2021:DFP

[LXBG21] Futian Li, Jiekai Xu, John Beardall, and Kunshan Gao. Diurnally fluctuating pCO_{2} enhances growth of a coastal strain of Emiliania huxleyi under future-projected ocean acidification conditions. ICES Journal of Marine Science, 78(4): 1301–1310, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1301/6158504.

Liu:2020:NME

[LZWH20] Chanjuan Liu, Shijie Zhou, You-Gan Wang, and Zhihua Hu. Natural mortality estimation using tree-based ensemble learning models. ICES Journal of Marine Science, 77(4): 1414–1426, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/4/1414/5854079.

Li:2021:PST

[LZZ^{+}21] Yali Li, Xinqing Zou, Shichun Zou, Pu Li, Ying Yang, and Junjie Wang. Pollution status and trophic transfer of polycyclic aromatic hydrocarbons in coral reef ecosystems of the South China Sea. ICES Journal of Marine Science, 78(6):2053–2064, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2053/6312006.

Murphy:2021:IPD

[MAC21] Hannah M. Murphy, Aaron T. Adamack, and Frédéric Cyr. Identifying possible drivers of the abrupt and persistent delay in capelin spawning timing following the 1991 stock collapse in Newfoundland, Canada. ICES Journal of Marine Science, 78(8):2709–2723, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2709/6356489.

Mackinson:2022:FRI

[Mac22] Steven Mackinson. The fall and rise of industry participation in fisheries science — a European story. ICES Journal of Marine Science, 79(4):1024–1033, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1024/6551737.

Martinez-Abrain:2022:FFE

[MACSR22] Alejandro Martínez-Abraín, Inés Castejón-Silvo, and Sergio Roiloa. Foreseeing the future of Posidonia oceanica meadows by accounting for the past evolution of the Mediterranean Sea. ICES Journal of Marine Science, 79(10):2597–2599, December 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/10/2597/6855639.

Mills:2023:CPK

[MAE^{+}23] Katherine E. Mills, Derek Armitage, Jacob G. Eurich, Kristin M. Kleisner, Gretta T. Pecl, and Kanae Tokunaga. Co-production of knowledge and strategies to support climate resilient fisheries. ICES Journal of Marine Science, 80 (2):358–361, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/2/358/6611665.

MacMillan:2022:STV

[MAI^{+}22] Isla MacMillan, Martin J. Attrill, Taha Imzilen, Christophe Lett, Simon Walmsley, Clarus Chu, and David M. Kaplan. Spatio-temporal variability in drifting Fish Aggregating Device (dFAD) beaching events in the Seychelles Archipelago. ICES Journal of Marine Science, 79(5):1687–1700, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1687/6591913.

Mangel:2021:SHP

[Man21] Marc Mangel. Sidney Holt on principles for the conservation of wild living resources, whaling in the Antarctic, and the Beverton–Holt stock–recruitment relationship. ICES Journal of Marine Science, 78(6):2211–2217, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/2211/6039465.

Maltby:2023:EFP

[MAT^{+}23] Katherine M. Maltby, Luis Acosta, Bryony Townhill, Julia Touza, Piran White, and Stephen C. Mangi. Exploring fishers’ perceptions of index insurance and coral reef health in the context of climate-driven changes in extreme events. ICES Journal of Marine Science, 80(8):2210–2221, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/8/2210/6516158.

Mullowney:2021:SMS

[MB21] Darrell R. J. Mullowney and Krista D. Baker. Size-at-maturity shift in a male-only fishery: factors affecting molttype outcomes in Newfoundland and Labrador snow crab (Chionoecetes opilio). ICES Journal of Marine Science, 78 (2):516–533, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/2/516/5974112.

Mahmood:2020:ADW

[MBA^{+}20] Ammar Mahmood, Mohammed Bennamoun, Senjian An, Ferdous Sohel, Farid Boussaid, Renae Hovey, and Gary Kendrick. Automatic detection of western rock lobster using synthetic data. ICES Journal of Marine Science, 77(4):1308– 1317, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/4/1308/5637822.

Mildenberger:2020:TVP

[MBP^{+}20] Tobias K. Mildenberger, Casper W. Berg, Martin W. Pedersen, Alexandros Kokkalis, and J. Rasmus Nielsen. Timevariant productivity in biomass dynamic models on seasonal and long-term scales. ICES Journal of Marine Science, 77(1):174–187, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/174/5572245.

McInturf:2023:UPD

[MBS^{+}23] A. G. McInturf, J. Bowman, J. M. Schulte, K. C. Newton, B. Vigil, M. Honig, S. Pelletier, N. Cox, O. Lester, M. Cantor, and T. K. Chapple. A unified paradigm for defining elasmobranch aggregations. ICES Journal of Marine Science, 80(6):1551–1566, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
6/1551/7210132.

Munch:2020:FAQ

[MBSR20] Stephan B. Munch, Antoine Brias, George Sugihara, and Tanya L. Rogers. Frequently asked questions about nonlinear dynamics and empirical dynamic modelling. ICES Journal of Marine Science, 77(4):1463–1479, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1463/5643857.

Martino:2022:SEO

[MCC^{+}22] Jasmin C. Martino, Ming-Tsung Chung, Chun-I Chiang, Chia-Hui Wang, Kotaro Shirai, and Zoë A. Doubleday. Systematic evaluation of oxygen isotopes in cephalopod statoliths as thermal proxies. ICES Journal of Marine Science, 79(6):1719–1729, August 2022.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1719/6608334.

McQueen:2020:RSS

[MCD^{+}20] Kate McQueen, Michele Casini, Bodo Dolk, Stefanie Haase, Jakob Hemmer-Hansen, Annelie Hilvarsson, Karin Hüssy, Monica Mion, Thomas Mohr, Krzysztof Radtke, Franziska Maria Schade, Norbert Schulz, and Uwe Krumme. Regional and stock-specific differences in contemporary growth of Baltic cod revealed through tag-recapture data. ICES Journal of Marine Science, 77(6):2078–2088, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2078/5880832.

Munnelly:2024:AAA

[MCH^{+}24] Ryan T Munnelly, Jose C Castillo, Nils Olav Handegard, Matthew E Kimball, Kevin M Boswell, and Guillaume Rieucau. Applications and analytical approaches using imaging sonar for quantifying behavioural interactions among aquatic organisms and their environment. ICES Journal of Marine Science, 81(2):207–251, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/2/207/7456026.

Molto:2022:MFF

[MCOÁ^{+}22] Vicenç Moltó, Ignacio A. Catalán, Andrés Ospina-Álvarez, Pilar Hernández, and Rubén H. Roa-Ureta. A multiannual five-fleet generalized depletion model for the stock assessment of the Mediterranean dolphinfish (Coryphaena hippurus) fishery. ICES Journal of Marine Science, 79(5):1481– 1496, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1481/6581598.

McKeown:2020:GPS

[MCS^{+}20] Niall J. McKeown, Piera Carpi, Joana F. Silva, Amy J. E. Healey, Paul W. Shaw, and Jeroen van der Kooij. Genetic population structure and tools for the management of European sprat (Sprattus sprattus). ICES Journal of Marine Science, 77(6):2134–2143, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2134/5885266.

Minello:2021:EIM

[MCX21] Murilo Minello, Leandro Calado, and Fabio C. Xavier. Ecoacoustic indices in marine ecosystems: a review on recent developments, challenges, and future directions. ICES Journal of Marine Science, 78(9):3066–3074, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3066/6381746.

Magnusson:2021:EDT

[MD21] A. K. Magnusson and W. Dekker. Economic development in times of population decline — a century of European eel fishing on the Swedish west coast. ICES Journal of Marine Science, 78(1):185–198, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/185/6026102.

Miethe:2022:TLB

[MD22] Tanja Miethe and Helen Dobby. Testing length-based reference points in a management strategy evaluation for cuckoo ray (Leucoraja naevus) and thornback ray (Raja clavata). ICES Journal of Marine Science, 79(1):129–146, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/129/6473230.

MaresMayagoitia:2021:NAP

[MDAMPE^{+}21] Jorge Alberto Mares Mayagoitia, Juan A. De-Anda-Montañez, Ricardo Perez-Enriquez, Paulina Mejía-Ruíz, José Luis Gutiérrez-González, and Fausto Valenzuela-Quiñonez. Neutral and adaptive population structure of pink abalone (Haliotis corrugata): fishery management implications. ICES Journal of Marine Science, 78(5):1909–1919, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/5/1909/6289743.

Miller:2021:AMP

[MDHB21] Sara E. Miller, Sherri C. Dressel, Christine M. Hinds, and Detlef Buettner. Age at maturation predicted from scale measurements in Pacific herring (Clupea pallasii). ICES Journal of Marine Science, 78(2):704–713, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/2/704/6139922.

Mesquita:2021:ASD

[MDP^{+}21] Carlos Mesquita, Helen Dobby, Graham J. Pierce, Catherine S. Jones, and Paul G. Fernandes. Abundance and spatial distribution of brown crab (Cancer pagurus) from fishery-independent dredge and trawl surveys in the North Sea. ICES Journal of Marine Science, 78(2):597–610, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/2/597/5880823.

Mahajan:2023:GBU

[MEOA23] Shauna L. Mahajan, Estradivari, Lenice Ojwang, and Gabby N. Ahmadia. The good, the bad, and the ugly: reflections on co-designing science for impact between the Global South and Global North. ICES Journal of Marine Science, 80(2):390–393, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/2/390/6613543.

Methratta:2020:MFR

[Met20] Elizabeth T. Methratta. Monitoring fisheries resources at offshore wind farms: BACI vs. BAG designs. ICES Journal of Marine Science, 77(3):890–900, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
890/5802590.

Mofokeng:2024:FOP

[MFA^{+}24] R. P. Mofokeng, A. Faltynkova, M. B. Alfonso, I. Boujmil, I. R. B. Carvalho, K. Lunzalu, N. B. Mohd Zanuri, E. S. Nyadjro, P. S. Puskic, D. J. Lindsay, K. Willis, T. M. Adyel, C. Serra-Gonçalves, A. Zolich, T. S. Eriksen, H-C Evans, D. Gabriel, S. Hajbane, G. Suaria, K. L. Law, and D. Lobelle. The future of ocean plastics: designing diverse collaboration frameworks. ICES Journal of Marine Science, 81(1):43–54, January 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/1/43/7110746.

March:2023:CWD

[MFB23] Antaya March, Pierre Failler, and Michael Bennett. Challenges when designing blue bond financing for Small Island Developing States. ICES Journal of Marine Science, 80 (8):2244–2251, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/8/2244/7036139.

Morley:2020:CUC

[MFP20] James W. Morley, Thomas L. Frölicher, and Malin L. Pinsky. Characterizing uncertainty in climate impact projections: a case study with seven marine species on the North American continental shelf. ICES Journal of Marine Science, 77(6):2118–2133, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2118/5885271.

McCormick:2023:NEP

[MFV^{+}23] M. I. McCormick, E. P. Fakan, G. Vamvounis, S. Bosshard, C. Moad, E. Smyth, P. Gatenby, J. Edmiston, K. Patel, and B. J. M. Allan. No effects of plasticized microplastics on the body condition and reproduction of a marine fish. ICES Journal of Marine Science, 80(5):1267–1276, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1267/7103487.

Mamoozadeh:2023:GDR

[MGB^{+}23] N. R. Mamoozadeh, J. E. Graves, R. Bealey, J. Schratwieser, J. C. Holdsworth, S. Ortega-Garcia, and J. R. McDowell. Genomic data resolve long-standing uncertainty by distinguishing white marlin (Kajikia albida) and striped marlin (K. audax) as separate species. ICES Journal of Marine Science, 80(6):1802–1813, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1802/7223542.

Machado:2021:ADS

[MGF^{+}21] Alexandre M. S. Machado, Eduardo L. Hettwer Giehl, Luiza Pacheco Fernandes, Simon N. Ingram, and Fábio G.

Daura-Jorge. Alternative data sources can fill the gaps in data-poor fisheries. ICES Journal of Marine Science, 78(5): 1663–1671, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/5/1663/6231554.

Matthews:2021:RIZ

[MGO21] Stephanie A. Matthews, Erica Goetze, and Mark D. Ohman. Recommendations for interpreting zooplankton metabarcoding and integrating molecular methods with morphological analyses. ICES Journal of Marine Science, 78(9):3387–3396, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3387/6295520.

Mosman:2023:SRF

[MGO^{+}23] J. D. Mosman, B. L. Gilby, A. D. Olds, H. P. Borland, A. J. Rummell, and C. J. Henderson. Scavenger richness and functional diversity modify carrion consumption in the surf zone of ocean beaches. ICES Journal of Marine Science, 80(7):2024–2035, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/2024/7249984.

Martinez:2021:EZD

[MHA^{+}21] Miguel A. Martinez, Clara María Hereu, María Clara Arteaga, Sylvia P. A. Jiménez-Rosenberg, Sharon Z. Herzka, Anaid Saavedra-Flores, Javier Robles-Flores, Ricardo Gomez-Reyes, Paola Gabriela Batta-Lona, Jaime Gasca-Pineda, and Clara E. Galindo-Sánchez. Epipelagic zooplankton diversity in the deep water region of the Gulf of Mexico: a metabarcoding survey. ICES Journal of Marine Science, 78(9):3317–3332, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3317/6277126.

Malde:2020:MID

[MHES20] Ketil Malde, Nils Olav Handegard, Line Eikvil, and Arnt-Børre Salberg. Machine intelligence and the datadriven future of marine science. ICES Journal of Marine Science, 77(4):1274–1285, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1274/5457719.

Marks:2020:TCG

[MHJ^{+}20] Rachel Marks, S. Alex Hesp, Danielle Johnston, Ainslie Denham, and Neil Loneragan. Temporal changes in the growth of a crustacean species, Portunus armatus, in a temperate marine embayment: evidence of density dependence. ICES Journal of Marine Science, 77(2):773–790, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/773/5680364.

Monkman:2020:AEF

[MHKV20] Graham G. Monkman, Kieran Hyder, Michel J. Kaiser, and Franck P. Vidal. Accurate estimation of fish length in single camera photogrammetry with a fiducial marker. ICES Journal of Marine Science, 77(6):2245–2254, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2245/5380578.

Mytilineou:2023:EDL

[MHMP^{+}23] Chryssi Mytilineou, Bent Herrmann, Danai Mantopoulou-Palouka, Antonello Sala, and Persefoni Megalofonou. Escape, discard, and landing probability in multispecies Mediterranean bottom-trawl fishery. ICES Journal of Marine Science, 80(3):542–555, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
3/542/6301182.

Mosman:2020:SCE

[MHO^{+}20] Jesse D. Mosman, Christopher J. Henderson, Andrew D. Olds, Ben L. Gilby, and Thomas A. Schlacher. Seascape connectivity exerts differing effects for fish assemblages in distinct habitats of the surf zones of ocean beaches. ICES Journal of Marine Science, 77(3):1033–1042, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1033/5775557.

Millar:2024:HED

[Mil24] Russell B. Millar. Hybrid experimental designs for joint estimation of absolute and relative retention curves for trawl gears. ICES Journal of Marine Science, 81(1):66–74, January 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/1/66/7335583.

Monsanto:2023:BFD

[MKB^{+}23] Maxine Monsanto, Peter Kohler, Umberto Binetti, Briony Silburn, Josie Russell, Chris Corbin, Brett Lyons, Silvana N. R. Birchenough, and Thomas Maes. A Blue Future: developing a national marine litter action plan in SIDS — lessons learnt in Belize. ICES Journal of Marine Science, 80(8):2171–2182, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
8/2171/6953704.

Muller-Karger:2023:MLB

[MKCA^{+}23] Frank E. Muller-Karger, Gabrielle Canonico, Claudia Baron Aguilar, Nicholas J. Bax, Ward Appeltans, Kristen Yarincik, Venus Leopardas, Isabel Sousa-Pinto, Masahiro Nakaoka, Akkeshi Aikappu, Jonatha Giddens, Emma Heslop, Enrique Montes, and J. Emmett Duffy. Marine Life 2030: building global knowledge of marine life for local action in the Ocean Decade. ICES Journal of Marine Science, 80(2):355–357, March 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/2/355/6589469.

Morse:2020:PSA

[MKGC20] Molly R. Morse, Lisa A. Kerr, Benjamin Galuardi, and Steven X. Cadrin. Performance of stock assessments for mixed-population fisheries: the illustrative case of Atlantic bluefin tuna. ICES Journal of Marine Science, 77(6):2043– 2055, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/6/2043/5865469.

Mills:2024:MPS

[MKK^{+}24] Katherine E Mills, Adam Kemberling, Lisa A Kerr, Sean M Lucey, Richard S McBride, Janet A Nye, Andrew J Pershing, Miguel Barajas, and Carly S Lovas. Multispecies populationscale emergence of climate change signals in an ocean warming hotspot. ICES Journal of Marine Science, 81(2):375– 389, March 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/2/375/7596727.

Machida:2021:CAZ

[MKN^{+}21] Ryuji J. Machida, Haruko Kurihara, Ryota Nakajima, Takashi Sakamaki, Ya-Ying Lin, and Kazushi Furusawa. Comparative analysis of zooplankton diversities and compositions estimated from complement DNA and genomic DNA amplicons, metatranscriptomics, and morphological identifications. ICES Journal of Marine Science, 78(9):3428–3443, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3428/6304979.

McDonald:2021:EIS

[MKS^{+}21] Raphaël R. McDonald, David M. Keith, Jessica A. Sameoto, Jeffrey A. Hutchings, and Joanna M. Flemming. Explicit incorporation of spatial variability in a biomass dynamics assessment model. ICES Journal of Marine Science, 78(9):3265–3280, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3265/6386085.

McDonald:2023:IHF

[MKSM23] Raphaël R. McDonald, David M. Keith, Jessica A. Sameoto, and Joanna Mills Flemming. Integrating habitat features into spatio-temporal biomass dynamics models for a better understanding of stock productivity: a case study of sea scallop in the Bay of Fundy. ICES Journal of Marine Science, 80(6):1710–1726, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1710/7206784.

Medellin-Mora:2020:CSP

[MMAE20] Johanna Medellín-Mora, Angus Atkinson, and Rubén Escribano. Community structured production of zooplankton in the eastern boundary upwelling system off central/ southern Chile (2003–2012). ICES Journal of Marine Science, 77(1):419–435, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/419/5612122.

Maltby:2023:NCS

[MMC^{+}23] K. M. Maltby, J. G. Mason, H. Cheng, G. Fay, R. L. Selden, L. Williams, and C. L. Alves. Navigating concepts of socialecological resilience in marine fisheries under climate change: shared challenges and recommendations from the northeast United States. ICES Journal of Marine Science, 80(9):2266– 2279, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/9/2266/7301219.

Mendo:2024:WSA

[MMCS^{+}24] T Mendo, A Mujal-Colilles, J Stounberg, G Glemarec, J Egekvist, E Mugerza, M Rufino, R Swift, and M James. A workflow for standardizing the analysis of highly resolved vessel tracking data. ICES Journal of Marine Science, 81(2): 390–401, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/81/2/390/7516127.

McGill:2023:PSC

[MMH^{+}23] L. McGill, A. D. McDevitt, B. Hellemans, F. Neat, H. Knutsen, S. Mariani, H. Christiansen, T. Johansen, F. A. M. Volckaert, and I. Coscia. Population structure and connectivity in the genus Molva in the Northeast Atlantic. ICES Journal of Marine Science, 80(4):1079–1086, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1079/7079686.

Murillas-Maza:2023:PBB

[MMMB^{+}23] A. Murillas-Maza, E. Mugerza, E. Bachiller, L. A. Errazkin, and M. Louzao. Participatory-based bio-economic activity mapping of small-scale fisheries: towards holistic management in the Bay of Biscay. ICES Journal of Marine Science, 80(5):1202–1217, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1202/7180101.

McQueen:2022:SAC

[MMN^{+}22] Kate McQueen, Justin J. Meager, Daniel Nyqvist, Jon Egil Skjæraasen, Esben Moland Olsen, Ørjan Karlsen, Petter H. Kvadsheim, Nils Olav Handegard, Tonje Nesse Forland, and Lise Doksæter Sivle. Spawning Atlantic cod (Gadus morhua L.) exposed to noise from seismic airguns do not abandon their spawning site. ICES Journal of Marine Science, 79(10):2697–2708, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/10/2697/6827581.

Miller:2023:EED

[MOC^{+}23] M. R. Miller, R. L. Oakes, P. A. Covert, D. Ianson, and J. F. Dower. Evidence for an effective defence against ocean acidification in the key bioindicator pteropod Limacina helicina. ICES Journal of Marine Science, 80(5):1329–1341, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1329/7160602.

Montgomery:2022:WWF

[Mon22] John C. Montgomery. What will the future make of our discoveries? ICES Journal of Marine Science, 79(5):1435– 1444, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1435/6585090.

Moore:2023:PEN

[Moo23a] Michael J. Moore. Policy enabling North Atlantic right whale reproductive health could save the species. ICES Journal of Marine Science, 80(2):237–242, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/2/237/6992550.

Moore:2023:CIS

[Moo23b] Sue E. Moore. Chasing inter-species communication: what marine mammals are telling us about our oceans. ICES Journal of Marine Science, 80(4):690–697, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/690/7070108.

Mousing:2023:QDS

[MPA^{+}23] Erik Askov Mousing, Benjamin Planque, Per Arneberg, Vilde Regine Bjørdal, Felicia Keulder-Stenevik, Penny Lee Liebig, Herdis Langøy Mørk, and Mette Skern-Mauritzen. Quantifying diets for small pelagic fish: effects of weight versus occurrence methods and sampling effort. ICES Journal of Marine Science, 80(2):317–328, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/2/317/6990985.

Mueter:2021:PFS

[MPH^{+}21] Franz J. Mueter, Benjamin Planque, George L. Hunt, Irene D. Alabia, Toru Hirawake, Lisa Eisner, Padmini Dalpadado, Melissa Chierici, Kenneth F. Drinkwater, Naomi Harada, Per Arneberg, and Sei-Ichi Saitoh. Possible future scenarios in the gateways to the Arctic for subarctic and Arctic marine systems: II. Prey resources, food webs, fish, and fisheries. ICES Journal of Marine Science, 78(9):3017–3045, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3017/6384790.

Munroe:2022:ASF

[MPK^{+}22] Daphne M. Munroe, Eric N. Powell, John M. Klinck, Andrew M. Scheld, Sarah Borsetti, Jennifer Beckensteiner, and Eileen E. Hofmann. The Atlantic surfclam fishery and offshore wind energy development: 1. Model development and verification. ICES Journal of Marine Science, 79(6):1787– 1800, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/6/1787/6611678.

Miret-Pastor:2020:CBR

[MPMGGAHR20] Lluís Miret-Pastor, Agustín Molina-García, César García-Aranda, and Paloma Herrera-Racionero. The connection between recreational fishing and the traditional fishing sector in the emerging area of marine tourism: challenges and opportunities for diversification with the European Fisheries Fund (EFF). ICES Journal of Marine Science, 77(6):2369–2374, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2369/5543495.

Millette:2020:WTD

[MPN20] N. C. Millette, J. J. Pierson, and E. W. North. Water temperature during winter may control striped bass recruitment during spring by affecting the development time of copepod nauplii. ICES Journal of Marine Science, 77(1):300–314, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/1/300/5638886.

Martin:2020:PTB

[MPQS20] Karen L. M. Martin, Emily A. Pierce, Vincent V. Quach, and Melissa Studer. Population trends of beach-spawning California grunion Leuresthes tenuis monitored by citizen scientists. ICES Journal of Marine Science, 77(6):2226–2233, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2226/5497988.

Maps:2020:MEM

[MR20a] Frédéric Maps and Nicholas R. Record. Marine ecosystems model development should be rooted in past experiences, not anchored in old habits. ICES Journal of Marine Science, 77(1):46–57, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/46/5651047.

Meyer-Rochow:2020:IBS

[MR20b] Victor Benno Meyer-Rochow. Ingredients to become a scientist: curiosity, enthusiasm, perseverance, opportunity, and a good pinch of luck. ICES Journal of Marine Science, 77(6):2013–2021, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2013/5866692.

Madigan:2021:WCS

[MRC^{+}21] Daniel J. Madigan, Andrew J. Richardson, Aaron B. Carlisle, Sam B. Weber, Judith Brown, and Nigel E. Hussey. Water column structure defines vertical habitat of twelve pelagic predators in the South Atlantic. ICES Journal of Marine Science, 78(3):867–883, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
867/6050574.

McQueen:2023:BRW

[MSN^{+}23] K. McQueen, J. E. Skjæraasen, D. Nyqvist, E. M. Olsen, Ø Karlsen, J. J. Meager, P. H. Kvadsheim, N. O. Handegard, T. N. Forland, K. de Jong, and L. D. Sivle. Behavioural responses of wild, spawning Atlantic cod (Gadus morhua L.) to seismic airgun exposure. ICES Journal of Marine Science, 80(4):1052–1065, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1052/7076237. See correction [Ano23b].

Moriarty:2020:CFS

[MSP^{+}20] Meadhbh Moriarty, Suresh A. Sethi, Debbi Pedreschi, T. Scott Smeltz, Chris McGonigle, Bradley P. Harris, Nathan Wolf, and Simon P. R. Greenstreet. Combining fisheries surveys to inform marine species distribution modelling. ICES Journal of Marine Science, 77(2):539–552, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/539/5709875.

Martinez-Silva:2022:UGE

[MSVSA22] Maria Angelica Martínez-Silva, Marie Vagner, Caroline Senay, and Céline Audet. Using gene expression to identify the most suitable environmental conditions for growth and metabolism of juvenile deepwater redfish (Sebastes mentella) in the Estuary and the Gulf of St. Lawrence. ICES Journal of Marine Science, 79(2):382–393, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/382/6503931.

Mello:2021:VFR

[MSW21] H. Lee Mello, Abigail M. Smith, and Anna C. L. Wood. Voluntary fishing restrictions alone do not promote growth of bryozoan-dominated biogenic habitat on the Otago shelf, southeastern New Zealand. ICES Journal of Marine Science, 78(4):1542–1553, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1542/6207638.

McIvor:2022:UAV

[MSWB22] Ashlie J. McIvor, Julia L. Y. Spaet, Collin T. Williams, and Michael L. Berumen. Unoccupied aerial video (UAV) surveys as alternatives to BRUV surveys for monitoring elasmobranch species in coastal waters. ICES Journal of Marine Science, 79(5):1604–1613, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
5/1604/6597023.

Machado:2021:IAS

[MTCC21] I. Machado, C. M. Teixeira, J. L. Costa, and H. Cabral. Identifying assessment scales for food web criteria in the NE Atlantic: implications for the marine strategy framework directive. ICES Journal of Marine Science, 78(1):246–263, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/1/246/6024435.

Monnahan:2021:IVD

[MTK^{+}21] Cole C. Monnahan, James T. Thorson, Stan Kotwicki, Nathan Lauffenburger, James N. Ianelli, and André E. Punt. Incorporating vertical distribution in index standardization accounts for spatiotemporal availability to acoustic and bottom trawl gear for semi-pelagic species. ICES Journal of Marine Science, 78(5):1826–1839, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1826/6278035.

Mazur:2022:ISH

[MTS^{+}22] Mackenzie D. Mazur, Kisei R. Tanaka, Burton Shank, Jui-Han Chang, Cameron T. Hodgdon, Kathleen M. Reardon, Kevin D. Friedland, and Yong Chen. Incorporating spatial heterogeneity and environmental impacts into stock-recruitment relationships for Gulf of Maine lobster. ICES Journal of Marine Science, 79(2):362–372, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/362/6511800.

Merkel:2023:QCS

[MTT^{+}23] B. Merkel, P. Trathan, S. Thorpe, E. J. Murphy, H. Pehlke, K. Teschke, and G. P. Griffith. Quantifying circumpolar summer habitat for Antarctic krill and ice krill, two key species of the Antarctic marine ecosystem. ICES Journal of Marine Science, 80(6):1773–1786, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1773/7224140.

Melbourne-Thomas:2023:IHD

[MTTG^{+}23] Jess Melbourne-Thomas, Desiree Tommasi, Marion Gehlen, Eugene J. Murphy, Jennifer Beckensteiner, Francisco Bravo, Tyler D. Eddy, Mibu Fischer, Elizabeth Fulton, Mayya Gogina, Eileen Hofmann, Maysa Ito, Sara Mynott, Kelly Ortega-Cisneros, Anna N. Osiecka, Mark R. Payne, Romeo Saldívar-Lucio, and Kim J. N. Scherrer. Integrating human dimensions in decadal-scale prediction for marine social–ecological systems: lighting the grey zone. ICES Journal of Marine Science, 80(1):16–30, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/16/6960601.

Munyandorero:2020:CPL

[Mun20] Joseph Munyandorero. Conflicting perceptions on the lifehistory covariates of density-dependent recruitment based on information-generating equations. ICES Journal of Marine Science, 77(3):878–889, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/3/878/5803789.

Murawski:2021:WYC

[Mur21] Steve Murawski. When you come to a fork in the road — take it! ICES Journal of Marine Science, 78(7):2359–2370, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/7/2359/6346868.

Marchal:2023:STD

[MV23] Paul Marchal and Youen Vermard. Species targeting and discarding in mixed fisheries. ICES Journal of Marine Science, 80(3):532–541, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/3/532/6594653.

Meier:2023:TFA

[MvdMS^{+}23] Sonnich Meier, Terje van der Meeren, Jon Egil Skjæraasen, Raymond J. Bannister, Josef Daniel Rasinger, and Ørjan Karlsen. Terrestrial fatty acids from feed oil in feed for farmed salmonids are transferred to the liver, gonads, and muscle of wild Atlantic cod (Gadus morhua). ICES Journal of Marine Science, 80(4):1143–1154, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1143/7097698.

Minte-Vera:2021:ADA

[MVMdS21] Carolina V. Minte-Vera, Mark N. Maunder, and Alexandre M. Aires da Silva. Auxiliary diagnostic analyses used to detect model misspecification and highlight potential solutions in stock assessments: application to yellowfin tuna in the eastern Pacific Ocean. ICES Journal of Marine Science, 78(10):3521–3537, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3521/6430633.

Marquez-Velasquez:2021:RPB

[MVNR^{+}21] Viviana Márquez-Velásquez, Andrés F. Navia, Ricardo S. Rosa, Paulo R. Guimarães, Jr., and Rafael L. G. Raimundo. Resource partitioning between fisheries and endangered sharks in a tropical marine food web. ICES Journal of Marine Science, 78(7):2518–2527, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2518/6320818.

Ma:2024:SDC

[MWJ^{+}24] Yongxing Ma, Yongsheng Wu, Nicholas W Jeffery, Rachel Horwitz, Jinshan Xu, Ed Horne, and Ryan R E Stanley. Simulating dispersal in a complex coastal environment: the Eastern Shore Islands archipelago. ICES Journal of Marine Science, 81(1):178–194, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/1/178/7473284.

Munsch:2023:DEM

[MWS^{+}23] S. H. Munsch, R. K. Walter, B. L. Sanderson, L. Reshitnyk, J. K. O’Leary, P. M. Kiffney, M. Hessing-Lewis, T. G. Gerwing, C. A. Endris, W. B. Chesney, K. M. Beheshti, and F. L. Beaty. Decades of eelgrass meadow dynamics across the northeast Pacific support seascape-scale conservation. ICES Journal of Marine Science, 80(10):2520–2529, December 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/10/2520/7330258.

Mason:2021:PCD

[MWT^{+}21] Julia G. Mason, Pamela J. Woods, Magnús Thorlacius, Kristinn Gudnason, Vincent S. Saba, Patrick J. Sullivan, and Kristin M. Kleisner. Projecting climate-driven shifts in demersal fish thermal habitat in Iceland’s waters. ICES Journal of Marine Science, 78(10):3793–3804, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3793/6429120.

Melnychuk:2021:TRD

[MYAH21] Michael C. Melnychuk, Erik G. Young, Leif E. Anderson, and Ray Hilborn. Trading off retained daily catch for longer seasons to maximize angler benefits in recreational fisheries management. ICES Journal of Marine Science, 78(5):1650– 1662, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/5/1650/6226982.

Mackay:2020:INC

[MYJ^{+}20] Mary Mackay, Satoshi Yamazaki, Sarah Jennings, Hugh Sibly, Ingrid E. van Putten, and Timothy J. Emery. The influence of nudges on compliance behaviour in recreational fisheries: a laboratory experiment. ICES Journal of Marine Science, 77(6):2319–2332, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2319/5363229.

Mestdagh:2020:DIT

[MYMV20] Sebastiaan Mestdagh, Tom Ysebaert, Tom Moens, and Carl Van Colen. Dredging-induced turbid plumes affect bio-irrigation and biogeochemistry in sediments inhabited by Lanice conchilega (Pallas, 1766). ICES Journal of Marine Science, 77(3):1219–1226, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1219/5094980.

Mcilvenny:2021:CAT

[MYW^{+}21] J. Mcilvenny, A. Youngson, B. J. Williamson, N. R. Gauld, L. Goddijn-Murphy, and D. Del Villar-Guerra. Combining acoustic tracking and hydrodynamic modelling to study migratory behaviour of Atlantic salmon (Salmo salar) smolts on entry into high-energy coastal waters. ICES Journal of Marine Science, 78(7):2409–2419, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2409/6297551.

Nogues:2022:SEN

[NAB^{+}22] Quentin Nogues, Emma Araignous, Pierre Bourdaud, Ghassen Halouani, Aurore Raoux, Éric Foucher, François Le Loc’h, Frédérique Loew-Turbout, Frida Ben Rais Lasram, Jean-Claude Dauvin, and Nathalie Niquil. Spatialized ecological network analysis for ecosystem-based management: effects of climate change, marine renewable energy, and fishing on ecosystem functioning in the Bay of Seine. ICES Journal of Marine Science, 79(4):1098–1112, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1098/6535870.

Nicastro:2020:CBF

[NAS^{+}20] Katy R. Nicastro, Jorge Assis, Ester A. Serrão, Gareth A. Pearson, João Neiva, Myriam Valero, Rita Jacinto, and Gerardo I. Zardi. Congruence between fine-scale genetic breaks and dispersal potential in an estuarine seaweed across multiple transition zones. ICES Journal of Marine Science, 77(1): 371–378, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/1/371/5637823.

Nogues:2023:EWA

[NBA^{+}23] Quentin Nogues, Pierre Bourdaud, Emma Araignous, Ghassen Halouani, Frida Ben Rais Lasram, Jean-Claude Dauvin, François Le Loc'h, and Nathalie Niquil. An ecosystemwide approach for assessing the spatialized cumulative effects of local and global changes on coastal ecosystem functioning. ICES Journal of Marine Science, 80(4):1129–1142, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/1129/7092991.

Newton:2021:CAD

[NBL^{+}21] Matthew Newton, James Barry, Angus Lothian, Robert Main, Hannele Honkanen, Simon Mckelvey, Paul Thompson, Ian Davies, Nick Brockie, Alastair Stephen, Rory O’Hara Murray, Ross Gardiner, Louise Campbell, Paul Stainer, and Colin Adams. Counterintuitive active directional swimming behaviour by Atlantic salmon during seaward migration in the coastal zone. ICES Journal of Marine Science, 78(5): 1730–1743, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/5/1730/6262049.

Neuenfeldt:2020:FGAa

[NBO^{+}20a] Stefan Neuenfeldt, Valerio Bartolino, Alessandro Orio, Ken H. Andersen, Niels G. Andersen, Susa Niiranen, Ulf Bergström, Didzis Ustups, Nataliia Kulatska, and Michele Casini. Feeding and growth of Atlantic cod (Gadus morhua L.) in the eastern Baltic Sea under environmental change. ICES Journal of Marine Science, 77(2):624–632, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/2/624/5643855.

Neuenfeldt:2020:FGAb

[NBO^{+}20b] Stefan Neuenfeldt, Valerio Bartolino, Alessandro Orio, Ken H. Andersen, Niels G. Andersen, Susa Niiranen, Ulf Bergström, Didzis Ustups, Nataliia Kulatska, and Michele Casini. Feeding and growth of Atlantic cod (Gadus morhua L.) in the eastern Baltic Sea under environmental change. ICES Journal of Marine Science, 77(2):858, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/2/858/5707320.

Neuenfeldt:2020:RRG

[NBO^{+}20c] Stefan Neuenfeldt, Valerio Bartolino, Alessandro Orio, Ken H. Andersen, Niels G. Andersen, Susa Niiranen, Ulf Bergström, Didzis Ustups, Nataliia Kulatska, and Michele Casini. Reply to “Reduced growth in Baltic Sea cod may be due to mild hypoxia” — a comment to Neuenfeldt et al. (2020). ICES Journal of Marine Science, 77(5):2006–2008, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/2006/5862933.

Ng:2021:PSC

[NDE^{+}21] Elizabeth L. Ng, Jonathan J. Deroba, Timothy E. Essington, Arnaud Grüss, Brian E. Smith, and James T. Thorson. Predator stomach contents can provide accurate indices of prey biomass. ICES Journal of Marine Science, 78(3):1146– 1159, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/1146/6145864.

Nakayama:2020:MFT

[NFN20] Shin-Ichiro Nakayama, Hiromu Fukuda, and Shuya Nakatsuka. Model-free time series analysis detected the contributions of middle-age spawner biomass and the environment on Pacific bluefin tuna recruitment. ICES Journal of Marine Science, 77(4):1480–1491, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1480/5531239.

Navarro:2020:DSP

[NHL^{+}20] Matthew Navarro, Atakelty Hailu, Tim Langlois, Karina L. Ryan, and Marit E. Kragt. Determining spatial patterns in recreational catch data: a comparison of generalized additive mixed models and boosted regression trees. ICES Journal of Marine Science, 77(6):2216–2225, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2216/5532211.

Navarro:2022:CSE

[NHL^{+}22] Matthew Navarro, Atakelty Hailu, Tim Langlois, Karina L. Ryan, Michael Burton, and Marit E. Kragt. Combining spatial ecology and economics to incorporate recreational fishing into marine spatial planning. ICES Journal of Marine Science, 79(1):147–157, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/147/6470668.

Nielsen:2021:MFS

[NHM^{+}21] Anders Nielsen, Niels T. Hintzen, Henrik Mosegaard, Vanessa Trijoulet, and Casper W. Berg. Multi-fleet statespace assessment model strengthens confidence in singlefleet SAM and provides fleet-specific forecast options. ICES Journal of Marine Science, 78(6):2043–2052, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/6/2043/6317566.

Nishizawa:2020:CAS

[NKA^{+}20] Bungo Nishizawa, Naoya Kanna, Yoshiyuki Abe, Yoshihiko Ohashi, Daiki Sakakibara, Izumi Asaji, Shin Sugiyama, Atsushi Yamaguchi, and Yutaka Watanuki. Contrasting assemblages of seabirds in the subglacial meltwater plume and oceanic water of Bowdoin Fjord, northwestern Greenland. ICES Journal of Marine Science, 77(2):711–720, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/711/5677500.

Nordio:2021:APC

[NKA^{+}21] Davide Nordio, Natalie Khtikian, Sean Andrews, Daniela Bertotto, Karen Leask, and Timothy Green. Adaption potential of Crassostrea gigas to ocean acidification and disease caused by Vibrio harveyi. ICES Journal of Marine Science, 78(1):360–367, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/360/5864916.

Norton:2023:SOF

[NKS^{+}23] Emily L. Norton, Isaac C. Kaplan, Samantha Siedlecki, Albert J. Hermann, Simone R. Alin, Jan Newton, Kelly Corbett, Daniel Ayres, Ervin Joe Schumacker, Nicholas A. Bond, Kate Richerson, and Michael A. Alexander. Seasonal ocean forecasts to improve predictions of Dungeness crab catch rates, co-developed with state and tribal fishery managers. ICES Journal of Marine Science, 80(4):823–835, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/823/7034403.

Nomura:2022:FCM

[NSJ^{+}22] Keiko Nomura, Jameal F. Samhouri, Andrew F. Johnson, Alfredo Giron-Nava, and James R. Watson. Fisheries connectivity measures of adaptive capacity in small-scale fisheries. ICES Journal of Marine Science, 79(2):519–531, March 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/2/519/6374668.

Nooteboom:2023:IBM

[NSS^{+}23] Peter D. Nooteboom, Joe Scutt Phillips, Inna Senina, Erik van Sebille, and Simon Nicol. Individual-based model simulations indicate a non-linear catch equation of drifting Fish Aggregating Device-associated tuna. ICES Journal of Marine Science, 80(6):1746–1757, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1746/7220136.

Olin:2022:STV

[OBJ^{+}22] Agnes B. Olin, Neil S. Banas, David G. Johns, Michael R. Heath, Peter J. Wright, and Ruedi G. Nager. Spatiotemporal variation in the zooplankton prey of lesser sandeels: species and community trait patterns from the Continuous Plankton Recorder. ICES Journal of Marine Science, 79(5): 1649–1661, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/5/1649/6605793.

Oyafuso:2021:ISV

[OBK21] Zack S. Oyafuso, Lewis A. K. Barnett, and Stan Kotwicki. Incorporating spatiotemporal variability in multispecies survey design optimization addresses trade-offs in uncertainty. ICES Journal of Marine Science, 78(4):1288–1300, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/4/1288/6157461.

Ochwada-Doyle:2023:QPI

[ODHF^{+}23] Faith Ochwada-Doyle, Julian Hughes, Ashley M. Fowler, Jeffrey Murphy, Kate Stark, Michael Lowry, Laurie West, and Matthew D. Taylor. Quantifying the potential impact of reducing recreational harvest limits on the catch of a key marine species using off-site angler survey data. ICES Journal of Marine Science, 80(2):295–307, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/2/295/6988646.

Olson:2021:SSN

[OdS21] Julia Olson and Patricia Pinto da Silva. Science, social networks, and collaboration: an analysis of publications in fisheries science from 1990 to 2018. ICES Journal of Marine Science, 78(3):810–820, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/3/810/6124058.

OLeary:2022:UTS

[ODT^{+}22] Cecilia A. O’Leary, Lukas B. DeFilippo, James T. Thorson, Stan Kotwicki, Gerald R. Hoff, Vladimir V. Kulik, James N. Ianelli, and André E. Punt. Understanding transboundary stocks’ availability by combining multiple fisheriesindependent surveys and oceanographic conditions in spatiotemporal models. ICES Journal of Marine Science, 79(4): 1063–1074, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/1063/6555703.

Ounsley:2020:RVD

[OGMA20] James P. Ounsley, Alejandro Gallego, David J. Morris, and John D. Armstrong. Regional variation in directed swimming by Atlantic salmon smolts leaving Scottish waters for their oceanic feeding grounds — a modelling study. ICES Journal of Marine Science, 77(1):315–325, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/1/315/5581804.

Otero:2023:LHT

[OH23] J. Otero and M. Hidalgo. Life-history traits and environment shape small pelagic fish demography and responses to fishing and climate across European Atlantic seas. ICES Journal of Marine Science, 80(5):1447–1461, July 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/5/1447/7152845.

Ostman:2023:SBI

[ÖHB^{+}23] Ö Östman, K. Hommik, E. Bolund, O. Heikinheimo, M. Olin, A. M. Lejk, R. Svirgsden, S. Smoliński, and J. Olsson. Size-based indicators for assessments of ecological status of coastal fish communities. ICES Journal of Marine Science, 80(10):2478–2489, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/10/2478/7341836.

Ohayon:2023:CEE

[OHM^{+}23] Sarah Ohayon, Hikaru Homma, Shahar Malamud, Ilia Ostrovsky, Ruthy Yahel, Thomas Mehner, Mor Kanari, and Jonathan Belmaker. Consistent edge effect patterns revealed using continuous surveys across an Eastern Mediterranean no-take marine protected area. ICES Journal of Marine Science, 80(6):1594–1605, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1594/7218790.

Orofino:2023:OCI

[OMM^{+}23] Sara Orofino, Gavin McDonald, Juan Mayorga, Christopher Costello, and Darcy Bradley. Opportunities and challenges for improving fisheries management through greater transparency in vessel tracking. ICES Journal of Marine Science, 80(4):675–689, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/4/675/7030574.

ONeill:2021:GDD

[ON21] F. G. O’Neill and T. Noack. The geometry and dynamics of Danish anchor seine ropes on the seabed. ICES Journal of Marine Science, 78(1):125–133, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
125/6017348.

OSullivan:2022:GSI

[OOB^{+}22] Ronan James O'Sullivan, Mikhail Ozerov, Geir H. Bolstad, John Gilbey, Jan Arge Jacobsen, Jaakko Erkinaro, Audun H. Rikardsen, Kjetil Hindar, and Tutku Aykanat. Genetic stock identification reveals greater use of an oceanic feeding ground around the Faroe Islands by multi-sea winter Atlantic salmon, with variation in use across reporting groups. ICES Journal of Marine Science, 79(9):2442–2452, November 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/9/2442/6762978.

Olin:2022:IOS

[OOE^{+}22] Agnes B. Olin, Jens Olsson, Johan S. Eklöf, Britas Klemens Eriksson, Olavi Kaljuste, Laura Briekmane, and Ulf Bergström. Increases of opportunistic species in response to ecosystem change: the case of the Baltic Sea three-spined stickleback. ICES Journal of Marine Science, 79(5):1419– 1434, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1419/6581452.

Ommer:2022:ICS

[OP22] Rosemary E. Ommer and R. Ian Perry. The importance of complexities of scale and context in fisheries research and management. ICES Journal of Marine Science, 79(8):2196– 2203, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/8/2196/6724457.

Orozco-Ruiz:2023:GAR

[ORGTO^{+}23] A. M. Orozco-Ruiz, C. Galván-Tirado, A. M. Orlov, S. Y. Orlova, and F. J. García DeLeón. Genetic analyses reveal a non-panmictic genetic structure in the sablefish Anoplopoma fimbria in the northern Pacific. ICES Journal of Marine Science, 80(5):1319–1328, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1319/7110751.

Omori:2022:ISC

[OT22] Kristen L. Omori and James T. Thorson. Identifying species complexes based on spatial and temporal clustering from joint dynamic species distribution models. ICES Journal of Marine Science, 79(3):677–688, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/677/6530390.

Olsen:2023:TMS

[OTL^{+}23] Erik Olsen, Maciej T. Tomczak, Christopher P. Lynam, Andrea Belgrano, and Andrew Kenny. Testing management scenarios for the North Sea ecosystem using qualitative and quantitative models. ICES Journal of Marine Science, 80(1): 218–234, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/1/218/6965515.

Obregon:2020:DOP

[OTLH20] Clara Obregón, James R. Tweedley, Neil R. Loneragan, and Michael Hughes. Different but not opposed: perceptions between fishing sectors on the status and management of a crab fishery. ICES Journal of Marine Science, 77(6):2354–2368, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2354/5658677.

OLeary:2020:CMA

[OTMN20] Cecilia A. O’Leary, James T. Thorson, Timothy J. Miller, and Janet A. Nye. Comparison of multiple approaches to calculate time-varying biological reference points in climate-linked population-dynamics models. ICES Journal of Marine Science, 77(3):930–941, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
930/5637330.

Pendleton:2023:CDM

[PAC^{+}23] Linwood H. Pendleton, Stella J. Alexandroff, Alison Clausen, Jörn O. Schmidt, and Howard I. Browman. Co-designing marine science for the ocean we want. ICES Journal of Marine Science, 80(2):342–346, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/

2/342/7037136. See correction [Ano23c].

Parry:2021:MCT

[PASL21] H. E. Parry, A. Atkinson, P. J. Somerfield, and P. K. Lindeque. A metabarcoding comparison of taxonomic richness and composition between the water column and the benthic boundary layer. ICES Journal of Marine Science, 78(9):3333–3341, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3333/6042356.

Pedersen:2022:CEA

[PAW22] Torstein Pedersen, Colin Amundsen, and Stephen Wickler. Characteristics of early Atlantic cod (Gadus morhua L.) catches based on otoliths recovered from archaeological excavations at medieval to early modern sites in northern Norway. ICES Journal of Marine Science, 79(10):2667–2681, December 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/10/2667/6831971.

Pernet:2021:FNM

[PB21] Fabrice Pernet and Howard I. Browman. The future is now: marine aquaculture in the Anthropocene. ICES Journal of Marine Science, 78(1):315–322, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
315/6151698.

Page:2022:RBC

[PBC^{+}22] Heather N. Page, Keisha D. Bahr, Tyler Cyronak, Elizabeth B. Jewett, Maggie D. Johnson, and Sophie J. Mc-Coy. Responses of benthic calcifying algae to ocean acidification differ between laboratory and field settings. ICES Journal of Marine Science, 79(1):1–11, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/1/6458584.

Pardo:2021:TMS

[PBD^{+}21] Sebastián A. Pardo, Geir H. Bolstad, J. Brian Dempson, Julien April, Ross A. Jones, Dustin Raab, and Jeffrey A. Hutchings. Trends in marine survival of Atlantic salmon populations in eastern Canada. ICES Journal of Marine Science, 78(7):2460–2473, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2460/6307378.

Pelage:2021:BHP

[PBF^{+}21] Latifa Pelage, Arnaud Bertrand, Beatrice Padovani Ferreira, Flávia Lucena-Frédou, Anne K. S. Justino, and Thierry Frédou. Balanced harvest as a potential management strategy for tropical small-scale fisheries. ICES Journal of Marine Science, 78(7):2547–2561, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2547/6325327.

Paine:2023:NEO

[PBS^{+}23] Ellie R. Paine, Damon Britton, Matthias Schmid, Elizabeth A. Brewer, Guillermo Diaz-Pulido, Philip W. Boyd, and Catriona L. Hurd. No effect of ocean acidification on growth, photosynthesis, or dissolved organic carbon release by three temperate seaweeds with different dissolved inorganic carbon uptake strategies. ICES Journal of Marine Science, 80(2): 272–281, March 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/2/272/6982891.

Pennino:2023:CMA

[PCC23] Maria Grazia Pennino, Marta Coll, and Santiago Cerviño. The challenges of modelling and assessing fisheries resources. ICES Journal of Marine Science, 80(10):2563–2566, December 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/10/2563/7420083.

Peterson:2021:RCS

[PCCL21] Cassidy D. Peterson, Dean L. Courtney, Enric Cortés, and Robert J. Latour. Reconciling conflicting survey indices of abundance prior to stock assessment. ICES Journal of Marine Science, 78(9):3101–3120, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3101/6372896.

Phillips:2022:SVE

[PCG^{+}22] Elizabeth M. Phillips, Dezhang Chu, Stéphane Gauthier, Sandra L. Parker-Stetter, Andrew O. Shelton, and Rebecca E. Thomas. Spatiotemporal variability of euphausiids in the California Current Ecosystem: insights from a recently developed time series. ICES Journal of Marine Science, 79(4):1312–1326, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
4/1312/6567585.

Pascoe:2024:UCE

[PCH^{+}24] S Pascoe, R Curtotti, E Hoshino, S McWhinnie, and P Schrobback. Use of catch and effort data to monitor trends in economic performance in fisheries. ICES Journal of Marine Science, 81(1):97–107, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/1/97/7424414.

Pappalardo:2021:RTE

[PCL^{+}21] Paula Pappalardo, Allen G. Collins, Katrina M. Pagenkopp Lohan, Kate M. Hanson, Sarit B. Truskey, William Jaeckle, Cheryl Lewis Ames, Jessica A. Goodheart, Stephanie L. Bush, Leann M. Biancani, Ellen E. Strong, Michael Vecchione, M. G. Harasewych, Karen Reed, Chan Lin, Elise C. Hartil, Jessica Whelpley, Jamie Blumberg, Kenan Matterson, Niamh E. Redmond, Allison Becker, Michael J. Boyle, and Karen J. Osborn. The role of taxonomic expertise in interpretation of metabarcoding studies. ICES Journal of Marine Science, 78(9):3397–3410, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3397/6277123.

Plaganyi:2021:PFU

[PDH^{+}21] Éva Plagányi, Roy Aijun Deng, Trevor Hutton, Rob Kenyon, Emma Lawrence, Judy Upston, Margaret Miller, Chris Moeseneder, Sean Pascoe, Laura Blamey, and Stephen Eves. From past to future: understanding and accounting for recruitment variability of Australia’s redleg banana prawn (Penaeus indicus) fishery. ICES Journal of Marine Science, 78 (2):680–693, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/2/680/5855998.

Perretti:2020:STM

[PDL20] Charles T. Perretti, Jonathan J. Deroba, and Christopher M. Legault. Simulation testing methods for estimating misreported catch in a state-space stock assessment model. ICES Journal of Marine Science, 77(3):911–920, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/3/911/5804131.

Prado:2023:IWR

[PDP^{+}23] Jonatas H. F. Prado, Nicholas W. Daudt, Martin S. Perez, Pedro V. Castilho, and Danielle S. Monteiro. Intensive and wide-ranging beach surveys uncover temporal and spatial stranding patterns of marine megafauna. ICES Journal of Marine Science, 80(3):492–506, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/492/6640477.

Pol:2021:BBT

[PE21] M. Pol and S. Eayrs. Behaviour in a bottom trawl: moving forward to limit catches of Atlantic cod. ICES Journal of Marine Science, 78(7):2317–2324, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2317/6295519.

Pechenik:2023:MGT

[Pec23] Jan A. Pechenik. Maybe it’s good to not have too much focus. ICES Journal of Marine Science, 80(10):2470–2477, December 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/10/2470/7416713.

Pepin:2023:FLF

[Pep23] Pierre Pepin. Feeding by larval fish: how taxonomy, body length, mouth size, and behaviour contribute to differences among individuals and species from a coastal ecosystem. ICES Journal of Marine Science, 80(1):91–106, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/91/6956805.

Parsa:2020:EBA

[PEWN20] Mahdi Parsa, Timothy J. Emery, Ashley J. Williams, and Simon Nicol. An empirical Bayesian approach for estimating fleet- and vessel-level bycatch rates in fisheries with effort heterogeneity and limited data: a prospective tool for measuring bycatch mitigation performance. ICES Journal of Marine Science, 77(3):921–929, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
921/5762486.

Pauly:2021:MNN

[PF21] Daniel Pauly and Rainer Froese. MSY needs no epitaph — but it was abused. ICES Journal of Marine Science, 78(6):2204–2210, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2204/6050569.

Planque:2022:QTI

[PFH^{+}22] Benjamin Planque, Aurélien Favreau, Bérengère Husson, Erik Askov Mousing, Cecilie Hansen, Cecilie Broms, Ulf Lindstrøm, and Elliot Sivel. Quantification of trophic interactions in the Norwegian Sea pelagic food-web over multiple decades. ICES Journal of Marine Science, 79(6):1815–1830, August 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/6/1815/6644750.

Portnoy:2022:GAR

[PFP^{+}22] David S. Portnoy, Andrew T. Fields, Jonathan B. Puritz, Christopher M. Hollenbeck, and William F. Patterson III. Genomic analysis of red snapper, Lutjanus campechanus, population structure in the U.S. Atlantic and Gulf of Mexico. ICES Journal of Marine Science, 79(1):12–21, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/1/12/6453068.

Pouso:2020:MVR

[PFT^{+}20] Sarai Pouso, Silvia Ferrini, R. Kerry Turner, Ángel Borja, and María C. Uyarra. Monetary valuation of recreational fishing in a restored estuary and implications for future management measures. ICES Journal of Marine Science, 77(6):2295–2303, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2295/5510568.

Peters:2023:SDB

[PGRP23] S. M. Peters, R. Guppy, D. Ramsewak, and A. Potts. Socioeconomic dimensions of the Buccoo Reef Marine Park, an assessment of stakeholder perceptions towards enhanced management through MSP. ICES Journal of Marine Science, 80(5):1399–1409, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1399/7130094.

Popper:2021:FHH

[PH21] Arthur N. Popper and Anthony D. Hawkins. Fish hearing and how it is best determined. ICES Journal of Marine Science, 78(7):2325–2336, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2325/6307377.

Passerotti:2020:AER

[PHB^{+}20] Michelle S. Passerotti, Thomas E. Helser, Irina M. Benson, Beverly K. Barnett, Joseph C. Ballenger, Walter J. Bubley, Marcel J. M. Reichert, and Joseph M. Quattro. Age estimation of red snapper (Lutjanus campechanus) using FT-NIR spectroscopy: feasibility of application to production ageing for management. ICES Journal of Marine Science, 77(6):2144–2156, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2144/5885268.

Paradinas:2023:CFD

[PIAF^{+}23] Iosu Paradinas, Janine B. Illian, Alexandre Alonso-Fernändez, Maria Grazia Pennino, and Sophie Smout. Combining fishery data through integrated species distribution models. ICES Journal of Marine Science, 80(10):2579–2590, December 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/10/2579/7147648.

LaScala-Gruenewald:2023:BHC

[PJCS^{+}23a] Maria Pachiadaki, Felix Janssen, Marina Carreiro-Silva, Telmo Morato, Gilberto P. Carreira, Helena C. Frazão, Patrick Heimbach, Isabel Iglesias, Frank E. Muller-Karger, Miguel M. Santos, Leslie M. Smith, Michael F. Vardaro, Fleur Visser, Joanna J. Waniek, Ann-Christine Zinkann, and Ana Colaço. Building on a human-centred, iterative, and agile co-design strategy to facilitate the availability of deep ocean data. ICES Journal of Marine Science, 80(2):347–351, March 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/2/347/6694272.

Pachiadaki:2023:CDM

[PJCS^{+}23b] Maria Pachiadaki, Felix Janssen, Marina Carreiro-Silva, Telmo Morato, Gilberto P. Carreira, Helena C. Frazão, Patrick Heimbach, Isabel Iglesias, Frank E. Muller-Karger, Miguel M. Santos, Leslie M. Smith, Michael F. Vardaro, Fleur Visser, Joanna J. Waniek, Ann-Christine Zinkann, and Ana Colaço. Co-designing a multidisciplinary deep-ocean observing programme at the Mid-Atlantic Ridge in the Azores region: a blueprint for synergy in deep ocean research and conservation. ICES Journal of Marine Science, 80(2):352– 354, March 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/2/352/6793824.

Privitera-Johnson:2020:LSU

[PJP20] Kristin M. Privitera-Johnson and André E. Punt. Leveraging scientific uncertainty in fisheries management for estimating among-assessment variation in overfishing limits. ICES Journal of Marine Science, 77(2):515–526, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
515/5675586.

Pedersen:2020:DRS

[PKAT20] Eric J. Pedersen, Mariano Koen-Alonso, and Tyler D. Tunney. Detecting regime shifts in communities using estimated rates of change. ICES Journal of Marine Science, 77(4): 1546–1555, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/4/1546/5835266.

Probst:2021:SSP

[PKT^{+}21] Wolfgang Nikolaus Probst, Alexander Kempf, Marc Taylor, Inigo Martinez, and David Miller. Six steps to produce stock assessments for the marine strategy framework directive compliant with descriptor 3. ICES Journal of Marine Science, 78(4):1229–1240, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1229/6151700.

Parnell:2020:GOR

[PLN20] P. Ed Parnell, Lisa A. Levin, and Michael O. Navarro. Gauging oxygen risk and tolerance for the megafauna of the Southern California shelf based on in situ observation, species mobility, and seascape. ICES Journal of Marine Science, 77(5):1941–1952, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1941/5859710.

Planque:2020:MCN

[PM20] Benjamin Planque and Christian Mullon. Modelling chance and necessity in natural systems. ICES Journal of Marine Science, 77(4):1573–1588, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1573/5584404.

Pol:2023:CCI

[PM23] Mike Pol and Christos D. Maravelias. Cracking the challenges of incentivizing avoidance of unwanted catch. ICES Journal of Marine Science, 80(3):403–406, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/403/7083379.

Perricone:2023:NBB

[PMM^{+}23] V. Perricone, M. Mutalipassi, A. Mele, M. Buono, D. Vicinanza, and P. Contestabile. Nature-based and bioinspired solutions for coastal protection: an overview among key ecosystems and a promising pathway for new functional and sustainable designs. ICES Journal of Marine Science, 80(5): 1218–1239, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/5/1218/7181089.

Pena:2021:CEI

[PMO^{+}21a] Héctor Peña, Gavin J. Macaulay, Egil Ona, Sindre Vatnehol, and Arne J. Holmin. Corrigendum to: Estimating individual fish school biomass using digital omnidirectional sonars, applied to mackerel and herring. ICES Journal of Marine Science, 78(3):1174, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
1174/6270966.

Pena:2021:EIF

[PMO^{+}21b] Héctor Peña, Gavin J. Macaulay, Egil Ona, Sindre Vatnehol, and Arne J. Holmin. Estimating individual fish school biomass using digital omnidirectional sonars, applied to mackerel and herring. ICES Journal of Marine Science, 78 (3):940–951, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/3/940/6104104.

Peschko:2020:WDR

[PMS^{+}20] Verena Peschko, Sabine Müller, Philipp Schwemmer, Moritz Mercker, Peter Lienau, Tanja Rosenberger, Janne Sundermeyer, and Stefan Garthe. Wide dispersal of recently weaned grey seal pups in the Southern North Sea. ICES Journal of Marine Science, 77(5):1762–1771, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1762/5824890.

Proud:2020:ACS

[PMSK^{+}20] Roland Proud, Richard Mangeni-Sande, Robert J. Kayanda, Martin J. Cox, Chrisphine Nyamweya, Collins Ongore, Vianny Natugonza, Inigo Everson, Mboni Elison, Laura Hobbs, Benedicto Boniphace Kashindye, Enock W. Mlaponi, Anthony Taabu-Munyaho, Venny M. Mwainge, Esther Kagoya, Antonio Pegado, Evarist Nduwayesu, and Andrew S. Brierley. Automated classification of schools of the silver cyprinid Rastrineobola argentea in Lake Victoria acoustic survey data using random forests. ICES Journal of Marine Science, 77(4):1379–1390, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1379/5835268.

Phang:2023:RBE

[PMTG^{+}23] Sui Phang, Antaya March, Gregoire Touron-Gardic, Kieran Deane, and Pierre Failler. A review of the blue economy, potential, and opportunities in seven Caribbean nations pre-COVID-19. ICES Journal of Marine Science, 80(8):2233– 2243, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/8/2233/6988643.

Precoda:2024:IFO

[PO24] Kristin Precoda and Christopher D. Orphanides. Impact of fishery observer protocol on estimated bycatch rates of marine mammals. ICES Journal of Marine Science, 81(2): 348–357, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/81/2/348/7516126.

Popper:2020:CF

[Pop20] Arthur N. Popper. Colleagues as friends. ICES Journal of Marine Science, 77(6):2033–2042, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2033/5880820.

Pala:2023:ACI

[POUH23] Ahmet Pala, Anna Oleynik, Ingrid Utseth, and Nils Olav Handegard. Addressing class imbalance in deep learning for acoustic target classification. ICES Journal of Marine Science, 80(10):2530–2544, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
10/2530/7329477.

Pourtois:2022:MEH

[PPMD22] Julie D. Pourtois, Mikaela M. Provost, Fiorenza Micheli, and Giulio A. De Leo. Modelling the effect of habitat and fishing heterogeneity on the performance of a Total Allowable Catch-regulated fishery. ICES Journal of Marine Science, 79(5):1467–1480, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
5/1467/6581604.

Polo:2022:EAD

[PPV^{+}22] J. Polo, A. Punzón, P. Vasilakopoulos, R. Somavilla, and M. Hidalgo. Environmental and anthropogenic driven transitions in the demersal ecosystem of Cantabrian Sea. ICES Journal of Marine Science, 79(7):2017–2031, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2017/6648917.

Phelan:2023:ESD

[PRAR23] Anna (Anya) Phelan, Helen Ross, Dedi S. Adhuri, and Russell Richards. Equity in a sea of debt: how better understanding of small-scale fisheries can help reel in sustainable seafood. ICES Journal of Marine Science, 80(8):2222–2232, October 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/8/2222/6532048.

Pascoe:2022:CTK

[PRCM22] Penelope Pascoe, Ben Raymond, Noel Carmichael, and Julie McInnes. The current trajectory of king penguin (Aptenodytes patagonicus) chick numbers on Macquarie Island in relation to environmental conditions. ICES Journal of Marine Science, 79(7):2084–2092, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2084/6656014.

Pezy:2020:EAS

[PRD20] Jean-Philippe Pezy, Aurore Raoux, and Jean-Claude Dauvin. An ecosystem approach for studying the impact of offshore wind farms: a French case study. ICES Journal of Marine Science, 77(3):1238–1246, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1238/5096674.

Peres:2022:IVS

[PRMS22] Henrique A. Peres, Dominique Robert, Julien Mainguy, and Pascal Sirois. Interannual variability in size-selective winter mortality of young-of-the-year striped bass. ICES Journal of Marine Science, 79(5):1614–1623, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1614/6597020.

Probst:2020:HED

[Pro20] Wolfgang Nikolaus Probst. How emerging data technologies can increase trust and transparency in fisheries. ICES Journal of Marine Science, 77(4):1286–1294, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1286/5380566.

Punzon:2020:HSD

[PRRB^{+}20] Antonio Punzón, Lucia Rueda, Augusto Rodríguez-Basalo, Manuel Hidalgo, Pere Oliver, José Castro, Juan Gil, Antonio Esteban, Luis Gil de Sola, and Enric Massutí. History of the Spanish demersal fishery in the Atlantic and Mediterranean Seas. ICES Journal of Marine Science, 77(2):553– 566, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/553/5671741.

Plowman:2020:SAS

[PTD^{+}20] Caitlin Q. Plowman, Cynthia D. Trowbridge, John Davenport, Colin Little, Luke Harman, and Rob McAllen. Stressed from above and stressed from below: dissolved oxygen fluctuations in Lough Hyne, a semi-enclosed marine lake. ICES Journal of Marine Science, 77(6):2106–2117, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2106/5880831.

Pierson:2022:CHS

[PTR22] James J. Pierson, Jeremy M. Testa, and Michael R. Roman. Copepod habitat suitability estimates vary among oxygen metrics in Chesapeake Bay. ICES Journal of Marine Science, 79(3):855–867, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/3/855/6530391.

Pereniguez:2022:ATA

[PVGC^{+}22] J. M. Pereñíguez, L. A. Venerus, C. Gutiérrez-Cánovas, D. Abecasis, J. E. Ciancio, P. Jiménez-Montalbán, and J. A. García-Charton. Acoustic telemetry and accelerometers: a field comparison of different proxies for activity in the marine environment. ICES Journal of Marine Science, 79(10):2600– 2613, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/10/2600/6822703.

Pascoe:2020:DKD

[PVHV20] Sean Pascoe, Ingrid Van Putten, Eriko Hoshino, and Simon Vieira. Determining key drivers of perceptions of performance of rights-based fisheries in Australia using a Bayesian belief network. ICES Journal of Marine Science, 77(2):803– 814, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/803/5643856.

Peterson:2021:DFA

[PWCL21] Cassidy D. Peterson, Michael J. Wilberg, Enric Cortés, and Robert J. Latour. Dynamic factor analysis to reconcile conflicting survey indices of abundance. ICES Journal of Marine Science, 78(5):1711–1729, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1711/6261976.

Prince:2023:HEL

[PWH23] Jeremy D. Prince, Chris Wilcox, and Norman Hall. How to estimate life history ratios to simplify data-poor fisheries assessment. ICES Journal of Marine Science, 80(10):2619– 2629, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/10/2619/7076233.

Pitois:2022:CMZ

[PY22] Sophie Pitois and Lidia Yebra. Contribution of marine zooplankton time series to the United Nations Decade of Ocean Science for Sustainable Development. ICES Journal of Marine Science, 79(3):722–726, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/722/6555701.

Queffelec:2021:MSP

[QBF^{+}21] Betty Queffelec, Marie Bonnin, Beatrice Ferreira, Sophie Bertrand, Solange Teles Da Silva, Fatou Diouf, Brice Trouillet, Annie Cudennec, Adrien Brunel, Odeline Billant, and Hilde Toonen. Marine spatial planning and the risk of ocean grabbing in the tropical Atlantic. ICES Journal of Marine Science, 78(4):1196–1208, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1196/6154827.

Quinzan:2020:DIF

[QCM^{+}20] M. Quinzán, J. Castro, E. Massutí, L. Rueda, and M. Hidalgo. Disentangling the influence of fishing, demography, and environment on population dynamics of Iberian Peninsula waters fish stocks. ICES Journal of Marine Science, 77 (1):1–11, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/1/1/5675582.

Quintana:2023:SDC

[QMMC^{+}23] Anastasia Quintana, Sergio Marcos, Luis Malpica-Cruz, Lizbeth Tamayo, José Ángel Canto Noh, Francisco Fernández-Rivera Melo, and Stuart Fulton. Socioeconomic dilemmas of commercial markets for invasive species: lessons from lionfish in Mexico. ICES Journal of Marine Science, 80(1): 31–39, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/1/31/6884606.

Qin:2021:GRR

[QSKL21] Le-Zheng Qin, Zhaxi Suonan, Seung Hyeon Kim, and Kun-Seop Lee. Growth and reproductive responses of the seagrass Zostera marina to sediment nutrient enrichment. ICES Journal of Marine Science, 78(3):1160–1173, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/3/1160/6145866.

Qiao:2021:DLM

[QWT^{+}21] Maoying Qiao, Dadong Wang, Geoffrey N. Tuck, L. Richard Little, André E. Punt, and Mike Gerner. Deep learning methods applied to electronic monitoring data: automated catch event detection for longline fishing. ICES Journal of Marine Science, 78(1):25–35, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
25/6053706.

Rijnsdorp:2022:FTE

[RAH^{+}22] A. D. Rijnsdorp, G. Aarts, N. T. Hintzen, J. C. van Rijssel, A. M. Winter, and J. J. Poos. Fishing tactics and the effect of resource depletion and interference during the exploitation of local patches of flatfish. ICES Journal of Marine Science, 79(7):2093–2106, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2093/6661568.

Rubbens:2023:MLM

[RBC^{+}23] Peter Rubbens, Stephanie Brodie, Tristan Cordier, Diogo Destro Barcellos, Paul Devos, Jose A. Fernandes-Salvador, Jennifer I. Fincham, Alessandra Gomes, Nils Olav Handegard, Kerry Howell, Cédric Jamet, Kyrre Heldal Kartveit, Hassan Moustahfid, Clea Parcerisas, Dimitris Politikos, Raphaëlle Sauzède, Maria Sokolova, Laura Uusitalo, Laure Van den Bulcke, Aloysius T. M. van Helmond, Jordan T. Watson, Heather Welch, Oscar Beltran-Perez, Samuel Chaffron, David S. Greenberg, Bernhard Kühn, Rainer Kiko, Madiop Lo, Rubens M. Lopes, Klas Ove Möller, William Michaels, Ahmet Pala, Jean-Baptiste Romagnan, Pia Schuchert, Vahid Seydi, Sebastian Villasante, Ketil Malde, and Jean-Olivier Irisson. Machine learning in marine ecology: an overview of techniques and applications. ICES Journal of Marine Science, 80(7):1829–1853, September 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/7/1829/7236451.

Ruzicka:2020:EER

[RBCD20] James Ruzicka, Richard D. Brodeur, Kristin Cieciel, and Mary Beth Decker. Examining the ecological role of jellyfish in the Eastern Bering Sea. ICES Journal of Marine Science, 77(2):791–802, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/2/791/5680029.

Receveur:2022:WBC

[RBF^{+}22] Aurore Receveur, Martina Bleil, Steffen Funk, Sven Stötera, Ulf Gräwe, Michael Naumann, Cyril Dutheil, and Uwe Krumme. Western Baltic cod in distress: decline in energy reserves since 1977. ICES Journal of Marine Science, 79(4): 1187–1201, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/1187/6547254.

Rasmusson:2020:ETH

[RBG^{+}20] Lina M. Rasmusson, Pimchanok Buapet, Rushingisha George, Martin Gullström, Pontus C. B. Gunnarsson, and Mats Björk. Effects of temperature and hypoxia on respiration, photorespiration, and photosynthesis of seagrass leaves from contrasting temperature regimes. ICES Journal of Marine Science, 77(6):2056–2065, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2056/5869942.

Riley:2022:LRE

[RBH22] Madeline Riley, Corey J. A. Bradshaw, and Charlie Huveneers. Long-range electric deterrents not as effective as personal deterrents for reducing risk of shark bite. ICES Journal of Marine Science, 79(10):2656–2666, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/10/2656/6798393.

Raicevich:2021:SHG

[RCB^{+}21] Saša Raicevich, Bryony A. Caswell, Valerio Bartolino, Massimiliano Cardinale, Tyler D. Eddy, Ioannis Giovos, Ann-Katrien Lescrauwaet, Ruth H. Thurstan, Georg H. Engelhard, and Emily S. Klein. Sidney Holt, a giant in the history of fisheries science who focused on the future: his legacy and challenges for present-day marine scientists. ICES Journal of Marine Science, 78(6):2182–2192, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2182/6157426.

Rossi:2021:FRR

[RCH^{+}21] Steven P. Rossi, Sean P. Cox, Mike O. Hammill, Cornelia E. den Heyer, Douglas P. Swain, Arnaud Mosnier, and Hugues P. Benoît. Forecasting the response of a recovered pinniped population to sustainable harvest strategies that reduce their impact as predators. ICES Journal of Marine Science, 78(5):1804–1814, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1804/6277125.

Rehren:2022:EEI

[RCJ^{+}22] Jennifer Rehren, Marta Coll, Narriman Jiddawi, Lotta Clara Kluger, Omar Omar, Villy Christensen, Maria Grazia Pennino, and Matthias Wolff. Evaluating ecosystem impacts of gear regulations in a data-limited fishery — comparing approaches to estimate predator–prey interactions in Ecopath with Ecosim. ICES Journal of Marine Science, 79(5):1624– 1636, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1624/6603645.

Rowell:2020:DMS

[RDAOE20] Timothy J. Rowell, Gerald L. D’Spain, Octavio Aburto-Oropeza, and Brad E. Erisman. Drivers of male sound production and effective communication distances at fish spawning aggregation sites. ICES Journal of Marine Science, 77 (2):730–745, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/2/730/5707355.

Rijnsdorp:2021:SMB

[RDM^{+}21] A. D. Rijnsdorp, J. Depestele, P. Molenaar, O. R. Eigaard, A. Ivanović, and F. G. O’Neill. Sediment mobilization by bottom trawls: a model approach applied to the Dutch North Sea beam trawl fishery. ICES Journal of Marine Science, 78(5):1574–1586, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1574/6214487.

Roscher:2023:DCS

[RES^{+}23] M. B. Roscher, H. Eriksson, M. Sharp, O. Menaouer, and N. Andrew. Decadal characteristics of small-scale fishing livelihoods in 13 Pacific Island Countries and Territories. ICES Journal of Marine Science, 80(7):1963–1975, September 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/7/1963/7231248.

Rasmuson:2022:CVH

[RFB^{+}22] Leif K. Rasmuson, Stephanie A. Fields, Matthew T. O. Blume, Kelly A. Lawrence, and Polly S. Rankin. Combined video–hydroacoustic survey of nearshore semi-pelagic rockfish in untrawlable habitats. ICES Journal of Marine Science, 79(1):100–116, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/100/6470671.

Rutterford:2023:FIA

[RGE^{+}23] Louise A. Rutterford, Martin J. Genner, Georg H. Engelhard, Stephen D. Simpson, and Ewan Hunter. Fishing impacts on age structure may conceal environmental drivers of body size in exploited fish populations. ICES Journal of Marine Science, 80(4):848–860, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/848/7034401.

Rubio:2022:SPA

[RHO22] Iratxe Rubio, Alistair J. Hobday, and Elena Ojea. Skippers’ preferred adaptation and transformation responses to catch declines in a large-scale tuna fishery. ICES Journal of Marine Science, 79(2):532–539, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/532/6270967.

Rijnsdorp:2020:DBT

[RHvD^{+}20] A. D. Rijnsdorp, J. G. Hiddink, P. D. van Denderen, N. T. Hintzen, O. R. Eigaard, S. Valanko, F. Bastardie, S. G. Bolam, P. Boulcott, J. Egekvist, C. Garcia, G. van Hoey, P. Jonsson, P. Laffargue, J. R. Nielsen, G. J. Piet, M. Sköld, and T. van Kooten. Different bottom trawl fisheries have a differential impact on the status of the North Sea seafloor habitats. ICES Journal of Marine Science, 77(5):1772–1786, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1772/5824898.

Rouse:2020:CFL

[RHW20] Sally Rouse, Peter Hayes, and Thomas A. Wilding. Commercial fisheries losses arising from interactions with offshore pipelines and other oil and gas infrastructure and activities. ICES Journal of Marine Science, 77(3):1148–1156, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/3/1148/5116261.

Rinkevich:2021:EEA

[Rin21] Baruch Rinkevich. Ecological engineering approaches in coral reef restoration. ICES Journal of Marine Science, 78(1):410–420, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/410/5780429.

Reum:2021:NAF

[RKH^{+}21] Jonathan C. P. Reum, Christopher R. Kelble, Chris J. Harvey, Robert P. Wildermuth, Neda Trifonova, Sean M. Lucey, P. Sean McDonald, and Howard Townsend. Network approaches for formalizing conceptual models in ecosystem-based management. ICES Journal of Marine Science, 78(10):3674–3686, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
10/3674/6414826.

Rasmuson:2022:INB

[RMF^{+}22] Leif K. Rasmuson, Scott R. Marion, Stephanie A. Fields, Matthew T. O. Blume, Kelly A. Lawrence, and Polly S.

Rankin. Influence of near bottom fish distribution on the efficacy of a combined hydroacoustic video survey. ICES Journal of Marine Science, 79(7):2069–2083, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2069/6656952.

Rooper:2021:PSG

[ROH^{+}21] Christopher N. Rooper, Ivonne Ortiz, Albert J. Hermann, Ned Laman, Wei Cheng, Kelly Kearney, and Kerim Aydin. Predicted shifts of groundfish distribution in the Eastern Bering Sea under climate change, with implications for fish populations and fisheries management. ICES Journal of Marine Science, 78(1):220–234, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
220/6024437.

Ribeiro:2021:ASM

[RPB^{+}21] Bárbara Ribeiro, André Padua, Adam Barno, Helena Villela, Gustavo Duarte, André Rossi, Flávio da Costa Fernandes, Raquel Peixoto, and Michelle Klautau. Assessing skeleton and microbiome responses of a calcareous sponge under thermal and pH stresses. ICES Journal of Marine Science, 78(3):855–866, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/3/855/6046590.

Rodriguez-Pena:2021:DPM

[RPVT^{+}21] E. Rodríguez-Pena, P. Verísimo, O. Tully, L. Fernández, and A. Martínez-Lage. Differences in prevalence of multiple paternity in the spiny spider crab Maja brachydactyla in two localities that differ in female fecundity, fishing intensity, and management measures. ICES Journal of Marine Science, 78(5):1697–1710, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1697/6257888.

Roach:2022:CEP

[RRJ22] Michael Roach, Andy Revill, and Magnus J. Johnson. Coexistence in practice: a collaborative study of the effects of the Westermost Rough offshore wind development on the size distribution and catch rates of a commercially important lobster (Homarus gammarus) population. ICES Journal of Marine Science, 79(4):1175–1186, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1175/6547885.

Renn:2024:LLB

[RRR^{+}24] Chloe Renn, Sian Rees, Adam Rees, Bede F R Davies, Amy Y Cartwright, Sam Fanshawe, Martin J Attrill, Luke A Holmes, and Emma V Sheehan. Lessons from Lyme Bay (UK) to inform policy, management, and monitoring of Marine Protected Areas. ICES Journal of Marine Science, 81 (2):276–292, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/81/2/276/7542447.

Rohtla:2021:CRI

[RST^{+}21] Mehis Rohtla, Maidu Silm, Jouni Tulonen, Päärn Paiste, Håkan Wickström, Melanie Kielman-Schmitt, Ellen Kooijman, Väino Vaino, Redik Eschbaum, Lauri Saks, Aare Verliin, and Markus Vetemaa. Conservation restocking of the imperilled European eel does not necessarily equal conservation. ICES Journal of Marine Science, 78(1):101–111, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/1/101/6026101.

Ryan:2022:RFP

[RSTL22] Karina L. Ryan, Jenny Shaw, Sean R. Tracey, and Jeremy M. Lyle. Recreational fishers’ perceptions of climate change. ICES Journal of Marine Science, 79(2):540–551, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/540/6380053.

Ruhl:2020:MLS

[RTQW20] Saskia Rühl, Charlie Thompson, Ana M. Queirós, and Stephen Widdicombe. Missing links in the study of solute and particle exchange between the sea floor and water column. ICES Journal of Marine Science, 77(5):1602–1616, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1602/5843856.

Roa-Ureta:2021:ESS

[RUdPFRA^{+}21] Rubén H. Roa-Ureta, M. del Pino Fernández-Rueda, José Luis Acuña, Antonella Rivera, Ricardo González-Gil, and Lucía García-Flórez. Estimation of the spawning stock and recruitment relationship of Octopus vulgaris in Asturias (Bay of Biscay) with generalized depletion models: implications for the applicability of MSY. ICES Journal of Marine Science, 78(6):2256–2270, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2256/6299661.

Roman:2024:SRS

[RVR^{+}24] Salvador Román, Elsa Vázquez, Marta Román, Rosa M Viejo, Nicolás Weidberg, Jesús S Troncoso, Sarah A Woodin, David S Wethey, and Celia Olabarria. The stress response of the seagrass Zostera noltei and three commercial clam species to low salinity associated with heavy rainfall. ICES Journal of Marine Science, 81(2):358–374, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/2/358/7477276. See correction [Ano24b].

Stoeckle:2022:CLP

[SAA^{+}22] Mark Y. Stoeckle, Jason Adolf, Jesse H. Ausubel, Zach Charlop-Powers, Keith J. Dunton, and Greg Hinks. Current laboratory protocols for detecting fish species with environmental DNA optimize sensitivity and reproducibility, especially for more abundant populations. ICES Journal of Marine Science, 79(2):403–412, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/403/6503935.

Satterthwaite:2020:ETF

[SAB^{+}20] William H. Satterthwaite, Kelly S. Andrews, Brian J. Burke, Jennifer L. Gosselin, Correigh M. Greene, Chris J. Harvey, Stuart H. Munsch, Michael R. O’Farrell, Jameal F. Samhouri, and Kathryn L. Sobocinski. Ecological thresholds in forecast performance for key United States West Coast Chinook salmon stocks. ICES Journal of Marine Science, 77(4):1503–1515, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1503/5588520.

Stoeckle:2021:TEA

[SACP^{+}21] Mark Y. Stoeckle, Jason Adolf, Zachary Charlop-Powers, Keith J. Dunton, Gregory Hinks, and Stacy M. VanMorter. Trawl and eDNA assessment of marine fish diversity, seasonality, and relative abundance in coastal New Jersey, USA. ICES Journal of Marine Science, 78(1):293–304, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/1/293/6017541.

Stockin:2022:UCD

[SAM^{+}22] Karen A. Stockin, Christophe Amiot, Laureline Meynier, Cameron Purvin, and Gabriel E. Machovsky-Capuska. Understanding common dolphin and Australasian gannet feeding associations from nutritional and ethological perspectives. ICES Journal of Marine Science, 79(7):2032–2042, September 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/7/2032/6660730.

Saevik:2022:MCR

[SASM22] Pål Næverlid Sævik, Ann-Lisbeth Agnalt, Ole Bent Samuelsen, and Mari Myksvoll. Modelling chemical releases from fish farms: impact zones, dissolution time, and exposure probability. ICES Journal of Marine Science, 79(1): 22–33, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/1/22/6453070.

Stesko:2021:BRK

[SB21] Aleksei V. Stesko and Sergey V. Bakanev. Bycatches of the red king crab in the bottom fish fishery in the Russian waters of the Barents Sea: assessment and regulations. ICES Journal of Marine Science, 78(2):575–583, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/2/
575/5900772.

Sheerin:2022:MMA

[SBA^{+}22] Edel Sheerin, Leigh Barnwall, Esther Abad, Angela Larivain, Daniel Oesterwind, Michael Petroni, Catalina Perales-Raya, Jean-Paul Robin, Ignacio Sobrino, Julio Valeiras, Denise O’Meara, Graham J. Pierce, A. Louise Allcock, and Anne Marie Power. Multi-method approach shows stock structure in Loligo forbesii squid. ICES Journal of Marine Science, 79(4):1159–1174, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
4/1159/6551738.

Strom:2022:BRW

[SBB^{+}22] John Fredrik Strøm, Pål Arne Bjørn, Eirik Emil Bygdnes, Lars Kristiansen, Bjørnar Skjold, and Thomas Bøhn. Behavioural responses of wild anadromous Arctic char experimentally infested in situ with salmon lice. ICES Journal of Marine Science, 79(6):1853–1863, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1853/6617289.

Silburn:2023:BSM

[SBB^{+}23] Briony Silburn, Adil Bakir, Umberto Binetti, Josie Russell, Peter Kohler, Fiona Preston-Whyte, Bryony Meakins, Nanne van Hoytema, Gilbert Andrews, Abel Carrias, and Thomas Maes. A baseline study of macro, meso and micro litter in the Belize River basin, from catchment to coast. ICES Journal of Marine Science, 80(8):2183–2196, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/8/2183/6514214.

Sparholt:2021:CEE

[SBC^{+}21a] Henrik Sparholt, Bjarte Bogstad, Villy Christensen, Jeremy Collie, Rob van Gemert, Ray Hilborn, Jan Horbowy, Daniel Howell, Michael C. Melnychuk, Søren Anker Pedersen, Claus Reedtz Sparrevohn, Gunnar Stefansson, and Petur Steingrund. Corrigendum: Estimating F_{msy} from an ensemble of data sources to account for density dependence in Northeast Atlantic fish stocks. ICES Journal of Marine Science, 78(3):1175, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/

1175/6273054. See [SBC^{+}21b].

Sparholt:2021:EED

[SBC^{+}21b] Henrik Sparholt, Bjarte Bogstad, Villy Christensen, Jeremy Collie, Rob van Gemert, Ray Hilborn, Jan Horbowy, Daniel Howell, Michael C. Melnychuk, Søren Anker Pedersen, Claus Reedtz Sparrevohn, Gunnar Stefansson, and Petur Steingrund. Estimating F_{msy} from an ensemble of data sources to account for density dependence in Northeast Atlantic fish stocks. ICES Journal of Marine Science, 78(1): 55–69, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/1/55/5923788. See corri-

gendum [SBC^{+}21a].

Sokolova:2021:ERT

[SBD^{+}21] Nadezhda Sokolova, Martin Butzin, Flemming Dahlke, Karl Michael Werner, Daniel Balting, Gerrit Lohmann, and Hans-Otto Pörtner. Exploring the role of temperature in observed inter-population differences of Atlantic cod (Gadus morhua) growth with a 4-dimensional modelling approach. ICES Journal of Marine Science, 78(4):1519–1529, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1519/6207631.

Sandvik:2021:DSA

[SBH^{+}21] Anne D. Sandvik, Samantha Bui, Mats Huserbråten, Ørjan Karlsen, Mari S. Myksvoll, Bjørn Ådlandsvik, and Ingrid A. Johnsen. The development of a sustainability assessment indicator and its response to management changes as derived from salmon lice dispersal modelling. ICES Journal of Marine Science, 78(5):1781–1792, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1781/6277122.

Schemmel:2022:ASA

[SBKO22] Eva Schemmel, Erin C. Bohaboy, Michael J. Kinney, and Joseph M. O’Malley. An assessment of sampling approaches for estimating growth from fishery-dependent biological samples. ICES Journal of Marine Science, 79(5):1497–1514, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1497/6581603.

Scheld:2022:ASF

[SBM^{+}22] Andrew M. Scheld, Jennifer Beckensteiner, Daphne M. Munroe, Eric N. Powell, Sarah Borsetti, Eileen E. Hofmann, and John M. Klinck. The Atlantic surfclam fishery and offshore wind energy development: 2. Assessing economic impacts. ICES Journal of Marine Science, 79(6):1801–1814, August 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/6/1801/6611672.

Sobradillo:2021:ESS

[SBPA^{+}21] B. Sobradillo, G. Boyra, I. Pérez-Arjona, U. Martinez, and V. Espinosa. Ex situ and in situ target strength measurements of European anchovy in the Bay of Biscay. ICES Journal of Marine Science, 78(3):782–796, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
782/6114705.

Skold:2022:NTZ

[SBW^{+}22] Mattias Sköld, Patrik Börjesson, Håkan Wennhage, Joakim Hjelm, Johan Lövgren, and Katja Ringdahl. A notake zone and partially protected areas are not enough to save the Kattegat cod, but enhance biomass and abundance of the local fish assemblage. ICES Journal of Marine Science, 79(8):2231–2246, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2231/6726660.

Shellock:2023:CDB

[SCB^{+}23] R. J. Shellock, C. Cvitanovic, N. Badullovich, D. Catto, J. A. DelBene, J. Duggan, D. B. Karcher, A. Ostwald, and P. Tuohy. Crossing disciplinary boundaries: motivations, challenges, and enablers for early career marine researchers moving from natural to social sciences. ICES Journal of Marine Science, 80(1):40–55, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/40/6872748.

Sbragaglia:2020:DMY

[SCCA20] Valerio Sbragaglia, Ricardo A. Correia, Salvatore Coco, and Robert Arlinghaus. Data mining on YouTube reveals fisher group-specific harvesting patterns and social engagement in recreational anglers and spearfishers. ICES Journal of Marine Science, 77(6):2234–2244, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2234/5519069.

Szuwalski:2021:CCF

[SCF^{+}21] Cody Szuwalski, Wei Cheng, Robert Foy, Albert J. Hermann, Anne Hollowed, Kirstin Holsman, Jiwoo Lee, William Stockhausen, and Jie Zheng. Climate change and the future productivity and distribution of crab in the Bering Sea. ICES Journal of Marine Science, 78(2):502–515, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/2/502/5920400.

Suhring:2020:PPF

[SCG^{+}20] R. Sühring, A. Cousins, L. Gregory, C. Moran, A. Papachlimitzou, C. Phillips, R. Rowles, S. Supple, M. Wilczynska, and S. N. R. Birchenough. The past, present, and future of the regulation of offshore chemicals in the North Sea — a United Kingdom perspective. ICES Journal of Marine Science, 77(3):1157–1166, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1157/5578498.

Schneider:2021:BOS

[Sch21a] David C. Schneider. Becoming an ocean scientist: learning from surprise. ICES Journal of Marine Science, 78 (10):3544–3551, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/78/10/3544/6409470.

Snyder:2021:DNI

[SCH^{+}21b] Hunter T. Snyder, Michael E. Cox, Signe Bork Hansen, Christopher Connors, and Sonia Eckstein. Deterrents and nudges improve compliance in Greenland’s Atlantic salmon (Salmo salar) fishery. ICES Journal of Marine Science, 78(8):2809–2817, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2809/6364351.

Shellock:2023:BLO

[SCM^{+}23] R. J. Shellock, C. Cvitanovic, M. C. McKinnon, M. Mackay, I. E. van Putten, J. Blythe, R. Kelly, P. Tuohy, K. M. Maltby, S. Mynott, N. Simmonds, M. Bailey, A. Begossi, B. Crona, K. A. Fakoya, B. P. Ferreira, A. J. G. Ferrer, K. Frangoudes, J. Gobin, H. C. Goh, P. Haapasaari, B. D. Hardesty, V. Häussermann, K. Hoareau, A-K Hornidge, M. Isaacs, M. Kraan, Y. Li, M. Liu, P. F. M. Lopes, M. Mlakar, T. H. Morrison, H. A. Oxenford, G. Pecl, J. Penca, C. Robinson, S. A. Selim, M. Skern-Mauritzen, K. Soejima, D. Soto, A. K. Spalding, A. Vadrot, N. Vaidianu, M. Webber, and M. S. Wisz. Building leaders for the UN Ocean Science Decade: a guide to supporting early career women researchers within academic marine research institutions. ICES Journal of Marine Science, 80(1):56–75, January 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/1/56/6916918.

Subramaniam:2022:SEF

[SCMT^{+}22] Roshni C. Subramaniam, Stuart P. Corney, Jessica Melbourne-Thomas, Clara Péron, Philippe Ziegler, and Kerrie M. Swadling. Spatially explicit food web modelling to consider fisheries impacts and ecosystem representation within Marine Protected Areas on the Kerguelen Plateau. ICES Journal of Marine Science, 79(4):1327–1339, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
4/1327/6572845.

Sokolova:2023:IEE

[SCN^{+}23] Maria Sokolova, Manuel Cordova, Henk Nap, Aloysius van Helmond, Michiel Mans, Arjan Vroegop, Angelo Mencarelli, and Gert Kootstra. An integrated end-to-end deep neural network for automated detection of discarded fish species and their weight estimation. ICES Journal of Marine Science, 80(7):1911–1922, September 2023.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1911/7240285.

Stamp:2021:IJS

[SCP^{+}21] Thomas Stamp, David Clarke, Shaun Plenty, Tim Robbins, James E. Stewart, Elizabeth West, and Emma Sheehan. Identifying juvenile and sub-adult movements to inform recovery strategies for a high value fishery — European bass (Dicentrarchus labrax). ICES Journal of Marine Science, 78(9):3121–3134, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3121/6370941.

Sheehan:2020:DEA

[SCW^{+}20] E. V. Sheehan, A. Y. Cartwright, M. J. Witt, M. J. Attrill, M. Vural, and L. A. Holmes. Development of epibenthic assemblages on artificial habitat associated with marine renewable infrastructure. ICES Journal of Marine Science, 77(3): 1178–1189, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/3/1178/5137117.

Sydeman:2020:SFY

[SDGR^{+}20] William J. Sydeman, Simon Dedman, Marisol García-Reyes, Sarah Ann Thompson, Julie A. Thayer, Andrew Bakun, and Alec D. MacCall. Sixty-five years of northern anchovy population studies in the southern California Current: a review and suggestion for sensible management. ICES Journal of Marine Science, 77(2):486–499, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
486/5734670.

Suca:2021:EDT

[SDR^{+}21] Justin J. Suca, Jonathan J. Deroba, David E. Richardson, Rubao Ji, and Joel K. Llopiz. Environmental drivers and trends in forage fish occupancy of the Northeast US shelf. ICES Journal of Marine Science, 78(10):3687–3708, December 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/10/3687/6419780.

Semmar:2020:HGR

[SdS20] Nabil Semmar and André M. Vaz dos Santos. Highlighting growth regulation processes in fish populations by a simplex simulation approach: application to Merluccius hubbsi stocks in the Southwestern Atlantic. ICES Journal of Marine Science, 77(4):1401–1413, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/4/
1401/5681886.

Sandvik:2021:EWC

[SDSMS21] Anne D. Sandvik, Sussie Dalvin, Rasmus Skern-Mauritzen, and Morten D. Skogen. The effect of a warmer climate on the salmon lice infection pressure from Norwegian aquaculture. ICES Journal of Marine Science, 78(5):1849–1859, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/5/1849/6280175.

Semmouri:2021:MRH

[SDW^{+}21] Ilias Semmouri, Karel A. C. De Schamphelaere, Stijn Willemse, Michiel B. Vandegehuchte, Colin R. Janssen, and Jana Asselman. Metabarcoding reveals hidden species and improves identification of marine zooplankton communities in the North Sea. ICES Journal of Marine Science, 78(9):3411–3427, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
9/3411/6154330.

Stokesbury:2022:FCI

[SFG22] Kevin D. E. Stokesbury, Gavin Fay, and Robert Griffin. A framework for categorizing the interactions of offshore windfarms and fisheries. ICES Journal of Marine Science, 79(6): 1711–1718, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/6/1711/6633757.

Schijns:2021:FCC

[SFHP21] Rebecca Schijns, Rainer Froese, Jeffrey A. Hutchings, and Daniel Pauly. Five centuries of cod catches in Eastern Canada. ICES Journal of Marine Science, 78(8):2675– 2683, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/8/2675/6359257. See com-

ment [HHN22] and reply [SFHP22].

Schijns:2022:RHE

[SFHP22] Rebecca Schijns, Rainer Froese, Jeffrey A. Hutchings, and Daniel Pauly. Reply to Holm et al. 2022, “Comment on ‘Five centuries of cod catches in eastern Canada,’ by Schijns et al.”. ICES Journal of Marine Science, 79(5):1708, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1708/6593173. See [SFHP21, HHN22].

Soto:2023:EML

[SFPP^{+}23] María Soto, Lourdes Fernández-Peralta, Maria Grazia Pennino, Alexandros Kokkalis, Javier Rey, Francisca Salmerón, María Liébana, Beyah Meissa, and Laurie Kell. Effects of misreporting landings, discards, and Catch Per Unit of Effort index in state-space production models: the case of black hake in northwest Africa. ICES Journal of Marine Science, 80(10):2591–2605, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/10/2591/6827054.

Schadeberg:2022:PPN

[SFW^{+}22] Amanda Schadeberg, Eleanor Ford, Alina M. Wieczorek, Louise C. Gammage, María López-Acosta, Ivana Buselic, Timotej Turk Dermastia, Marcos Fontela, Cristina Galobart, Natalia Llopis Monferrer, Marek Lubośny, Stefania Piarulli, and Giuseppe Suaria. Productivity, pressure, and new perspectives: impacts of the COVID-19 pandemic on marine early-career researchers. ICES Journal of Marine Science, 79(8):2298–2310, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/8/2298/6748269.

Shephard:2021:WAS

[SG21] Samuel Shephard and Patrick Gargan. Wild Atlantic salmon exposed to sea lice from aquaculture show reduced marine survival and modified response to ocean climate. ICES Journal of Marine Science, 78(1):368–376, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/1/368/5864917.

Sciascia:2022:LDS

[SGM22] Roberta Sciascia, Katell Guizien, and Marcello G. Magaldi. Larval dispersal simulations and connectivity predictions for Mediterranean gorgonian species: sensitivity to flow representation and biological traits. ICES Journal of Marine Science, 79(7):2043–2054, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2043/6660729.

Strand:2024:RTV

[SGM^{+}24] Nina Strand, Kevin A Glover, Sonnich Meier, Fernando Ayllon, Vidar Wennevik, Abdullah Madhun, Øystein Skaala, Kristin Hamre, Per Gunnar Fjelldal, Tom Hansen, Eero Niemelä, Sofie Knutar, Per Tommy Fjeldheim, and Monica F Solberg. Regional and temporal variation in escape history of Norwegian farmed Atlantic salmon. ICES Journal of Marine Science, 81(1):119–129, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/81/1/119/7456025.

Saborowski:2022:HPB

[SH22] Reinhard Saborowski and Kim Hünerlage. Hatching phenology of the brown shrimp Crangon crangon in the southern North Sea: inter-annual temperature variations and climate change effects. ICES Journal of Marine Science, 79(4):1302– 1311, May 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/4/1302/6561889.

Sharov:2021:UBF

[Sha21] Alexei Sharov. The unknown baranov. forty years of polemics over the formal theory of the life of fishes. ICES Journal of Marine Science, 78(2):743–754, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/2/743/5859740.

Stenevik:2022:PDD

[SHE^{+}22] Erling Kåre Stenevik, Sondre Hølleland, Katja Enberg, Åge Høines, Are Salthaug, Aril Slotte, Sindre Vatnehol, and Sondre Aanes. Predicting density-dependent somatic growth in Norwegian spring-spawning herring. ICES Journal of Marine Science, 79(4):1340–1352, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1340/6565776.

Savina:2022:NME

[SHFK22] Esther Savina, Bent Herrmann, Rikke P. Frandsen, and Ludvig A. Krag. A new method for estimating length-dependent capture modes in gillnets: a case study in the Danish cod (Gadus morhua) fishery. ICES Journal of Marine Science, 79 (2):373–381, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/2/373/6503932.

Skov:2021:EOU

[SHG^{+}21] Christian Skov, Kieran Hyder, Casper Gundelund, Anssi Ahvonen, Jérôme Baudrier, Trude Borch, Sara deCarvalho, Karim Erzini, Keno Ferter, Fabio Grati, Tessa van der-Hammen, Jan Hinriksson, Rob Houtman, Anders Kagervall, Kostas Kapiris, Martin Karlsson, Adam M. Lejk, Jeremy M. Lyle, Roi Martinez-Escauriaza, Pentti Moilanen, Estanis Mugerza, Hans Jakob Olesen, Anastasios Papadopoulos, Pablo Pita, João Pontes, Zachary Radford, Krzysztof Radtke, Mafalda Rangel, Oscar Sagué, Hege A. Sande, Harry V. Strehlow, Rūdolfs Tutiņš, Pedro Veiga, Thomas Verleye, Jon Helge Vølstad, Joseph W. Watson, Marc Simon Weltersbach, Didzis Ustups, and Paul A. Venturelli. Expert opinion on using angler smartphone apps to inform marine fisheries management: status, prospects, and needs. ICES Journal of Marine Science, 78(3):967– 978, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/967/6104102.

Synnes:2021:LRA

[SHK^{+}21] Ann-Elin Wårøy Synnes, Mats Huserbråten, Halvor Knutsen, Per Erik Jorde, Marte Sodeland, and Even Moland. Local recruitment of Atlantic cod and putative source spawning areas in a coastal seascape. ICES Journal of Marine Science, 78(10):3767–3779, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/78/
10/3767/6440309.

Sakamoto:2023:SIE

[SHK^{+}23] Tatsuya Sakamoto, Sachiko Horii, Taketoshi Kodama, Kazutaka Takahashi, Atsushi Tawa, Yosuke Tanaka, and Seiji Ohshmio. Stable isotopes in eye lenses reveal migration and mixing patterns of diamond squid in the western North Pacific and its marginal seas. ICES Journal of Marine Science, 80(9):2313–2328, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/9/2313/7286231. See correction [Ano24a].

Sydeman:2021:SAE

[SHP^{+}21] William J. Sydeman, George L. Hunt, Ellen K. Pikitch, Julia K. Parrish, John F. Piatt, P. Dee Boersma, Les Kaufman, Daniel W. Anderson, Sarah Ann Thompson, and Richard B. Sherley. South Africa’s experimental fisheries closures and recovery of the endangered African penguin. ICES Journal of Marine Science, 78(10):3538–3543, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3538/6429113. See comment [BRG22].

Sydeman:2022:APL

[SHP^{+}22] W. J. Sydeman, G. L. Hunt, Jr., E. K. Pikitch, J. K. Parrish, J. F. Piatt, P. D. Boersma, L. Kaufman, D. W. Anderson, S. A. Thompson, and R. B. Sherley. African penguins and localized fisheries management: Response to Butterworth and Ross-Gillespie. ICES Journal of Marine Science, 79(6):1972–1978, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/

6/1972/6640472. See [SHP^{+}21, BRG22].

Stock:2021:PRR

[SHW^{+}21] Brian C. Stock, Scott A. Heppell, Lynn Waterhouse, India C. Dove, Christy V. Pattengill-Semmens, Croy M. Mc-Coy, Phillippe G. Bush, Gina Ebanks-Petrie, and Brice X. Semmens. Pulse recruitment and recovery of Cayman Islands Nassau Grouper (Epinephelus striatus) spawning aggregations revealed by in situ length-frequency data. ICES Journal of Marine Science, 78(1):277–292, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/1/277/6071467.

Sandvik:2020:PSL

[SJM^{+}20] Anne D. Sandvik, Ingrid A. Johnsen, Mari S. Myksvoll, Pål N. Sævik, and Morten D. Skogen. Prediction of the salmon lice infestation pressure in a Norwegian fjord. ICES Journal of Marine Science, 77(2):746–756, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/2/746/5704435.

Slesinger:2021:SPR

[SJS21] Emily Slesinger, Olaf P. Jensen, and Grace Saba. Spawning phenology of a rapidly shifting marine fish species throughout its range. ICES Journal of Marine Science, 78(3):1010– 1022, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/1010/6118442.

Song:2020:SHS

[SJX^{+}20] Hongjun Song, Rubao Ji, Ming Xin, Ping Liu, Zhaohui Zhang, and Zongling Wang. Spatial heterogeneity of seasonal phytoplankton blooms in a marginal sea: physical drivers and biological responses. ICES Journal of Marine Science, 77(1):408–418, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/77/1/408/5572610.

Scoulding:2021:ICT

[SK21] Ben Scoulding and Rudy Kloser. Industry-collected target strength of high seas orange roughy in the Indian Ocean. ICES Journal of Marine Science, 78(6):2120–2131, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/2120/5880834.

Sivle:2015:PPL

[SKA15] L. D. Sivle, P. H. Kvadsheim, and M. A. Ainslie. Potential for population-level disturbance by active sonar in herring. ICES Journal of Marine Science, 72(2):558–567, January 2015. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/72/2/558/2801451. See erratum [SKA22].

Sivle:2022:EPP

[SKA22] L. D. Sivle, P. H. Kvadsheim, and M. A. Ainslie. Erratum to: Potential for population-level disturbance by active sonar in herring. ICES Journal of Marine Science, 79(1):244, January 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/1/244/6459379. See [SKA15].

Schadeberg:2021:BMS

[SKH21] Amanda Schadeberg, Marloes Kraan, and Katell G. Hamon. Beyond métiers: social factors influence fisher behaviour. ICES Journal of Marine Science, 78(4):1530–1541, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1530/6207634.

Smith:2022:SLP

[SKM^{+}22] Jennifer E. Smith, John Keane, Craig Mundy, Caleb Gardner, and Michael Oellermann. Spiny lobsters prefer native prey over range-extending invasive urchins. ICES Journal of Marine Science, 79(4):1353–1362, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1353/6565266.

Schaal:2021:CTP

[SL21] Sara M. Schaal and Katie E. Lotterhos. Comparative thermal performance among four young-of-the-year temperate reef fish species. ICES Journal of Marine Science, 78(5): 1684–1696, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/5/1684/6237190.

Sabadin:2020:TRC

[SLB^{+}20] D. E. Sabadin, L. O. Lucifora, S. A. Barbini, D. E. Figueroa, and M. Kittlein. Towards regionalization of the chondrichthyan fauna of the Southwest Atlantic: a spatial framework for conservation planning. ICES Journal of Marine Science, 77(5):1893–1905, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1893/5837621.

Sun:2021:RDF

[SLRC21] Ming Sun, Yunzhou Li, Yiping Ren, and Yong Chen. Rebuilding depleted fisheries towards BMSY under uncertainty: harvest control rules outperform combined management measures. ICES Journal of Marine Science, 78(6):2218– 2232, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/6/2218/5868120.

Sando:2021:BSP

[SMB^{+}21] Anne Britt Sandø, Erik Askov Mousing, W. P. Budgell, Solfrid S. Hjøllo, Morten D. Skogen, and B. Ådlandsvik. Barents Sea plankton production and controlling factors in a fluctuating climate. ICES Journal of Marine Science, 78(6):1999– 2016, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/6/1999/6297834.

Sathianandan:2021:SIM

[SMJ^{+}21] T. V. Sathianandan, Kolliyil Sunil Mohamed, Jayaraman Jayasankar, Somy Kuriakose, K. G. Mini, Eldho Varghese, P. U. Zacharia, P. Kaladharan, T. M. Najmudeen, Mohammed K. Koya, Geetha Sasikumar, Vivekanand Bharti, Prathibha Rohit, G. Maheswarudu, K. Augustine Sindhu, V. Sreepriya, Joseph Alphonsa, and A. Deepthi. Status of Indian marine fish stocks: modelling stock biomass dynamics in multigear fisheries. ICES Journal of Marine Science, 78(5):1744–1757, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1744/6268978.

Steins:2023:BAN

[SMK23] Nathalie A. Steins, Alessa L. Mattens, and Marloes Kraan. Being able is not necessarily being willing: governance implications of social, policy, and science-related factors influencing uptake of selective gear. ICES Journal of Marine Science, 80(3):469–482, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/3/469/6530388.

Skern-Mauritzen:2022:MMC

[SMLB^{+}22] Mette Skern-Mauritzen, Ulf Lindstrøm, Martin Biuw, Bjarki Elvarsson, Thorvaldur Gunnlaugsson, Tore Haug, Kit M. Kovacs, Christian Lydersen, Margaret M. McBride, Bjarni Mikkelsen, Nils Øien, and Gísli Víkingsson. Marine mammal consumption and fisheries removals in the Nordic and Barents Seas. ICES Journal of Marine Science, 79(5):1583– 1603, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1583/6594656.

Scoulding:2022:EAB

[SMO22] Ben Scoulding, Kylie Maguire, and Eric C. Orenstein. Evaluating automated benthic fish detection under variable conditions. ICES Journal of Marine Science, 79(8):2204–2216, October 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/8/2204/6724460.

Sainte-Marie:2021:DOM

[SMOD21] Bernard Sainte-Marie, Jean-François Ouellet, and Hélène Dionne. Direction, ontogeny, and mechanism of the temperature-size rule operating in a large marine crab, Chionoecetes opilio. ICES Journal of Marine Science, 78(2):557– 574, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/2/557/5998278.

Skaret:2023:DBE

[SMP^{+}23] G. Skaret, G. J. Macaulay, R. Pedersen, X. Wang, T. A. Klevjer, L. A. Krag, and B. A. Krafft. Distribution and biomass estimation of Antarctic krill (Euphausia superba) off the South Orkney Islands during 2011–2020. ICES Journal of Marine Science, 80(5):1472–1486, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1472/7161638.

Snyder:2021:CMN

[SMS^{+}21] Richard A. Snyder, Joseph A. Moss, Luciana Santoferrara, Marie Head, and Wade H. Jeffrey. Ciliate microzooplankton from the Northeastern Gulf of Mexico. ICES Journal of Marine Science, 78(9):3356–3371, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3356/6122863.

Schram:2022:EES

[SMS^{+}22] Edward Schram, Pieke Molenaar, Maarten Soetaert, Dirk Burggraaf, Pim G. Boute, Martin J. Lankheet, and Adriaan D. Rijnsdorp. Effect of electrical stimulation used in the pulse trawl fishery for common sole on internal injuries in sandeels. ICES Journal of Marine Science, 79(5):1561– 1568, July 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/5/1561/6585624.

Silva:2023:ASD

[SMSP^{+}23] D. Silva, R. Menezes, B. Serra-Pereira, M. Azevedo, and I. Figueiredo. Adapting the sampling design of research surveys to improve the biomass estimation of secondary species — the case study of Raja clavata. ICES Journal of Marine Science, 80(10):2567–2578, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/10/2567/7190299.

Souza:2023:PPA

[SOB23] P. M. Souza, Jr., Z. Olsen, and S. J. Brandl. Paired passive acoustic and gillnet sampling reveal the utility of bioacoustics for monitoring fish populations in a turbid estuary. ICES Journal of Marine Science, 80(5):1240–1255, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1240/7179496.

Sguotti:2020:NLS

[SOC^{+}20] Camilla Sguotti, Saskia A. Otto, Xochitl Cormon, Karl M. Werner, Ethan Deyle, George Sugihara, and Christian Möllmann. Non-linearity in stock–recruitment relationships of Atlantic cod: insights from a multi-model approach. ICES Journal of Marine Science, 77(4):1492–1502, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/4/1492/5522545.

Strand:2022:TTO

[SOCN^{+}22] Mia Strand, Kelly Ortega-Cisneros, Holly J. Niner, Michel Wahome, James Bell, Jock C. Currie, Hashali Hamukuaya, Giulia La Bianca, Alana M. S. N. Lancaster, Ntemesha Maseka, Lisa McDonald, Kirsty McQuaid, Marly M. Samuel, and Alexander Winkler. Transdisciplinarity in transformative ocean governance research — reflections of early career researchers. ICES Journal of Marine Science, 79(8):2163– 2177, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/8/2163/6698563.

Smolinski:2022:CSC

[SOO^{+}22] Szymon Smoliński, Daniel Ottmann, Okko Outinen, Amanda Schadeberg, Valentina Melli, Lara Funk, Côme Denechaud, Alina Wieczorek, Alessandro Orio, Robert Mussgnug, Rasa Morkūnė, Emilie Hernes Vereide, Maja Zdulska, Genevieve Phillips, Fedor Lishchenko, and Greta Srėbalienė. Counting stars: contribution of early career scientists to marine and fisheries sciences. ICES Journal of Marine Science, 79(9):2351–2361, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
9/2351/6795149.

Sorochan:2021:ASA

[SPBJ21] K. A. Sorochan, S. Plourde, M. F. Baumgartner, and C. L. Johnson. Availability, supply, and aggregation of prey (Calanus spp.) in foraging areas of the North Atlantic right whale (Eubalaena glacialis). ICES Journal of Marine Science, 78(10):3498–3520, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3498/6410484.

Sande:2022:FFR

[SPd^{+}22] Hege Sande, Nuno Prista, Annica de Groote, Michele Casini, Cynthia Jones, and Andreas Sundelöf. Frameless — finding and refining a sampling frame for surveying recreational fisheries: lessons from estimating Swedish harvest of western Baltic cod. ICES Journal of Marine Science, 79(4):1217– 1231, May 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/4/1217/6554382.

Sorochan:2023:NBA

[SPJ23] K. A. Sorochan, S. Plourde, and C. L. Johnson. Near-bottom aggregations of Calanus spp. copepods in the southern Gulf of St. Lawrence in summer: significance for North Atlantic right whale foraging. ICES Journal of Marine Science, 80 (4):787–802, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/4/787/7034396.

Surma:2021:TOU

[SPP21] Szymon Surma, Tony J. Pitcher, and Evgeny A. Pakhomov. Trade-offs and uncertainties in Northeast Pacific herring fisheries: ecosystem modelling and management strategy evaluation. ICES Journal of Marine Science, 78(6):2280–2297, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/2280/6324220.

Stephenson:2021:POH

[SRA^{+}21] Fabrice Stephenson, Ashley A. Rowden, Owen F. Anderson, C. Roland Pitcher, Matt H. Pinkerton, Grady Petersen, and David A. Bowden. Presence-only habitat suitability models for vulnerable marine ecosystem indicator taxa in the South Pacific have reached their predictive limit. ICES Journal of Marine Science, 78(8):2830–2843, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2830/6355116.

SalesHenriques:2023:AMQ

[SRB^{+}23] Nuno Sales Henriques, Tommaso Russo, Luis Bentes, Pedro Monteiro, Antonio Parisi, Ramiro Magno, Frederico Oliveira, Karim Erzini, and Jorge M. S. Gonçalves. An approach to map and quantify the fishing effort of polyvalent passive gear fishing fleets using geospatial data. ICES Journal of Marine Science, 80(6):1658–1669, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1658/7197498.

Shirk:2023:PBC

[SRBT23] Philip L. Shirk, Kate Richerson, Michael Banks, and Vanessa Tuttle. Predicting bycatch of Chinook salmon in the Pacific hake fishery using spatiotemporal models. ICES Journal of Marine Science, 80(1):133–144, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/133/6891089.

Speir:2020:MHC

[SRM^{+}20] Cameron Speir, Corey Ridings, Jennifer Marcum, Michael Drexler, and Karma Norman. Measuring health conditions and behaviours in fishing industry participants and fishing communities using the Behavioral Risk Factor Surveillance Survey (BRFSS). ICES Journal of Marine Science, 77(5):1830–1840, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1830/5835263.

Suthers:2021:NFI

[SRMS21] Iain M. Suthers, Dennis D. Reid, Erlend Moksness, and Hayden T. Schilling. Novel fisheries investigations by harald dannevig: some parallels with Johan Hjort on the other side of the world. ICES Journal of Marine Science, 78(2):755–764, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/2/755/5715944.

Shephard:2021:ULE

[SROR21] Samuel Shephard, Diarmuid Ryan, Paul O’Reilly, and Willie Roche. Using local ecological knowledge to inform semi-quantitative fishery surveillance indicators: an example in marine recreational angling. ICES Journal of Marine Science, 78(10):3805–3816, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3805/6430631.

Szymkowiak:2021:EEQ

[SRR21] Marysia Szymkowiak and Melissa Rhodes-Reese. Examining the expansion of qualitative network models towards integrating multifaceted human dimensions. ICES Journal of Marine Science, 78(7):2395–2408, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2395/6297552.

Saavedra:2021:CCD

[SSBV21] Luisa M. Saavedra, Gonzalo S. Saldías, Bernardo R. Broitman, and Cristian A. Vargas. Carbonate chemistry dynamics in shellfish farming areas along the Chilean coast: natural ranges and biological implications. ICES Journal of Marine Science, 78(1):323–339, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
323/5902158.

Suca:2022:TUD

[SSF^{+}22] Justin J. Suca, Jarrod A. Santora, John C. Field, K. Alexandra Curtis, Barbara A. Muhling, Megan A. Cimino, Elliott L. Hazen, and Steven J. Bograd. Temperature and upwelling dynamics drive market squid (Doryteuthis opalescens) distribution and abundance in the California Current. ICES Journal of Marine Science, 79(9):2489–2509, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/9/2489/6780182.

Sun:2023:EBH

[SSF^{+}23] Runlong Sun, Peng Sun, Caihong Fu, Guankui Liu, Zhenlin Liang, Yunne-Jai Shin, Nicolas Barrier, and Yongjun Tian. Exploring balanced harvest as a potential strategy for highly exploited multispecies fisheries. ICES Journal of Marine Science, 80(4):897–910, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/4/897/7048540.

Schulte:2020:STD

[SSH^{+}20] K. F. Schulte, V. Siegel, M. Hufnagl, T. Schulze, and A. Temming. Spatial and temporal distribution patterns of brown shrimp (Crangon crangon) derived from commercial logbook, landings, and vessel monitoring data. ICES Journal of Marine Science, 77(3):1017–1032, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1017/5780427.

Seri:2023:CCC

[SSP^{+}23] Sai Geetha Seri, Matthew Edward Schinault, Seth Michael Penna, Chenyang Zhu, Lise Doksæter Sivle, Karen de Jong, Nils Olav Handegard, and Purnima Ratilal. Characterizing coastal cod vocalization using a towed hydrophone array. ICES Journal of Marine Science, 80(6):1727–1745, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1727/7218794.

Salman:2020:AFD

[SSS^{+}20] Ahmad Salman, Shoaib Ahmad Siddiqui, Faisal Shafait, Ajmal Mian, Mark R. Shortis, Khawar Khurshid, Adrian Ulges, and Ulrich Schwanecke. Automatic fish detection in underwater videos by a deep neural network-based hybrid motion learning system. ICES Journal of Marine Science, 77(4): 1295–1307, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/4/1295/5366225.

Swieca:2023:GDL

[SSS^{+}23] K. Swieca, S. Sponaugle, M. S. Schmid, J. Ivory, M. Corrales-Ugalde, K. R. Sutherland, and R. K. Cowen. Growth and diet of a larval myctophid across distinct upwelling regimes in the California Current. ICES Journal of Marine Science, 80(5):1431–1446, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1431/7151267.

Svedang:2022:DEB

[SSV^{+}22a] Henrik Svedäng, Oleg Savchuk, Anna Villnäs, Alf Norkko, Bo G. Gustafsson, Sofia A. Wikström, and Christoph Humborg. On the decline of eastern Baltic cod: we need to take more holistic views into account. reply to Brander (2022) comment on Svedäng et al. (2022). ICES Journal of Marine Science, 79(7):2157–2158, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/7/2157/6571567. See [SSV^{+}22b, Bra22].

Svedang:2022:RTE

[SSV^{+}22b] Henrik Svedäng, Oleg Savchuk, Anna Villnäs, Alf Norkko, Bo G. Gustafsson, Sofia A. Wikström, and Christoph Humborg. Re-thinking the “ecological envelope” of Eastern Baltic cod (Gadus morhua): conditions for productivity, reproduction, and feeding over time. ICES Journal of Marine Science, 79(3):689–708, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/689/6532053. See comment [Bra22] and re-

ply [SSV^{+}22a].

Stafford:2022:CSE

[Sta22] R. Stafford. Comment on “Seaweed ecosystems may not mitigate CO_{2} emissions” by Gallagher et al. (2022). ICES Journal of Marine Science, 79(5):1701–1702, July 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/5/1701/6589472. See [GSL22] and reply

[GS22].

Stevens:2021:UDT

[Ste21] Bradley G. Stevens. The ups and downs of traps: environmental impacts, entanglement, mitigation, and the future of trap fishing for crustaceans and fish. ICES Journal of Marine Science, 78(2):584–596, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/2/
584/5896070.

Sassa:2020:REYb

[STK^{+}20a] Chiyuki Sassa, Motomitsu Takahashi, Yoshinobu Konishi, Yoshimasa Aonuma, and Youichi Tsukamoto. The rapid expansion of yellowtail (Seriola quinqueradiata) spawning ground in the East China Sea is linked to increasing recruitment and spawning stock biomass. ICES Journal of Marine Science, 77(3):1249, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1249/5801147.

Sassa:2020:REYa

[STK^{+}20b] Chiyuki Sassa, Motomitsu Takahashi, Yoshinobu Konishi, Aonuma Yoshimasa, and Youichi Tsukamoto. The rapid expansion of yellowtail (Seriola quinqueradiata) spawning ground in the East China Sea is linked to increasing recruitment and spawning stock biomass. ICES Journal of Marine Science, 77(2):581–592, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
581/5662382.

Selden:2020:CCB

[STS^{+}20] Rebecca L. Selden, James T. Thorson, Jameal F. Samhouri, Steven J. Bograd, Stephanie Brodie, Gemma Carroll, Melissa A. Haltuch, Elliott L. Hazen, Kirstin K. Holsman, Malin L. Pinsky, Nick Tolimieri, and Ellen Willis-Norton. Coupled changes in biomass and distribution drive trends in availability of fish stocks to US West Coast ports. ICES Journal of Marine Science, 77(1):188–199, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/1/188/5614897.

Suuronen:2022:UPB

[Suu22] Petri Suuronen. Understanding perspectives and barriers that affect fishers’ responses to bycatch reduction technologies. ICES Journal of Marine Science, 79(4):1015–1023, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1015/6555700.

Silvar-Viladomiu:2022:ERI

[SVBM^{+}22] Paula Silvar-Viladomiu, Luke Batts, Cóilín Minto, David Miller, and Colm Lordan. An empirical review of ICES reference points. ICES Journal of Marine Science, 79(10):2563– 2578, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/10/2563/6812839. See cor-

rection [Ano23a].

Silvar-Viladomiu:2022:PPM

[SVMBR22] Paula Silvar-Viladomiu, Cóilín Minto, Deirdre Brophy, and David G. Reid. Peterman’s productivity method for estimating dynamic reference points in changing ecosystems. ICES Journal of Marine Science, 79(4):1034–1047, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1034/6549169.

Silvar-Viladomiu:2023:SMS

[SVML^{+}23] Paula Silvar-Viladomiu, Cóilín Minto, Colm Lordan, Deirdre Brophy, Rich Bell, Jeremy Collie, and David Reid. Stochastic modelling and synthesis of dynamic fish recruitment productivity in the Celtic Seas ecoregion. ICES Journal of Marine Science, 80(9):2329–2341, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/9/2329/7282467.

Strong:2023:MHM

[SWH^{+}23] J. A. Strong, C. Wardell, A. Haïssoune, A. L. Jones, and L. Coals. Marine habitat mapping to support the use of conservation and anti-trawl structures in Kep Province, Cambodia. ICES Journal of Marine Science, 80(8):2197–2209, October 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/8/2197/6511220.

Sojitra:2022:MDC

[SWLW22] Milan Sojitra, Eric J. Woehler, Mary-Anne Lea, and Simon Wotherspoon. Multi-decadal changes in the at-sea distribution and abundance of black-browed and light-mantled sooty albatrosses in the southwest Pacific Ocean. ICES Journal of Marine Science, 79(10):2630–2642, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/10/2630/6808343.

Stamp:2022:LSH

[SWR^{+}22] Thomas Stamp, Elizabeth West, Tim Robbins, Shaun Plenty, and Emma Sheehan. Large-scale historic habitat loss in estuaries and its implications for commercial and recreational fin fisheries. ICES Journal of Marine Science, 79(7):1981–1991, September 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
7/1981/6687832.

Suca:2021:SSL

[SWS^{+}21] Justin J. Suca, David N. Wiley, Tammy L. Silva, Anna R. Robuck, David E. Richardson, Sarah G. Glancy, Emily Clancey, Teresa Giandonato, Andrew R. Solow, Michael A. Thompson, Peter Hong, Hannes Baumann, Les Kaufman, and Joel K. Llopiz. Sensitivity of sand lance to shifting prey and hydrography indicates forthcoming change to the northeast US shelf forage fish complex. ICES Journal of Marine Science, 78(3):1023–1037, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
1023/6120244.

Thorson:2020:SIV

[TAB^{+}20] James T. Thorson, Charles F. Adams, Elizabeth N. Brooks, Lisa B. Eisner, David G. Kimmel, Christopher M. Legault, Lauren A. Rogers, and Ellen M. Yasumiishi. Seasonal and interannual variation in spatio-temporal models for index standardization and phenology studies. ICES Journal of Marine Science, 77(5):1879–1892, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1879/5837191.

Townhill:2022:ECE

[TAPB22] B. L. Townhill, Y. Artioli, J. K. Pinnegar, and S. N. R. Birchenough. Exposure of commercially exploited shellfish to changing pH levels: how to scale-up experimental evidence to regional impacts. ICES Journal of Marine Science, 79(9):2362–2372, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
9/2362/6761941.

Trochta:2021:ABM

[TB21] John T. Trochta and Trevor A. Branch. Applying Bayesian model selection to determine ecological covariates for recruitment and natural mortality in stock assessment. ICES Journal of Marine Science, 78(8):2875–2894, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/8/2875/6368095.

Thorson:2020:STN

[TBH^{+}20] James T. Thorson, Meaghan D. Bryan, Peter-John F. Hulson, Haikun Xu, and André E. Punt. Simulation testing a new multi-stage process to measure the effect of increased sampling effort on effective sample size for age and length data. ICES Journal of Marine Science, 77(5):1728–1737, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/5/1728/5814325.

Trijoulet:2022:TRP

[TBM^{+}22] Vanessa Trijoulet, Casper W. Berg, David C. M. Miller, Anders Nielsen, Anna Rindorf, and Christoffer Moesgaard Albertsen. Turning reference points inside out: comparing MSY reference points estimated inside and outside the assessment model. ICES Journal of Marine Science, 79(4): 1232–1244, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/4/1232/6552728.

Tewfik:2020:SLF

[TBP20] Alexander Tewfik, Elizabeth A. Babcock, and Myles Phillips. Spiny lobster fisheries status across time and a mosaic of spatial management regimes. ICES Journal of Marine Science, 77(3):1002–1016, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/3/
1002/5728667.

Trenkel:2022:CKM

[TCLB22] Verena M. Trenkel, Grégory Charrier, Pascal Lorance, and Mark V. Bravington. Close-kin mark–recapture abundance estimation: practical insights and lessons learned. ICES Journal of Marine Science, 79(2):413–422, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/2/413/6513634.

Taillebois:2021:IAS

[TDB^{+}21] Laura Taillebois, Danielle Davenport, Diane P. Barton, David A. Crook, Thor Saunders, Mark Hearnden, Richard J. Saunders, Stephen J. Newman, Michael J. Travers, Christine L. Dudgeon, Safia L. Maher, and Jennifer R. Ovenden. Integrated analyses of SNP-genotype and environmental data in a continuously distributed snapper species (Lutjanus johnii, Bloch, 1792) reveals a mosaic of populations and a challenge for sustainable management. ICES Journal of Marine Science, 78(9):3212–3229, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/9/3212/6380056.

Tamini:2023:BSL

[TDC^{+}23] L. L. Tamini, R. F. Dellacasa, L. N. Chavez, C. J. Marinao, M. E. Góngora, R. Crawford, and E. Frere. Bird scaring lines reduce seabird mortality in mid-water and bottom trawlers in Argentina. ICES Journal of Marine Science, 80(9):2393– 2404, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/9/2393/7223059.

Trathan:2021:EEA

[TFH^{+}21] P. N. Trathan, S. Fielding, P. R. Hollyman, E. J. Murphy, V. Warwick-Evans, and M. A. Collins. Enhancing the ecosystem approach for the fishery for Antarctic krill within the complex, variable, and changing ecosystem at South Georgia. ICES Journal of Marine Science, 78(6):2065–2081, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/2065/6295716.

Trathan:2022:SSR

[TFWE^{+}22] P. N. Trathan, S. Fielding, V. Warwick-Evans, J. Freer, and F. Perry. Seabird and seal responses to the physical environment and to spatio-temporal variation in the distribution and abundance of Antarctic krill at South Georgia, with implications for local fisheries management. ICES Journal of Marine Science, 79(9):2373–2388, November 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/9/2373/6759148.

Thingstad:2020:HTC

[Thi20] T. Frede Thingstad. How trophic cascades and photic zone nutrient content interact to generate basin-scale differences in the microbial food web. ICES Journal of Marine Science, 77(5):1639–1647, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1639/5827722.

Thorson:2022:DST

[Tho22] James T. Thorson. Development and simulation testing for a new approach to density dependence in species distribution models. ICES Journal of Marine Science, 79(1):117–128, January 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/1/117/6463441.

Thorson:2021:GCH

[THSZ21] James T. Thorson, Albert J. Hermann, Kevin Siwicke, and Mark Zimmermann. Grand challenge for habitat science: stage-structured responses, nonlocal drivers, and mechanistic associations among habitat variables affecting fishery productivity. ICES Journal of Marine Science, 78(6):1956–1968, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/6/1956/6297833.

Tseng:2020:DCH

[TK20] Chi-Hsuan Tseng and Yan-Fu Kuo. Detecting and counting harvested fish and identifying fish types in electronic monitoring system videos using deep convolutional neural networks. ICES Journal of Marine Science, 77(4):1367– 1378, July 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/4/1367/5847960.

Tate:2020:SHR

[TLM^{+}20] Alissa Tate, Johnny Lo, Ute Mueller, Glenn A. Hyndes, Karina L. Ryan, and Stephen M. Taylor. Standardizing harvest rates of finfish caught by shore-based recreational fishers. ICES Journal of Marine Science, 77(6):2207–2215, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2207/5650356.

Titelboim:2021:TSR

[TLS21] Danna Titelboim, Oliver T. Lord, and Daniela N. Schmidt. Thermal stress reduces carbonate production of benthic foraminifera and changes the material properties of their shells. ICES Journal of Marine Science, 78(9):3202–3211, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3202/6375121.

Thiebot:2020:WDM

[TM20] Jean-Baptiste Thiebot and Julie C. McInnes. Why do marine endotherms eat gelatinous prey? ICES Journal of Marine Science, 77(1):58–71, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http:/

/academic.oup.com/icesjms/article/77/1/58/5613760.

Taormina:2020:OIB

[TMD^{+}20] Bastien Taormina, Martin P. Marzloff, Nicolas Desroy, Xavier Caisey, Olivier Dugornay, Emmanuelle Metral Thiesse, Aurélien Tancray, and Antoine Carlier. Optimizing image-based protocol to monitor macroepibenthic communities colonizing artificial structures. ICES Journal of Marine Science, 77(2):835–845, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
835/5695542.

Thomas:2020:SBA

[TMG20] Alyssa S. Thomas, Taciano L. Milfont, and Michael C. Gavin. Scenario-based analyses evaluate potential outcomes of proposed regulatory changes in recreational fishery. ICES Journal of Marine Science, 77(6):2333–2343, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2333/5536334.

Tidbury:2021:BBE

[TMLH21] Hannah J. Tidbury, Angela Muench, Philip D. Lamb, and Kieran Hyder. Balancing biological and economic goals in commercial and recreational fisheries: systems modelling of sea bass fisheries. ICES Journal of Marine Science, 78(5): 1793–1803, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/5/1793/6277124.

Thorson:2023:MTS

[TMS23] James T. Thorson, Timothy J. Miller, and Brian C. Stock. The multivariate-Tweedie: a self-weighting likelihood for age and length composition data arising from hierarchical sampling designs. ICES Journal of Marine Science, 80(10):2630– 2641, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/10/2630/6710216.

Toledo:2020:TEP

[TND^{+}20] Pamela Toledo, Edwin J. Niklitschek, Audrey M. Darnaude, Félix P. Leiva, Chris Harrod, Sergio Lillo, Vilma Ojeda, Sebastián Klarian, Blanca E. Molina-Burgos, Patricio Gálvez, and Cristian B. Canales-Aguirre. The trophic ecology of partial migration: insights from Merluccius australis off NW Patagonia. ICES Journal of Marine Science, 77(5):1927– 1940, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/5/1927/5859702.

Thebaud:2023:IEF

[TNM^{+}23] O. Thébaud, J. R. Nielsen, A. Motova, H. Curtis, F. Bastardie, G. E. Blomqvist, F. Daurès, L. Goti, J. Holzer, J. Innes, A. Muench, A. Murillas, R. Nielsen, R. Rosa, E. Thunberg, S. Villasante, J. Virtanen, S. Waldo, S. Agnarsson, D. Castilla Espino, R. Curtin, G. DePiper, R. Doering, H. Ellefsen, J. J. García del Hoyo1, S. Gourguet, P. Greene, K. G. Hamon, A. Haynie, J. B. Kellner, S. Kuikka, B. Le Gallic, C. Macher, R. Prellezo, J. Santiago Castro-Rial, K. Sys, H. van Oostenbrugge, and B. M. J. Vastenhoud. Integrating economics into fisheries science and advice: progress, needs, and future opportunities. ICES Journal of Marine Science, 80(4):647–663, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/4/647/7036153.

Tiedemann:2021:EIN

[TNS^{+}21] Maik Tiedemann, Richard D. M. Nash, Erling K. Stenevik, Martina H. Stiasny, Aril Slotte, and Olav S. Kjesbu. Environmental influences on Norwegian spring-spawning herring (Clupea harengus L.) larvae reveal recent constraints in recruitment success. ICES Journal of Marine Science, 78(2): 640–652, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/2/640/5828450.

Tam:2024:ESR

[TPCM^{+}24] Jamie C. Tam, Courtenay E. Parlee, Jill Campbell-Miller, Manuel Bellanger, Jacob Bentley, Vahab Pourfaraj, Evan J. Andrews, Sondra L. Eger, Adam M. Cook, and Gabrielle Beaulieu. Expanding the scope and roles of social sciences and humanities to support integrated ecosystem assessments and ecosystem-based management. ICES Journal of Marine Science, 81(1):22–42, January 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/81/1/22/7337019.

Taylor:2020:CWA

[TR20] S. M. Taylor and K. L. Ryan. Concurrent Western Australian telephone surveys highlight the advantages of sampling from a registry of recreational fishers. ICES Journal of Marine Science, 77(6):2181–2191, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2181/5529303.

Tutar:2022:HLG

[TRD^{+}22] Ozge Tutar, Miriam Ruocco, Emanuela Dattolo, Guglielmo Lacorata, Raffaele Corrado, Romain Watteaux, Daniele Iudicone, Bettina Fach, and Gabriele Procaccini. High levels of genetic diversity and population structure in the Mediterranean seagrass Posidonia oceanica at its easternmost distribution limit. ICES Journal of Marine Science, 79(8):2286– 2297, October 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/8/2286/6698561.

Thomas:2021:CEL

[TRP^{+}21] J.-B. E. Thomas, M. Sodré Ribeiro, J. Potting, G. Cervin, G. M. Nylund, J. Olsson, E. Albers, I. Undeland, H. Pavia, and F. Gröndahl. A comparative environmental life cycle assessment of hatchery, cultivation, and preservation of the kelp Saccharina latissima. ICES Journal of Marine Science, 78(1):451–467, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/451/5892999.

Taylor:2021:CCE

[TSD^{+}21] S. M. Taylor, C. B. Smallwood, C. J. Desfosses, K. L. Ryan, and G. Jackson. Corroborating catch estimates to inform monitoring of a small-scale marine recreational fishery in a world heritage property. ICES Journal of Marine Science, 78(5):1887–1899, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/5/
1887/6290961.

Tillotson:2021:NSE

[TSE^{+}21] Michael D. Tillotson, Timothy F. Sheehan, Brandon Ellingson, Ruth E. Haas-Castro, Maxime Olmos, and Katherine E. Mills. Non-stationary effects of growth on the survival of North American Atlantic salmon (Salmo salar). ICES Journal of Marine Science, 78(8):2967–2982, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2967/6370947.

Todd:2021:CSW

[TSW^{+}21] Victoria L. G. Todd, Irene Susini, Laura D. Williamson, Ian B. Todd, Dianne L. McLean, and Peter I. Macreadie. Characterizing the second wave of fish and invertebrate colonization of an offshore petroleum platform. ICES Journal of Marine Science, 78(3):1131–1145, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
1131/6145865.

Thatcher:2023:RHU

[TSWM23] H. Thatcher, T. Stamp, D. Wilcockson, and P. J. Moore. Residency and habitat use of European lobster (Homarus gammarus) within an offshore wind farm. ICES Journal of Marine Science, 80(5):1410–1421, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1410/7145793.

Todd:2020:CFW

[TWC^{+}20] Victoria L. G. Todd, Laura D. Williamson, Sophie E. Cox, Ian B. Todd, and Peter I. Macreadie. Characterizing the first wave of fish and invertebrate colonization on a new offshore petroleum platform. ICES Journal of Marine Science, 77(3): 1127–1136, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/3/1127/5485757.

Tookes:2023:RFE

[TYF23] Jennifer Sweeney Tookes, Tracy Yandle, and Bryan Fluech. The role of fisher engagement in the acceptance of turtle excluder devices in Georgia’s shrimping industry. ICES Journal of Marine Science, 80(3):407–416, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/407/6568362.

Tenningen:2021:PCD

[TZE21] Maria Tenningen, Fabian Zimmermann, and Katja Enberg. Pre-catch and discard mortality in Northeast Atlantic herring and mackerel fisheries: consequences for stock estimates and advice. ICES Journal of Marine Science, 78(7):2603– 2614, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/7/2603/6346871.

Tang:2021:IRR

[TZR^{+}21] Xiaozhuo Tang, Nan Zheng, Rick M. Rideout, Shijia Wang, and Fan Zhang. Identification of recruitment regime shifts with a hidden Markov stock-recruitment model. ICES Journal of Marine Science, 78(7):2591–2602, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/7/2591/6332553.

Urban:2023:SEB

[UBD^{+}23] P. Urban, D. Bekkevold, H. Degel, B. K. Hansen, M. W. Jacobsen, A. Nielsen, and E. E. Nielsen. Scaling from eDNA to biomass: controlling allometric relationships improves precision in bycatch estimation. ICES Journal of Marine Science, 80(4):1066–1078, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
4/1066/7077003.

Utne:2021:PFO

[UPH^{+}21] Kjell Rong Utne, Beatriz Diaz Pauli, Monika Haugland, Jan Arge Jacobsen, Niall Maoileidigh, Webjørn Melle, Cecilie Thorsen Broms, Leif Nøttestad, Marianne Holm, Katie Thomas, and Vidar Wennevik. Poor feeding opportunities and reduced condition factor for salmon post-smolts in the Northeast Atlantic Ocean. ICES Journal of Marine Science, 78(8):2844–2857, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/
2844/6357781.

Uriarte:2021:COP

[UVI^{+}21a] Ibon Uriarte, Fernando Villate, Arantza Iriarte, Álvaro Fanjul, Angus Atkinson, and Kathryn Cook. Corrigendum to: Opposite phenological responses of zooplankton to climate along a latitudinal gradient through the European Shelf. ICES Journal of Marine Science, 78(7):2643, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/7/2643/6322254. See [UVI^{+}21b].

Uriarte:2021:OPR

[UVI^{+}21b] Ibon Uriarte, Fernando Villate, Arantza Iriarte, Álvaro Fanjul, Angus Atkinson, and Kathryn Cook. Opposite phenological responses of zooplankton to climate along a latitudinal gradient through the European Shelf. ICES Journal of Marine Science, 78(3):1090–1107, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/

1090/6132743. See corrigendum [UVI^{+}21a].

Varanasi:2021:CWN

[Var21] Usha Varanasi. Casting a wide net and making the most of the catch. ICES Journal of Marine Science, 78(3):832– 847, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/3/832/6140879.

Vieira:2020:UNR

[VAS^{+}20] Manuel Vieira, M. Clara P. Amorim, Andreas Sundelöf, Nuno Prista, and Paulo J. Fonseca. Underwater noise recognition of marine vessels passages: two case studies using hidden Markov models. ICES Journal of Marine Science, 77(6):2157–2170, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2157/5609037.

Vighi:2021:MWR

[VBJ^{+}21] Morgana Vighi, Asunción Borrell, Jennifer A. Jackson, Emma L. Carroll, Maria Grazia Pennino, and Alex Aguilar. The missing whales: relevance of “struck and lost” rates for the impact assessment of historical whaling in the southwestern Atlantic Ocean. ICES Journal of Marine Science, 78(1): 14–24, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/1/14/6026100.

Vaughan:2021:CGR

[VBO^{+}21a] Louise Vaughan, Deirdre Brophy, Ciar O’Toole, Conor Graham, Niall Ó Maoiléidigh, and Russell Poole. Corrigendum to: Growth rates in a European eel Anguilla anguilla (L., 1758) population show a complex relationship with temperature over a seven-decade otolith biochronology. ICES Journal of Marine Science, 78(8):3012, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/8/

3012/6355110. See [VBO^{+}21b].

Vaughan:2021:GRE

[VBO^{+}21b] Louise Vaughan, Deirdre Brophy, Ciar O’Toole, Conor Graham, Niall Ó Maoiléidigh, and Russell Poole. Growth rates in a European eel Anguilla anguilla (L., 1758) population show a complex relationship with temperature over a seven-decade otolith biochronology. ICES Journal of Marine Science, 78 (3):994–1009, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/3/994/6104621. See corri-

gendum [VBO^{+}21a].

Vanermen:2020:AOM

[VCD^{+}20] Nicolas Vanermen, Wouter Courtens, Robin Daelemans, Luc Lens, Wendt Müller, Marc Van de walle, Hilbran Verstraete, and Eric W. M. Stienen. Attracted to the outside: a mesoscale response pattern of lesser black-backed gulls at an offshore wind farm revealed by GPS telemetry. ICES Journal of Marine Science, 77(2):701–710, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
701/5652052.

Volstad:2020:FSM

[VCF^{+}20] Jon Helge Vølstad, Mary Christman, Keno Ferter, Alf Ring Kleiven, Håkon Otterå, Øystein Aas, Robert Arlinghaus, Trude Borch, Jonathan Colman, Bruce Hartill, Thrond O. Haugen, Kieran Hyder, Jeremy M. Lyle, Martin Junker Ohldieck, Christian Skov, Harry V. Strehlow, Dave van Voorhees, Marc Simon Weltersbach, and Edward D. Weber. Field surveying of marine recreational fisheries in Norway using a novel spatial sampling frame reveals striking undercoverage of alternative sampling frames. ICES Journal of Marine Science, 77(6):2192–2205, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/6/
2192/5525341.

Vieira:2020:OIW

[VDAT20] Ana Rita Vieira, Sandra Dores, Manuela Azevedo, and Susanne E. Tanner. Otolith increment width-based chronologies disclose temperature and density-dependent effects on demersal fish growth. ICES Journal of Marine Science, 77 (2):633–644, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/2/633/5680030.

vandenBurg:2021:TSE

[vdBDH21] S. W. K. van den Burg, H. Dagevos, and R. J. K. Helmes. Towards sustainable European seaweed value chains: a triple P perspective. ICES Journal of Marine Science, 78(1):443– 450, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/1/443/5580339.

vanDenderen:2020:EIB

[vDBF^{+}20] P. D. van Denderen, S. G. Bolam, R. Friedland, J. G. Hiddink, K. Norén, A. D. Rijnsdorp, M. Sköld, A. Törnroos, E. A. Virtanen, and S. Valanko. Evaluating impacts of bottom trawling and hypoxia on benthic communities at the local, habitat, and regional scale using a modelling approach. ICES Journal of Marine Science, 77(1):278–289, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/1/278/5634219.

vanDenderen:2022:PBF

[vDHR^{+}22] P. Daniël van Denderen, Helen Holah, Laura M. Robson, Jan Geert Hiddink, Lénaick Menot, Debbi Pedreschi, Georgios Kazanidis, Marcos Llope, Phillip J. Turner, David Stirling, F. Javier Murillo, Andrew Kenny, Neil Campbell, A. Louise Allcock, Andreia Braga-Henriques, Jose M. González-Irusta, Graham Johnston, Covadonga Orejas, Alberto Serrano, Joana R. Xavier, Peter Hopkins, Ellen Kenchington, Eugene Nixon, and Sebastian Valanko. A policybased framework for the determination of management options to protect vulnerable marine ecosystems under the EU deep-sea access regulations. ICES Journal of Marine Science, 79(1):34–49, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/34/6453073.

vandeWolfshaar:2022:SFY

[vdWBL22] K. E. van de Wolfshaar, L. Barbut, and G. Lacroix. From spawning to first-year recruitment: the fate of juvenile sole growth and survival under future climate conditions in the North Sea. ICES Journal of Marine Science, 79(2):495–505, March 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/2/495/6140865.

van:2021:ADR

[vEMvH^{+}21] Rick van Essen, Angelo Mencarelli, Aloysius van Helmond, Linh Nguyen, Jurgen Batsleer, Jan-Jaap Poos, and Gert Kootstra. Automatic discard registration in cluttered environments using deep learning and object tracking: class imbalance, occlusion, and a comparison to human review. ICES Journal of Marine Science, 78(10):3834–3846, December 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/10/3834/6444891.

Vihtakari:2022:MPG

[VET^{+}22] Mikko Vihtakari, Bjarki ór Elvarsson, Margaret Treble, Adriana Nogueira, Kevin Hedges, Nigel E. Hussey, Laura Wheeland, Denis Roy, Lise Helen Ofstad, Elvar H. Hallfredsson, Amanda Barkley, Daniel Estévez-Barcia, Rasmus Nygaard, Brian Healey, Petur Steingrund, Torild Johansen, Ole Thomas Albert, and Jesper Boje. Migration patterns of Greenland halibut in the North Atlantic revealed by a compiled mark–recapture dataset. ICES Journal of Marine Science, 79(6):1902–1917, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/6/1902/6646040.

Vihtakari:2021:PAS

[VHT^{+}21] Mikko Vihtakari, Robinson Hordoir, Margaret Treble, Meaghan D. Bryan, Bjarki Elvarsson, Adriana Nogueira, Elvar H. Hallfredsson, Jørgen Schou Christiansen, and Ole Thomas Albert. Pan-Arctic suitable habitat model for Greenland halibut. ICES Journal of Marine Science, 78(4): 1340–1356, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1340/6168597.

Vad:2020:ECAa

[VKH^{+}20a] J. Vad, G. Kazanidis, L-A. Henry, D. O. B. Jones, A. R. Gates, and J. M. Roberts. Environmental controls and anthropogenic impacts on deep-sea sponge grounds in the Faroe–Shetland Channel, NE Atlantic: the importance of considering spatial scale to distinguish drivers of change. ICES Journal of Marine Science, 77(1):451–461, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/1/451/5599858.

Vad:2020:ECAb

[VKH^{+}20b] J. Vad, G. Kazanidis, L.-A. Henry, D. O. B. Jones, A. R. Gates, and J. M. Roberts. Environmental controls and anthropogenic impacts on deep-sea sponge grounds in the Faroe–Shetland Channel, NE Atlantic: the importance of considering spatial scale to distinguish drivers of change. ICES Journal of Marine Science, 77(5):2009, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/5/2009/5810164.

Vollset:2021:WSR

[VLD^{+}21] Knut Wiik Vollset, Robert J. Lennox, Jan Grimsrud Davidsen, Sindre Håvarstein Eldøy, Trond E. Isaksen, Abdullah Madhun, Sten Karlsson, and Kristina M. Miller. Wild salmonids are running the gauntlet of pathogens and climate as fish farms expand northwards. ICES Journal of Marine Science, 78(1):388–401, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/1/
388/5920397.

vonLeesen:2020:SOI

[vLNC20] Gotje von Leesen, Ulysses S. Ninnemann, and Steven E. Campana. Stable oxygen isotope reconstruction of temperature exposure of the Icelandic cod (Gadus morhua) stock over the last 100 years. ICES Journal of Marine Science, 77 (3):942–952, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/77/3/942/5736261.

Valdes:2022:SIV

[VLUB^{+}22] Luis Valdés, Angel López-Urrutia, Gregory Beaugrand, Roger P. Harris, and Xabier Irigoien. Seasonality and interannual variability of copepods in the Western English Channel, Celtic Sea, Bay of Biscay, and Cantabrian Sea with a special emphasis to Calanus helgolandicus and Acartia clausi. ICES Journal of Marine Science, 79(3):727–740, April 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/3/727/6565659.

Vilalta-Navas:2023:CSL

[VNBLZR^{+}23] Ainoa Vilalta-Navas, Rodrigo Beas-Luna, Manuel J. Zetina-Rejón, Gladis A. López-Ibarra, Luis Malpica-Cruz, Julio Lorda, José Alberto Zepeda-Dominguez, and Lotta Clara Kluger. Coupling scientific and local ecological knowledge network models for temperate coastal ecosystems. ICES Journal of Marine Science, 80(1):185–196, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/1/185/6958702.

Vigo:2023:BAC

[VNR^{+}23] M. Vigo, J. Navarro, G. Rotllant, N. Bahamon, M. Carretón, J. Quevedo, A. Rojas, and J. B. Company. Before– after control–impact (BACI) assessment of the effects of a deep-water no-take fishery reserve to recover Norway lobster (Nephrops norvegicus) overfished populations and coexisting megafauna. ICES Journal of Marine Science, 80(7):2008– 2023, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/7/2008/7248747.

Varnes:2023:FCD

[VO23] B. K. Varnes and E. M. Olsen. Fish community dynamics in a coastal no-take marine protected area compared to a harvested area before and after protection from fishing. ICES Journal of Marine Science, 80(5):1462–1471, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1462/7164138.

Vilas:2020:SSP

[VPB^{+}20] Daniel Vilas, Maria Grazia Pennino, Jose Maria Bellido, Joan Navarro, Isabel Palomera, and Marta Coll. Seasonality of spatial patterns of abundance, biomass, and biodiversity in a demersal community of the NW Mediterranean Sea. ICES Journal of Marine Science, 77(2):567–580, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/567/5644315.

Valle-Pereira:2022:RBV

[VPCM^{+}22] João V. S. Valle-Pereira, Mauricio Cantor, Alexandre M. S. Machado, Damien R. Farine, and Fábio G. Daura-Jorge. The role of behavioural variation in the success of artisanal fishers who interact with dolphins. ICES Journal of Marine Science, 79(4):1150–1158, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1150/6549170.

VanBeveren:2023:RIV

[VPP^{+}23] Elisabeth Van Beveren, Stéphane Plourde, Pierre Pepin, Karen Cogliati, and Martin Castonguay. A review of the importance of various areas for northern contingent West-Atlantic mackerel spawning. ICES Journal of Marine Science, 80(1):1–15, January 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/1/1/6891092.

Voss:2022:REE

[VQN22] Rudi Voss, Martin Quaas, and Stefan Neuenfeldt. Robust, ecological–economic multispecies management of Central Baltic fishery resources. ICES Journal of Marine Science, 79(1):169–181, January 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/1/169/6479779.

Vary:2023:TDS

[VRH^{+}23] L. Vary, L. Rogers, M. Harte, R. Howard, and L. Ciannelli. Temperature-dependent spawning behaviour and larval thermohaline associations of Bering Sea groundfish. ICES Journal of Marine Science, 80(7):1881–1898, September 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/7/1881/7231243.

Veron:2023:EDC

[VRM^{+}23] Pierre Veron, Romane Rozanski, Virginie Marques, Stéphane Joost, Marie Emilie Deschez, Verena M. Trenkel, Pascal Lorance, Alice Valentini, Andrea Polanco F., Loïc Pellissier, David Eme, and Camille Albouy. Environmental DNA complements scientific trawling in surveys of marine fish biodiversity. ICES Journal of Marine Science, 80(8):2150–2165, October 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/8/2150/7280042.

Vereshchaka:2022:SST

[VSL22] A. L. Vereshchaka, A. V. Shatravin, and A. A. Lunina. Shifting seasonal timing of peak abundance of two invading ctenophore populations in the Black Sea during the period 1991–2017. ICES Journal of Marine Science, 79(3): 954–962, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/3/954/6529071.

Venello:2022:DVC

[VSS^{+}22] Theresa A. Venello, Akash R. Sastri, Karyn D. Suchy, Moira D. Galbraith, and John F. Dower. Drivers of variation in crustacean zooplankton production rates differ across regions off the west coast of Vancouver Island and in the subarctic NE Pacific. ICES Journal of Marine Science, 79(3): 741–760, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/3/741/6444893.

Votier:2023:OIF

[VSS^{+}23] S. C. Votier, R. B. Sherley, K. L. Scales, K. Camphuysen, and R. A. Phillips. An overview of the impacts of fishing on seabirds, including identifying future research directions. ICES Journal of Marine Science, 80(9):2380–2392, November 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/9/2380/7453018.

Waples:2020:SM

[Wap20] Robin S. Waples. Serendipity and me. ICES Journal of Marine Science, 77(5):1658–1665, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/5/
1658/5841731.

Ward:2022:INS

[WBA^{+}22] Eric J. Ward, Lewis A. K. Barnett, Sean C. Anderson, Christian J. C. Commander, and Timothy E. Essington. Incorporating non-stationary spatial variability into dynamic species distribution models. ICES Journal of Marine Science, 79(9):2422–2429, November 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/9/2422/6758105.

Whitmarsh:2021:NDE

[WBB^{+}21] Sasha K. Whitmarsh, Greg M. Barbara, James Brook, Dimitri Colella, Peter G. Fairweather, Tim Kildea, and Charlie Huveneers. No detrimental effects of desalination waste on temperate fish assemblages. ICES Journal of Marine Science, 78(1):45–54, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/1/45/6032770.

Waller:2023:RRH

[WBB^{+}23] J. Waller, J. Bartlett, E. Bates, H. Bray, M. Brown, M. Cieri, C. Clark, W. DeVoe, B. Donahue, D. Frechette, H. Glon, M. Hunter, C. Huntsberger, K. Kanwit, S. Ledwin, B. Lewis, R. Peters, K. Reardon, R. Russell, M. Smith, C. Uraneck, R. Watts, and C. Wilson. Reflecting on the recent history of coastal Maine fisheries and marine resource monitoring: the value of collaborative research, changing ecosystems, and thoughts on preparing for the future. ICES Journal of Marine Science, 80(8):2074–2086, October 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/8/2074/7257039.

Wang:2021:ESS

[WBC^{+}21] Jintao Wang, Robert Boenish, Xinjun Chen, Siquan Tian, and JiangFeng Zhu. The effects of spatiotemporal scale on commercial fishery abundance index suitability. ICES Journal of Marine Science, 78(7):2506–2517, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/7/2506/6318895.

Wells:2022:EDD

[WBC^{+}22] Seòna R. Wells, Eileen Bresnan, Kathryn Cook, Dafne Eerkes-Medrano, Margarita Machairopoulou, Daniel J. Mayor, Berit Rabe, and Peter J. Wright. Environmental drivers of a decline in a coastal zooplankton community. ICES Journal of Marine Science, 79(3):844–854, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/3/844/6368415.

Wilber:2022:DFI

[WBG^{+}22] Dara H. Wilber, Lorraine Brown, Matthew Griffin, Gregory R. DeCelles, and Drew A. Carey. Demersal fish and invertebrate catches relative to construction and operation of North America’s first offshore wind farm. ICES Journal of Marine Science, 79(4):1274–1288, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1274/6555702.

Williamson:2021:AME

[WBWS21] Benjamin J. Williamson, Philippe Blondel, Laura D. Williamson, and Beth E. Scott. Application of a multibeam echosounder to document changes in animal movement and behaviour around a tidal turbine structure. ICES Journal of Marine Science, 78(4):1253–1266, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1253/6154329.

Wang:2020:QLT

[WCCY20] Na Wang, Anthony J. Courtney, Matthew J. Campbell, and Wen-Hsi Yang. Quantifying long-term discards from Queensland’s (Australia) east Coast otter trawl fishery. ICES Journal of Marine Science, 77(2):680–691, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
680/5644309.

Wu:2023:RTM

[WCL^{+}23] X. H. Wu, Y. Chang, T. Y. Liao, M. M. Ding, and C. C. Ke. Real-time multi-month forecasting of skipjack tuna (Katsuwonus pelamis) habitat in the western and central Pacific Ocean for improved fishing efficiency and fisheries management. ICES Journal of Marine Science, 80(10):2490–2503, December 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/10/2490/7310994.

Wang:2023:ECI

[WCLB23] Jintao Wang, Xinjun Chen, YunKai Li, and Robert Boenish. The effects of climate-induced environmental variability on Pacific Ocean squids. ICES Journal of Marine Science, 80 (4):878–888, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/4/878/7041963. See cor-

rection [Ano23l].

Wolfe:2023:HDR

[WDPM23] Kennedy Wolfe, Amelia A. Desbiens, Emma Pietsch, and Peter J. Mumby. Habitat and distribution of the red decorator crab, Schizophrys aspera, a cryptic crown-of-thorns seastar predator. ICES Journal of Marine Science, 80(8):2114–2124, October 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/8/2114/7280038.

Warwick-Evans:2021:MSA

[WESWT21] V. Warwick-Evans, J. A. Santora, J. J. Waggitt, and P. N. Trathan. Multi-scale assessment of distribution and density of procellariiform seabirds within the Northern Antarctic Peninsula marine ecosystem. ICES Journal of Marine Science, 78(4):1324–1339, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/4/
1324/6161357.

White:2023:ICL

[WFG^{+}23] Jonathan White, Colm Fitzgerald, Patrick Gargan, Elvira de Eyto, Michael Millane, Gerald Chaput, Paddy Boylan, Walter W. Crozier, Dennis Doherty, Bryan Kennedy, Ian Lawler, David Lyons, Ferdia Marnell, Phil McGinnity, Kealan O'Higgins, William K. Roche, Hugo Maxwell, and Niall Ó Maoiléidigh. Incorporating conservation limit variability and stock risk assessment in precautionary salmon catch advice at the river scale. ICES Journal of Marine Science, 80(4):803–822, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/803/7034404. See correction [Ano23e].

Williams:2022:HGA

[WGCC^{+}22] Megan J. Peterson Williams, Becca Robbins Gisclair, Elizabeth Cerny-Chipman, Michael LeVine, and Theresa Peterson. The heat is on: Gulf of Alaska Pacific cod and climateready fisheries. ICES Journal of Marine Science, 79(2):573– 583, March 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/2/573/6154493.

Ward:2021:IPD

[WGI^{+}21] Timothy M. Ward, Gretchen L. Grammer, Alex R. Ivey, Jonathan J. Smart, and Richard McGarvey. Increasing the precision of the daily egg production method; 2020’s remix of a 1980’s classic. ICES Journal of Marine Science, 78(4): 1177–1195, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1177/6199409.

Wain:2021:QIF

[WGKG21] Gwenaëlle Wain, Loreleï Guéry, David Michael Kaplan, and Daniel Gaertner. Quantifying the increase in fishing efficiency due to the use of drifting FADs equipped with echosounders in tropical tuna purse seine fisheries. ICES Journal of Marine Science, 78(1):235–245, January 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/78/1/235/6032769.

Wosnick:2023:OER

[WGL^{+}23] Natascha Wosnick, Eloísa Pinheiro Giareta, Renata Daldin Leite, Ingrid Hyrycena, and Patricia Charvet. An overview on elasmobranch release as a bycatch mitigation strategy. ICES Journal of Marine Science, 80(3):591–604, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/591/6711584.

Weston:2022:EGH

[WJ22] Johanna N. J. Weston and Alan J. Jamieson. Exponential growth of hadal science: perspectives and future directions identified using topic modelling. ICES Journal of Marine Science, 79(4):1048–1062, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1048/6576454.

Weber:2021:IOM

[WJB^{+}21] D. Nick Weber, Michael G. Janech, Louis E. Burnett, Gorka Sancho, and Bryan S. Frazier. Insights into the origin and magnitude of capture and handling-related stress in a coastal elasmobranch Carcharhinus limbatus. ICES Journal of Marine Science, 78(3):910–921, July 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/3/
910/6054606.

Williamson:2024:SNA

[WJB^{+}24] Michael J Williamson, David M P Jacoby, Tea Bašić, Alan Walker, and Adam T Piper. Social network analysis as a tool to inform anguillid eel conservation and management. ICES Journal of Marine Science, 81(2):402–410, March 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/2/402/7587675.

Wright:2020:SSS

[WLR^{+}20] Serena R. Wright, Christopher P. Lynam, David A. Righton, Julian Metcalfe, Ewan Hunter, Ainsley Riley, Luz Garcia, Paulette Posen, and Kieran Hyder. Structure in a sea of sand: fish abundance in relation to man-made structures in the North Sea. ICES Journal of Marine Science, 77(3):1206– 1218, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/3/1206/5145713.

Wheeland:2020:ASS

[WM20] Laura J. Wheeland and M. Joanne Morgan. Age-specific shifts in Greenland halibut (Reinhardtius hippoglossoides) distribution in response to changing ocean climate. ICES Journal of Marine Science, 77(1):230–240, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/1/230/5569489.

Woods:2022:RAO

[WMB^{+}22] P. J. Woods, J. I. Macdonald, H. Bárdarson, S. Bonanomi, W. J. Boonstra, G. Cornell, G. Cripps, R. Danielsen, L. Färber, A. S. A. Ferreira, K. Ferguson, M. Holma, R. E. Holt, K. L. Hunter, A. Kokkalis, T. J. Langbehn, G. Ljungström, E. Nieminen, M. C. Nordström, M. Oostdijk, A. Richter, G. Romagnoni, C. Sguotti, A. Simons, N. L. Shackell, M. Snickars, J. D. Whittington, H. Wootton, and J. Yletyinen. A review of adaptation options in fisheries management to support resilience and transition under socioecological change. ICES Journal of Marine Science, 79(2): 463–479, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/2/463/6354503.

Woodings:2021:PGE

[WML^{+}21] Laura N. Woodings, Nicholas P. Murphy, Geoffrey W. Liggins, Marcus E. Miller, Giles M. Ballinger, Sally C. Y. Lau, and Jan M. Strugnell. Population genomics of the Eastern rock lobster, Sagmariasus verreauxi, during spawning stock recovery from over-exploitation. ICES Journal of Marine Science, 78(7):2448–2459, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2448/6307375.

Walker:2023:DSP

[WOB^{+}23] N. D. Walker, R. Ouréns, J. E. Ball, J. van der Kooij, A. Uriarte, J. White, P. Carpi, P. Schuchert, and R. D. M. Nash. Defining sustainable and precautionary harvest rates for data-limited short-lived stocks: a case study of sprat (Sprattus sprattus) in the English Channel. ICES Journal of Marine Science, 80(10):2606–2618, December 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/10/2606/7125894.

Weijerman:2021:SEB

[WOL^{+}21] Mariska Weijerman, Zack S. Oyafuso, Kirsten M. Leong, Kirsten L. L. Oleson, and Morgan Winston. Supporting ecosystem-based fisheries management in meeting multiple objectives for sustainable use of coral reef ecosystems. ICES Journal of Marine Science, 78(8):2999–3011, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2999/5999229.

Woods:2022:AIE

[Woo22] Pamela J. Woods. Aligning integrated ecosystem assessment with adaptation planning in support of ecosystembased management. ICES Journal of Marine Science, 79(2): 480–494, March 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/79/2/480/6332557.

Waldo:2020:ECS

[WPB20] Staffan Waldo, Anton Paulrud, and Johan Blomquist. The economic costs of seal presence in Swedish small-scale fisheries. ICES Journal of Marine Science, 77(2):815–825, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/2/815/5643854.

Walter:2023:WCW

[WPM^{+}23] J. F. Walter III, C. D. Peterson, K. Marshall, J. J. Deroba, S. Gaichas, B. C. Williams, S. Stohs, D. Tommasi, and R. Ahrens. When to conduct, and when not to conduct, management strategy evaluations. ICES Journal of Marine Science, 80(4):719–727, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/80/4/719/7092988.

Wegner:2021:PRS

[WPN^{+}21] Nicholas C. Wegner, Elan J. Portner, Drew T. Nguyen, Lyall Bellquist, Andrew P. Nosal, Alena L. Pribyl, Kevin L. Stierhoff, Paul Fischer, Ken Franke, Russell D. Vetter, Philip A. Hastings, Brice X. Semmens, and John R. Hyde.

Post-release survival and prolonged sublethal effects of capture and barotrauma on deep-dwelling rockfishes (genus Sebastes): implications for fish management and conservation. ICES Journal of Marine Science, 78(9):3230–3244, November 2021. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/78/9/3230/6378054.

Whitlock:2021:TOA

[WPP^{+}21] R. E. Whitlock, T. Pakarinen, S. Palm, M. L. Koljonen, J. Östergren, and J. Dannewitz. Trade-offs among spatio-temporal management actions for a mixed-stock fishery revealed by Bayesian decision analysis. ICES Journal of Marine Science, 78(10):3625–3638, December 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/78/10/3625/6425787.

Waller:2021:CSM

[WRC^{+}21] Jesica D. Waller, Kathleen M. Reardon, Sarah E. Caron, Blaise P. Jenner, Erin L. Summers, and Carl J. Wilson. A comparison of the size at maturity of female American lobsters (Homarus americanus) over three decades and across coastal areas of the Gulf of Maine using ovarian staging. ICES Journal of Marine Science, 78(4):1267–1277, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/4/1267/6154828.

Westgaard:2023:GPS

[WSJ23] Jon-Ivar Westgaard, Guldborg Søvik, and Torild Johansen. Genetic population structure in Norway lobster (Nephrops norvegicus): management regime under panmixia. ICES Journal of Marine Science, 80(4):766–774, May 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/4/766/7023837.

Watson:2023:ERB

[WSL^{+}23] James R. Watson, Claire M. Spillman, L. Richard Little, Alistair J. Hobday, and Phillip S. Levin. Enhancing the resilience of blue foods to climate shocks using insurance.

ICES Journal of Marine Science, 80(10):2457–2469, December 2023. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/80/10/2457/7424415.

Wo:2022:MTA

[WZJ^{+}22] Jia Wo, Chongliang Zhang, Yupeng Ji, Binduo Xu, Ying Xue, and Yiping Ren. A multispecies TAC approach to achieving long-term sustainability in multispecies mixed fisheries. ICES Journal of Marine Science, 79(1):218–229, January 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/1/218/6482878.

Xia:2023:QBR

[XCK^{+}23] Meng Xia, Tom Carruthers, Richard Kindong, Libin Dai, Zhe Geng, Xiaojie Dai, and Feng Wu. Quantifying bycatch risk factors for the Chinese distant water fishery. ICES Journal of Marine Science, 80(3):507–517, April 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/3/507/6414825.

Yadav:2020:KTI

[YABM20] Shreya Yadav, Ameer Abdulla, Ned Bertz, and Alexander Mawyer. King tuna: Indian Ocean trade, offshore fishing, and coral reef resilience in the Maldives archipelago. ICES Journal of Marine Science, 77(1):398–407, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/1/398/5584406.

Yalcin:2023:ELS

[YARE23] Semra Yalcin, Sean C. Anderson, Paul M. Regular, and Philina A. English. Exploring the limits of spatiotemporal and design-based index standardization under reduced survey coverage. ICES Journal of Marine Science, 80(9):2368– 2379, November 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://academic.

oup.com/icesjms/article/80/9/2368/7310990.

Yeung:2020:CVS

[YC20] Cynthia Yeung and Daniel W. Cooper. Contrasting the variability in spatial distribution of two juvenile flatfishes in relation to thermal stanzas in the eastern Bering Sea. ICES Journal of Marine Science, 77(3):953–963, May 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/3/953/5637324.

Yan:2023:SMG

[YCF^{+}23] Yuan Yan, Eva Cantoni, Chris Field, Margaret Treble, Hugues P. Benoît, Rick M. Rideout, and Joanna Mills Flemming. Spatiotemporal modelling of Greenland halibut maturation across the Northwest Atlantic. ICES Journal of Marine Science, 80(6):1787–1801, August 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/6/1787/7225125.

Yasumiishi:2020:DNS

[YFM^{+}20] Ellen M. Yasumiishi, Edward V. Farley, Jacek Maselko, Kerim Y. Aydin, Kelly A. Kearney, Albert J. Hermann, Gregory T. Ruggerone, Katherine G. Howard, and Wesley W. Strasburger. Differential north–south response of juvenile Chinook salmon (Oncorhynchus tshawytscha) marine growth to ecosystem change in the eastern Bering Sea, 1974–2010. ICES Journal of Marine Science, 77(1):216–229, January 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/1/216/5561470.

Yang:2022:CLB

[YLC^{+}22] Zhenyu Yang, Jianping Li, Tao Chen, Yuchun Pu, and Zhenghui Feng. Contrastive learning-based image retrieval for automatic recognition of in situ marine plankton images. ICES Journal of Marine Science, 79(10):2643–2655, December 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/10/2643/6812844.

Young:2020:IFO

[YMAH20] Erik G. Young, Michael C. Melnychuk, Leif E. Anderson, and Ray Hilborn. The importance of fishing opportunity to angler utility analysis in marine recreational fisheries. ICES Journal of Marine Science, 77(6):2344–2353, November 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/77/6/2344/5680028.

Yan:2024:BTS

[YMH^{+}24] Naizheng Yan, Tohru Mukai, Kohei Hasegawa, Jun Yamamoto, and Yoshiaki Fukuda. Broadband target strength of arabesque greenling, Pacific sand lance, and pointhead flounder. ICES Journal of Marine Science, 81(1):195–203, January 2024. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/81/1/195/7469915.

Ma:2021:HSM

[yMlZlL^{+}21] Chao yi Ma, Xiao lu Zhu, Ming ling Liao, Shuang lin Dong, and Yun wei Dong. Heat sensitivity of mariculture species in China. ICES Journal of Marine Science, 78(8):2922–2930, November 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/8/2922/6360555.

Yaragina:2022:BDI

[YSL22] Natalia A. Yaragina, Leif Chr Stige, and Øystein Langangen. Bycatch data from ichthyoplankton surveys reveal longterm trends in gelatinous zooplankton in the Norwegian and Barents Seas. ICES Journal of Marine Science, 79(3):868– 881, April 2022. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL https://academic.oup.com/
icesjms/article/79/3/868/6426104.

Yang:2022:RES

[YY22] Yingxi Yang and Takashi Yamakawa. Re-examination of stock–recruitment relationships: a meta-analysis. ICES Journal of Marine Science, 79(4):1380–1393, May 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/4/1380/6565775.

Zhou:2021:IFS

[ZB21] Can Zhou and Nigel Brothers. Interaction frequency of seabirds with longline fisheries: risk factors and implications for management. ICES Journal of Marine Science, 78(4): 1278–1287, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/78/4/1278/6154494.

Zahner:2024:MSE

[ZB24] Joshua A. Zahner and Trevor A. Branch. Management strategy evaluation of harvest control rules for Pacific herring in Prince William Sound, Alaska. ICES Journal of Marine Science, 81(2):317–333, March 2024. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/81/2/317/7511708.

Zhang:2020:EIS

[ZCX^{+}20] Chongliang Zhang, Yong Chen, Binduo Xu, Ying Xue, and Yiping Ren. Evaluating the influence of spatially varying catchability on multispecies distribution modelling. ICES Journal of Marine Science, 77(5):1841–1853, September 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/
article/77/5/1841/5835269.

Zimmermann:2021:DIM

[ZEM21] Fabian Zimmermann, Katja Enberg, and Marc Mangel. Density-independent mortality at early life stages increases the probability of overlooking an underlying stock–recruitment relationship. ICES Journal of Marine Science, 78(6):2193–2203, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2193/6130109.

Zhou:2022:EGL

[ZHL^{+}22] Shijie Zhou, Trevor Hutton, Yeming Lei, Margaret Miller, Tonya van Der Velde, and Roy Aijun Deng. Estimating growth from length frequency distribution: comparison of ELEFAN and Bayesian approaches for red endeavour prawns (Metapenaeus ensis). ICES Journal of Marine Science, 79(6):1942–1953, August 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/79/
6/1942/6646042.

Zhou:2021:MDR

[Zho21] Can Zhou. Multidecadal daily resolved growth increments reveal climate effect on the growth of a highly migratory shark in the North Atlantic. ICES Journal of Marine Science, 78(7):2496–2505, October 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/7/
2496/6317418.

Zang:2021:SVP

[ZJF^{+}21] Zhengchen Zang, Rubao Ji, Zhixuan Feng, Changsheng Chen, Siqi Li, and Cabell S. Davis. Spatially varying phytoplankton seasonality on the Northwest Atlantic shelf: a model-based assessment of patterns, drivers, and implications. ICES Journal of Marine Science, 78(5):1920–1934, August 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/78/5/1920/6291728.

Zang:2023:EWF

[ZJH^{+}23] Z. Zang, R. Ji, D. R. Hart, D. Jin, C. Chen, Y. Liu, and C. S. Davis. Effects of warming and fishing on Atlantic sea scallop (Placopecten magellanicus) size structure in the Mid-Atlantic rotationally closed areas. ICES Journal of Marine Science, 80(5):1351–1366, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/80/5/1351/7125901.

Zhao:2022:IFF

[ZL22] Youzhu Zhao and Yangfan Li. Impact of fisheries footprint on an early warning indicator of resilience reduction in marine net primary productivity. ICES Journal of Marine Science, 79(10):2741–2751, December 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL https://academic.oup.com/icesjms/
article/79/10/2741/6862926.

Zhou:2020:BHA

[ZMFS20] Shijie Zhou, Sarah Martin, Dan Fu, and Rishi Sharma. A Bayesian hierarchical approach to estimate growth parameters from length data of narrow spread. ICES Journal of Marine Science, 77(2):613–623, March 2020. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/77/2/
613/5675587.

Zuercher:2022:NGB

[ZMMF22] Rachel Zuercher, Nicole Motzer, Rafael A. Magris, and Wesley Flannery. Narrowing the gap between marine spatial planning aspirations and realities. ICES Journal of Marine Science, 79(3):600–608, April 2022. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL https://

academic.oup.com/icesjms/article/79/3/600/6533689.

Zakharov:2021:DSC

[ZMN^{+}21] Denis V. Zakharov, Igor E. Manushin, Tatiana B. Nosova, Natalya A. Strelkova, and Valery A. Pavlov. Diet of snow crab in the Barents Sea and macrozoobenthic communities in its area of distribution. ICES Journal of Marine Science, 78(2):545–556, March 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/78/2/545/5920409.

Zhang:2021:ANS

[ZRW^{+}21] Fan Zhang, Paul M. Regular, Laura Wheeland, Rick M. Rideout, and M. Joanne Morgan. Accounting for nonstationary stock–recruitment relationships in the development of MSY-based reference points. ICES Journal of Marine Science, 78(6):2233–2243, September 2021. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/78/6/
2233/5944179.

Zhou:2023:ICA

[ZYJZ23] Z. Zhou, X. Yang, H. Ji, and Z. Zhu. Improving the classification accuracy of fishes and invertebrates using residual convolutional neural networks. ICES Journal of Marine Science, 80(5):1256–1266, July 2023. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL https://academic.oup.com/icesjms/article/80/
5/1256/7100606.